A Complete Bibliography of Publications in The Computer Journal: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 October 2025 Version 1.63

Title word cross-reference

(k, t) [EMYRV21]. {1, 2, 3} [YYL^{+}20]. ^{∗} [GLXP23, GLX25, ZZC24a]. ^{2} [SMK23]. ^{∗} [RW20]. _{n} [WL23]. a [YZMC23]. C [KRR20]. ℓ_{∞} [CC22]. g [DWLF25b, qLlHG20, YX22, YLW21, ZM21]. h [HZ25, LZM^{+}22]. i∗ [SK22]. K [CLX^{+}24, KK23b, LQCM23, PG22, TZW^{+}24, ADFS24, BZZ23, GCH21, HZ25, LZL^{+}21, LFZ^{+}22, YFA20]. L [SZS23]. m [KA20]. μ^{2} [SLL24]. N [YÖ23, BZZ23, GCH21, HZ25, LFZ^{+}22, SG23, WZ20, WX25]. p [BJ22, HJK20]. Q [LdXZ^{+}21, LdXZ^{+}22, ZZ21]. |(n, k)| [YYL^{+}20].

-Adic [BJ22]. -Anonymity [YFA20]. -Ary [LFZ^{+}22, BZZ23, GCH21, HZ25]. -average [YZMC23]. -Cube [LFZ^{+}22]. -Cubes [WZ20, BZZ23, GCH21, HZ25]. -Enhanced [YYL^{+}20]. -Extra [YX22, YLW21, ZM21]. -faulty-block [HZ25]. -Good-Neighbor [qLlHG20, DWLF25b]. -graph [SZS23]. -Learning [LdXZ^{+}21, LdXZ^{+}22]. -Mean [KRR20]. -means [LZL^{+}21, LQCM23]. -Means-SMOTE-ENN [PG22]. -Metric [EMYRV21]. -nearest [KK23b, TZW^{+}24]. -Norm [CC22]. -Planar [ADFS24]. -Restricted [LZM^{+}22, YYL^{+}20]. -Star [HJK20]. -Subdivision [ZZ21]. -th [WX25]. -wave [KA20].

19 [AASG23, AP22, Ano23b, BY24, ETA22, MST23, MKB23, RJ23, SAK22, TH23, VSS^{+}22]. 1K [CWD^{+}23]. 1K-AES [CWD^{+}23].

2-Interval [PPR23]. 2-Layer [ADFS24]. 256 [LSG^{+}20, YXLS25]. 2D [CC24, MMK22]. 2D&3D [MK20]. 2L [WZG^{+}24]. 2L-LSH [WZG^{+}24].

3-Connectivity [WL23]. 3-Extra [YFC^{+}22]. 32 [JSBV22, JSB24]. 3D [ANA^{+}22, HCZ23, LZ23, MMK22, NCL22, SJC23, SZL22, ZWZW23]. 3G [SJ23].

5G [HHX^{+}23, HXCH24, Hsu24].

6 [YYLX21]. 64 [JSBV22, JSB24].

80 [WLLM23]. 802.11ah [CLH22, HH24]. 802.11ah-based [CLH22].

AAL [SA24]. ABE [LSX^{+}21, SG23, ZKQH24]. Abnormal [CCZ^{+}22, GEZ24]. Abnormalities [Ano24c, RK22, VKM21]. Abort [ZJZ^{+}22]. Abstract [NAD21]. Abstracts [KS24]. Accelerate [MKI20]. Accelerated [ZCH23, MXH^{+}25]. Accelerating [MP22]. Access [ACK^{+}24, BK23, CLH22, PCMPCA^{+}20, SVD^{+}24, WZMZ23, WZL^{+}24, YNZ^{+}22, SLWL25]. Accident [KAMA22]. account [RGB^{+}24]. Accountable [SVD^{+}24, ZWG^{+}20]. Accumulator [SLW24b]. Accumulators [ZYWH23]. Accuracy [MWH^{+}24, SR20, SD20, WTWZ25]. Accurate [HJH^{+}24, PiKT21, SM23, YHX^{+}24, LLS^{+}25]. ACD [DM23]. achieving [PCS25]. Acoustic [LLY^{+}21, SLC^{+}20]. Acquisition [DFS^{+}21a, WG21, WZQ^{+}22, KXLL25]. across [Fat22, LLS^{+}25]. Action [BIZA21, OB21]. Active [SB20]. Activities [HYTG24]. Activity [KKBK24, TYTZSE24]. ADAF [DM23]. ADAF-Adjusted [DM23]. Adaptability [BYYS21]. Adaptation [FS24]. Adaptive [AKI20, AKA^{+}21, BTT^{+}22, HXLX18, HXLX22, HZW21, HZLW24, KK23c, LZL20, MS20, NM23, RD20, SS22d, WWY^{+}20b, XCC20, YWY21, Y^{+}23c, YZN24, Z^{+}23a, ZHL^{+}24a, ZYG23, LZY^{+}25, WY25, YLW^{+}25a, Ano23a]. Adaptively [LZ20b, SG23]. Addition [Per24]. Additional [PMMS22]. Additive [XWH^{+}23]. Address [WLW24]. Adic [BJ22]. adjacent [CHQM25]. Adjusted [DM23]. Adjustment [Gan20, RA22b]. Admission [SLWL25]. Admitting [CGY22]. Adopting [AHI22]. Advanced [Als25, MFHhK21, SRS^{+}23, WPF^{+}25, PCS25]. advancing [LLL^{+}25]. Adversarial [ASPB22, Ano23b, GT22, JMATQ23, Liu24b, PLY^{+}23, SS22a, SAK22, WZH^{+}22, WCC^{+}22, ZZL^{+}22, ZZZ^{+}22, LDFD23]. Adversary [Ala20]. AE [ZLD^{+}20]. AEGIS [JLD22]. AES [CWD^{+}23, ZZW^{+}24a]. AES-Like [ZZW^{+}24a]. AFDX [TBH21]. Affect [WSA22]. Affecting [BÖI22]. Affine [ZYS^{+}23]. Affinity [CGY22]. After [JMR23]. Against [Ala20, LWS^{+}21, LMH^{+}21, LHW21, TZS^{+}23, WLLM23, YWHY20, ZQZ24, ZZC24b, LZ22, LZX^{+}22, LLS^{+}25, YCL^{+}20]. Age [SAR^{+}22a, SAR^{+}22b]. Age-invariant [SAR^{+}22a, SAR^{+}22b]. Agenda [Ang24]. agents [SKE^{+}25]. Aggregate [BRL24, Q^{+}23, YQL24]. Aggregation [Hsu24, MWQ^{+}24, CLY^{+}25, SZZ^{+}25]. Aging [SS22d]. Agricultural [DFS^{+}21a]. Agriculture [SP20a]. AI [Ang24, GT22, WZH^{+}22]. Aided [ASSA21, C^{+}23b, Gök21, KK23b, MJQ23, WDM^{+}23, ML20]. Air [BS24a]. airborne [PL25]. AKGF [DJD24]. Algebra [CKR22, HPV^{+}21]. Algeria [DH21]. Algorithm [ASSA21, AKA^{+}21, BRA21, BGHG22, CP23, CZH^{+}21, CKA21, DFS^{+}21b, DLW24, EC22, EO21, FSN21, GS22, GG22, Hsu24, JB23, JZD21, JDH23, KT24, KA20, KSG22, KKK21, KSKM23, LC20, LYK^{+}20, LdXZ^{+}21, LHL21, LL21, LdXZ^{+}22, LQCM23, LLY^{+}23, LPP21, LZL^{+}21, LRG^{+}24, LLAL22, LL24b, MGB23, MBA23, Nar22, OMG22, PLY^{+}23, Per24, PTH^{+}21, PiKT21, PFT24, RP22a, RFAY22, RG22, Rat20, SRC20, SP20b, SJ22, SRS^{+}23, SR23, SLC^{+}20, SMK23, TYTZSE24, TTPD21, VB21, WCF^{+}22, XZXY25, XQYA23, YSH^{+}22, YLG^{+}24, ZWXL24, DWLF25a, KK23b, LXW^{+}25, QLH24, SLWL25, WY25, ZYL^{+}25]. Algorithm-Based [JB23, MGB23, MBA23, RP22a, Rat20, SP20b]. Algorithm-Switching-Based [LYK^{+}20]. Algorithms [DO22, GC22, GC23, GGŞ23, HYTG24, Kam24, LLGC22, LWL^{+}22, OS23, QGL^{+}22, SZX20, TT24, YCZ^{+}22, YTXW24, YFA20, HNR24, ADJ23]. Alignment [ZYG23, ZWZ^{+}24]. Alliance [ZSQS24]. Allocation [BK23, DY21, JCY^{+}20, JTGJ20, NM23, PP22, LZ25, QSC^{+}24, SGGM21, ZYLW25]. Almost [FGS24, MH21]. Along [LZN^{+}24]. Alternating [GHC21, HLX22, HLCX23, SLL25]. Alternating-update-strategy [SLL25]. Aluminum [CRA23]. Aluminum-Titanium [CRA23]. Always [QXZZ21]. AmBC [ZZL20]. Among [KS24]. Amount [AA20, BÖI22]. amounts [YZC^{+}24]. Amplification [QYZ^{+}21]. Amplified [SJC23, ZQY^{+}22]. Amygdala [SJC23]. Analogy [WBG21]. Analogy-Centered [WBG21]. Analysis [AÖ21, AASG23, AlS23a, ARAA^{+}22, AP22, BRA21, ÇÖİ21, DH21, Fat22, GCR^{+}22, GKK22, GHC21, HJH^{+}24, HCZ^{+}21, HZY21, JZ23, KAMA22, KTK23, LQZ^{+}24, LWS^{+}21, LHLY20, LZLZ24, LHH23, MK20, MKS^{+}22, MPP23, Mou22, NRA23, Noo23, PC22, PGV^{+}22, RJ23, SR20, SDW24, SZH^{+}24, Suc24, SK22, SYC^{+}23, TTCCMR^{+}23, WPY^{+}23, WhLY23, XG22, XZYZ24, YS22, Zar20, ZZL20, ZTK21, AAL24, AAL25, Eld25, KNWT24, LDFD23, Man24a, SGGM21, WZ25, WX25, YJY^{+}24, ZFY^{+}25]. Analysis-Based [Noo23]. Analytics [MKB23, MAAJ24, MP22, SV22b, VSS^{+}22]. Analytics-inspired [MKB23]. Analyzing [uHLH22]. Anchoring [WLZ22]. Android [ErEE20, L^{+}23c, NPF^{+}25, OEAA23, YAHVC20]. Angle [LZ23]. Anisotropic [BSM21]. Anomalies [CLX^{+}22, LL24a, LHH23]. Anomaly [LGC^{+}24, LKA24, PVFP22a, PVFP22b, SB22, WHL22, WGL^{+}22]. Anomaly-Based [LKA24]. Anonymity [H^{+}23c, YFA20]. Anonymization [XCHY24]. Anonymous [FNLW23, MH21, Oks24]. Ant [NCRS22, SS22a]. Antennas [WWY20a]. Anti [LGC^{+}24, SR23, Z^{+}23b, LUW^{+}25]. Anti-Forensic [SR23]. Anti-Occlusion [Z^{+}23b]. Anti-Spoofing [LGC^{+}24, LUW^{+}25]. Antioxidant [GR21]. AODV [KSG20b]. Aperture [BT23b]. API [AAWX24, LCC^{+}24]. APIs [Ala23]. Apollonian [SK25]. Application [Far24, GZY^{+}23, HJH^{+}24, HJ20, JWR24, LZL20, MA20, PS20, RJ23, ZSM24, ZZW^{+}24a, ZCW^{+}24, BMA^{+}25]. Applications [AYW^{+}22, AEGA22, AEGA23, CGY22, CNL^{+}23, ErEE20, GHC20, HLJW22, JZD21, KCLH21, LWX22, MHG22, Oks24, RK21, RMSMAH^{+}20, SZS23, SD23, WtZLS20, WZXX20, XWL23, ZZC24b, XZXY25]. applied [VTZX25]. Applying [BFM22]. Appraisal [TEGB22]. Approach [AKA^{+}22, ASM^{+}21, ASM^{+}22, AKD^{+}21, AS24a, AR21, AEGA22, AEGA23, BFG^{+}21, BS24c, CCC^{+}23, Çal22, CLM^{+}22, DAR22, FFH22, HLML21, HYTG24, HPV^{+}21, HAWA^{+}22, HH22, KGA22, KSG20b, KESZ22, KP20, KS24, LFZ24, LJ23, MHG22, MNN20, MAEK23, NM23, PDO22, PAG^{+}22, PCLZ23, RP22b, RW23a, RMSMAH^{+}20, Sha22, SA22, SSMS22, SB20, TH23, VSS^{+}22, WLZ20, XQYA23, Y^{+}23a, YFA20, ZZC24b, CTL^{+}25, HHL^{+}24, LXW^{+}25]. Approaches [AM23, HZKF23, SCS24]. Approximate [FPT22]. Approximating [CT23, CC22]. Approximation [HNR24]. Arabia [AA23, Fat22]. Arabic [TEGB22]. Arbitrary [SIT20]. Arbitrary-Shaped [SIT20]. Architecture [DTP24, KKM21, LYK^{+}20, PAO20, RP22a, SK24a, SR20, ZCH23, FMX^{+}25, HBI^{+}25]. Architectures [MNN20]. Area [BRA21, JMR23, PMMS22, SB22]. Areas [M^{+}23]. Argumentation [NAD21]. Arguments [YD21]. Arithmetic [HHX^{+}23]. Armed [AV20]. Arming [ÇA25]. Array [HCZ^{+}21]. Arrays [BKS21]. Arrhythmia [Çal22, Eld25]. Artifact [ANA^{+}22]. Artifacts [SRS^{+}23]. Artificial [Ang24, Ano24e, CP23, EO21, Gok22, HH22, KS20a, KKK21, KTK21, LH20, OMG22, SD20]. ARX [HXW22]. Ary [LFZ^{+}22, WZ20, BZZ23, GCH21, HZ25]. ASCON [ZZC24b]. ASCUE [WCC^{+}22]. ASLR [XJWX24]. Aspect [KF24, LHLY20, LH21, WPY^{+}23]. Aspect-Based [WPY^{+}23, LHLY20]. Aspects [SA24]. Asperger [aBWH^{+}22]. Assess [Meg20, PK20, PK22]. Assessment [AlS23a, JMR23, MG22, NCL22, SWFW24, VG23, ZHP21, DWLF25b]. Assets [PAO20]. Assignment [BC22]. Assisted [MM21, SSMS22, YNZ^{+}22, ZY23, SS22b]. Associated [PCMPCA^{+}20]. Association [DMG^{+}23, ETA22]. Assortativity [Meg20]. Assumption [LLGC20, ZWQ^{+}23]. Assumptions [WLW24]. Asymmetric [WZL^{+}24]. Asymptotic [ZY21]. Asynchronous [DLP^{+}21, KT24, ZZW^{+}24b, CL25]. Attack [AGA24, Ano24b, C^{+}23b, DWZS24, JLH20, LMH^{+}21, LRG^{+}24, LSS24, LSG^{+}23, SLL24, SDW24, XWCZ24, ZZW22, ZY23, ZWL^{+}25b, ZQZ24, ZWZ^{+}24, PZL^{+}25, SK25, SWZW25, WJHL25, ZMC25, PL25]. Attack-Oblivious [PL25]. Attacks [AAJ^{+}20, CWD^{+}23, DWGC23, FQZL24, GT22, GHZ24, HCZ23, H^{+}23b, HWW^{+}24, JLD22, LC22, LWS^{+}21, LSG^{+}20, LZX^{+}22, Liu24b, LCZ^{+}22, MG21, SS22c, Sha24, SJ23, TZS^{+}23, WCD21, WLLM23, WZJ^{+}24, WDM^{+}23, XWL23, XZYZ24, YWHY20, YCL^{+}20, YQDJ24, ZLD^{+}20, ZZD^{+}21]. Attempts [BM23]. Attention [HHLL22, JWR24, MST23, SM23, SLW^{+}24a, WPY^{+}23, XZZ^{+}24, ZWM22, BY24, CLY^{+}25, HZS^{+}25, LXP25, LWJ25, LXWD24, ZJLM25, uhLM^{+}25]. Attentions [ZLCW24]. Attestation [CEN24]. Attribute [BG24, CLG^{+}20, LHW21, YX23, ZC23b, CZPL25]. Attribute-Based [CLG^{+}20, LHW21, YX23, ZC23b, CZPL25]. Attributed [LSH^{+}22]. Auction [HWC^{+}24]. Audit [CLWH24, LLT^{+}24]. auditing [HMG25]. Auditor [MHAR20]. Auditory [AMA^{+}24]. augmentation [SSN25]. Augmented [BZZ23, GCH21, WZ20, HHL^{+}24]. Auscultation [Gök21]. Authentic [HAWA^{+}22]. Authenticated [HM23, HHX^{+}23, HXCH24, JZWN23, LZ22, LTT^{+}22, LRG^{+}24, SJHL21]. Authenticating [SLT^{+}22]. Authentication [CHWM21, SKB^{+}22, TLD^{+}24, ZYA^{+}22, ANCLHD25, QXX^{+}25, ZLZZ25]. Authority [ZWG^{+}20]. Authorization [CDR^{+}24, LZG^{+}21b]. Authorship [FYH20]. Autism [GCD21, GCD22]. Auto [CGY22, SII22, BS24b]. auto-encoder [BS24b]. Auto-Estimating [CGY22]. Auto-Scale [SII22]. Autoencoder [YLY24a, WJHL25, SLL25]. AutoEncoders [CLYZ25]. Automata [MMR22, PFT24, SZS23, DCV25]. Automated [KK23b, SSS^{+}22, WCD^{+}24, XJWX24]. Automatic [Ano24c, CLWH24, DJD24, GCD22, HXW22, JMV22, LSG^{+}23, MA20, RK22, RMSMAH^{+}20, SJ23, WZG25]. Automating [ZHL^{+}24b]. Automation [CDCN21, CKR22]. Automaton [SZA24]. Autotuning [TADD23]. Auxiliary [LZ22]. Availability [GQL^{+}20]. average [YZMC23]. AVL [WZQ^{+}23]. Avoidance [CLH22, Hsu25]. Aware [AKD^{+}21, AS24a, BA23b, BK23, BS23, CHWM21, CLH22, Gan20, HH24, LZL20, LZY20, LL24b, LZL^{+}24, LSC^{+}22, MWH^{+}24, NJ21, PFT24, SKB^{+}22, SVD^{+}24, WWW^{+}22, WWSW24, YLY24a, Y^{+}23b, YLX^{+}24, ZYC^{+}24, ZeWSL24, CLY^{+}25, DY21, LDCX25, SGGM21]. Awareness [PJ24]. AWFC [LCZ^{+}22]. Axioms [BT23a].

B [MHG22]. Backdoor [FQZL24, ZQZ24]. Backoff [CLH22, HH24]. Backward [KK23a, L^{+}23e]. Bag [PG22]. Bag-Boost [PG22]. Balance [HH24]. Balanced [GCH20, LWZR24, YX24, PCS25, WX25]. Balancing [BA23b, BK23, CLM^{+}20, HZZ20, LSF25, LWL^{+}22]. Band [WWY^{+}20b]. Bandit [AR21]. Bandits [AV20]. Bandwidth [CCY^{+}24, VD20, VC23]. Bandwidth-Efficient [CCY^{+}24]. Banking [MAEK23]. Bankruptcy [ASM^{+}21]. Barrier [LRG^{+}24]. Base [JWR24]. Based [ASPB22, ACK^{+}24, ANA^{+}22, AASG23, AEZ20, AMA^{+}24, ASSA21, AlS23a, AHI22, ASK^{+}23, AA23, AVA21, Ano24c, Ano24g, Ano24f, AS24b, AEGA22, AEGA23, AA20, Bad24, BDFP23, BA23b, aBWH^{+}22, BGHG22, BS24d, BRL24, BG24, BS23, CCC^{+}23, CRA23, CHT20, CLG^{+}20, CLX^{+}22, CZC22, CLWH24, CLZ^{+}24, CCZ^{+}22, CDCN21, DM23, DFS^{+}21b, DYZY24, DTP24, ETA22, EO21, FZH21, FNLW23, FWC24, FGM21, GK21, GZL^{+}21, GSFS21, GS22, GLY^{+}24, GÖ23, GG22, GK23, HKA24, HLML21, HWS21, HFT22, HYTG24, HDK24, H^{+}23b, HM23, Hsu23, Hsu24, H^{+}23c, HSL^{+}22, HH24, JB23, JWR24, JCY^{+}20, JMR23, JZ23, KS21, K.24, Kal23, KGA22, KSG22, KS20a, KK23b, KYAN21, KKM21, KP20, KRR20, KASV23, KGA23, KC23, KKBK24, LHHW22, LQZ^{+}24, LYK^{+}20, LLG^{+}20, LLGC20, LHL21, LL21, LZQL22, LLGC22, LZM^{+}22, LOGC22, LS23, LLW24, LZN^{+}24, LSQ20, LWM^{+}22, LSW^{+}23]. Based [LLF^{+}24, LL24a, LMH^{+}21, L^{+}23d, LH20, LPP21, LTH21, LHW21, LZL^{+}21, LDC^{+}21, LYW^{+}22, LL22, LRG^{+}24, LWZR24, LLT^{+}24, LGX^{+}24, LZLZ24, LJ23, LKA24, LLS^{+}24, L^{+}23f, LC24b, LFZ^{+}22, MWH^{+}24, MST23, MS20, MYB20, MM20, MJQ23, MRVC23, MHG22, MMK22, MAAJ24, MS23, M^{+}23, MGB23, MBA23, NB21, NCRS22, NLD^{+}23, Noo23, NCL22, PLY^{+}23, PK22, PS20, PCMPCNHR22, Per24, PTH^{+}21, PP22, PFT24, PJ20, QGL^{+}22, QZTZ21, RP22a, RG23, RSSJ23, RFAY22, RCK22, Rat20, RK22, RMW24, RAD20, RKA^{+}22, SS22a, SB22, SK24a, SP20b, SCH^{+}20, SP23, SKA23, SLT^{+}22, SCS24, SJHL21, SM23, SLW^{+}24a, SLW24b, SV22b, SD20, SJC23, SSMS22, SS20, SD23, SZX20, SR23, SJ20, SLY^{+}24, SMK23, TR24, TYTZSE24, TAA23, TJB23, T^{+}23, TLD^{+}24, TKV23, TEHG22, TH23, VKM21, WG21, WQZ^{+}22, WZQ^{+}22, WZH^{+}22, WGL^{+}22, WPY^{+}23, WZQ^{+}23, WZPZ24, WZG^{+}24]. Based [WLW24, WBG21, WZG^{+}20, WCC^{+}22, WCD^{+}24, XDZ22, XQYA23, XZYZ24, YXLZ23, YDS^{+}20, YCZ^{+}22, YTXW24, YHW^{+}24, YX23, YLC^{+}24, YFA20, YX24, Y^{+}23c, YLX^{+}24, YZX^{+}24, YLG^{+}24, ZHYH22, ZHYH23, ZSM24, ZM21, ZJZ^{+}22, ZC23b, ZOLX23, ZWG^{+}20, ZZL20, ZYG23, ZCC^{+}23, ZYWH23, ZYW^{+}20, Z^{+}23b, ZWQ^{+}23, ZZ24, ZWC^{+}24, ZWXL24, ANCLHD25, ASCJ25, BS24b, CZLY21, CGQ^{+}25, CZPL25, CLH22, CTL^{+}25, ÇÖİ21, ÇA25, CKA21, DFS^{+}21a, DWLF25b, GLW^{+}24, GZF25, HZZ20, HY25, HHL^{+}24, Hsu25, HMG25, ICAAFSS25, JMV22, KXLL25, LLW22, LW23b, LLM^{+}24, LSF25, LHLY20, LWL^{+}25, LFDW24, Ma25, MLX^{+}25, MFSH^{+}25, MG22, PZL^{+}25, PFGW25, QLH24, QWC24, RD20, RGB^{+}24, SS22b, SLL25, Sha21, SIT20, SZCY25, TZwSX24, WLZ^{+}25, WLTY25, WZB^{+}25, WJHL25, WG25, WY25, YZS25, YJY^{+}24, YLWL25, ZLM^{+}24, ZJLM25, ZHS^{+}25, ZFY^{+}25, ZBQ^{+}25, ZTT24, Zou25]. Bat [SS22a]. Batch [K^{+}23, YLC^{+}24]. Battering [KZH^{+}22]. Battery [GGŞ23]. Bayes [MAB^{+}24]. Bayesian [AR21, ETA22, MM20, PJ20, RNJRLL^{+}22]. Bayesian-Based [MM20]. BC [LSG^{+}20, LC22]. BCDC [LCF^{+}21]. BCM [YLC^{+}24]. BCube [YFC^{+}22]. BD [PJ20]. BDEA [HMG25]. Be [GGJM21, LHW21]. Beam [CRA23]. Beamforming [WWY20a]. Bee [EO21, KKK21, OMG22]. Beetle [Ano23d, KB22]. Behavior [CLC^{+}19, CLJ^{+}22, GW24, GKK22, LQJ^{+}24, Noo23, PC22]. Behavior-Obfuscation [CLC^{+}19, CLJ^{+}22]. Behavioral [SCH^{+}20]. Behaviors [MHG22]. Behaviour [BJ22, XZYZ24]. Being [BM23]. Belief [BT23b, BGHG22, PXWL22, Sha21]. Benchmark [BXZ22]. Benefit [HWC^{+}24]. BERT [BS24c, LCC^{+}24, SM23]. BERT-BiLSTM [SM23]. BERT-CNN [BS24c]. Best [JZD21, NJ21]. Best-Worst [NJ21]. Better [ASM^{+}21, CPN^{+}21]. Between [LLF^{+}21, SFC^{+}23, YLW21, BMA^{+}25]. Bi [TYTZSE24]. Bi-LSTM [TYTZSE24]. Biased [AM20]. Bidirectional [LH22]. Big [AEGA22, AEGA23, HDK24, NCRS22, RP22a, WQZ^{+}22, SGGM21]. Bigdata [BG24]. Bijective [YZMC23]. Bike [HSL^{+}22]. Bike-Sharing [HSL^{+}22]. Bilateral [WZL^{+}24]. Bilinear [ZYW^{+}20]. BiLSTM [uhLM^{+}25, SM23, SLW^{+}24a]. Binary [Çal22, FGC22, KP20, LW23b, MP22, S^{+}23, TT24, WSA22, GZF25]. Binary-Coded [TT24]. Bind [CEN24]. Binding [WZXX20]. biochips [BG25]. BioElectra [uhLM^{+}25]. BioElectra-BiLSTM-Dual [uhLM^{+}25].

Biogeography [SR23].

Biogeography-Based [SR23]. Biomarkers [KD21]. Biomedical [KGA22, KS24, SKA23]. Biometric [ANG20, Ano24e, Gok22, PCLZ23, WSA22]. Biometrics [TLD^{+}24]. Bird [SHJA24]. Birds [HHLL22]. Bit [XWQ24]. Bitcoin [MQL23, WZH^{+}22, YZW23]. Biterm [ZZC^{+}22]. Black [RP22b, ZWG^{+}20]. Black-Box [ZWG^{+}20]. BLFair [FMX^{+}25]. Blind [CZC22, DST20, H^{+}23a, LLG^{+}20, Nar22, SLC^{+}20, XWY^{+}24, QLH24]. Block [AMT23, BGGBN24, DWGC23, HLJW22, HLC^{+}23, LC22, LW23b, LLCL24, LLS^{+}24, LSH^{+}24, SLL24, YLZ20, HZ25]. Blockchain [ATZ^{+}21, AVA21, AS24b, CLWH24, GCR^{+}22, LLW22, LWL^{+}22, LYW^{+}22, M^{+}23, Oks24, PCMPCNHR22, SS25, T^{+}23, WZH^{+}22, YLC^{+}24, YLZ20, ZHYH22, ZHYH23, ZSM24, ZC23b, ZYC^{+}24, CZPL25, HMG25, ZWL^{+}25a, LLT^{+}24, Alm25]. Blockchain-Based [AVA21, AS24b, CLWH24, LYW^{+}22, M^{+}23, PCMPCNHR22, WZH^{+}22, ZSM24, LLW22, HMG25]. blockchain-enabled [ZWL^{+}25a]. blockchains [LLM^{+}24]. Blog [BS21]. Bloom [KSS^{+}24]. BLS [FHWH24]. Boat [CRRY23]. Body [BRA21, KESZ22, SB22]. Boolean [Sha22, WZL^{+}24, YX23, ZYS^{+}23]. Boolean-Widths [Sha22]. Boost [PG22]. Boosting [HB23, KTK23]. Booth [Per24]. Border [CR22]. Borehole [RMW24]. BORON [LSH^{+}24]. Bot [WLY24]. BotGSL [WLY24]. Both [Ano23c, LWX22, LCC23]. Bound [ZZW^{+}24a]. boundary [YZLL25]. Bounded [LZH^{+}20]. Bower [SHJA24]. Box [LLCL24, ZWG^{+}20, CLL^{+}24, LZX^{+}22]. boxes [HLJW22]. BR [ZNMZ24]. Brain [Ano23d, BTT^{+}22, CP23, CKA21, KB22, NBTB20, RPP24]. Brainstem [AMA^{+}24]. Break [LRG^{+}24]. Breast [AAR^{+}24, CP22, CLZ^{+}24, GK23, RA22b, SRL20, SRLM22, TT24]. Broad [LWM^{+}22].

Broadcast [CDR^{+}24]. broadcasting [KXLL25]. broadcasting-based [KXLL25]. Broken [GLXP23]. Broken-Down [GLXP23]. BSO [KYAN21]. Bubble [GC22]. Buckets [CC24]. Budget [LZY20, QSC^{+}24]. Budget-Feasible [LZY20]. Building [RNJRLL^{+}22, SD20, HTJ25, Tan24]. Buildings [LHL21]. Bumps [SZL22]. Bunching [BPNdQ22]. Bundle [ZNMZ24]. Bus [BPNdQ22]. bypassing [XJWX24]. Byte [JZ23]. Byte2vec [YAHVC20]. Bytecode [OEAA23]. byzantine [SZZ^{+}25, ZWL^{+}25a, CL25, ZZW^{+}24b]. byzantine-robust [SZZ^{+}25, ZWL^{+}25a].

Cache [LYK^{+}20, SV22b]. Cache-Enabled [SV22b]. Caching [Y^{+}23b]. Cactus [LZLZ24]. Cactus-Based [LZLZ24]. CAIMP [DTP24]. CalBehav [SCH^{+}20]. Calculation [LWZR24]. Calculus [BT23a, BT24]. Calendar [SCH^{+}20]. Call [Ala23, DFS^{+}21a, LCC^{+}24]. Camera [ZZJZ20]. Can [ACLA23, SZY^{+}25]. Cancer [ASPB22, AAR^{+}24, CP22, CLZ^{+}24, GG22, GK23, KM23, MRVC23, RA22b, SRL20, SRLM22, SSS^{+}22, TT24]. Cancerous [SSS^{+}22]. Canny [Ano24f, LGX^{+}24]. Capabilities [PMMS22, Alm25]. Capability [LLY^{+}21, WhLY23, ZWLP24]. Capacity [CZZ23]. Capsule [YHW^{+}24]. Captions [PK20]. Capture [LH22]. Card [RFAY22]. Cardinality [BCS22, LZY^{+}25]. Care [Kat23]. Carrier [Hsu24]. Carrying [YCZ^{+}22]. Cartesian [FLPS23, WX25, Yan24]. Case [ÇÖİ21, DFS^{+}21b, Fat22, KF24, MAEK23, Suc24, TTCCMR^{+}23, WZH^{+}22]. Case-Based [DFS^{+}21b]. Cash [H^{+}23a]. Cat [CR22, NCRS22]. Catching [PLY^{+}23]. Categorization [VS21]. Categorized [LFDF24]. Categorizing [Wan21]. Category [KF24]. CATI [CHQM25]. Cauliflower [ZTT24]. Causal [TJB23].

Caviar [JB23]. Caviar-Sunflower [JB23]. Cavity [SSS^{+}22]. Cayley [LX20, RW20, RW23b, XDL23]. CBAM [YHW^{+}24]. CBAM-Capsule [YHW^{+}24]. CCA [MH20, MH21, ZQY^{+}22, ZWQ^{+}23]. CCA-Almost-Full [MH21]. CCA2 [LSX^{+}21]. CCESHP [TJB23]. CD [YLC^{+}24]. CD-BCM [YLC^{+}24]. CDNM [WCD^{+}24]. Cell [FSN21]. Cement [Sha21]. Center [AT24, DFS^{+}21a, DFC^{+}23, HZZ20, LCF^{+}21, Y^{+}23a, YFC^{+}22, DWLF25a, DWLF25b]. Centered [WBG21]. Central [HSL^{+}22]. Centralities [Dan24]. Centrality [BXZ22, JMV22, LZ20a, QZTZ21]. Centrality-based [JMV22]. Centric [Zar20]. centroid [BS24b]. Cerebrospinal [BTT^{+}22]. Certificate [SYC^{+}23, ZYW^{+}20]. Certificate-Based [ZYW^{+}20]. Certificateless [CDR^{+}24, DYZY24, LSY^{+}20, LWS^{+}21, Q^{+}23, RS22, SD23, YQL24, YWHY20]. Certificates [BM23, YLC^{+}24]. Cervical [GG22]. CFAuditChain [LLT^{+}24]. CFS [YZS25]. CFS-COA [YZS25]. CFTP [VD20, VC23]. CGH [TTPD21]. Chain [KSG20b, LTH21, LWZR24, WLZ22, YXZ23]. Chameleon [T^{+}23, LLM^{+}24]. Change [KRR20]. Changes [MYB20]. Channel [ANA^{+}22, HZS^{+}23, JZ23, LWZR24, MM20, RM22, ZeWSL24]. chaos [Eld25]. Chaotic [ANG20, K.24, SR23]. Character [GZL^{+}21, L^{+}23d, LPP21, RKA^{+}22]. Character-Based [L^{+}23d]. Characteristics [MA20, ZZ24]. Characterization [LZH^{+}20]. Characters [DMP^{+}23]. Charging [CHT20, LL24a]. Chaskey [XWL23]. ChatGPT [NPF^{+}25]. ChatGPT-driven [NPF^{+}25]. Checking [DO22, IOS^{+}22, MKI20, PDO22, WCC^{+}22]. Chest [MST23, PXWL22]. Chicken [CKA21, RP22a]. Children [GCD21, GCD22]. China [DFS^{+}21a]. Chinese [FLZ^{+}22, LLF^{+}24, LL23]. Choice [SWJkL20]. Chromatic [WZ23]. Chronic [Gök21]. Chronological [Kal23]. Chronos [ZZW^{+}24b]. Chunk [TKV23]. Cipher [AMT23, DWGC23, HLC^{+}23, JLH20, LC22, LLCL24, LSH^{+}24, MG21, SLL24]. Ciphers [AGA24, Ano24b, DWZS24, HLJW22, HXW22, JSBV22, JSB24, LW23b, LLAL22, LLS^{+}24, WLLM23, ZZW^{+}24a, Zha24]. Ciphertext [CSYY23]. Circuit [B^{+}23, XWW23, ZYS^{+}23]. Circuits [SG23]. Cities [HJ20]. City [AKA^{+}22, HJ20]. Class [KM23, LZM^{+}22, LGC^{+}24, LHH23, PG22, RW23b, WCW^{+}24]. Classes [PPR23]. Classical [LRG^{+}24]. Classically [DCV25]. Classification [ASPB22, ATZ^{+}21, AMA^{+}24, AP22, Ano23d, AAR^{+}24, BA23a, BT23b, Çal22, CP23, CC24, CZH^{+}21, CKA21, ESC24, GLW^{+}24, GG22, GCD21, JMR23, KF24, KK23b, KM23, KB22, KK23c, LHL21, LS23, LL23, LJ23, LH21, MK20, NB21, NCRS22, NBTB20, OMG22, PKLK21, RP22a, SV22a, SHJA24, SP23, SCS24, SDM23, TT24, TH23, WQZ^{+}22, YWTL23, LXWD24, LWZ^{+}25, SSN25, Tan24, TZW^{+}24, uhLM^{+}25, ZWL^{+}25b]. classified [SZCY25]. Classifier [BIZA21, CP23, JB23, LQJ^{+}24, SK24b, uhLM^{+}25]. ClassRoom [PCLZ23]. Clauses [Wil22]. Cleaning [YZX^{+}24, TDZ^{+}24]. click [LXP25]. click-through [LXP25]. Climate [YLX^{+}24]. Climate-Season-Based [YLX^{+}24]. Climatic [BÖI22]. Climbing [LLY^{+}23]. Clinical [SHJA24]. CLIP [CGQ^{+}25]. CLIP-based [CGQ^{+}25]. clipping [WZG25]. Clique [YZW23]. Cliques [S^{+}23]. Closed [PK20, PXWL22]. Closeness [Dan24, WZ23, XLZ24]. Closest [CC22]. Closure [MMR22]. Clothoid [L^{+}23f]. Cloud [ANKZ^{+}22, AT24, AHI22, AS24a, AEGA22, AEGA23, BS23, FQZL24, GS22, GQL^{+}20, JCY^{+}20, Kal23, LZL20, LZG^{+}21b, LZQ^{+}23, LAKL^{+}22, MHAR20, NJ21, PAO20, PP22, PAG^{+}22, QMR^{+}20, RK21, SRC20, SSMS22, SZX20, SK24b, TEHG22, WZMZ23, WZG^{+}24, YSTD24, YMD^{+}24a, YLX^{+}24, LWJ25, SGGM21, Mou22]. cloud-aware [SGGM21]. Cloud-Based [AHI22, AEGA22, AEGA23]. Cloud-Edge [JCY^{+}20]. Clouds [CLM^{+}20, LZQL22, SII22, TADD23, HBI^{+}25]. CLS [DMG^{+}23]. CLS-MMS [DMG^{+}23]. Cluster [KYAN21, KP20, SMK23, ZZL20, ZZLY23, LFDW24, ZYL^{+}25]. Cluster-Based [KYAN21, KP20]. Clustered [LZL^{+}23]. Clustering [aBWH^{+}22, GÖ23, KAMA22, KRR20, KASV23, LQCM23, MFSH^{+}25, PS20, PJ20, RSSJ23, RP22b, VB21, VSS^{+}22, WtZLS20, WZC^{+}21, WCD^{+}24, YFA20, ZWXL24, BS24b, SZCY25, Wan25, ZYL^{+}25]. Clustering-Based [WCD^{+}24, MFSH^{+}25]. Clustering-Evolutionary [aBWH^{+}22]. Clusters [LZL^{+}24]. CM [GLW^{+}24]. CM-UTC [GLW^{+}24]. CMA [SLC^{+}20]. CMT [CZH^{+}21]. CNFs [Wil22]. CNN [BS24c, JWR24, KSN24, MSH22, SK24a, SLW^{+}24a]. CNN-BiLSTM [SLW^{+}24a]. CNN-LSTM [JWR24]. Co [BDFP23, RD20]. CO-OFDM [RD20]. Co-operative [BDFP23]. Co-simulation [BDFP23]. COA [YZS25]. Coalitional [HWC^{+}24]. Coarse [LSC^{+}22]. Coarse-Grained [LSC^{+}22]. Code [CZC22, Thi24, YYLX21, ZWC^{+}24, ASCJ25, NPF^{+}25, WPF^{+}25]. Code-Based [CZC22, ASCJ25]. Coded [TT24]. Codes [BKS22, KP20, S^{+}23, ZLZ22]. Coding [DYZY24, LC24b]. Coding-Based [LC24b]. Cognitive [MM20, QAA^{+}22, RG22]. Coherence [SC22]. Coherent [DMG^{+}23, XZXY25]. Coin [BS24d, YZC^{+}24]. CoinFA [YZC^{+}24]. Cold [JGJ^{+}24, LFDF24, YTXW24, YÖ23]. Cold-Start [LFDF24, YTXW24, YÖ23]. Collaborative [ASSA21, DY21, GGD22, LZL20, PP22, WG25]. Collie [CR22]. Collision [CLH22, L^{+}23f, SWW^{+}22, Hsu25]. collision-avoidance [Hsu25]. Collusive [XZYZ24]. Colony [EO21, KKK21, OMG22]. Colony-Based [EO21]. Color [PS20, XXW^{+}24]. Colored [DLW24]. Combination [MMR22, NCL22, SS24a, SWW^{+}22, XCC20]. Combinatorial [JXZ^{+}22]. Combined [DFS^{+}21b, MAB^{+}24, VG23]. Combining [ANG20, L^{+}23c, LZ23, RM22, XZLZ22, ZLX24, LDFD23, ZMC25]. COMET [XLW22]. Commerce [GÖ23, ZL24b]. Commitment [WZXX20, XWQ24]. Common [BT23a, BT24, FLPS23]. Communication [FNLW23, GAK20, LW23a, LX23, LLW24, SLC^{+}20, WZH^{+}22, WWSW24, YVCC23, ZZJZ20, CPLW25, CL25, SLWL25]. Communication-Aware [WWSW24]. Communication-Efficiency [LW23a]. Communication-Efficient [LX23]. Communications [WWY20a]. Communities [HZY21]. Community [LDW^{+}24, MYB20, MPP23, Nar21, YXLZ23]. Compact [C^{+}23a, JM24, BATN24, ICAAFSS25]. Companies [AÖ21]. Comparative [MPP23, PGV^{+}22, SAR^{+}22a, SAR^{+}22b, SCS24]. Comparison [JCY^{+}20, MMR22, Sha22, SB24, VTZX25, Zou25]. Compatibility [PPR23, MS25]. Competency [RM20]. Competitive [OMG22]. Compiled [ZLCW24]. Compiler [GCG^{+}22, XWW23]. Complete [FGS24, RW20, YX24, WX25]. Completely [WCF^{+}22]. completion [HY25, LWJ25]. Complex [LFZ24, LGFD24, Meg20, QZTZ21, QLZ24, WYXZ22, XTLM24, SZY^{+}25, ZYW25]. Complexity [FLZ^{+}22, YD21, ZWQ^{+}23]. Compliance [SA24]. Component [GCH20, HLCX23, LLF^{+}21, LFZ^{+}22, MHG22, SFC^{+}23, XG22, ZLC^{+}22, WPF^{+}25]. Component-Based [MHG22]. Composite [LQZ^{+}24, SZX20]. Composition [BS23, JDH23]. Compound [LZM^{+}22].

Comprehensive [AAWX24, MK20, RPP24]. Compressed [BKS21, CC24, SLW24b]. Compression [FGM21, GKO^{+}25, K.24, OS23, QGL^{+}22, RG23, Wil22, MFSH^{+}25]. Compressive [K.24, KGA23, RG23, Sha21, Wan25]. Compressor [OS23]. Computation [HLML21, HHX^{+}23, ICAAFSS25, YXLS25]. Computational [AKA^{+}22, GG22, GG23, Hsu23, TTCCMR^{+}23]. Computations [K^{+}23]. Compute [AR21]. Computed [JCG23]. Computer [Gök21, KK23b, TTPD21, XXW^{+}24]. Computer-Aided [Gök21, KK23b]. Computer-Generated [TTPD21, XXW^{+}24]. Computing [AS24a, BK23, GCH20, JCY^{+}20, JTGJ20, LZQ^{+}23, LAKL^{+}22, MHAR20, MFHhK21, NAD21, PAO20, PFT24, SRC20, SK24b, TEHG22, WG21, WWSW24, WYC^{+}23, Ant24, LZ25, TDZ^{+}24]. Concentrating [ARAA^{+}22]. Concept [AlS23a, GCMRGM24, Mou22, RCK22, SCS24]. Concurrence [LH21]. Concurrence-Words [LH21]. Concurrent [TKV23]. Condition [SWW^{+}22, YLZ24]. Conditional [EI21, JMATQ23, qLlHG20, LX20, LCF^{+}21, RW23b, ZM21, ZL24a, GLX25]. Conditioning [LHL21]. Conditions [BÖI22, DWLF25b]. Confidence [Z^{+}23b, ZYL^{+}25]. Confidential [HSC^{+}24, PCMPCNHR22]. confidentiality [ZLZZ25]. Conflicting [SK22]. Confluence [WTWZ25]. Conformance [WCC^{+}22]. Congestion [AKA^{+}22, AKA^{+}21, BA23b, VD20, VC23]. Congestion-Aware [BA23b]. Congestion-Free [VD20]. Congestive [Çal22]. Conjunctions [ZZ23]. Conjunctive [CLG^{+}20, JZWN23, L^{+}23e]. Connected [HDK24, WhLY23, QWC24]. Connectedness [GCH21, WCW^{+}24]. Connection [YZMC23]. Connectivities [GCH20]. Connectivity [BZ22, BZZ23, JQ24, LLF^{+}21, LZM^{+}22, LX20, LFZ^{+}22, SFC^{+}23, WZ20, WL23, XG22, YZMC23, YFC^{+}22, YX22, YYL^{+}20, YLW21, Zar20, ZL24a, ZZLY23, BZ25, HZ25, ZXDL24]. Conquer [GC22, PDO22]. Consecutive [HYZ^{+}20]. Consensus [Sha24, ZZW^{+}24b]. Considering [CSYY23, LSC^{+}22, SA24]. Consistency [TJB23]. Consortium [GCR^{+}22]. Constant [FGC22, HXCH24, S^{+}23]. Constant-Round [HXCH24]. Constant-Size [FGC22]. Constants [SLL24]. Constrained [HXCH24, ZC23a]. Constraints [BCS22, CCC^{+}23, HWC^{+}24, RAD20, YSTD24, SLWL25]. Construct [BM23, WCF^{+}22]. Constructing [CX22, LW23b]. Construction [AGA24, Ano24b, AA20, BS24d, DM23, HHX^{+}23, LW23a, LTT^{+}22, Q^{+}23, SZH^{+}24, TLD^{+}20, WZG^{+}20, YQL24, DWLF25a]. Constructions [AGA24, Ano24b]. Constructors [HCZ^{+}21]. Consumption [LQJ^{+}24, SKB^{+}22, SHL22, WWSW24]. Contact [Nar21, BATN24]. Container [CGY22]. Containers [Liu24b]. Content [Ano24a, DK22, GSFS21, HZY21, KSS^{+}24, PS20, R^{+}23, TR24, ZLL^{+}22, KXLL25]. Content-Based [GSFS21, PS20, TR24]. Contention [BRA21, RS21]. Context [AKD^{+}21, BKS22, BYYS21, FGM21, GZL^{+}21, LSC^{+}22, SKB^{+}22, TGZ^{+}21]. Context-Aware [AKD^{+}21, LSC^{+}22, SKB^{+}22]. Contextualised [KF24]. Continual [HYZH22]. Continuous [LLW24, LTH21, QYZ^{+}21, ZQY^{+}22, ZYW^{+}20, ZWQ^{+}23, ANCLHD25, MLX^{+}25]. Continuous-Time [LTH21]. Contour [LZ23]. Contourlet [WWY^{+}20b]. Contract [ACK^{+}24, PCMPCNHR22, LSF25, TZW^{+}24]. Contract-Based [ACK^{+}24, LSF25]. Contraction [YCZ^{+}22]. Contracts [WLZ22]. Contradiction [PS22a, PS22b]. Contrast [CYH^{+}20, ZL24b, CLYZ25]. Contrastive [XCHY24, ZNMZ24]. Control [AKA^{+}22, ACK^{+}24, AKA^{+}21, BRA21, BDFP23, BHJ20, DO22, PTH^{+}21, SVD^{+}24, WZL^{+}24, ZLCW24, SLWL25, ZZZL25]. Control-Flow [ZLCW24]. Controllable [WZMZ23]. controlled [DCV25]. Controller [HXLX18, HXLX22]. Controllers [TBH21]. Controlling [SSMS22]. Convergence [KSKM23, PMMS22]. Convex [YWZB24, ZYLW25]. Convolution [GGD22, KGA23, LJLQ24, RJ23, ZWY^{+}24]. Convolutional [AAM^{+}22, BTT^{+}22, FZH21, JCG23, KKBK24, OB21, PKLK21, PVFP22a, PVFP22b, RA22a, Rat20, SRLM22, WZQ^{+}22, ASCJ25, LYZ^{+}25, QWC24]. Cooperative [ZYC^{+}24]. Coordinate [BS23, LZZZ21]. Copy [Ano24f, KSG22, LGX^{+}24, YTZ^{+}24]. Copy-Move [KSG22]. Core [TADD23, TK22]. Corona [Ano24d, ZXL^{+}24]. Corporate [ASM^{+}21]. Corpus [CBEBPPZM20, KNWT24]. Correction [Ano23a, Ano23b, Ano23c, Ano24c, Ano24b, Ano24d, Ano24e, Ano24g, Ano24f, Ano24a, DAR22, LLF^{+}24]. Correlated [DMG^{+}23, WW22]. Correlation [GYY^{+}20, WWW^{+}22, WLLM23, ZZC^{+}22, BMA^{+}25]. Correlations [BG24]. Corrigendum [Ano23e, Ano23d, HXLX22, LdXZ^{+}22, PVFP22a, PS22a, SAR^{+}22a, VC23, ZHYH23]. Corset [LW23a]. Cosine [ASPB22]. Cost [GLW^{+}24, LZQL22, WBG21, Y^{+}23b]. Cost- [LZQL22]. Cost-Aware [Y^{+}23b]. Cost-sensitive [GLW^{+}24]. Cotton [DFS^{+}21b]. Counting [FGS24, YSH^{+}22]. Coupled [SZL22]. Couplet [GZL^{+}21]. Coupling [CZZW21]. Cover [BT23b, KCLH21]. Coverage [HCZ^{+}21, WWY20a]. Covering [HCZ^{+}21]. Coverless [TR24]. COVID [AASG23, AP22, Ano23b, BY24, ETA22, MST23, MKB23, SAK22, TH23, VSS^{+}22, RJ23]. COVID-19 [AASG23, AP22, Ano23b, BY24, ETA22, MST23, MKB23, SAK22, TH23, VSS^{+}22, RJ23]. CP [LSX^{+}21, ZKQH24]. CP-ABE [LSX^{+}21, ZKQH24]. CPLM [GK21]. CPLM-Based [GK21]. CPU [ACLA23, DJD24, ZCH23]. CPU-FPGA [DJD24]. cQA [KC23]. CR [PCLZ23]. Crack [SP23, PFGW25]. CRAFT [GZY^{+}23, MA20]. Creation [AAWX24, LZL20]. Credentials [SKB^{+}22]. Credit [RFAY22]. Crick [MMR22]. Crime [AA23, Als25]. Crisis [Ano23b, SAK22]. criteria [TEHG22]. Critical [VG23]. CRLPPM [SZCY25]. Crop [SJ22]. Cross [DTP24, HHLL22, LFDF24, RMSMAH^{+}20, SC22, WLZ22, YLC^{+}24, YXZ23, YWY21, BZ25]. Cross-Architecture [DTP24]. Cross-Chain [WLZ22, YXZ23]. Cross-Domain [LFDF24, SC22, YLC^{+}24]. Cross-Lingual [YWY21]. Cross-Modal [HHLL22]. Cross-Platform [RMSMAH^{+}20]. Crossed [Liu24a]. Crossing [ADFS24, BFG^{+}21]. Crowd [LZY20, Noo23, Rat20]. Crowdsourcing [DY21, LLGC22]. Cryptanalysis [BMV22, CZZ23, C^{+}23b, HY24a, HLC^{+}23, HXW22, LLCL24, LSH^{+}24, WLLM23, ZWZW23, ZY23, ZZC24b]. Cryptographic [CHWM21, ZYWH23, Als25]. Cryptography [PCMPCNHR22]. Cryptology [AM23]. Cryptosystem [GKO^{+}25, YCL^{+}20]. Cryptosystems [WZJ^{+}24]. CSEA [YLX^{+}24]. CSFL [ZYC^{+}24]. CSI [YHX^{+}24]. CT [ASPB22, KM23, TH23]. CT-Based [TH23]. Cube [LFZ^{+}22]. Cubes [KCLH21, qLlHG20, Liu24a, WZ20, BZZ23, BZ25, GCH21, HZ25, ZXDL24]. CubeSat [Alm25]. cubic [LFDW24]. Cuckoo [LdXZ^{+}21, LdXZ^{+}22, LLT^{+}24, RCK22]. Current [ATZ^{+}21]. curriculum [LWZ^{+}25]. Curve [BJ22, JM24, L^{+}23f, PCMPCNHR22]. Curve-Based [L^{+}23f]. Customer [LQCM23, LQJ^{+}24, NRA23]. Customization [YMD24b]. CutTheTail [PiKT21]. CWOA [PAG^{+}22]. CWSOGG [PLY^{+}23]. Cyber [BHJ20, GW24, SS22b, YWTL23, ZHL^{+}24b, HTJ25]. cyber-physical [HTJ25]. Cyberbullying [Bad24, EO21, Fat22]. CyberCIEGE [Sae24]. CyberEyes [FZH21]. Cyberincidents [GCMRGM24]. Cybersecurity [FZH21]. Cycle [BÖI22, KCLH21]. Cycles [Yan24]. Cyclic [ZZC24a].

D2D [GAK20]. DA-VAPR [Gan20]. DaaS [LC24b]. Data [AT24, AVA21, Ano23e, AGZ24, ALS23b, AEGA22, AEGA23, BS24c, BC22, BJ22, CZLY21, CCC^{+}23, CYH^{+}20, CLX^{+}22, CLX^{+}24, ÇÖİ21, CPN^{+}21, DFC^{+}23, GAK20, GKK22, GQL^{+}20, GCD22, HYTG24, HZZ20, HDK24, HLC^{+}23, HJ20, Kam24, KSA20, Kat23, KTK23, LH22, LLW22, LZQL22, L^{+}23b, LLW24, LWL^{+}22, LYW^{+}22, LZQ^{+}23, LJ23, LZL^{+}24, LAKL^{+}22, LCF^{+}21, MRVC23, MMMZ23, MAAJ24, MUA23, MWQ^{+}24, MFHhK21, MM21, Mou22, M^{+}23, MG22, NCRS22, Oks24, PKLK21, PG22, RP22a, RP22b, R^{+}23, SRC20, SV22a, SRL20, SCH^{+}20, SJ22, SLT^{+}22, SCS24, SKB^{+}22, TBH21, TTCCMR^{+}23, TZS^{+}23, TKV23, WQZ^{+}22, WW22, WZMZ23, WLTY25, WDC^{+}23, WCD^{+}24, Y^{+}23a, YFC^{+}22, YNZ^{+}22, YYLX21, Y^{+}23c, YZN24, YZX^{+}24, Z^{+}23a, ZOLX23, ZSQS24, ZHL^{+}24a, ZJ23, ZTK21, ASYLAM25, Als25, BATN24, CTL^{+}25, DWLF25a, DWLF25b, GSL25, HBI^{+}25, HHL^{+}24, HMG25, KXLL25, LSF25, LDFD23, SGGM21, SS25]. data [SZCY25, TDZ^{+}24, WTWZ25, Wan25, ZLZZ25, ZCZ^{+}24, zLB^{+}25]. Data-Based [HDK24]. Database [CBEBPPZM20, QLH24]. Datacenter [BA23b]. Dataset [ASM^{+}22, RK21, SHJA24, ZCW^{+}24]. Datasets [CP22, KZH^{+}22, YÖ23, PCS25]. DCAF [WLZ22]. DCIAExchange [ZLZZ25]. DCT [TCY^{+}20]. DD LMS [SLC^{+}20]. DDoS [WJHL25, ZMC25]. decay [Ma25]. Decentralized [ALS23b, LLW22, ZKQH24, ZSQS24]. Decision [ASK^{+}23, BPNdQ22, B^{+}23, ETA22, ESC24, MAEK23, RJ23, SHL22, TEHG22]. Decision-Making [ETA22, TEHG22]. Declarative [CFR23]. Decoder [GZL^{+}21]. Decoders [XWY^{+}24]. Decoding [ZLX24]. Decomposition [HJK20, KSG22, KSG20a]. deconvolution [SSN25]. Decoupling [FGM21]. Decreasing [FHWH24, XTLM24]. Decryption [CPN^{+}21]. deduplication [HMG25]. Deep [AÖ21, AAJ^{+}20, AAM^{+}22, AAR^{+}24, Bad24, BT23b, CLM^{+}22, CR22, ÇÖİ21, DV22, DAR22, FQZL24, GEZ24, GGD22, GCD22, HZKF23, HYTG24, H^{+}23b, JB23, JZ23, KESZ22, KASV23, KGA23, KKBK24, LPP21, LFZ24, LRP22, LJ23, LLDX23, MSZ^{+}20, MRVC23, MSH22, MGB23, NCRS22, PXWL22, PVFP22a, PVFP22b, PFT24, RA22a, RCK22, Rat20, RKA^{+}22, SP20a, SV22a, SAR^{+}22b, SSS^{+}22, Sha21, SS22d, SSB23, TT24, TH23, VSAR24, WYC^{+}23, XCC20, YS22, ZLZS22, AK24, Als25, Eld25, GSL25, HHL^{+}24, SS25, WLTY25, YLY^{+}24b, ZYL^{+}25, Zha25, ZMC25, SAR^{+}22a]. Deep-Learning [JZ23]. DeepDetect [AAJ^{+}20]. deeply [SZY^{+}25]. DeepSTF [LLDX23]. Deer [MRVC23]. Defending [TZS^{+}23]. defense [XJWX24]. Defenses [GT22]. Defined [GP21, WJHL25, ZMC25]. definition [DCV25]. Defog [SS24b]. Degree [YZMC23]. dehazing [GZ25]. Delaunay [YLG^{+}24]. Delay [YSTD24, SLWL25]. deletion [CTL^{+}25, ZCZ^{+}24]. Delivery [LC24b, MM21]. Demand [HSL^{+}22]. Demirci [LSG^{+}23]. Dendritic [CP23, FSN21]. Dendritic-Squirrel [CP23]. Dengue [SSMS22]. Denial [AAJ^{+}20]. Denoising [MS20, WWY^{+}20b, ZeWSL24]. Density [ADFS24, KAMA22, MW21, RSSJ23, ZWXL24]. Density-Based [RSSJ23]. Deoxys [LC22, LSG^{+}20]. Deoxys-BC [LC22]. Deoxys-BC-256 [LSG^{+}20]. Dependences [RLIK21]. dependency [WG25]. dependency-based [WG25]. Dependent [CZZ23, ZYLW25]. Deployment [LZQL22, LOGC22, QMR^{+}20, ZHL^{+}24b]. Depression [BS24c]. Depth [Gan20, RMW24]. Depthwise [SK24a]. Derived [CSYY23, OMG22]. Deriving [BHJ20, TY23]. Descent [Rat20]. Description [WDC^{+}23]. Descriptors [EI21]. Design [AMT23, PAO20, WLZ20]. Designated [HSC^{+}24]. Designed [DAR22]. Designing [Sae24]. Desired [HWC^{+}24]. DESL [JZD21]. Despeckling [BSM21]. Destination [HZY21]. detect [SZY^{+}25]. Detecting [Ala23, AGZ24, CLX^{+}22, CCZ^{+}22, Fat22, LHH23, QAA^{+}22, XXW^{+}24, XZYZ24, ZYS^{+}23]. Detection [ASPB22, Ano24g, Ano24f, AAJ^{+}20, Bad24, BS24c, BG24, CCC^{+}23, CLM^{+}22, CR22, CLC^{+}19, CLJ^{+}22, DV22, DTP24, FSN21, GEZ24, GCD21, GCD22, HCZ^{+}21, JCG23, KSA20, KF24, KSG20a, KSG22, KS20b, KRR20, KC23, Lee20, L^{+}23c, LYL24, LWM^{+}22, LL24a, LCC^{+}24, Liu24b, LDW^{+}24, LGX^{+}24, LKA24, MMK22, MPP23, OEAA23, PVFP22a, PVFP22b, PTH^{+}21, RA21, RA22a, RFAY22, RPP24, RW23a, RMW24, SB22, SK24a, SCS24, SSS^{+}22, SWW^{+}22, SLW^{+}24a, SE23, SJL20, TAA23, VKM21, VS21, WHL22, WGL^{+}22, WLY24, WCD^{+}24, XWCZ24, YXLZ23, YFW^{+}23, YHW^{+}24, YTZ^{+}24, ZZL20, ZL24b, BG25, LLL^{+}25, LLGP25, NPF^{+}25, PCS25, RGB^{+}24, SS24b, WCLY24, WPF^{+}25, WJHL25, YLW^{+}25a, YZS25, YLW^{+}25b, ZLM^{+}24, ZJLM25, ZHS^{+}25, ZBQ^{+}25, ZYW25, ZMC25, ZSH25]. Detector [HWS21, ZLZ^{+}24]. Determination [BÖI22]. Determine [DWZS24, JLH20, JLD22]. Determining [BÖI22, HSL^{+}22, HH22, MW21]. DEV [YFW^{+}23]. DEV-ETA [YFW^{+}23]. developer [YLWL25]. Developing [Alm25, GR22, RMSMAH^{+}20]. Development [HLML21, KK23c, RM20]. Device [HZS^{+}25]. Devices [AS24a, GGŞ23, ZLZS22, QXX^{+}25]. DGA [YLY^{+}24b]. DHCL [ZNMZ24]. DHCL-BR [ZNMZ24]. Diabetes [HH22, NGS22]. Diabetic [Ano24c, KK23b, RK22]. Diagnosabilities [LZH^{+}20]. Diagnosability [GLXP23, LZC^{+}21, qLlHG20, RW20, RW23b, SZL^{+}20, SFC^{+}23, WZ20, YLW21, ZLC^{+}22, ZZC24a, GLX25]. Diagnosis [ASK^{+}23, Ano24c, aBWH^{+}22, DFS^{+}21b, FWC24, Gök21, GG22, GG23, HLCX23, KK23b, LZL^{+}23, RK22, RA22b, Sar20, SDK^{+}22, SSS^{+}22, WQZ^{+}22, WhLY23, ZTT24, KSN24, SS25]. Diagnostic [AP22, LZC^{+}21]. Diagrams [B^{+}23]. Diamonds [FGS24]. Different [ZLCW24]. Differential [CZLY21, HWW^{+}24, JZD21, LRG^{+}24, LLS^{+}24, MA20, SLL24, WBG21, WDM^{+}23, XWL23, ZZC24b, KSG20a, SZCY25]. Differential-Aided [WDM^{+}23]. Differential-Based [LLS^{+}24]. Differential-Linear [XWL23]. Differential/Linear [JZD21]. Differentially [WL20, WW22, QYH^{+}25, QSC^{+}24]. Differentials [GZY^{+}23, ZZW^{+}24a]. Differentiated [YNZ^{+}22]. Diffusion [BSM21]. digit [AAM^{+}22]. Digital [SR23, SHL22, TZ23, BG25, LWL^{+}25]. Dimension [EMYRV21]. Dimensional [BJ22, BGGBN24, Çal22, CC24, SRLM22, TGZ^{+}21, PFGW25, WZC^{+}21, XZYZ24]. Direction [SIT20]. Direction-based [SIT20]. Directional [YLY24a, LRP22]. Directions [AYW^{+}22]. Disaster [KS20b]. Disasters [SS22b]. Discharge [GGŞ23]. Disclosure [GKK22]. discourse [WY25]. Discover [GC23]. Discovery [NGS22, XXM^{+}22]. Discrete [LDW^{+}24, MKI20, MP22, XLL24]. Discrete-Time [LDW^{+}24, MKI20]. Discriminating [BXZ22]. Discriminative [GYY^{+}20, HZLW24]. Disease [ASK^{+}23, DV22, DFS^{+}21b, Gök21, RJ23, Sar20, ZTT24, KSN24]. Diseases [KM23]. disinformation [YLW^{+}25a]. Disjoint [DWLF25a, KCLH21]. Disjunctive [Ano23c, ÇA23, LCC23]. Disk [CZZW21, YYLX21]. Dissemination [LLGC22]. Dissimilar [CRA23]. Distance [KASV23, LPH21, WKW^{+}24, GPR20]. Distances [AEZ20, ZZ21, ZXZ^{+}23]. Distillation [LL23, ZFY^{+}25]. Distinguisher [CSYY23, YLZ24, ZY21]. Distinguishers [LLS^{+}24, WWZZ24]. Distinguishing [LCC^{+}24]. Distributed [AYS^{+}24, AAJ^{+}20, CvDRV24, DO22, KSKM23, LZQL22, TZS^{+}23, WYC^{+}23, YLZ20, ZLZZ25]. Distributing [LYW^{+}22]. Distribution [LYW^{+}22, TBH21, Y^{+}23c]. District [KK23c]. Diverse [NLD^{+}23]. Diversity [LGFD24]. Divide [GC22, PDO22]. Divide-and-Conquer [GC22]. Division [HLJW22]. DNFs [Wil22]. DNN [DJD24, WCLY24]. DNS [ARDP21, LYW^{+}22, LCX^{+}22]. Document [AASG23, BFM22, JB23, VB21]. Does [AV22]. DolLion [GR22]. Domain [BSM21, CCZ^{+}22, FS24, GÖ23, LFDF24, LDW^{+}24, SC22, SYC^{+}23, YLC^{+}24, ZYG23, YLY^{+}24b]. Dominating [KSKM23]. Domination [ÇA23, LZG21a]. Double [HWC^{+}24, LL21, LTH21, R^{+}23, QYH^{+}25]. Double-Layer [LL21, LTH21]. Double-Lead [R^{+}23]. Doubling [GHZ24].

Down [GLXP23]. downloads [BMA^{+}25]. DPCell [DWLF25a, DWLF25b, LZYT25]. DQcube [ZM21, ZZLY23]. DQN [Hsu23]. DQN-Based [Hsu23]. Dragonfly [HY24b, KRR20, MGB23, MBA23, PK21]. Drawings [BFG^{+}21]. Drift [SCS24]. Drift-Based [SCS24]. Drive [CZZW21]. Driven [ANKZ^{+}22, TTCCMR^{+}23, NPF^{+}25, RA22a]. Driving [HHX^{+}20]. DRL [HHL^{+}24]. DRL-Tomo [HHL^{+}24]. Drone [BDFP23]. Dropout [MSZ^{+}20]. Drought [KS20a]. Drug [GR21, YS22, JYW^{+}25]. DSDV [KSG20b]. DSR [KSG20b]. DT [YLG^{+}24]. Dual [ARDP21, CZZW21, JWR24, SS24a, SLY^{+}24, XWY^{+}24, YLW^{+}25a, ZNMZ24, ZeWSL24, ZLCW24, QLH24, WLZ^{+}25, uhLM^{+}25]. Dual-Channel [ZeWSL24]. Dual-feature [YLW^{+}25a]. Dual-Mode [CZZW21]. Dual-Stack [ARDP21]. Dueling [Hsu23]. DuoSQL [zLB^{+}25]. Duplications [ZLX24]. During [DH21, GGJM21, VD20, SS22b, VC23]. Dynamic [AT24, BATN24, CvDRV24, EI21, GP21, KKM21, KK23c, LJLQ24, L^{+}23e, MNN20, NJ21, PKLK21, Sha24, SKB^{+}22, SJL20, TK22, WLZ22, ZHT22, ZWM22, QWHC25, SKE^{+}25, ZLZ^{+}24, ZYW25, MNN20]. Dynamic-SoS [MNN20]. Dynamically [RSSJ23].

e-Cash [H^{+}23a]. e-Coin [BS24d]. E-Commerce [GÖ23, ZL24b]. E-Learning [AHI22, RW23a]. e-Voting [H^{+}23a]. Eager [BT24]. Early [DM23, HYTG24, MKB23, NGS22, SSS^{+}22]. Earthworm [SJ20]. eBiBa [LLW24]. ECDSA [CCY^{+}24]. eCK [LZ22]. eCK-Secure [LZ22]. Economic [Ano23b, SDW24, SAK22]. Ecosystems [Ano24a, ZLL^{+}22]. Edge [Ano24d, BXZ22, BA23b, BK23, BS23, GCH20, JCY^{+}20, LZL20, LZ20a, MJQ23, TJB23, WYXZ22, WYC^{+}23, YZMC23, Y^{+}23a, Y^{+}23b, ZL24a, ZYA^{+}22, ZXL^{+}24, Ant24, GLX25, LZ25, TDZ^{+}24, Zou25]. Edge-Aided [MJQ23]. Edge-Based [BA23b, Zou25]. Edge-Connectivity [YZMC23]. Edge-enabled [Y^{+}23b]. Edge-Fault-Tolerant [ZL24a]. Edges [GLXP23, JQ24, RW20, RW23b]. Editing [SLT^{+}22]. Editing-Enabled [SLT^{+}22]. Editorial [Ang24, Ano22, Kam24, Man24b, Ant24, Man24a]. Education [LS23]. Educational [MUA23]. EEG [ANA^{+}22, MK20, MMK22, QAA^{+}22, SRS^{+}23]. Effect [LLAL22, TA23]. Effective [BK23, LFZ24, MPL21, OMG22, WtZLS20, SZZ^{+}25, WZG25]. Effectively [CLX^{+}22]. Effectiveness [AHI22]. Effects [KD21, LZ20a, MS25, ZYLW25]. Efficiency [AT24, LZQL22, LWX22, LW23a, YZW23, CL25]. Efficient [AMT23, CLM^{+}22, CZH^{+}21, CCY^{+}24, GAK20, GSL25, IOS^{+}22, JSB24, KSD22, KSS^{+}24, KYAN21, KKM21, Lee20, LC20, LX23, LMH^{+}21, LWL^{+}22, LC24b, MHAR20, MWQ^{+}24, OFMH22, PTH^{+}21, PiKT21, RFAY22, SS22b, SB24, SMK23, WYC^{+}23, XDZ22, YSH^{+}22, YXZ23, ZLZS22, ZJZ^{+}22, ZC23b, ZZW^{+}24b, ZLZ^{+}24, Zha24, JYW^{+}25, Wan25, YZC^{+}24]. effort [YMD24b]. EFTA [MWQ^{+}24]. Elastic [ANKZ^{+}22, zLB^{+}25]. Electric [LL24a]. Electricity [SKB^{+}22]. Electronic [ZHYH22, ZHYH23, ZSQS24]. Electrostatic [WSCL22]. Electrovibration [SZL22]. Element [GR21]. Elements [DFC^{+}23]. Elephant [ZZD^{+}21]. Elliptic [JM24, PCMPCNHR22]. ELSM [HKA24]. Email [FYH20]. EmailDetective [FYH20]. Embedding [Ano24a, HFT22, KCLH21, LL21, LL24b, MPL21, Nar22, ZHT22, ZLL^{+}22, QLH24, YLY^{+}24b]. embeddings [ZYL^{+}25]. Emojis [Wan21]. Emotion [AEZ20, DH21, HZY21, LL23, RW23a, Rat20, LWZ^{+}25]. Emotional [KESZ22]. Emotions [AEZ21, NRA23]. Empirical [Gök21, CHQM25]. Employing [TYTZSE24]. Empowered [AAM^{+}22, AKA^{+}21]. Emulation [YXLZ23]. Enable [LLY^{+}21]. Enabled [Ano24e, GK21, Gok22, Hsu23, NCRS22, PVFP22a, PVFP22b, PP22, RP22b, RM22, SLT^{+}22, SV22b, VB21, Y^{+}23b, ZWL^{+}25a]. Enabling [AKA^{+}22, LHW21, CL25, FMX^{+}25]. Encoder [GZL^{+}21, XWY^{+}24, BS24b]. Encoding [CLL^{+}24, P^{+}23, SS24a]. Encodings [KC23]. Encrypted [GLW^{+}24, KSA20, RK21, WZMZ23, YFW^{+}23, MXH^{+}25]. Encryption [Ala20, CNL^{+}23, CDR^{+}24, CLX^{+}24, FGC22, HYZH22, JZWN23, K.24, LZG^{+}21b, LMH^{+}21, LZ20b, LHW21, LTT^{+}22, LZQ^{+}23, LRG^{+}24, L^{+}23e, MH20, MAAJ24, QYZ^{+}21, WCD21, WZQ^{+}23, WZL^{+}24, XWH^{+}23, X^{+}23, XTW^{+}23, YMD^{+}24a, ZQY^{+}22, ZYW^{+}20, ZWQ^{+}23, Als25, ASCJ25]. End [YSTD24]. End-to-End [YSTD24]. Endorsement [CEN24]. Energy [EC22, ECE20, KYAN21, NM23, PFT24, RG22, SS22b, SMK23, WWSW24, Wan25, YSH^{+}22, YLX^{+}24]. Energy-Aware [PFT24, YLX^{+}24]. Energy-Efficient [KYAN21, YSH^{+}22]. Energy-Throughput [RG22]. enforcement [XLL24, ZZZL25]. Engine [Ang24, SP20b]. English [HAWA^{+}22]. Enhance [ANG20, KGA22]. Enhanced [Ano24a, BKS22, HHLL22, Hsu25, LKA24, NBTB20, WSCL22, WM22, XG22, XXW^{+}24, YX22, YYL^{+}20, ZZL^{+}22, ZLL^{+}22, CHQM25, CLYZ25, JWJ24, KSN24]. Enhancement [AK20, BSM21, CYH^{+}20, HWS21, PK21, LXP25, LDFD23, TZW^{+}24, ZLM^{+}24]. Enhancements [Far24]. Enhancing [AMA^{+}24, ASCJ25, JZ23, KF24, MUA23, RCK22]. ENN [PG22]. Enriching [SC22]. Ensemble [AAR^{+}24, AA20, aBWH^{+}22, BPNdQ22, Çal22, CK21, CP22, GR21, GG22, GK23, KZH^{+}22, L^{+}23c, PG22, RFAY22, SV22a, SRL20, SK24b, TT24, ZWL^{+}25b, ZYL^{+}25]. Ensemble-Based [GG22, GK23, RFAY22]. Ensembles [BIZA21]. Entities [KS24]. Entity [FZH21, GLY^{+}24, HZKF23, WZB^{+}25]. Entropic [RP22b]. Entropy [LZL^{+}21, QZTZ21, XZLZ22, ZWLP24, HMG25]. Enumerating [YCZ^{+}22]. Enumeration [SLY^{+}24]. Environment [BC22, BS23, GS22, Kat23, PAG^{+}22, TEHG22, TDZ^{+}24]. environmental [Alm25]. Environments [HXCH24, MHAR20]. Episode [KK23a]. Epistemic [CvDRV24]. Equality [LSX^{+}21, LZG^{+}21b, LMH^{+}21, YMD^{+}24a]. Equalization [SLC^{+}20]. Equalizer [RD20]. equations [XZXY25]. Equilibrium [AR21, ZCC^{+}23]. equipment [GZ25]. Equivalence [ZYS^{+}23, CLL^{+}24]. Equivalent [KC23]. Eradication [SRS^{+}23]. ERP [YMD24b]. Erratum [AEGA23, CLJ^{+}22]. Error [HB23, Lee20, LLF^{+}24, SK25, ZZ24]. Errors [LSW^{+}23, LLS^{+}25]. Establishment [HXCH24]. Estimate [ErEE20]. Estimating [CGY22, SA22]. Estimation [DM23, KAMA22, L^{+}23b, MM20, PP22, RM22, WKW^{+}24, WBG21, XLLG22, XZLZ22, ZY21, LZY^{+}25, YMD24b]. ETA [YFW^{+}23]. Ethereum [RGB^{+}24, ZSM24, ZWL^{+}25b]. Etiological [ASM^{+}22]. Evacuation [SS22b]. Evaluate [PK22]. Evaluating [FLZ^{+}22, ZZZ^{+}22, ZZC24b]. Evaluation [Ano24g, CK22, CLX^{+}24, DAR22, Gan20, LZL^{+}23, LCF^{+}21, PCMPCA^{+}20, SDK^{+}22, TAA23, YX24, ZWLP24, CHQM25]. Even [Mer20]. Event [KK23a, MHG22, TTCCMR^{+}23, TGZ^{+}21, WCC^{+}22, XLL24]. Event-B-Based [MHG22]. Event-Driven [TTCCMR^{+}23]. Events [GEZ24, HTJ25]. Evidence [HKA24, HZY21].

Evidence-Based [HKA24]. Evolution [LH20, WGL^{+}22, WBG21, KSG20a]. Evolution-Based [WGL^{+}22]. Evolutionary [aBWH^{+}22, YFA20]. Evolving [SCS24, WHL22, WGL^{+}22]. EWA [SJ20]. Exact [Lem24, LPH21, MW21, Sha22]. Exactly [Suc24]. Examples [Ano23c, LCC23]. Exchange [HM23, LZ22, SJHL21, ZKQH24, ZLZZ25]. Exchanged [qLlHG20, LC24a]. Excitation [LYL24]. Expansion [HHC22, LLAL22]. Experience [HXLX18, HXLX22]. Experimental [HCZ^{+}21, Sha22]. Expert [DFS^{+}21a]. Explicit [BM23, ZZW^{+}24a]. exploitation [XJWX24]. Exploiting [KAMA22]. Exploration [DLP^{+}21, WBG21]. Exploratory [KTK23]. Exponential [SR23, ZYLW25]. Exponentiation [BMV22, SZX20]. Exposing [GCG^{+}22]. exposure [HZS^{+}23]. Expression [LRP22, MRVC23, WDC^{+}23, ZCH23]. Expressions [Ano23c, LCC23, ZMWL20]. extended [XZXY25]. Extensions [NAD21]. Extra [LLF^{+}21, YFC^{+}22, YX22, YLW21, ZM21]. Extract [PTH^{+}21]. Extraction [GW24, JMV22, Kat23, KASV23, MST23, Mou22, SP20b, Sar20, SS22d, WCC^{+}22, CLH^{+}25, YZS25]. Extraction-Based [MST23]. Extractive [BFM22]. Extractor [LLGC20, MQL23]. Extreme [RKA^{+}22]. Extremely [HXCH24].

F [SMK23]. F2PQNN [LYZ^{+}25]. Fabric [ZSM24]. Face [AK24, JMATQ23, LGC^{+}24, MJQ23, SAR^{+}22a, SAR^{+}22b, ZZJZ20, LUW^{+}25, XCHY24]. Facebook [ÇÖİ21, Wan21]. Facial [LRP22, SS22d]. Factor [CPN^{+}21, LHH23, QXX^{+}25]. Factoring [WLW24]. Factorization [GLLZ21, L^{+}23d]. Factors [ASM^{+}22, ASYLAM25]. Factual [KTK21].

Failed [AM23, BM23]. Failing [BBD^{+}24]. Failure [Çal22, ZZ24]. Failures [VD20, VC23]. Fair [LLW22, ZKQH24, ZLZZ25]. fairness [WG25]. Fall [HYTG24, ZBQ^{+}25, ZBQ^{+}25]. False [CNL^{+}23]. Families [SLY^{+}24]. Family [L^{+}23c, PJ24]. Fan [SG23]. Fan-In [SG23]. Fan-Out [SG23]. FANETS [KSG20b]. Far [ACLA23]. Fast [AÖ21, HJH^{+}24, LPH21, MSH22, QXZZ21, WLLM23, WLW24, Y^{+}23c, ZCH23, LYZ^{+}25]. Fault [DWLF25b, FWC24, HCZ^{+}21, HLX22, HLCX23, Lee20, LRG^{+}24, LC24a, LFDW24, MWQ^{+}24, SM21, SDK^{+}22, SWFW24, SZL^{+}20, WhLY23, XDL23, ZM21, ZL24a, BG25, CL25, ZZZL25]. Fault-Free [HLX22, ZZZL25]. Fault-Tolerance [SWFW24]. Fault-Tolerant [MWQ^{+}24, XDL23]. Faults [GHC20, LZL^{+}23]. Faulty [DFC^{+}23, JQ24, RW23b, ZLC^{+}22, HZ25]. FCM [KRR20, RA22a]. FDP [QSC^{+}24]. FDP-FL [QSC^{+}24]. Feasible [LZY20]. Feature [Ano24c, CCZ^{+}22, DTP24, EO21, GK21, GCD21, KKK21, KASV23, LXP25, LDC^{+}21, LRP22, MW21, MST23, MP22, Mou22, RK22, SHJA24, SP20b, SS22d, SE23, WHL22, YAHVC20, ZJ23, ZWC^{+}24, ÇA25, LLGP25, TZW^{+}24, WZB^{+}25, YLW^{+}25a, YZS25, ZLZ^{+}24, ZHS^{+}25]. Feature-Based [CCZ^{+}22]. Feature-Evolving [WHL22]. Features [AEZ20, Ano24e, AAR^{+}24, AA20, BIZA21, CSYY23, DV22, Gok22, LZ23, PTH^{+}21, QAA^{+}22, SDM23, WSA22, XZYZ24, HY25, PCS25, YZLL25]. Federated [AKD^{+}21, LW23a, LX23, LL24a, LCZ^{+}22, TZS^{+}23, ZZL^{+}22, Z^{+}23a, ZYC^{+}24, LSF25, QYH^{+}25, QSC^{+}24, SZZ^{+}25, WZG25, ZWL^{+}25a, ZMC25]. FedEVCP [LL24a]. Feedback [AM20, AR21, MKS^{+}22, WSCL22]. Feedback-Guided [AM20]. Few [ZWZ^{+}24]. Few-Shot [ZWZ^{+}24]. Fewer [Mer20]. FHE [AMT23]. FHE-Friendly [AMT23].

Fidelity [ADF22]. Fields [EI21, LSS24, Per24]. Fighting [ETA22]. File [WCD21]. File-injection [WCD21]. Filling [BJ22]. Filter [CLYZ25, GYY^{+}20, GGD22, KSS^{+}24, LLG^{+}20, LLT^{+}24, MS20]. Filter-enhanced [CLYZ25]. Filtering [ANA^{+}22, ÇA25, WG25]. Filters [CBEBPPZM20, WWW^{+}22, QWC24]. FINB [WZB^{+}25]. Find [CRRY23, GZY^{+}23]. Finding [LFZ24]. Fine [HJH^{+}24, LSC^{+}22, SLT^{+}22, YLX^{+}24, WPF^{+}25]. Fine-Grained [HJH^{+}24, LSC^{+}22, SLT^{+}22, YLX^{+}24]. fine-tuned [WPF^{+}25]. Fingerprint [LMMR22, LZX^{+}22]. Fingerprinting [XWCZ24, ZWZ^{+}24, SWZW25]. Finite [CKR22, MMR22, Per24, SZA24, TY23, WCD21]. Firebug [SMK23]. Firefly [KSG22, SV22a]. First [LWS^{+}21, ZZ24]. First-Error [ZZ24]. Fish [ASPB22]. Fixed [XLZ24]. FL [QSC^{+}24]. FLAKE [JMV22]. Flash [BKS22, CZZW21, CTL^{+}25, ZLX24]. Flash-based [CTL^{+}25]. FlashFox [CTL^{+}25]. Flattening [BT23a]. Flattering [SJC23]. FLDS [QYH^{+}25]. Fleet [CT23, SDK^{+}22]. Flexible [WZMZ23, QSC^{+}24, YZC^{+}24]. Flipping [LCZ^{+}22]. Floor [WLZ20]. Flow [BGHG22, Ken20, LJLQ24, Liu24b, LLDX23, PTH^{+}21, TTCCMR^{+}23, YTZ^{+}24, ZWY^{+}24, ZLCW24, MXH^{+}25, MLX^{+}25, PZL^{+}25]. Flow-Based [PTH^{+}21]. Flower [KSN24, RG23]. Flows [ALS23b]. Fluid [BTT^{+}22]. Flush [SS22c]. fMRI [aBWH^{+}22]. fMRI-Based [aBWH^{+}22]. Fog [AS24a, JTGJ20, PFT24, SS22b, SSMS22, XQYA23, YNZ^{+}22, Ant24]. Fog-Assisted [YNZ^{+}22, SS22b]. Fog-Based [XQYA23]. foggy [SS24b]. Folded [BZ22, Liu24a, WL23, ZL24a, BZ25]. follower [Hsu25]. Food [AÖ21]. Footprint [SHL22]. Force [SA22]. Forecast [Ano23b, HJ20, LLDX23, SAK22]. Forecasting [KS20a, MLX^{+}25, ZWL^{+}25a].

Forensic [GCMRGM24, SR23]. Forest [KD21, LHL21, MAB^{+}24, SK24b, SLY^{+}24]. Foreword [AM23]. Forgery [Ano24f, KSG22, LGX^{+}24, SK24a, VS21, KSG20a, YZLL25]. Formal [AlS23a, BDFP23, Mou22]. Formalizing [SDW24]. format [XJWX24]. FormatAEG [XJWX24]. Forms [ZLCW24]. Forward [BG21, JZWN23, KK23a, L^{+}23e, ZYA^{+}22]. Forward-Secure [ZYA^{+}22]. FOX [DWGC23]. FPGA [Ano24g, DJD24, TAA23, ZCH23]. FPGA-Based [Ano24g, TAA23]. FPGA-CPU [ZCH23]. FPGAs [ESC24]. FQ [WL23]. FR [NCL22]. FRACE [ZWL^{+}25b]. Fracterm [BT23a, BT24]. Fractional [CKA21, KGA23]. Fractional-Chicken [CKA21]. Framework [ACK^{+}24, AVA21, Ano24e, GC23, GP21, Gok22, HJH^{+}24, JTGJ20, KS20b, L^{+}23c, LW23a, MH20, MHAR20, NCRS22, RM20, SB22, Sar20, SV22b, SVD^{+}24, TEHG22, WLZ22, YFW^{+}23, YLX^{+}24, Alm25, LWZ^{+}25, MXH^{+}25, SGGM21, SZY^{+}25, Tan24, XJWX24, YLW^{+}25a, YJY^{+}24, ZWL^{+}25a, Zha25]. Frameworks [NAD21, YS22]. Fraud [RFAY22, ZL24b]. Free [HLX22, RS22, VC23, ZC23b, LSQ20, ZZZL25, VD20]. Frequency [FGM21, L^{+}23b, TK22]. Frequent [KK23a]. Friendly [AMT23]. Friends [LLGC22]. Friends-Based [LLGC22]. Front [ZWL^{+}25b]. Front-Running [ZWL^{+}25b]. Fruit [WLLM23]. Fruit-80 [WLLM23]. Fruit-v2 [WLLM23]. Full [MH21]. Fully [dAJS22, SS22d, XWH^{+}23, X^{+}23, QWC24]. Function [WZG^{+}24, YSTD24]. Function-Based [WZG^{+}24]. Functional [ALS23b, B^{+}23, LZ20b, SA24]. Functions [CX22, HYZ^{+}20, LWX22, LZ20b, ZYS^{+}23, ZLCW24]. fungible [LWL^{+}25]. Fusion [Ano24e, Gok22, HZS^{+}23, LQZ^{+}24, LPP21, LDC^{+}21, YLG^{+}24, ZWC^{+}24, LLGP25, ZLZ^{+}24]. Future [AYW^{+}22, SJ22].

Fuzzing [WZPZ24]. Fuzzy [ASK^{+}23, AS24a, DFS^{+}21b, HHB24, HXLX18, HXLX22, JMV22, JMR23, KRR20, KASV23, LLGC20, LWM^{+}22, MQL23, MS20, PJ20, RD20, RP22b, SZA24, SD20, SMK23, VSS^{+}22]. Fuzzy-Logic [AS24a].

Gambling [SYC^{+}23]. Game [DAR22, Sae24, VTZX25, YTXW24]. Games [AR21]. GAN [CLX^{+}22, LUW^{+}25]. GAN-Based [CLX^{+}22]. GANs [MJQ23]. gated [WJHL25]. Gathering [AA23, CBM24, SKE^{+}25]. Gaze [XLLG22]. GBE [TBH21]. GBO [SS25]. GEMMs [LDCX25]. gender [ÇA25]. Gene [MRVC23]. General [CLWH24, SZL^{+}20, WZG^{+}20, YLZ24]. Generalization [LLF^{+}24, LL22]. Generalize [Ano23c, LCC23]. Generalized [JQ24, KSKM23, SWFW24, WGL^{+}22, WL23]. Generated [Ano24a, HZY21, LX20, RW20, Shu23, TTPD21, XXW^{+}24, XDL23, ZLL^{+}22]. Generating [HTJ25]. Generation [DJD24, GC23, GZL^{+}21, HHLL22, Lee20, NLD^{+}23, YWY21, ZMWL20, HHL^{+}24, LWJ25, NPF^{+}25, ZWL^{+}25a]. Generative [ASPB22, Ano23b, JMATQ23, PLY^{+}23, SS22a, SAK22, ZZL^{+}22]. Generic [LTT^{+}22, WBG21, Tan24]. Genesis [ADF22]. Genetic [Hsu24, LQJ^{+}24, PLY^{+}23, SRC20, TT24, LXW^{+}25]. Genetic-XGBoost [LQJ^{+}24]. Genetically [NCL22]. Geo [TJB23]. Geo-Replication [TJB23]. Geolocation [ZLWZ24]. Geometric [CNL^{+}23, SZL22]. Geospatial [LH22]. Gestational [HH22]. Gesture [KESZ22]. Gestures [WSCL22]. GIFT [JZD21, YQDJ24]. GIS [KAMA22]. GitHub [YLWL25]. Glaucoma [SE23]. Global [LXWD24, Nar21, WPY^{+}23, WZC^{+}21]. Global-local [LXWD24]. GNN [GLY^{+}24, YTXW24]. GNN-Based [GLY^{+}24]. Go [ACLA23]. Goal [BYYS21, SK22]. Goal-oriented [BYYS21]. Goals [BYYS21, SK22]. Good [LW23b, qLlHG20, WZ20, DWLF25b]. Google [BMA^{+}25]. GPT [LLL^{+}25]. GPUs [LDCX25]. Grade [PK22]. Gradient [Rat20, WZG25]. Grain [WLLM23]. Grain-Like [WLLM23]. Grained [HJH^{+}24, LSC^{+}22, SLT^{+}22, YLX^{+}24]. Grammar [LZZZ21, WLZ20, Zou25]. Grammatical [LLF^{+}24]. GRAN [ZJLM25]. Graph [AYS^{+}24, FZH21, HPV^{+}21, HLX22, HZLW24, JMV22, LLY^{+}23, L^{+}23a, LJLQ24, LZZZ21, LPH21, LH21, PKLK21, RGB^{+}24, RW20, SWJkL20, WLZ20, WLY24, ZeWSL24, ZL24b, HZS^{+}25, HY25, JWJ24, LXWD24, MLX^{+}25, MS25, SZS23, SLL25, WG25, WX25, YLWL25, ZJLM25, Zou25, SLL25]. Graphs [ADFS24, Ano24d, BFG^{+}21, BBD^{+}24, ÇA23, Dan24, EMYRV21, GHC21, HLCX23, JQ24, JXZ^{+}22, LZG21a, LZ20a, LZZZ21, LPH21, LX20, PPR23, QDZZ20, RW23b, WCF^{+}22, WYXZ22, WhLY23, XLZ24, XDL23, YCZ^{+}22, YX24, ZZ21, ZLCW24, ZYA^{+}22, ZXL^{+}24, B^{+}23, HNR24, ICAAFSS25, MS25]. Gravity [XTLM24]. Gray [BTT^{+}22]. Green [Ang24]. Grey [PP22, VB21, WZ25]. Grid [DLP^{+}21, PPR23, MWQ^{+}24]. Grids [KT24]. GRNet [JWJ24]. Grounded [NAD21]. Group [GHC21, HM23, HXCH24, HLX22, HLCX23, LSH^{+}22, LMH^{+}21, MH21, MQL23, XZYZ24, YLZ20]. Group-Shared [MQL23]. Grouping [HH24]. Groups [LLGC22]. growth [WZ25]. Guess [AM20, DWZS24, JLH20, JLD22]. Guess-And-Determine [JLH20, JLD22]. GUI [ZLWZ24]. Guided [AM20, SSY^{+}25, ZOLX23, LYL24]. Guiding [NB21]. Gustafson [VSS^{+}22].

Hamiltonian

[DFC^{+}23, HY24b, XDL23, Yan24]. Handle [Ano23b, SAK22]. Handling [FFH22, PG22, RG22, SS25, ZYL^{+}25]. Handshake [TLD^{+}20, TLMY21]. Haptic [CDCN21]. Harary [ÇA23]. Hard [CZZW21, Kam24]. Hardness [CRA23, YLZY22]. Hardware [Per24, SWJkL20, SJHL21, VG23, Zha24, WLZ^{+}25, YXLS25]. Hardware/Software [SWJkL20]. Harmony [BFM22]. Harvesting [ECE20]. Hash [BBD^{+}24, LLW24, TJB23, T^{+}23, WZG^{+}24, YLZ20, LLM^{+}24]. Hash-Based [LLW24]. Hashing [Ano24f, C^{+}23a, GEZ24, LZ23, LGX^{+}24, TCY^{+}20, TYY^{+}21, YTL^{+}23, YTZ^{+}24]. Health [CDR^{+}24, Kat23, MM21, MBA23, SS22a]. Healthcare [ACK^{+}24, AS24b, BS24a, CR22, MM21, Oks24, Q^{+}23, YQL24, SS25]. Heapsort [Suc24]. Heart [Çal22]. Heat [MGB23, MBA23]. Heat-Level [MBA23]. Hedged [HYZH22]. Hercules [CZZW21]. Heterogeneous [BC22, CBM24, GS22, LJLQ24, LZL^{+}24, NM23, WWSW24, WDC^{+}23, Y^{+}23a, ZHT22]. HetNets [BK23]. Heuristic [BFG^{+}21, ICAAFSS25, PiKT21, RM22]. Heuristic-based [ICAAFSS25]. Heuristic-Enabled [RM22]. Heuristics [BCS22]. Hidden [GCG^{+}22, HZZ20]. Hiding [CYH^{+}20, Y^{+}23c, YZN24, ZOLX23]. Hierarchical [BK23, BIZA21, GÖ23, LQZ^{+}24, MW21, MST23, QDZZ20, SSMS22, VS21, YDS^{+}20, CLY^{+}25, LFDW24]. Hierarchy [ARDP21, LHLY20]. High [BFG^{+}21, BJ22, GZY^{+}23, MWH^{+}24, WZC^{+}21, WDC^{+}23, YZX^{+}24, ZBQ^{+}25]. High-Accuracy [MWH^{+}24]. High-Dimensional [BJ22, WZC^{+}21]. High-Probability [GZY^{+}23]. Higher [WYXZ22, ZY21]. Higher-Order [WYXZ22]. Highly [WhLY23, ZYW25]. Hill [LLY^{+}23]. hinge [ZZF25]. Hippocampus [SJC23]. Hirak [DH21]. Histogram [PK21, YTZ^{+}24]. Histograms [Y^{+}23c, ZOLX23]. Histograms-Based [Y^{+}23c]. Histopathological [AAR^{+}24]. Hitting [QDZZ20, ZZ21, ZXL^{+}24, Ano24d]. Hoax [KTK21]. Hole [RP22b]. Hologram [TTPD21]. homing [TY23]. Homogeneous [ADF22]. Homomorphic [DYZY24, LWX22, LXY^{+}20, MAAJ24, XWH^{+}23, X^{+}23, XTW^{+}23]. Homomorphism [XWH^{+}23]. Hop [LOGC22]. Hops [AKA^{+}22]. Hotspots [KAMA22]. HotStuff [GHZ24]. HPC [CLM^{+}20, QWHC25]. HPMC [GLX25]. HSDC [DFC^{+}23]. Human [AASG23, BIZA21, HYTG24, KTK21, KKBK24, LH22, TYTZSE24, ZOLX23]. Hunting [KRR20, MRVC23]. Hunting-Based [KRR20, MRVC23]. Hybrid [AKA^{+}22, ANA^{+}22, AEZ20, AEGA22, AEGA23, BS24c, BC22, DLW24, FFH22, GAK20, GR22, GCD21, HAWA^{+}22, JGJ^{+}24, JDH23, KYAN21, LYK^{+}20, LQCM23, LZC^{+}21, MAEK23, MG22, PAG^{+}22, PG22, RM22, SKA23, Sha21, SK24b, WSA22, WhLY23, XQYA23, ZWY^{+}24, ZLX24, GPR20, GLX25, HBI^{+}25, LXW^{+}25, ZHS^{+}25, ZMC25]. Hybrid-based [Sha21]. Hybrid-Convolution [ZWY^{+}24]. Hybridised [KF24]. Hybridized [PK21, RCK22]. Hydrographical [Ken20]. Hyper [ANG20, XDL23, Yan24]. Hyper-Chaotic [ANG20]. Hyper-Hamiltonian [XDL23, Yan24]. Hypercube [LZM^{+}22, LC24a, SWFW24, WL23]. Hypercube-Based [LZM^{+}22]. Hypercubes [BZ22, GCH20, JQ24, S^{+}23, WM22, XG22, YX22, YYL^{+}20, ZLC^{+}22, ZZC24a, ZL24a]. Hypergraph [ZNMZ24]. Hyperledger [ZSM24]. hypotheses [WX25].

I/O [FMX^{+}25]. IaaS [SII22]. IBE [ML20]. icons [BMA^{+}25]. ID [LSQ20, LMH^{+}21]. ID-Based [LSQ20, LMH^{+}21]. Ideation [CK21]. Identification [ASM^{+}22, CK22, DK22, FYH20, JMATQ23, LGFD24, MJQ23, ZZJZ20, ZCW^{+}24, CLY^{+}25, Eld25]. Identify [MKS^{+}22]. Identifying [ASYLAM25, KS24, QZTZ21, QLZ24, XTLM24]. Identity [LHHW22, ML20, WZG^{+}20, YDS^{+}20, ZWG^{+}20, ZWQ^{+}23]. Identity-Based [LHHW22, WZG^{+}20, YDS^{+}20, ZWG^{+}20, ZWQ^{+}23]. IDF [OEAA23]. IDS [SCS24]. IEC [TZ23]. IEEE [CLH22, HH24]. IID [WL20, Z^{+}23a]. IIoT [Hsu23]. Im [CX22]. Image [AlS23a, AP22, Ano24f, AAR^{+}24, BSM21, BS24b, CBEBPPZM20, CHWM21, EI21, GR22, HWS21, HHLL22, HZS^{+}23, HZY21, KS21, KP20, KASV23, KGA23, LZ23, LGX^{+}24, MS20, MJQ23, NB21, NBTB20, NLD^{+}23, OS23, PS20, PK21, RG23, RJ23, RMSMAH^{+}20, SSN25, SR23, TR24, TYY^{+}21, VS21, WWY^{+}20b, XWY^{+}24, XZLZ22, XZZ^{+}24, YZLL25, YTL^{+}23, ZFY^{+}25, ZZH24, CGQ^{+}25, GZ25, ZLM^{+}24]. Image-Based [NB21]. image-text [CGQ^{+}25]. Images [ASPB22, AAM^{+}22, BT23b, CKA21, JCG23, K.24, KM23, KRR20, KK23c, MST23, NGS22, PDRC22, SS24a, SSB23, XXW^{+}24, XWCZ24, ASCJ25, KK23b, LLGP25, XCHY24]. Imagine [Kam24]. Imaging [BTT^{+}22]. ImagIngDev [RMSMAH^{+}20]. Imbalanced [ASM^{+}22, KZH^{+}22, PG22, ZHL^{+}24a]. Immune [CP23]. Impact [AASG23, Meg20]. Impaired [NB21]. Imperialist [OMG22]. Implementation [CLL^{+}24, LW23b, P^{+}23, PCMPCA^{+}20, TBH21, YXLS25]. Implementations [LLCL24, ZY23]. Implicit [BM23]. ImposeSVD [YÖ23]. Imposing [Wil22]. Impossible [ZZW^{+}24a]. Improve [FS24, KSG20b, SLW^{+}24a, XWL23]. Improved [Ano24c, FSN21, GQL^{+}20, H^{+}23c, JDH23, LLG^{+}20, LC22, LLY^{+}23, LQJ^{+}24, LSG^{+}20, LZX^{+}22, LLS^{+}24, LSH^{+}24, MG21, MP22, PS20, PP22, PG22, RK22, RMW24, SWW^{+}22, SM23, SR20, SDM23, SZL22, WCD21, WZ24, YQDJ24, YLW^{+}25b, ZWXL24, TDZ^{+}24, WCLY24, WJHL25]. Improvement [CYH^{+}20, GCR^{+}22, OB21, ZCC^{+}23]. Improving [BIZA21, JZD21, LZQL22, PCLZ23, QWHC25, Sar20, Thi24, WG25, YZW23, ZJ23, WTWZ25]. In-Organization [Sae24]. In/Out [WSCL22]. Incentive [LZY20]. Incident [HDK24]. inclination [CLY^{+}25]. inclination-aware [CLY^{+}25]. Incompleteness [BT23a]. Incorporating [ZZC^{+}22]. Incorrectly [Shu23]. Increasing [MK20]. Incremental [GÖ23]. Incrementing [YÖ23]. IND-CCA [MH20]. Independence [SB24]. Independent [KT24, WCF^{+}22]. Index [HLML21, Meg20]. Indexing [KGA22, KP20, KASV23, WZG^{+}24]. India [BA23a]. Indices [BJ22]. Indistinguishable [XWW23]. Individual [PC22]. Indoor [WWY20a]. Induced [S^{+}23]. Industrial [GCR^{+}22]. Industry [NRA23]. Infection [SSMS22]. Inference [ASK^{+}23, RD20, ZZZ^{+}22, CHQM25, ÇA25, GZF25, LYZ^{+}25]. Inference-Based [ASK^{+}23]. Influence [PiKT21, WZ24]. Influencers [LFZ24]. Influential [QZTZ21, QLZ24, XTLM24]. Information [AA20, CP23, GSFS21, HZS^{+}23, KGA22, L^{+}23c, LZN^{+}24, LSY^{+}20, LH20, LTH21, LZL^{+}21, LGFD24, LSC^{+}22, MPL21, SKA23, ZHT22, PZL^{+}25, PFGW25]. Information-Theoretic [HZS^{+}23]. Infrastructure [AS24b, QMR^{+}20]. Infrastructures [CLX^{+}22]. Initiative [SSB23]. injection [WCD21]. Ink [FLZ^{+}22]. Inpainting [RCK22]. Input [LZ22]. Inscription [DMP^{+}23]. Insertion [GC22, WLZ^{+}25]. insertions [BATN24]. Insider [LMH^{+}21]. Insightful [AM23]. Inspection [RSSJ23]. inspired [MKB23]. Instance [GW24]. Instantiation [LTT^{+}22]. instruction [CHQM25]. Integer [BC22]. Integral [WWZZ24]. Integrated [KSG20b, RM20]. Integrating [BYYS21, LH22, LGFD24, LXW^{+}25, Wan25, YZLL25]. integration [WZB^{+}25]. Intelligence [AA23, Ang24, GW24, HH22, KS20a, KTK21, YWTL23, YJY^{+}24]. Intelligence-Based [KS20a]. Intelligent [AKA^{+}22, BS24a, CZZW21, HAWA^{+}22, LKA24, LCZ^{+}22, OS23, RNJRLL^{+}22, SAK22, WG21, Alm25, Ano23b]. Intensional [CFR23]. Intensities [AEZ21]. intent [SZY^{+}25]. Interaction [HCZ^{+}21, WPY^{+}23, JYW^{+}25]. Interactions [LH22, WYXZ22]. Interactive [TADD23, WSCL22, WZL^{+}24]. Interconnection [GLXP23, SM21]. Interest [BA23a, KSS^{+}24]. Interesting [PJ20]. Interfaces [CDCN21]. Interference [WWY20a, ZYG23]. Intermediate [ZWC^{+}24]. Intermittent [FWC24, SZL^{+}20, ASCJ25]. internal [HNR24]. International [HHC22]. internet [Ant24, AYW^{+}22, Far20, GCR^{+}22, HAWA^{+}22, L^{+}23f, M^{+}23, MBA23, SD23, TTCCMR^{+}23, TZwSX24, WCC^{+}22, YZS25, YX23, YNZ^{+}22, ZZL^{+}22]. Internet-of-Things [GCR^{+}22, M^{+}23]. Interopera [YXZ23]. Interpolation [ANA^{+}22, MS20, ZZD^{+}21, VTZX25]. Interpretable [YFW^{+}23]. Interrelation [AA20]. Intersection [L^{+}23f]. Interval [PPR23]. Interview [KESZ22]. Intra [RLIK21]. Intra-Tile [RLIK21]. Intrusion [CLM^{+}22, CR22, FSN21, KSA20, Liu24b, SCS24, SLW^{+}24a, YHW^{+}24, ZHS^{+}25, WCLY24, YZS25, ZJLM25]. Invariant [C^{+}23a, LRP22, SAR^{+}22a, SAR^{+}22b]. Inventory [HSL^{+}22]. Inversion [JM24]. Investigation [GCMRGM24, PZL^{+}25]. Ion [GGŞ23]. IoT [AVA21, AS24a, Ano23e, AS24b, ALS23b, AKA^{+}21, BG24, CLM^{+}22, CR22, CLX^{+}22, CLH22, DTP24, GT22, GGŞ23, GCMRGM24, HHX^{+}23, HH24, Kat23, LCZ^{+}22, MWH^{+}24, MKB23, MMMZ23, MGB23, Noo23, PTH^{+}21, SS22a, SSMS22, SJ20, WZPZ24, XQYA23, ZLZS22, uHLH22]. IoT-Based [AS24b, SSMS22]. IP [ZLWZ24]. IPOD2 [ZCZ^{+}24]. IPv4 [AAL25]. IPv4-as-a-Service [AAL25]. IPv6 [AAL24, AAL25]. irrecoverable [ZCZ^{+}24]. Irreducible [SZA24]. Irreversible [LMMR22]. Irrigation [BÖI22]. ISO [TZ23]. ISO/IEC [TZ23]. Isogeny [BBD^{+}24]. Issue [AM23, GT22]. Issues [ATZ^{+}21, Ano23e, MMMZ23]. Item [ADF22]. ITÖ [DLW24].

jailbreaking [SZY^{+}25]. Japanese

[WZB^{+}25]. Jaro [KASV23]. Java [CFR23]. Jaya [Ano24c, RK22, VB21]. Jaya-Based [Ano24c, RK22]. jitter [SLWL25]. Job [KESZ22, TA23]. join [HPV^{+}21]. Joining [Mou22]. Joint [HHX^{+}23, LZ25, Z^{+}23b, ZWZ^{+}24]. Jr [ZLD^{+}20]. JSetL [CFR23]. Jumping [MMR22].

K2 [MG21]. Karatsuba [Per24]. Keccak [WDM^{+}23]. KEFSAR [MWH^{+}24]. Kernel [DJD24, KAMA22, MS20]. Kernels [OFMH22]. Kessel [VSS^{+}22]. Ketje [ZLD^{+}20]. Key [Ala20, AS24b, CZZ23, CEN24, FGC22, HWW^{+}24, HM23, HXCH24, HYZH22, Kat23, LZ22, LZG^{+}21b, LTT^{+}22, LGFD24, LLAL22, LLS^{+}24, MH20, MG21, QYZ^{+}21, SJHL21, WZPZ24, WZJ^{+}24, X^{+}23, YMD^{+}24a, YQDJ24, ZQY^{+}22, ZZW22, ZLD^{+}20, JZWN23, XWH^{+}23]. Key-Dependent [CZZ23]. Key-Expansion [LLAL22]. Key-Recovery [ZLD^{+}20]. Key-Value-Based [WZPZ24]. Keyframe [LYL24]. Keyframe-guided [LYL24]. Keys [CEN24, CRRY23, Shu23]. keystroke [PFGW25]. Keyword [CLG^{+}20, JMV22, JZWN23, LTT^{+}22, LZQ^{+}23, MH20, YX23, CZPL25]. keywords [GZF25]. KGC [LWS^{+}21, YWHY20]. Khatri [Kal23]. Kind [LX20]. Knee [SA22]. Knowledge [CCY^{+}24, DFS^{+}21a, Sar20, SLW24b, SZH^{+}24, WZXX20, WG21, YD21, ZZC^{+}22, ZeWSL24, HY25, YLWL25, ZLZZ25]. Knowledge-Aware [ZeWSL24]. Known [LSY^{+}20]. KORGAN [KKM21]. KP [SG23]. KP-ABE [SG23]. KPCA [NBTB20]. Kravatte [ZZD^{+}21]. Kurdish [Bad24]. KVFL [WZPZ24].

Label [KF24, LCZ^{+}22, PJ24, LWZ^{+}25]. Laceability [XDL23, Yan24]. Lai [LW23b]. LAM [Ala23]. LAMMPS [QWHC25]. Land [BT23b]. Landsat [BT23b]. Landslide [MG22]. Landslide-Occurring [MG22]. Language [Bad24, BS21, HAWA^{+}22, KF24, LLL^{+}25]. Languages [RKA^{+}22]. Laplacian [Ano23a, HZS^{+}25, HZW21]. Laptops [LLY^{+}21]. Large [BRL24, CLX^{+}22, FGS24, HLJW22, LSX^{+}21, LLW24, LPH21, LSS24, TADD23, YXLZ23, ZLC^{+}22, ZCW^{+}24, LLL^{+}25, QWHC25]. Large-Scale [CLX^{+}22, LLW24, YXLZ23, ZLC^{+}22, ZCW^{+}24, QWHC25]. Laser [CRA23]. Last [LYK^{+}20]. Last-Level [LYK^{+}20]. Latency [LZL20, LW23b, ZBQ^{+}25]. Latency-Aware [LZL20]. Latent [EI21]. Latent-Dynamic [EI21]. Lattice [DYZY24, H^{+}23c, KSD22, MAAJ24, SLW24b, ZJZ^{+}22, ZYWH23]. Lattice-Based [H^{+}23c, MAAJ24, SLW24b, ZJZ^{+}22, ZYWH23]. Lattices [BM23, BRL24, LXY^{+}20, LAKL^{+}22]. Law [JXZ^{+}22]. Laws [CFRS24]. Layer [ADFS24, Far24, LL21, LTH21, MM20, SK24a, SR20, ZLC^{+}25]. layers [LUW^{+}25, QWC24]. LDPC [ZLX24]. Lead [R^{+}23]. leader [Hsu25]. Leadership [KKK21]. Leading [Ala23]. Leads [PDO22]. Leads-to [PDO22]. Leakage [Ala20, HYZ^{+}20, HYZH22, LZ22, LSQ20, LZX^{+}22, QYZ^{+}21, XWW23, ZQY^{+}22, ZYW^{+}20, ZWQ^{+}23]. Leakage-Amplified [ZQY^{+}22]. Leakage-free [LSQ20]. Leakage-Resilient [HYZH22, XWW23, ZYW^{+}20, ZWQ^{+}23]. Learned [Shu23]. Learners [SRL20]. Learning [AÖ21, AASG23, ASSA21, AHI22, ASM^{+}22, AKD^{+}21, Ang24, Ano23a, AAJ^{+}20, AKA^{+}21, AEGA22, AEGA23, Bad24, CLM^{+}22, CBEBPPZM20, CK21, CK22, ÇÖİ21, FFH22, FS24, GW24, GGŞ23, GK23, GCD22, HZKF23, HYTG24, H^{+}23b, HHC22, HZW21, HZLW24, JB23, JZ23, Kat23, KTK23, KZH^{+}22, LdXZ^{+}21, LL21, LdXZ^{+}22, LW23a, LCC23, LX23, LWM^{+}22, LSW^{+}23, LL24a, LPP21, LFZ24, LRP22, LJ23, LH21, LCZ^{+}22, Ma25, MS23, MSH22, NLD^{+}23, PK22, PFT24, PDRC22, PGV^{+}22, PCLZ23, RCK22, RW23a, RKA^{+}22, SAR^{+}22a, SAR^{+}22b, SCH^{+}20, SDK^{+}22, SSS^{+}22, SS22d, SR20, SA22, SB20, SSB23, TK22, TZS^{+}23, TT24, TA23, TEGB22, TH23, VSAR24, WQZ^{+}22, WWW^{+}22, WLY24, WYC^{+}23, XCC20, XCHY24, XXW^{+}24, XZL^{+}24, YLZY22, YS22, YZX^{+}24, Z^{+}23a, ZYC^{+}24, ZNMZ24, ZHL^{+}24a, ZWL^{+}25b, ZL24b, AK24, Als25, Eld25, GSL25, HHL^{+}24, JWJ24, LSF25, LWZ^{+}25, LDFD23, LXW^{+}25]. learning [NPF^{+}25, QYH^{+}25, QSC^{+}24, SS25, SZZ^{+}25, WLTY25, WZG25, YLY^{+}24b, YJY^{+}24, ZJLM25, ZYLW25, ZWL^{+}25a, Zha25, ZMC25, Ano23c]. Learning-Based [ASSA21, Bad24, LL24a, PK22, SCH^{+}20, YZX^{+}24, ÇÖİ21, HHL^{+}24]. LED [ZZJZ20]. LEDet [HWS21]. Lemma [ADFS24]. Lemuria [SJ22]. Lesion [TH23]. Lesions [SSS^{+}22]. Level [ASPB22, CKA21, LYK^{+}20, MGB23, MBA23, OEAA23, PK22, PJ20, RS22, RLIK21, SP23, WDC^{+}23, DO22, Tan24]. Levels [DAR22]. Leveraging [FS24, LLY^{+}21]. LFA [LZX^{+}22]. LH [RAD20]. Li [GGŞ23]. Li-Ion [GGŞ23]. Life [BÖI22]. Lifelong [RM20]. Lifetime [CHT20]. Light [HWS21, SB22, SK24a, SLWL25]. Light-Weight [SB22]. Lightweight [HHX^{+}23, HXCH24, LZG^{+}21b, Zha24, ZTT24, ZSH25]. Like [RKA^{+}22, WLLM23, ZZW^{+}24a, LLAL22, ZSH25]. Line [HKA24, JQ24, PXWL22, WCF^{+}22]. Linear [BC22, CZZ23, HWW^{+}24, JZD21, LSH^{+}24, SJ23, SB24, XWL23, ZY23, ZZC24b, XZXY25]. Linearly [DYZY24, GHC20, LXY^{+}20]. Lingual [YWY21]. Link [LZN^{+}24, LL22, MYB20, Meg20, MPL21, VD20, XLZ24, ZC23a, VC23]. Linkable [LSW^{+}23, TLD^{+}20]. Linkage [GÖ23, LLS^{+}25]. Lion [SS22a]. Listing [FGS24]. LiteMixer [ZTT24]. Literacy [PK20]. Literature [Ano23e, MMMZ23, uhLM^{+}25]. Liveness [GHZ24]. Liver [KM23]. LLMs [SZY^{+}25]. Load [AK20, BA23b, BK23, HZZ20, HH24, LWL^{+}22]. Loading [Y^{+}23b]. Lobe [MST23]. Local [Ano24e, BTT^{+}22, Çal22, GS22, Gok22, LC20, LZN^{+}24, LRP22, MH21, RW23b, WZC^{+}21, YZW23, LXWD24]. Locality [WZG^{+}24]. Locality-Sensitive [WZG^{+}24]. Localization [RA21, RA22a, VKM21, YHX^{+}24, ZWC^{+}24, TDZ^{+}24, YZLL25, ZFY^{+}25]. Localized [XTLM24]. Locally [KCLH21, qLlHG20]. Location [Lee20, LLG^{+}20, R^{+}23, YYLX21, LLS^{+}25, SZCY25]. Location-Based [LLG^{+}20]. Log [SB24]. Log-Linear [SB24]. Logarithmic [GCD22]. Logic [AS24a, BDFP23, CvDRV24, DFS^{+}21b, HXLX18, HXLX22, JMR23, SD20]. Logic-Based [BDFP23, JMR23, SD20]. Logistic [CLZ^{+}24, XLLG22]. Lollipop [Dan24]. London [WKW^{+}24]. Long [AAM^{+}22, CGY22, Çal22, GYY^{+}20, LH22, MGB23]. Long-Running [CGY22]. Long-Short [LH22]. Long-Time [GYY^{+}20]. Longest [FLPS23]. Lookback [AM20]. Lookback-Guess-Next [AM20]. Loops [RLIK21]. Loss [PCLZ23, ZZF25]. Losses [AM23]. Lossy [HYZ^{+}20]. Lost [GCG^{+}22, Shu23]. Low [HWS21, HLC^{+}23, LW23b, PK20, ZJ23, HMG25, ZBQ^{+}25]. Low-Data [HLC^{+}23]. low-entropy [HMG25]. Low-Latency [LW23b]. Low-Light [HWS21]. Low-Sample-Size [ZJ23]. Lower [YD21]. LPN [LLGC20, LSS24]. LraSched [CGY22]. LSH [WZG^{+}24]. LSTM [Ano24g, BY24, JWR24, LCC^{+}24, TYTZSE24, TAA23]. LTL [IOS^{+}22]. LUISA [FGM21]. Luminous [KT24]. Lung [ASPB22, JCG23, MST23, PXWL22, KSN24]. LWE [HM23, YCL^{+}20].

MAC [MM20, RS21, ZYG23]. Machine [AASG23, AT24, ASM^{+}22, Ang24, Ano23a, AKA^{+}21, AEGA22, AEGA23, aBWH^{+}22, CBEBPPZM20, CK21, GGŞ23, GK23, HHC22, HZW21, Kat23, KTK23, LHH23, MS23, NLD^{+}23, PK22, PDRC22, PGV^{+}22, RKA^{+}22, SCH^{+}20, SDK^{+}22, SR20, SA22, TK22, TT24, TEGB22, WSA22, YS22, NPF^{+}25, ZZF25]. Machine-Based [RKA^{+}22]. Machines [CFRS24, TY23]. MACs [LWX22]. MADRL [L^{+}23d, ZCC^{+}23]. Magnetic [BTT^{+}22, GGD22, KGA23]. Main [LYK^{+}20]. Majority [AA20]. Majority-Voting [AA20]. Making [ETA22, TEHG22]. Malicious [CCZ^{+}22, LWS^{+}21, Sha24, YWHY20, YFW^{+}23, YZX^{+}24, SZY^{+}25]. Malicious-But-Passive [LWS^{+}21, YWHY20]. Malignant [ASM^{+}22, GG23]. Malleable [YD21]. Malware [CLC^{+}19, CLJ^{+}22, DTP24, L^{+}23c, OEAA23, PJ24, PTH^{+}21, YAHVC20, NPF^{+}25, WPF^{+}25]. Management [AK20, KS20b, Noo23, RM20, WG21, YLC^{+}24, SZCY25, zLB^{+}25]. Managers [KKK21]. MANETs [KXLL25]. Manually [DAR22]. Many [AV22, GHC20]. Map [ANG20, NCRS22, YTL^{+}23, YLG^{+}24, AAL24]. MAP-T [AAL24]. MAPBFT [WY25]. Mapping [AM20]. MapReduce [SP20a]. Marine [PAO20]. Market [GK21]. Markets [Ano23b, SAK22]. Markov [HZZ20, LTH21, MKI20]. Mashup [AAWX24]. Masking [LZX^{+}22, Zha24, LDFD23]. Mass [TT24, Wan21]. Massey [LW23b]. Massive [RM22]. Master [LYW^{+}22]. Match [PJ20]. Matching [AYS^{+}24, DMP^{+}23, FPT22, HHB24, LZG21a, RCK22, WM22, WZ23, XDZ22, ZCH23, ZLZ22]. Material [RSSJ23]. Matrices [LW23b]. Matrix [GLLZ21, GLW^{+}24, WPY^{+}23]. Matsui [JZD21]. Matter [BTT^{+}22]. Matting [SSY^{+}25]. Maximization [HHX^{+}20, PiKT21, WZ24]. maximizing [CZLY21]. Maximum [DLW24, KT24, OFMH22, S^{+}23, ZZZL25, HNR24]. MaxSAT [LC20]. MBSO [RG22]. mCityPASS [PCMPCA^{+}20]. MD5 [SWW^{+}22]. MDER [TGZ^{+}21]. Me [ZZJZ20]. Meadows [BT23a, BT24]. Meal [HH22]. Mean [HB23, KRR20]. Means [LQCM23, PG22, LZL^{+}21]. Measurements [YHX^{+}24]. Measures [PS22a, PS22b]. Measuring [FLZ^{+}22]. MEC [Hsu23]. MEC-Enabled [Hsu23]. Mechanical [SZL22]. Mechanism [CZLY21, GHZ24, HH24, JWR24, LZY20, SM23, SLW^{+}24a]. Mechanisms [Lee20, RPP24, BY24]. Media [BS24c, CLH22, DK22, Fat22, LH22, MPP23, NRA23, SV22b, TGZ^{+}21, Wan21, Als25]. Medical [AP22, M^{+}23, PDRC22, Q^{+}23, RP22b, SD23, WQZ^{+}22, YQL24, ZSQS24, ASCJ25]. Medium [HJ20]. Medium-Sized [HJ20]. Meet [DWGC23, LC22, LSG^{+}20, LSG^{+}23]. Meet-In-The-Middle [LSG^{+}23, DWGC23, LC22, LSG^{+}20]. Meeting [CBM24]. Mellitus [HH22].

Membership

[HXCH24, WLW24, ZZZ^{+}22, ÇA25].

Membership-Authenticated [HXCH24]. Memory [Çal22, CZZW21, LH22, LYK^{+}20, MGB23, XWQ24]. MENA [ARAA^{+}22]. Menger [GCH21, JQ24, WCW^{+}24, ZL24a]. Menger-Type [JQ24]. Merging [HKA24, Mer20]. Mesh [NCL22, KGA22]. MeSH-Based [KGA22]. Mesoscopic [Zar20]. Mesothelioma [ASM^{+}22, GG23]. Messaging [SV22b]. Meta [DO22, FS24, XZL^{+}24]. Meta-Learning [FS24, XZL^{+}24]. Meta-level [DO22]. Metaheuristic [Gok22, ZHS^{+}25, Ano24e]. Metaheuristic-Enabled [Gok22, Ano24e]. Metals [CRA23]. Meteorological [HJ20]. Meter [MAAJ24]. Method [ANA^{+}22, AP22, B^{+}23, CLH22, DFS^{+}21b, ErEE20, GLW^{+}24, GZY^{+}23, HB23, Hsu23, Hsu24, LL22, LGFD24, MYB20, MA20, MG22, NJ21, NBTB20, NLD^{+}23, QGL^{+}22, QLZ24, RNJRLL^{+}22, SLW^{+}24a, SHL22, WG21, WZQ^{+}22, WGL^{+}22, WZG^{+}24, WCC^{+}22, WCD^{+}24, XZLZ22, XZL^{+}24, ZLWZ24, CGQ^{+}25, CLH^{+}25, GZ25, LDFD23, LDCX25, WZB^{+}25, WZ25, YZS25, ZMC25]. Methodology [AAL25, GCMRGM24]. Methods [ETA22, MKI20, MFHhK21, OMG22, TEGB22, TEHG22, ZJ23, Hsu25]. Metric [EMYRV21, Meg20, PK20, YLY^{+}24b]. Metrics [BXZ22, DM23, LGC^{+}24, LL24b]. Metrics-Aware [LL24b]. MFAS [LUW^{+}25]. Microarray [SV22a]. microfluidic [BG25]. Microsoft [YMD24b]. Middle [DWGC23, LC22, LSW^{+}23, LSG^{+}20, LSG^{+}23]. Middle-Product [LSW^{+}23]. Midori [GZY^{+}23, MA20, ZZW^{+}24a]. Migrating [ZSM24]. Migration [YYLX21]. Millimeter [RM22]. MIMO [RM22]. min [CLX^{+}24]. Minimal [CCC^{+}23, YWY21]. Minimum [dAJS22, LZG21a, ZC23a]. Mining [ASYLAM25, BCS22, DMG^{+}23, HLML21, KK23a, MUA23, MFHhK21, SRL20, SJ22, YS22, YJY^{+}24]. Mining-Based [HLML21]. Minority [ZHL^{+}24a]. MinRank [SZH^{+}24]. Minutiae [LMMR22]. Misinformation [SST23]. Mismatch [ZZW22]. Missing [GLXP23, RW20]. Mitigation [SST23]. Mixing [QXZZ21, YZC^{+}24]. MKSIFT [KP20]. ML [Ang24]. MLP [RA22b, SRL20]. MM [GLXP23, RW20, ZZC24a]. MM* [YLW21]. MMS [DMG^{+}23]. MNVPCS [CHT20]. Mobile [MM21, RMSMAH^{+}20, Sha24, SK24b, ZCW^{+}24, QXX^{+}25, SKE^{+}25]. Mobile-Assisted [MM21]. Mobility [AS24a, KSG20b, R^{+}23]. Möbius [ZXDL24]. Modal [HHLL22, ZZH24, JWJ24, LUW^{+}25]. Mode [CZZW21]. Model [ANKZ^{+}22, Ala20, ADF22, AKI20, Ano24a, BSM21, BPNdQ22, CRA23, CLX^{+}22, C^{+}23a, DMP^{+}23, DO22, DY21, DH21, FYH20, FZH21, FGC22, FWC24, FGM21, GK21, GZL^{+}21, GR22, GSFS21, GG23, GK23, HHLL22, HZZ20, HHB24, HHC22, H^{+}23c, JCY^{+}20, KS20a, KSG20b, KYAN21, LSX^{+}21, LdXZ^{+}21, LHL21, LLW22, LdXZ^{+}22, LZL^{+}23, LZH^{+}20, LZC^{+}21, LSY^{+}20, LLF^{+}24, LH20, LHLY20, LPP21, LRP22, LLDX23, MH21, MHG22, MKI20, MSH22, Nar22, PK22, PDO22, PFT24, PDRC22, RW20, RW23b, SCH^{+}20, Sha21, SM23, SLW^{+}24a, SA22, SR23, SK22, TJB23, VSAR24, WQZ^{+}22, WWSW24, XTLM24, X^{+}23, YTXW24, Y^{+}23b, YLZ20, YTL^{+}23, YLW21, ZLL^{+}22, ZYC^{+}24, ZZC24a, GLX25, LSF25, LZY^{+}25, SS25, SZCY25, WCLY24, WPF^{+}25, WZB^{+}25, WJHL25, WZ25, YJY^{+}24, ZJLM25, ZBQ^{+}25]. Model-Based [JCY^{+}20, HZZ20]. Model-Driven [ANKZ^{+}22]. Modeling [AASG23, Ano24a, AEGA22, AEGA23, CT23, LTH21, WWY20a, WYXZ22, Y^{+}23a, ZZC^{+}22, ZLL^{+}22, uHLH22]. Modelling [Ken20, VSS^{+}22]. Models [AÖ21, CvDRV24, CK22, DM23, GLXP23, KF24, LH22, Lee20, LCC^{+}24, MKI20, PGV^{+}22, SB24, LLL^{+}25, WY25]. Modern [BG21]. Modification [SLT^{+}22, ZOLX23]. Modified [Ano23b, BSM21, SV22a, SAK22, X^{+}23]. Modular [BMV22, SZX20]. Modularization [IOS^{+}22]. Modules [SDK^{+}22]. Modulo [RAD20]. Modulus [ZLZ22]. MoG [CLC^{+}19, CLJ^{+}22]. Moisture [MGB23, MBA23]. molecular [QWHC25]. Molecules [GR21]. Monarch [SJ20]. Monarch-Earthworm-Based [SJ20]. Monarch-EWA [SJ20]. Monetization [AVA21]. Monitoring [BS24a, JMR23, MKB23, MBA23, SS22a, Alm25, WTWZ25]. Monoculus [PMMS22]. Monotonic [CCC^{+}23]. Mostly [Ang24]. Moth [GS22]. Motion [SRS^{+}23]. mountaineering [GZ25]. Move [Ano24f, KSG22, LGX^{+}24]. Moves [Mer20]. MPSoC [SWJkL20]. MR [NBTB20]. MRI [CKA21]. MSVNN [GR22]. Multi [ANA^{+}22, AV20, Ano23d, AAM^{+}22, BFM22, BK23, CC24, DLM20, GW24, HZS^{+}23, JB23, JZ23, KSD22, KF24, KESZ22, KM23, KB22, LdXZ^{+}21, L^{+}23a, LYL24, LZQ^{+}23, LDW^{+}24, LUW^{+}25, MSH22, PKLK21, PDRC22, RCK22, RA22b, TLD^{+}24, TEHG22, TGZ^{+}21, WWY^{+}20b, WZ24, XWH^{+}23, X^{+}23, XZYZ24, XWCZ24, YLY24a, ZHT22, ZZH24, CPLW25, CZPL25, HBI^{+}25, JWJ24, LLGP25, LWZ^{+}25, PFGW25, QXX^{+}25, Tan24, WTWZ25, WLTY25, WZB^{+}25, YZLL25, LdXZ^{+}22]. Multi-Access [BK23]. Multi-Armed [AV20]. Multi-Band [WWY^{+}20b]. Multi-Biometrics [TLD^{+}24]. Multi-Byte [JZ23]. Multi-Channel [ANA^{+}22]. Multi-Class [KM23]. multi-clouds [HBI^{+}25]. Multi-criteria [TEHG22]. Multi-digit [AAM^{+}22]. Multi-Dimensional [CC24, TGZ^{+}21, XZYZ24, PFGW25].

Multi-Directional [YLY24a]. Multi-Document [BFM22, JB23]. Multi-Domain [LDW^{+}24]. Multi-exposure [HZS^{+}23]. multi-factor [QXX^{+}25]. multi-feature [WZB^{+}25]. Multi-Instance [GW24]. Multi-Key [X^{+}23, XWH^{+}23]. Multi-Keyword [LZQ^{+}23, CZPL25]. Multi-Label [KF24, LWZ^{+}25]. multi-level [Tan24]. Multi-Modal [ZZH24, JWJ24, LUW^{+}25]. Multi-Net [MSH22]. Multi-Objective [Ano23d, KB22, WZ24, BFM22, WLTY25]. multi-party [CPLW25]. Multi-Passenger [DLM20]. Multi-path [LYL24]. multi-pipeline [WTWZ25]. multi-scale [LLGP25, YZLL25]. Multi-semantic [LUW^{+}25]. Multi-Signature [KSD22]. Multi-Stage [KESZ22, RA22b]. Multi-Swarm [LdXZ^{+}21, LdXZ^{+}22]. Multi-Tab [XWCZ24]. Multi-Verse [Ano23d, KB22, RCK22]. Multi-View [L^{+}23a, PDRC22, PKLK21, ZHT22]. Multiantenna [ZZL20]. Multichannel [RS21]. Multidimensional [CZZ23]. multilabel [uhLM^{+}25]. Multilayered [WSA22]. multilevel [WY25]. Multilingual [HY25, LHLY20]. Multimedia [OS23, ZTK21, SGGM21]. Multimodal [Ano24e, BS24c, Gok22, GLY^{+}24, LQZ^{+}24, SGGM21, YLW^{+}25a]. Multinode [CHT20]. Multipartite [YX24, WX25]. Multiparty [PCMPCNHR22]. Multipath [TKV23]. Multiple [BIZA21, CP22, CSYY23, DLM20, H^{+}23a, LL24b, SWJkL20, SR20, SE23, Y^{+}23c, ZOLX23]. Multiple-Choice [SWJkL20]. Multiple-Layer [SR20]. Multiplication [JM24]. Multiplicity [Ano23c, LCC23]. Multipliers [Lem24]. Multiprocessor [YX24]. Multiresolution [TADD23]. Multiscale [BSM21, LWM^{+}22, SWZW25]. Multiset [SZA24]. Multispectral [GR22, WZQ^{+}22]. Multivariate [DST20, SD23].

Multivariate-Based [SD23]. Muscle [SA22]. Muscular [ANA^{+}22]. MuST [LUW^{+}25]. MuST-GAN [LUW^{+}25]. mutation [BY24]. Mutual [CP23, HHLL22]. My [GGJM21]. Myopic [KT24]. mysteries [ZLC^{+}25].

Nagapattinam [JMR23]. Naïve [MAB^{+}24]. Name [K.24, SYC^{+}23]. Name-Based [K.24]. Named [GLY^{+}24, HZKF23, R^{+}23, KXLL25, WZB^{+}25]. Names [CCZ^{+}22]. NAND [ZLX24]. NARX [GK21, GPR20]. Nash [AR21]. Nastalique [RKA^{+}22]. Natural [BS21, HAWA^{+}22]. Nature [LH22]. NDN [KSS^{+}24, SVD^{+}24]. NDN-RBE [SVD^{+}24]. Near [WKW^{+}24]. nearest [KK23b, TZW^{+}24]. Need [BM23]. Needed [BÖI22]. Negative [Ano23c, LCC23, LTH21, GLLZ21]. Neighbor [qLlHG20, WZ20, DWLF25b, KK23b, TZW^{+}24, ZZJZ20]. Neighborhood [QZTZ21]. NER [YJY^{+}24]. Nested [RLIK21]. Net [BTT^{+}22, MSH22, SSY^{+}25]. NetFlow [LJ23]. nets [ZZZL25]. Network [ASPB22, ARAA^{+}22, Ano24c, Ano24e, BTT^{+}22, BT23b, CR22, CRA23, DV22, DFC^{+}23, DYZY24, FZH21, Far24, FWC24, Gok22, GGD22, Hsu23, HH24, JCG23, JMATQ23, KS21, KSA20, KGA23, KKBK24, LL21, L^{+}23a, LJLQ24, LH20, LHLY20, LTH21, LWL^{+}22, LL22, Liu24b, LWZR24, LGFD24, LJ23, LL24b, LC24b, LCF^{+}21, MST23, MRVC23, Man24b, MPL21, NCRS22, PLY^{+}23, PKLK21, PXWL22, PVFP22a, PVFP22b, RP22a, RA21, RA22a, Rat20, RK22, RJ23, RA22b, SP20a, SS22a, SRLM22, SP23, Sha21, SE23, SD20, SJC23, VKM21, WZQ^{+}22, WPY^{+}23, WCC^{+}22, XLLG22, YZW23, YXLZ23, YFC^{+}22, YLG^{+}24, ZHT22, ZWY^{+}24, ZTT24, ASCJ25, CLY^{+}25, ÇA25, DWLF25a, DWLF25b, GPR20, HHL^{+}24, JWJ24, MLX^{+}25, RGB^{+}24, SLL25, SS20, SLWL25, WTWZ25, WJHL25, ZJLM25, ZFY^{+}25, ZYW25, ZSH25].

Network-Based [SJC23, WCC^{+}22, YLG^{+}24]. Networking [R^{+}23, ZMC25]. Networks [ASSA21, AV20, Ano23b, Ano24g, AAM^{+}22, BRA21, CHT20, CLH22, DAR22, Far24, FQZL24, FGS24, Gan20, GLXP23, HZZ20, HHX^{+}23, Hsu24, HY24b, KA20, LLF^{+}21, LSH^{+}22, LZM^{+}22, LOGC22, LZL^{+}23, LZC^{+}21, LDW^{+}24, LFZ24, LZLZ24, MSZ^{+}20, MW21, MYB20, MM20, Meg20, NM23, OB21, QXZZ21, Q^{+}23, QZTZ21, QLZ24, RG22, RNJRLL^{+}22, RS21, SB22, SM21, Sha24, SST23, SAK22, SZL^{+}20, SFC^{+}23, SLY^{+}24, TAA23, WYXZ22, WGL^{+}22, WZ24, WCW^{+}24, XTLM24, XZZ^{+}24, YQL24, YZMC23, YHW^{+}24, Y^{+}23b, YFA20, YLW21, ZZL^{+}22, ZLZS22, ZeWSL24, ZHP21, ZQZ24, ZWZ^{+}24, ZXZ^{+}23, HZS^{+}25, Hsu25, LYZ^{+}25, LFDW24, PL25, QWC24, SK25, Wan25]. Neural [AV20, Ano24c, Ano24e, Ano24g, AAM^{+}22, BTT^{+}22, CR22, CRA23, C^{+}23b, CSYY23, DV22, DAR22, Gok22, GGD22, JCG23, KS21, KF24, KGA23, KKBK24, L^{+}23a, LH20, LLS^{+}24, MSZ^{+}20, MRVC23, MPL21, NCRS22, OB21, PVFP22a, PVFP22b, RP22a, RA22a, Rat20, RK22, RJ23, RA22b, SRLM22, SS22d, SE23, SD20, SJC23, SS20, TAA23, VKM21, WZQ^{+}22, WPY^{+}23, XLLG22, ZLZS22, ZeWSL24, ZLCW24, ZTT24, ZQZ24, ASCJ25, GPR20, LYZ^{+}25, MLX^{+}25, QWC24, RGB^{+}24, SLL25]. Neural-Aided [C^{+}23b]. Neuro [RD20]. Neuro-Fuzzy [RD20]. neuropathy [ASYLAM25]. NewHope [ZZW22]. Next [AM20]. NMF [TCY^{+}20]. Node [GLX25, LZN^{+}24, LGFD24, PL25, YLZ20, LXWD24]. Node-edge [GLX25]. Nodes [CBM24, GLXP23, QZTZ21, ZLC^{+}22]. NODSTAC [BG24]. Nodule [JCG23]. NOEKEON [ZZC24b]. Noise [WL20, YLZY22, YZX^{+}24]. Noisy [PG22]. NoisyOffice [CBEBPPZM20]. Non [BSM21, GLLZ21, LZX^{+}22, OFMH22, RLIK21, SA24, VG23, WW22, WZL^{+}24, YD21, Z^{+}23a, LWL^{+}25]. Non-functional [SA24]. non-fungible [LWL^{+}25]. Non-IID [Z^{+}23a]. Non-interactive [WZL^{+}24]. Non-Malleable [YD21]. Non-Maximum [OFMH22]. Non-negative [GLLZ21]. Non-Numerical [WW22]. Non-profiled [LZX^{+}22]. Non-repairable [VG23]. Non-Subsampled [BSM21]. Non-Uniform [RLIK21]. Nonce [LRG^{+}24]. Nonce-Based [LRG^{+}24]. Nonlinear [RD20, SDM23, WZC^{+}21]. Nonuniform [HLX22]. Norm [CC22]. Normalization [WCD^{+}24]. Note [GHC20]. Notion [LZM^{+}20]. Novel [AlS23a, AP22, BG24, ETA22, GAK20, HDK24, HH22, KS24, LHLY20, MAEK23, NM23, NLD^{+}23, QYZ^{+}21, Q^{+}23, SRL20, SP23, SJ22, SLC^{+}20, SLY^{+}24, SSB23, TH23, XZLZ22, XWCZ24, YQL24, ZTT24, LWZ^{+}25, MXH^{+}25]. NP [LZX^{+}22]. NP-LFA [LZX^{+}22]. NTRU [XTW^{+}23]. NuDist [LC20]. Number [AAM^{+}22, DLP^{+}21, WZ23]. Numerical [WW22]. NVMe [FMX^{+}25].

O [FMX^{+}25]. obfuscating [SZY^{+}25].

Obfuscation

[CLC^{+}19, CLJ^{+}22, PLY^{+}23, ZZ23]. Object [HWS21, LWM^{+}22, NB21, RA21, RA22a, SK24a, XWCZ24, YWZB24, GPR20, LLGP25, SS24b, TDZ^{+}24, ZLZ^{+}24, ZYW25]. Object-Based [SK24a]. Objective [Ano23d, KB22, WZ24, BFM22, WLTY25]. Objects [SIT20]. Oblivious [BS24d, DLP^{+}21, PL25]. observability [WLZ^{+}25]. Observation [SB20]. Observations [XLW22]. observed [XLL24]. Obstructive [Gök21]. OCCI [ANKZ^{+}22]. Occlusion [Z^{+}23b]. Occurring [MG22]. Ocular [ANA^{+}22]. ODBN [SS25]. OFDM [RD20]. Off [LWZR24, RG22]. Off-Chain [LWZR24]. Office [LHL21]. Offline [LHHW22, YX23, ZJZ^{+}22]. Offloading [ASSA21, GAK20, Hsu23, SK24b, LZ25, LXW^{+}25]. Offset [LWJ25]. One [Çal22, LHH23, SRLM22, Wil22]. One-by-One [Wil22]. One-Dimensional [Çal22, SRLM22]. Online [ESC24, HB23, HZY21, LHHW22, LS23, LH20, LHLY20, MYB20, XZL^{+}24, YX23, YYLX21, ZJZ^{+}22, ÇA25]. Online/Offline [LHHW22, YX23, ZJZ^{+}22]. Only [MG22]. Ontology [SA24]. opacity [XLL24]. Open [ATZ^{+}21, JTGJ20, LCX^{+}22, WPF^{+}25, WZ25]. open-source [WPF^{+}25, WZ25]. OpenStack [ZHL^{+}24b]. operation [SSN25]. Operational [CLX^{+}22]. Operations [RK21, MS25]. operative [BDFP23]. Operator [Ano24f, LGX^{+}24]. Operators [AGZ24]. Opinion [BA23a, DH21, LH20, LTH21, TEGB22, YS22]. Opportunities [GCG^{+}22]. optic [ASYLAM25]. Optical [BGHG22, RKA^{+}22, YTZ^{+}24]. Optimal [DLP^{+}21, EC22, ECE20, GCD21, MM20, PK21, SK24b, VB21, WZ25, YVCC23, ZY21, CPLW25, CL25, SGGM21]. Optimistic [ZKQH24]. Optimization [ASPB22, AM20, AK20, Ano23d, AR21, BFM22, BC22, BÖI22, CR22, CKA21, GCG^{+}22, GGD22, HPV^{+}21, JB23, JDH23, Ken20, KYAN21, KB22, LQCM23, LZL^{+}24, MS20, MP22, Nar22, NCRS22, OMG22, PVFP22a, PVFP22b, PP22, RP22a, RG23, RA22a, RP22b, RM22, RCK22, Rat20, SRC20, SS22a, SS20, SR23, SMK23, VB21, WZ24, XZL^{+}24, YSTD24, Y^{+}23c, ZCC^{+}23, LXW^{+}25, WLTY25]. Optimization-Based [MS20, RG23, RCK22, SS22a, SS20]. Optimization-driven [RA22a]. Optimization-Enabled [NCRS22, PVFP22a, PVFP22b]. optimize [Ma25]. Optimized [BT23b, GCD22, JM24, LZL^{+}21, LRP22, SP20a, SHJA24, SK24a, SM23, YXLS25]. Optimizer [AM20, Kal23, KGA23, MGB23, MBA23]. Optimizing [Eld25, YYLX21, uhLM^{+}25].

Optimum [AT24, BÖI22]. Oracle [RS22]. Oral [SSS^{+}22]. Orbits [Liu24a]. Orchestration [QMR^{+}20]. Order [LZM^{+}20, WYXZ22, ZY21, MLX^{+}25]. Ordered [DLM20, RS21, ZZW^{+}24b]. Organization [Sae24]. Orientation [SS24a]. Oriented [ALS23b, GZY^{+}23, HXLX18, HXLX22, VB21, BYYS21]. Orthogonal [Wil22]. Orthros [HWW^{+}24]. Other [Sha22]. Out-Of-Core [TADD23]. Outer [ZY23]. Outlier [BG24, CCC^{+}23]. Outsourced [LWL^{+}22, ZCZ^{+}24]. Outsourcing [BMV22, K^{+}23, SZX20]. Overfull [BRL24]. Overlap [MPP23]. Oversampling [ZHL^{+}24a].

P [YJY^{+}24]. P-TIMA [YJY^{+}24]. Paced [ZL24b]. Packet [CC24, CZH^{+}21]. Pages [PK22, PJ20]. Paintings [FLZ^{+}22]. Pairing [K^{+}23, YMD^{+}24a]. Pairings [ZYW^{+}20]. Pairs [CSYY23]. Pairwise [PPR23, MS25]. Palm [WZQ^{+}22]. Palmprint [PC22]. PAN [MRVC23]. PAN-Cancer [MRVC23]. Panama [ARAA^{+}22]. Pancyclic [HY24b]. Pancyclicity [KCLH21]. Panic [SS22b]. Panic-based [SS22b]. Panoptic [YWZB24]. Papers [ARAA^{+}22]. para [FMX^{+}25]. para-virtualization [FMX^{+}25]. Paradigms [TK22]. Parallel [BC22, GC22, OFMH22, SV22b]. Parallelism [WDC^{+}23]. Parallelization [RLIK21]. Parameter [LX23, LLF^{+}24, WYC^{+}23]. Parameterized [RSSJ23]. Parameters [KD21, MW21, Sha22, XLZ24]. Pareto [ZCC^{+}23]. Parity [YLZY22]. Part [RNJRLL^{+}22, VSAR24]. Part-of-Speech [VSAR24]. Partial [AYS^{+}24, LC20, LZG^{+}21b, SKE^{+}25, TJB23, YLZ20]. partially [XLL24]. Particle [BÖI22, CP23, Ken20, LQCM23, MP22, PK21, WZ24]. Partition [LZL^{+}24, WDC^{+}23, YXLZ23, ÇA25]. Partitioned [CC24]. Partitioning [LLY^{+}23, SWJkL20, ZHL^{+}24a]. Party [MHAR20, TZ23, CPLW25, LYZ^{+}25]. Parzen [XZLZ22]. Passenger [DLM20]. Passive [LWS^{+}21, YWHY20]. Password [SJHL21]. Password-Authenticated [SJHL21]. Patch [RCK22, TR24]. Path [BZZ23, LGFD24, TKV23, YCZ^{+}22, LYL24]. Path-Rank-Based [TKV23]. Path-Structure [BZZ23]. Paths [LZN^{+}24, Yan24, ZC23a, DWLF25a]. Pathway [GR21]. Pathwidth [ADFS24]. Patient [Oks24]. Pattern [AYS^{+}24, CCC^{+}23, CvDRV24, DMP^{+}23, WDC^{+}23, YZN24, LFDW24]. Pattern-Based [CCC^{+}23]. Patterns [Çal22, GGŞ23, LRP22, PJ20, XXW^{+}24]. Payment [LWZR24]. PBFL [LX23]. PBFT [KKM21, WY25]. PCA [HZS^{+}25, YTZ^{+}24]. PCDP [SZCY25]. PCDP-CRLPPM [SZCY25]. Peaks [ZWXL24]. PEAR [SZZ^{+}25]. Pending [KSS^{+}24]. Perception [KKK21, SZL22]. Perceptual [LGX^{+}24, CLY^{+}25, Ano24f]. Perfectly [RLIK21]. Performance [ARDP21, AEGA22, AEGA23, BRA21, CK22, CLM^{+}20, CYH^{+}20, Gan20, MAB^{+}24, OB21, RJ23, SDK^{+}22, TAA23, WSCL22, Y^{+}23a, Y^{+}23c, ZJ23, LSF25, QWHC25, VTZX25, Ano24g]. Periodic [AGZ24]. Permission [BCS22]. Permission-Role-Usage [BCS22]. Permutation [AV22, GC23, WWZZ24]. Persistence [LZC^{+}21]. Persistent [WZPZ24]. Personal [CDR^{+}24, LHL21]. Personality [AASG23, PC22]. Personalized [CZLY21, LHL21, SCH^{+}20, SZCY25]. Persons [PK20]. Perspective [BS24a, ZJZ^{+}22]. Perspectives [AK20]. Perturbation [SRC20, WZG25]. Petri [ZZZL25]. PHISH ATTENTION [PCS25]. phishing [PCS25, RGB^{+}24]. Phonocardiogram [SDM23]. photovoltaic [ZWL^{+}25a]. Physical [BHJ20, SS22b, HTJ25]. Pi [AK24]. Pile [LL24a]. Pilsung [CRRY23]. pipeline [WTWZ25]. Pipelined [ESC24]. Pixel [KS21, SP23, Y^{+}23c]. Pixel-Level [SP23]. Pizza [NRA23]. PKI [KKM21]. Placement [WYC^{+}23, YSTD24, HZS^{+}25]. Places [BA23a]. Plain [X^{+}23]. Plaintexts [Shu23]. Plan [HH22]. Planar [ADFS24, ICAAFSS25]. Plane [VD20, VC23, WTWZ25]. Plans [WLZ20]. Plant [DV22, MBA23]. Plantlet [WLLM23]. Plants [BÖI22]. Platform [MPP23, RMSMAH^{+}20, WZQ^{+}22, LWL^{+}25]. Platforms [Fat22, LLS^{+}25]. Play [BMA^{+}25]. PLSR [MK20]. PMC [FWC24, GLXP23, LZH^{+}20, RW23b, ZZC24a]. PNCTS [LC24b]. Point [CHT20, FQZL24, LDC^{+}21, PS20, TADD23, WZG^{+}24, ZC23a, LWJ25]. Poisoning [TZS^{+}23]. Polarity [SM23]. Policies [JCY^{+}20]. Policy [AK20, SLT^{+}22, T^{+}23, TLD^{+}24, WZMZ23, LLM^{+}24]. Policy-Based [SLT^{+}22, T^{+}23, TLD^{+}24, LLM^{+}24]. Political [MGB23, MBA23]. Pollination [GS22, KSN24]. Pollination-Based [GS22]. pollination-enhanced [KSN24]. Pollution [BS24a]. Polygonal [PXWL22]. Polygons [ZC23a]. Polynomial [HJK20, LWX22, YDS^{+}20]. Pop [AV22]. Pop-Stacks [AV22]. Pornographic [ZLM^{+}24]. Positioning [KK23a, Wan21]. Positive [Ano23c, CNL^{+}23, LCC23]. Possession [LAKL^{+}22]. Possibility [CX22]. Post [LLW24, WZJ^{+}24]. Post-Quantum [LLW24, WZJ^{+}24]. Potential [Ano23e, AR21, MMMZ23, MS23]. Power [AT24, BXZ22, CLM^{+}20, Far24, JXZ^{+}22, MMK22, OCPM20, ZWL^{+}25a]. Power-Efficiency [AT24]. Power-Law [JXZ^{+}22]. Powers [S^{+}23]. Practical [CLG^{+}20, HCZ23, H^{+}23b, LLM^{+}24, P^{+}23, QXX^{+}25, WZH^{+}22, ZLD^{+}20]. Pre [OB21, SSS^{+}22, YTXW24, YLY24a, WPF^{+}25].

Pre-Cancerous [SSS^{+}22]. Pre-trained [OB21, WPF^{+}25]. Pre-Training [YLY24a, YTXW24]. Precise [LDC^{+}21]. precision [ZBQ^{+}25]. Preclusion [WM22]. Precoding [RM22]. Predict [CP22, KKK21]. Predicting [BPNdQ22, LX23, MUA23, PJ20]. Prediction [ASM^{+}21, AA23, BS24a, CK21, CRA23, CLZ^{+}24, DH21, DTP24, FGM21, GK21, GGŞ23, GR21, GG22, GG23, GK23, HFT22, HYTG24, HSL^{+}22, JWR24, KS21, KS20a, KS20b, L^{+}23a, LJLQ24, LQJ^{+}24, LZN^{+}24, LL22, LC24b, MWH^{+}24, MST23, MYB20, MRVC23, MKB23, MAB^{+}24, Meg20, MPL21, MS23, MGB23, MBA23, PJ24, PGV^{+}22, RJ23, R^{+}23, SP20a, SRLM22, SJ22, Sha21, SM23, SHL22, SSB23, YZN24, ZWY^{+}24, ZZ24, BY24, LXP25]. Prediction-Based [FGM21, KS21]. Predictive [BS24a, CLX^{+}22, VSS^{+}22]. Preference [LFDF24, SHL22, WtZLS20]. Preferences [BYYS21]. Preimage [WDM^{+}23]. Preprocessing [ZCH23]. Preservation [Kat23, WZC^{+}21, WZG25]. Preserved [SRC20]. Preserving [CLZ^{+}24, HSC^{+}24, Kal23, LLG^{+}20, MAAJ24, ZZH24, CLX^{+}24, CLH^{+}25, SZZ^{+}25, ZTK21]. Pressure [Gan20]. Preventing [LCZ^{+}22]. prevention [Als25]. Priced [BS24d]. Prices [PGV^{+}22]. Primary [AEZ21]. Primitives [CHWM21, HWW^{+}24]. PRINCE [CWD^{+}23]. Print [WZQ^{+}22]. prior [ZLZ^{+}24]. Prioritization [BRA21]. Prioritizing [MKI20]. Privacy [ATZ^{+}21, AKD^{+}21, CZLY21, CHWM21, CLZ^{+}24, CLX^{+}24, CPN^{+}21, GT22, HSC^{+}24, Kal23, Kat23, LLG^{+}20, LWX22, L^{+}23e, MAAJ24, NLD^{+}23, SRC20, SVD^{+}24, ZZL^{+}22, ZZH24, ZTK21, LSF25, QSC^{+}24, SZCY25, SZZ^{+}25, WZG25]. Privacy-Aware [CHWM21]. Privacy-Enhanced [ZZL^{+}22]. Privacy-Preserving [CLZ^{+}24, HSC^{+}24, LLG^{+}20, MAAJ24, ZZH24, CLX^{+}24, ZTK21, SZZ^{+}25]. privacy-protection [SZCY25]. Private [GKK22, PCMPCA^{+}20, RK21, WL20, WW22, QYH^{+}25, QSC^{+}24]. Probabilistic [LZL^{+}23, MKI20, SRLM22, GZF25]. Probabilities [HLX22]. Probability [GZY^{+}23, HXW22, LPP21, LGFD24]. Probing [BA23b]. Problem [CC22, DLW24, FLPS23, JGJ^{+}24, dAJS22, KSKM23, LSW^{+}23, RAD20, SZH^{+}24]. Problems [AM20, JXZ^{+}22, TA23, YÖ23]. Process [ETA22, WBG21, XZL^{+}24]. Processing [BS21, BC22, CBEBPPZM20, HAWA^{+}22, RJ23, RMSMAH^{+}20, SV22b, XZZ^{+}24, zLB^{+}25]. Processor [TK22]. Producer [R^{+}23]. Product [Ano24d, FWC24, Kal23, LSW^{+}23, SRC20, WYXZ22, ZXL^{+}24, WX25]. Products [Lem24, Yan24]. profiled [LZX^{+}22]. Profit [CZLY21, HHX^{+}20]. Profit-maximizing [CZLY21]. Prognosis [GK23]. Program [ZZ24]. Programming [BC22, CFR23, MUA23]. Programs [BHJ20]. ProInfer [GZF25]. project [YLWL25]. projects [YMD24b]. Prolong [CHT20]. prompt [YJY^{+}24]. prompt-learning [YJY^{+}24]. Proof [CKR22, FHWH24, SJHL21, ZLZZ25]. Proofs [CLWH24, CCY^{+}24, GQL^{+}20, SLW24b, SZH^{+}24]. Propagation [BGHG22, HLJW22, LH20, LTH21, YZW23]. Properties [DFC^{+}23, HY24b, LW23b, PPR23, DCV25]. Property [HLJW22]. proportional [FMX^{+}25]. Proposals [AGA24, Ano24b]. Proposed [AGA24, Ano24b, HFT22, KKK21, RM22, SGGM21]. Proprietary [HHC22]. Protect [CPN^{+}21]. Protected [ZY23, ZLZZ25]. Protection [AKD^{+}21, GT22, LMMR22, PL25, SZCY25, WLZ^{+}25]. Protocol [Far20, Far24, Gan20, KYAN21, LLG^{+}20, MM20, PCMPCA^{+}20, PCMPCNHR22, RS21, SJ20, SMK23, YXZ23, ZKQH24, ZYG23, GZF25].

Protocols [BMV22, KSG20b]. Provable [LAKL^{+}22]. Provably [LHHW22, SD23, XTW^{+}23, YWHY20]. Providing [HYTG24]. Provisioning [SII22]. Proxies [OCPM20]. Proxy [LWS^{+}21, YWHY20]. Pruning [AV20, CC24, MM21, Mou22]. Psychology [MBA23]. Public [Ala20, AS24b, FGC22, HYZH22, JZWN23, LZG^{+}21b, LH20, LTH21, LTT^{+}22, MH20, QYZ^{+}21, YMD^{+}24a, ZQY^{+}22, HMG25]. Public-Key [Ala20, HYZH22, LTT^{+}22, QYZ^{+}21, ZQY^{+}22, JZWN23]. Publication [WW22]. Publicly [ZKQH24]. Publishing [SRC20, WL20, TZwSX24]. Pulmonary [Gök21, JCG23]. Pulse [Ano24g, TAA23]. PureSVD [YÖ23]. Pursuit [Ang24].

Qassim [AHI22]. QoE [HWC^{+}24]. QoS [AK20, BK23, GP21, KSG20b, LL24b, NJ21]. QoS-Aware [BK23, NJ21]. Qualitative [RNJRLL^{+}22]. Quality [AlS23a, Ano23e, MMMZ23, NCL22, PP22]. Quantitative [Meg20, RNJRLL^{+}22]. quantized [LYZ^{+}25]. Quantum [CWD^{+}23, DWZS24, LLW24, LTT^{+}22, LDW^{+}24, XWQ24, DCV25, QWC24, WZJ^{+}24, XZXY25]. Quantum-Resistant [LTT^{+}22]. Quaternion [TYY^{+}21]. Queries [ARDP21, P^{+}23, ZC23a, SZY^{+}25]. Query [AGZ24, CZLY21, HPV^{+}21, LPH21, SKA23, ZZH24, GPR20]. Query-based [CZLY21]. query-specific [GPR20]. Querying [LSH^{+}22]. Question [YWY21]. Questions [KC23]. Queue [SV22b].

R [HM23]. R-LWE [HM23]. Radar [Ano24g, BT23b, TAA23]. Radiation [MS23]. Radical [LS23]. Radical-Based [LS23]. Radio [MM20, RG22]. Radiographs [PXWL22]. radix [TZwSX24]. RADNet [ZYW25]. RAID [YYLX21]. RAID-6 [YYLX21]. Rainbow [ACLA23]. Rainfall [SP20a]. Random [AM20, Ano24d, aBWH^{+}22, CX22, EI21, GGŞ23, KD21, LHL21, LZ20a, MAB^{+}24, QDZZ20, RS22, Sha24, Sha22, ZXL^{+}24, SLWL25]. Random-Oracle-Free [RS22]. Randomized [LZ20b]. Randomly [Ano23d, KB22]. Range [P^{+}23, ZHL^{+}24b]. Rank [SP20b, TKV23]. Ranked [LZQ^{+}23, CZPL25]. Ranking [SHJA24, XTLM24, SP20b]. Ransomware [LCC^{+}24]. Rao [Kal23]. Rapid [WZG^{+}24, PZL^{+}25]. Rare [CCC^{+}23, DMG^{+}23]. Raspberry [AK24]. raster [MFSH^{+}25]. Ratcheting [YVCC23]. Rate [HB23, ZY21, LXP25, Ma25]. Rating [HFT22]. Ratings [ADF22]. Ratio [ZWXL24]. Ray [MST23]. RBE [SVD^{+}24]. RCR [JSBV22, JSB24]. RCR-32 [JSBV22, JSB24]. RCR-64 [JSBV22, JSB24]. RDH [SS24a]. Re [HH24, SP20b, CLY^{+}25]. Re-Grouping [HH24]. re-identification [CLY^{+}25]. Re-ranking [SP20b]. reachability [BATN24]. Reaction [Wan21]. Readability [PK20, PK22]. Real [ANCLHD25, AS24a, Ano24g, EC22, ECE20, HWS21, MYB20, MKB23, QXZZ21, SJL20, TAA23, WSA22, WKW^{+}24, SZZ^{+}25]. Real-Time [AS24a, Ano24g, EC22, ECE20, HWS21, MYB20, SJL20, TAA23, WSA22, WKW^{+}24, MKB23]. Real-World [QXZZ21, ANCLHD25, SZZ^{+}25]. Realistic [MJQ23]. Realizability [IOS^{+}22]. Reasoning [DFS^{+}21b, JWJ24]. receive [MXH^{+}25]. Receiver [ZZL20]. Receivers [Ano24g, TAA23]. Receiving [HAWA^{+}22]. Rechargeable [CHT20, MWH^{+}24]. Recognition [AEZ20, AEZ21, Ano24e, AAM^{+}22, BIZA21, DMP^{+}23, FFH22, FZH21, Gok22, GLY^{+}24, HZKF23, HZY21, KESZ22, KKBK24, LPP21, LRP22, MSH22, OB21, RMW24, RKA^{+}22, SAR^{+}22a, SAR^{+}22b, SS22d, SS20, SYC^{+}23, TYTZSE24, WSA22, WZQ^{+}22, XDZ22, ZWM22, AK24, JWJ24, PFGW25, WZB^{+}25].

Recognization [FQZL24]. Recommendation [ADF22, AAWX24, GSFS21, GGD22, HHX^{+}20, JGJ^{+}24, LFDF24, SC22, SD20, TGZ^{+}21, XQYA23, YTXW24, ZNMZ24, ZeWSL24, CLYZ25, WG25, YLWL25]. Recommendations [DLM20, LSC^{+}22, YÖ23]. Recommender [ATZ^{+}21, AKD^{+}21, AYW^{+}22, HFT22, XZYZ24]. Reconfigurable [NM23]. Reconfiguration [GGJM21]. Reconstruction [KGA23, TTPD21]. Record [GÖ23]. Records [CDR^{+}24, CCZ^{+}22]. Recover [Shu23]. Recovery [MG21, RG23, ZLD^{+}20]. Rectangle [YQDJ24]. Recurrent [Ano24c, CR22, MRVC23, NCRS22, RP22a, RK22, SS22d, SE23, ZWY^{+}24, WJHL25]. Recursive [SM21, WCW^{+}24]. Redactable [GCR^{+}22, ZYWH23, LLM^{+}24]. Redis [SV22b]. Redis-Based [SV22b]. Reduce [NCRS22, ZLX24]. Reduced [AlS23a, HY24a, HCZ23, H^{+}23b, LC22, LSG^{+}20, WDM^{+}23, ZWZW23, ZLD^{+}20, ZZD^{+}21]. Reduced-Round [HY24a, HCZ23, LC22, LSG^{+}20, ZWZW23]. Redundancy [LZY20]. Redundancy-Aware [LZY20]. Reference [AlS23a]. Referenced [XZZ^{+}24]. Refinement [ZZW22, CGQ^{+}25, LXP25]. Reformulation [SKA23]. REFS [MXH^{+}25]. Regenerative [EC22]. Region [ARAA^{+}22, XXM^{+}22, ZFY^{+}25]. regional [SZCY25]. Registered [CLH22, HH24]. Registered-Backoff-Time [CLH22]. Registration [LDC^{+}21]. Regression [CLZ^{+}24, MS23, XLLG22, ZLCW24]. Regret [GKK22]. Regular [LZL^{+}23, LZC^{+}21, SZL^{+}20, SFC^{+}23, ZHP21, ZMWL20, ZCH23, HNR24]. Regularization [HZLW24]. Regularizing [MSZ^{+}20]. Reinforcement [LL21, LFZ24, XCC20, HHL^{+}24, LXW^{+}25, WLTY25, Zha25]. Rejection [TA23].

Related [BS21, LLS^{+}24, WLW24, YQDJ24]. Related-Key [LLS^{+}24, YQDJ24]. Relation [BÖI22, CLH^{+}25, CKR22, HPV^{+}21]. relational [CLH^{+}25, WZ25]. Relationship [LLF^{+}21, SFC^{+}23, YLW21]. Relationships [ADFS24, KS24]. Relative [Meg20]. Relay [LOGC22]. release [LWL^{+}25]. Reliability [GHC21, HLX22, LZL^{+}23, LZYT25, LZLZ24, LCF^{+}21, LFZ^{+}22, VG23, WX25, YX24, ZM21, ZHP21, WZ25]. Reliable [KXLL25]. Relinearization [LSS24]. Reload [SS22c]. Relying [SB20]. Remote [GZ25, MM21, SSB23, TLD^{+}24, WWY^{+}20b, LLGP25]. Removal [ANA^{+}22]. Removing [XZZ^{+}24]. repairable [VG23]. Replica [LZL20]. Replication [AK20, GQL^{+}20, TJB23]. Reporters [HSC^{+}24]. Reporting [HSC^{+}24]. Representation [GEZ24, HFT22, LH21, SSB23, YWZB24, YAHVC20, ZWC^{+}24]. reputation [WY25]. Request [Y^{+}23b]. Requirement [AKI20]. Requirements [SA24]. rescaled [ZZF25]. Rescue [GGJM21]. Research [LZYT25, LCX^{+}22, QGL^{+}22, QLH24, ZYLW25, uHLH22]. reserved [Zou25]. Residual [AT24, WZ23, XLZ24, ZJLM25]. Resilient [HYZH22, Sha24, XWW23, ZYW^{+}20, ZWQ^{+}23]. Resistance [CLC^{+}19, CLJ^{+}22, WL20, ZZ21, ZXZ^{+}23]. Resistant [LTT^{+}22]. Resolution [BFG^{+}21, GR22]. Resolvers [LCX^{+}22]. Resonance [BTT^{+}22, KGA23]. Resource [BK23, HXCH24, JCY^{+}20, JTGJ20, NM23, PP22, SII22, XQYA23, LZ25, LDCX25, SGGM21, ZYLW25]. resource-aware [LDCX25]. Resource-Constrained [HXCH24]. Resources [ANKZ^{+}22, TTCCMR^{+}23, YWY21]. Response [AMA^{+}24, GR21, LCX^{+}22]. RESTFUL [HDK24]. Restricted [LZM^{+}22, SLT^{+}22, YYL^{+}20]. Results [KTK23, MK20]. Retinal [Ano24c, RK22, KK23b]. Retinopathy [Ano24c, KK23b, RK22]. Retrievability [CLWH24, GQL^{+}20, WLW24]. Retrievable [SRC20]. Retrieval [AP22, KGA22, KP20, KASV23, PS20, SKA23, TR24, ZZH24, BS24b, CGQ^{+}25, GPR20]. Retrieving [EI21, SIT20]. Retroactive [dAJS22]. Reusable [LLGC20, MQL23]. Reuse [ML20, WZJ^{+}24]. Reverse [SST23]. Reversible [B^{+}23, CYH^{+}20, QGL^{+}22, Y^{+}23c, YZN24, ZOLX23, QLH24]. Review [Ano23e, HFT22, KZH^{+}22, MMMZ23, RPP24, VS21, ZL24b]. Review-Based [HFT22]. Reviews [BA23a, HFT22, LS23, SC22, YS22]. Revisited [CC22, LZM^{+}20, LSQ20, WZJ^{+}24]. Revisiting [ARDP21, JSB24]. Revocable [ML20, TLMY21, T^{+}23, YMD^{+}24a]. Revocation [MH21, ZC23b, ZWG^{+}20]. Rewriting [BT24, T^{+}23]. RGIM [KSG20b]. Rice [DV22]. Ride [HHX^{+}20]. Rider [CP23, MGB23, Nar22, RP22a, SP20b]. Rider-Rank [SP20b]. Ridesharing [DLM20, HWC^{+}24]. Ring [DST20, H^{+}23c, LSW^{+}23, TJB23]. Rings [YDS^{+}20]. RIoT [Far20]. Risk [ErEE20, ASYLAM25]. River [Ken20]. RNN [Hsu25]. Road [AKA^{+}21, KAMA22, SP23, SS24b, ZYW25]. RoBERTa [LL23]. Robots [DLP^{+}21, KT24, PMMS22]. Robust [C^{+}23a, KSG22, LYW^{+}22, LZ23, PKLK21, RA21, TYY^{+}21, WHL22, XWY^{+}24, YTL^{+}23, YTZ^{+}24, PCS25, SWZW25, SZZ^{+}25, ZWL^{+}25a, ZYW25]. Robustness [ZZZ^{+}22, ZLZS22]. Role [BCS22, Kat23]. Roman [KNWT24]. Room [Ang24]. Root [LYW^{+}22]. Rotating [LZX^{+}22]. Rotation [C^{+}23a]. Rotation- [C^{+}23a]. Rotational [HXW22]. Rotational-XOR [HXW22]. Roughness [NCL22]. Round [HY24a, HCZ23, H^{+}23b, HXCH24, LC22, LSG^{+}20, WDM^{+}23, YD21, ZWZW23, ZLD^{+}20, ZZD^{+}21, YXLS25].

Round-Reduced

[H^{+}23b, WDM^{+}23, ZLD^{+}20, ZZD^{+}21]. Rounds [ZY23]. Route [HHX^{+}20]. Routing [Far20, FNLW23, Gan20, KSG20b, KYAN21, LWZR24, MWH^{+}24, RAD20, SJ20, SMK23, Y^{+}23b, PL25]. Row [CRRY23]. RSA [Shu23]. RSCOEWR [LS23]. RTIM [C^{+}23a]. Rule [DMG^{+}23]. Rules [CCZ^{+}22, ETA22, KK23a]. Rumor [DK22]. Running [CGY22, ZWL^{+}25b]. Rural [M^{+}23].

S [GR22, HLJW22, LZX^{+}22]. S-box

[LZX^{+}22]. S-boxes [HLJW22].

S-DolLion-MSVNN [GR22]. SA [SS25]. SA-GBO-ODBN [SS25]. SAAChain [LWL^{+}25]. SAC [Hsu25]. SAC-RNN-based [Hsu25]. Safe [KSKM23]. Sail [ASPB22]. Sailfish [Kal23, KGA23]. Salesman [DLW24]. Saliency [NCL22, WWW^{+}22, YTL^{+}23]. Saliency-Aware [WWW^{+}22]. Salient [LWM^{+}22]. Salinity [JMR23]. Sample [ZJ23]. Sampling [SST23]. SAR [KRR20]. sarcasm [LLL^{+}25, LLL^{+}25]. Sarcasm-GPT [LLL^{+}25]. sasa [ADJ23]. Satellite [GR22, KK23c, MS20]. Satin [SHJA24]. Satisfaction [NJ21]. Satisfiability [RAD20]. Saturnin [HCZ23, ZWZW23]. Saudi [AA23, Fat22]. Saving [NM23, OCPM20]. Scalability [ALS23b]. Scalable [CZH^{+}21, GGJM21, SST23, YLZ20]. Scalar [JM24]. Scale [CLX^{+}22, LLW24, LRP22, SII22, WG21, YXLZ23, ZLC^{+}22, ZCW^{+}24, LLGP25, QWHC25, YZLL25]. Scale-Invariant [LRP22]. Scaled [BJ22]. Scaling [TK22, YYLX21]. scan [YLW^{+}25b]. Scatter [DLW24]. Scattering [SDM23]. scenario [Hsu25]. Scenarios [Sae24, SZZ^{+}25]. scene [JWJ24]. scenes [ZYW25]. Schedule [BÖI22, HWW^{+}24]. Scheduling [AT24, AS24a, EC22, GS22, Hsu24, LZQL22, LZL^{+}24, PFT24, PAG^{+}22, TA23, TKV23, LDCX25, ZYLW25]. Scheme [CHT20, CLG^{+}20, CZC22, CLWH24, CLH22, DYZY24, DST20, FHWH24, GAK20, HFT22, KSD22, LHHW22, LLG^{+}20, LWS^{+}21, LZX^{+}22, LKA24, L^{+}23f, LC24b, MWQ^{+}24, Q^{+}23, R^{+}23, SD23, WZQ^{+}23, X^{+}23, YQL24, YWHY20, ZZ23, ZLX24, ZYWH23, ZYW^{+}20, CZPL25, TDZ^{+}24, TZwSX24, Wan25, YZC^{+}24, ZCZ^{+}24]. Schemes [LAKL^{+}22, RS22, WL20, WLW24, WZG^{+}20, Tan24]. Science [Kam24, Thi24]. scientific [uhLM^{+}25]. Sclera [XDZ22]. Scores [NJ21]. Screen [LC24b]. Script [RKA^{+}22]. Scrutinizing [Ala23]. SDN [FNLW23, VD20, VC23, ZJLM25]. SDTA [ZSQS24]. SDVoIP [GP21]. Search [AM20, BFM22, CP23, CLG^{+}20, GS22, HXW22, JZD21, JZWN23, KK23a, KSS^{+}24, KP20, LC20, LdXZ^{+}21, LdXZ^{+}22, LTT^{+}22, MH20, MRVC23, MA20, RCK22, R^{+}23, SP20b, SJ23, YX23, CZPL25]. Searchable [CNL^{+}23, LZQ^{+}23, L^{+}23e, WCD21, WZQ^{+}23, WZL^{+}24]. Season [YLX^{+}24]. second [MLX^{+}25]. second-order [MLX^{+}25]. Secret [TLD^{+}20, TLMY21, CTL^{+}25, VTZX25]. secret-sharing [CTL^{+}25]. Sectors [Kat23]. Secure [AKI20, CPLW25, JTGJ20, JZWN23, JM24, LHHW22, LZ22, LZ20b, MH20, M^{+}23, PCMPCA^{+}20, RS22, SD23, SZX20, SJ20, SG23, WZG^{+}20, XTW^{+}23, YVCC23, YWHY20, YNZ^{+}22, ZSQS24, Zha24, ZYA^{+}22, BG25, HBI^{+}25, HMG25, LLM^{+}24, LYZ^{+}25, SS25]. securing [CTL^{+}25]. Security [ANG20, AAL24, AS24a, BG21, CPN^{+}21, ErEE20, GCR^{+}22, GT22, JSBV22, LSY^{+}20, LWS^{+}21, LKA24, XLW22, YQL24, YCL^{+}20, ZQY^{+}22, ZYC^{+}24, ZWQ^{+}23, ZZC24b, AAL25, ASCJ25, GSL25, HTJ25]. SEDD [XWY^{+}24]. Seed [PS20, LWJ25, SS22c]. Segment [CC24, FNLW23, HKA24]. Segmentation [Ano23d, BTT^{+}22, KK23b, KB22, LQCM23, LWM^{+}22, MST23, MAEK23, PXWL22, RA21, SP23, SJC23, TH23, XDZ22, YWZB24]. Selcuk [LSG^{+}23]. Selection [Ano24c, EO21, GAK20, GC22, GK21, GCD21, HHC22, KD21, LZL20, MP22, NJ21, PS20, RK22, YAHVC20, ZJ23, ÇA25, WZ25, YZS25, ZHS^{+}25]. Selection-Based [PS20]. Selector [KKK21]. Self [ADJ23, CLL^{+}24, HJK20, KSKM23, LZY^{+}25, RSSJ23, SM23, SLY^{+}24, ZHYH22, ZHYH23, ZL24b]. Self-adaptive [LZY^{+}25]. Self-equivalence [CLL^{+}24]. Self-Paced [ZL24b]. Self-Parameterized [RSSJ23]. Self-Similar [SLY^{+}24]. Self-Stabilizing [HJK20, KSKM23, ADJ23]. Self-Tallying [ZHYH22, ZHYH23]. Seljuk [AMT23]. Semantic [Ano24a, EI21, KGA22, MKS^{+}22, PS22a, PS22b, RA21, RM20, SKA23, ZLL^{+}22, CGQ^{+}25, CLH^{+}25, LUW^{+}25, TZW^{+}24, ZLM^{+}24, ZLC^{+}25, SSY^{+}25]. Semantic-Based [SKA23]. semantic-preserving [CLH^{+}25]. Semantical [WCC^{+}22]. Semantically [KC23]. Semi [Ano23a, HPV^{+}21, HZW21, WQZ^{+}22]. Semi-join [HPV^{+}21]. Semi-Supervised [WQZ^{+}22, Ano23a, HZW21]. Sensation [LHL21]. Sensing [K.24, KGA23, LZY20, MM20, RG23, SSB23, WWY^{+}20b, GZ25, LLGP25, TDZ^{+}24, Wan25]. Sensitive [BKS22, WZG^{+}24, GLW^{+}24]. Sensitivity [SK22]. Sensor [ASSA21, CHT20, Gan20, HYTG24, MM20, Q^{+}23, RS21, TYTZSE24, YQL24, Y^{+}23b, YLG^{+}24, Wan25]. Sensor-Based [TYTZSE24]. Sensors [EC22, ZWLP24]. Sentences [PS22a, PS22b]. Sentiment [AÖ21, BA23a, ÇÖİ21, Fat22, LQZ^{+}24, LS23, LHLY20, LH21, OMG22, SR20, WPY^{+}23, XXM^{+}22, KNWT24, LDFD23, Man24a, ZFY^{+}25]. Sentiment-Specific [XXM^{+}22]. Separable [SK24a]. separated [zLB^{+}25]. SeqMask [GW24]. Sequences [Ala23, KK23a, LCC^{+}24, TY23]. Sequential [PJ20, CLYZ25]. Serial [KK23a]. Series [KK23c, SCH^{+}20, WHL22, MFSH^{+}25]. Serpent [ZZC24b]. Server [ML20, WYC^{+}23]. Server-aided [ML20]. Servers [OCPM20, WZPZ24]. Service [AAL25, ALS23b, AAJ^{+}20, HDK24, JDH23, MM21, NJ21, PP22, QMR^{+}20, TBH21, WtZLS20, YSTD24]. Service-Oriented [ALS23b]. Services [AVA21, BS23, LLG^{+}20, Noo23, PCMPCA^{+}20]. Servicing [ECE20]. Session [LSY^{+}20]. Session-Specific [LSY^{+}20]. Set [CT23, CC24, CKR22, KT24, KSKM23, LDC^{+}21, RK21, SE23, WCD21, ZHL^{+}24a]. Set-Pruning [CC24]. Sets [CKR22, CFR23]. Severity [ASPB22, CKA21]. Severn [Ken20]. SGD [SP23]. SGD-U-Network-Based [SP23]. SGX [ZLZZ25]. SHA [YXLS25]. SHA-256 [YXLS25]. Shamir [VTZX25]. Shape [EI21]. Shaped [SIT20]. Shapes [AEZ20]. Shared [B^{+}23, MQL23]. Sharing [HHX^{+}20, HSL^{+}22, LLF^{+}24, LZQ^{+}23, VD20, WZMZ23, CTL^{+}25, FMX^{+}25, HBI^{+}25, VC23, VTZX25]. Shearlet [BSM21]. Sheet [CRA23]. Shell [PLY^{+}23]. Shifted [Çal22]. Shilling [XZYZ24]. ship [ZSH25]. Short [Çal22, LH22, Lem24, MGB23, YYLX21, ZZC^{+}22]. Short-Code [YYLX21]. Short-Term [Çal22, MGB23]. Shorter [LSW^{+}23]. Shot [HWS21, ZWZ^{+}24]. shrunk [ZBQ^{+}25]. shufflers [QYH^{+}25]. shuffling [CPLW25]. Side [JZ23, YLW^{+}25b]. Side-Channel [JZ23]. Sierpiński [QDZZ20]. SIFT [BGHG22]. SIFT-Based [BGHG22]. Signal [Ano24g, QAA^{+}22, TAA23, ZZL20]. Signalling [GR21]. Signals [LLY^{+}21, MK20]. Signature [CZC22, DYZY24, DST20, H^{+}23a, H^{+}23c, KSD22, LHHW22, LLW24, LSQ20, LWS^{+}21, LSW^{+}23, MH21, Q^{+}23, RS22, SD23, TZ23, WZG^{+}20, YQL24, YDS^{+}20, YWHY20, ZJZ^{+}22, ZC23b, ZYWH23]. Signatures [BRL24, KKM21, LXY^{+}20, SLT^{+}22]. Signcryption [LSY^{+}20]. Significant [XDZ22]. Signing [PCMPCNHR22, TZ23]. Silent [HJK20]. SIMD [FPT22]. SIMECK [LLS^{+}24]. Similar [AGZ24, SLY^{+}24]. Similarity [PS22a, PS22b, VB21, YXLZ23, ZLCW24, YLWL25, PS22a]. SIMON [H^{+}23b, LLS^{+}24, LSG^{+}23, LLAL22]. Simon-like [LLAL22]. Simple [ZC23a]. Simplicial [ZXZ^{+}23]. Simplification [LPH21]. Simplified [MG21, WDC^{+}23]. Simulation [ANKZ^{+}22, AYS^{+}24, MNN20, BDFP23, QWHC25]. SimulAtor [ADJ23]. Simulink [TBH21]. Simultaneous [Lee20]. Simultaneously [JZ23]. Sine [ASPB22]. Singh [SRC20]. Single [HWS21, SZX20, TK22, XWY^{+}24, ZC23a]. Single-Core [TK22]. Single-Point [ZC23a]. Single-Shot [HWS21]. Singular [TYY^{+}21]. SIP [GP21]. SipHash [HY24a]. Site [KKK21]. Sites [CT23, GSFS21]. Situation [CLH22, FFH22]. Situation-Aware [CLH22]. Size [CGY22, DM23, FHWH24, FGC22, LLW24, ZJ23, ZBQ^{+}25]. Sized [HJ20]. Sketch [HJH^{+}24, L^{+}23b]. Skill [MUA23]. Skills [PK20, HTJ25]. SKINNY [GZY^{+}23, HLC^{+}23, MA20]. Skyline [SIT20]. SkySlide [MG22]. Slack [ECE20]. Slices [GGJM21]. slicing [WPF^{+}25]. SM4 [CLL^{+}24, LLCL24]. SM9 [LHHW22]. Small [AGA24, Ano24b, GR21, HJ20, LLW24, LLGP25, WLLM23, ZWM22]. Small-State [AGA24, Ano24b]. Smaller [BKS21]. Smart [AKA^{+}22, ACK^{+}24, AKA^{+}21, BA23b, HXCH24, KS20b, MAAJ24, MWQ^{+}24, OS23, WSA22, WLZ22, ZLZS22, TZW^{+}24]. Smartphone [SCH^{+}20, YHX^{+}24, ZWLP24, PFGW25]. Smartphones [YSH^{+}22]. smartwatch [ANCLHD25]. smartwatch-based [ANCLHD25]. Smoothing [HHB24, LZY^{+}25]. SMOTE [PG22, ZHL^{+}24a]. SMT [RAD20]. SMT-LH [RAD20]. Snap [KA20].

Snap-Stabilizing [KA20]. SnorkelPlus [KS24]. SNOW [JLH20, SJ23]. SNOW-V [JLH20, SJ23]. Social [ARAA^{+}22, Ano24a, BS24c, DK22, Fat22, LH22, LLGC22, LH20, LHLY20, LL22, MYB20, MPP23, MFHhK21, NRA23, SST23, SV22b, TGZ^{+}21, WKW^{+}24, WGL^{+}22, WZ24, WCC^{+}22, YFA20, ZLL^{+}22, Als25, ÇA25, LLS^{+}25]. Software [BYYS21, CDCN21, DM23, GP21, GSFS21, JSB24, SWJkL20, VG23, WBG21, Zha24, ZZ24, WJHL25, WZ25, ZMC25]. Software-Defined [GP21, ZMC25]. Software-Efficient [JSB24]. Soil [MGB23, MBA23]. Soils [JMR23]. Solapur [KK23c]. Solar [MWH^{+}24, MS23]. Solar-Aware [MWH^{+}24]. Solution [JGJ^{+}24, Nar21]. Solutions [ATZ^{+}21, Ano23e, MW21, MMMZ23]. Solving [AM20, RAD20, XZXY25]. Some [JXZ^{+}22, SZS23, SZL^{+}20, SFC^{+}23, TA23, ZHP21]. Sonar [YLG^{+}24, YLW^{+}25b]. Sort [AV22, GC22]. SoS [MNN20]. Sound [HHLL22]. Sounds [Gök21]. source [WPF^{+}25, WZ25]. South [BA23a]. Space [BJ22, LZ23, PiKT21, ZHL^{+}24a]. Space-Efficient [PiKT21]. Spam [KZH^{+}22]. Spanning [dAJS22, WCF^{+}22, HNR24, ICAAFSS25]. Spark [LZL^{+}24, RP22a]. Sparse [GEZ24, RA22a, YÖ23, YTL^{+}23, ZQZ24]. Sparse-FCM [RA22a]. Sparsely [YLZY22]. Spatial [BG24, L^{+}23a, LZZZ21, LLDX23, SIT20, YLY24a, ZWY^{+}24]. Spatial-Aware [YLY24a]. Spatio [MLX^{+}25, YZS25]. Spatio-temporal [MLX^{+}25, YZS25]. Spatiotemporal [KAMA22]. SPAW [ZHL^{+}24a]. SPAW-SMOTE [ZHL^{+}24a]. SPDK [FMX^{+}25]. SPDPOA [MBA23]. Speaker [SS20]. Special [AM23, GT22]. Specific [LSY^{+}20, XXM^{+}22, GPR20, LDFD23]. Specification [LZZZ21, SA24]. Specifications [BDFP23, BHJ20, IOS^{+}22]. Specifying [AKI20, DO22]. Spectra [ZZ21].

Spectral [WZC^{+}21]. Spectrum [MM20]. Speech [DFS^{+}21a, VSAR24]. speed [ICAAFSS25]. Spiral [Nar22]. splicing [KSG20a]. Spline [ANA^{+}22]. Split [GHC21]. Split-Stars [GHC21]. spoof [LUW^{+}25]. Spoofing [LGC^{+}24, LUW^{+}25]. Spread [DH21]. Spreaders [QLZ24, XTLM24]. Squirrel [CP23, MRVC23]. SS6 [YYLX21]. SSD [CTL^{+}25, FMX^{+}25]. Stabilizing [HJK20, KA20, KSKM23, ADJ23]. Stable [KSS^{+}24]. Stack [ARDP21, Far24]. Stacked [LL22, WJHL25]. Stacks [AV22]. Stage [JCG23, KESZ22, RA22b]. Standard [Ala20, FGC22, H^{+}23c, LSX^{+}21, LSY^{+}20, LAKL^{+}22, MH21]. Standards [TZ23]. Star [BZ22, HJK20, MS25]. Star-Structure [BZ22]. Star-Substructure [BZ22]. Stars [FGS24, GHC21]. Start [JGJ^{+}24, LFDF24, YTXW24, YÖ23]. State [AGA24, Ano24b, CFRS24, TY23, WLLM23, ZZ24]. Static [DTP24, ZWQ^{+}23]. Station [HSL^{+}22, JWR24]. Station-Based [HSL^{+}22]. Statistical [C^{+}23b, NCL22, PGV^{+}22, SB22, TEGB22]. Statistical-Based [SB22]. statistics [GZF25]. Status [ATZ^{+}21, LZG21a]. Stealing [ECE20]. Stealth [WLW24]. Stealthy [FQZL24]. steering [MXH^{+}25]. Steganographic [WZH^{+}22]. Steganography [KS21, TR24]. Stego [SS24a]. Step [Sar20, YSH^{+}22]. Step-Counting [YSH^{+}22]. Stereo [HHB24]. Stimuli [SZL22]. Stochastic [Rat20]. Stock [Ano23b, GK21, PGV^{+}22, SAK22]. Stone [DMP^{+}23]. Stop [Shu23]. Stops [DLM20]. Storage [GQL^{+}20, LZG^{+}21b, LWL^{+}22, TJB23, YMD^{+}24a, YLZ20, YLX^{+}24, LWL^{+}25]. Store [BMA^{+}25]. Stored [ANG20]. Storing [CDR^{+}24, Oks24]. Strategies [LZL20, LZL^{+}24, XWL23]. Strategy [JDH23, LWZR24, MSZ^{+}20, MWH^{+}24, RFAY22, TADD23, QLH24, SLL25, PL25]. Stream [AGA24, Ano24b, DWZS24, JLH20, JSBV22, JSB24, MG21, SCS24, WLLM23]. Streaming [HXLX18, HXLX22, TTCCMR^{+}23]. Streams [CCC^{+}23, L^{+}23b]. Strength [CRA23, Sha21]. String [FPT22, ZMWL20, XJWX24]. Strong [Ala20, FNLW23, GCH21, ZL24a]. Strongly [WCW^{+}24, XDL23]. Structural [MW21, SS22a, WGL^{+}22, YLZ24, HY25]. Structure [BZ22, BZZ23, BZ25, LYK^{+}20, LW23b, L^{+}23b, LZN^{+}24, LC24a, SM21, SWFW24, WLY24, WZC^{+}21, ZXDL24, BATN24, LZYT25]. Structured [WSA22]. Structures [SB24]. Student [MAB^{+}24, MUA23, MKS^{+}22, MBA23]. Study [Fat22, HJ20, KF24, MAEK23, NCL22, SAR^{+}22b, SCS24, TTCCMR^{+}23, WZH^{+}22, CHQM25, SAR^{+}22a]. Sub [B^{+}23]. Sub-graphs [B^{+}23]. Subdivision [ZZ21]. subdivisions [HNR24]. Subgraph [HLX22]. Subgraphs [FGS24, S^{+}23]. Subgroup [WLW24]. Subsampled [BSM21]. Subspace [HZLW24, XZZ^{+}24]. Substring [FLPS23]. Substructure [BZ22, ZXDL24]. Subsystem [LZLZ24]. Subtrees [SLY^{+}24, YCZ^{+}22]. Subversion [YCL^{+}20]. Success [ZY21]. Sufficient [SWW^{+}22]. Suffix [BKS21]. Suicidal [CK21, CK22]. Suitable [CDR^{+}24]. Sum [WZXX20]. Summarization [BS21, BFM22, JB23, KKBK24, TEGB22]. Sun [RG23]. Sunflower [JB23]. Super [GR22, SWFW24, ZZLY23]. Super-Resolution [GR22]. superposition [XZXY25]. Supersingular [BBD^{+}24]. Supervised [AKA^{+}21, CBEBPPZM20, WQZ^{+}22, XXM^{+}22, Ano23a, HZW21]. Supervisor [XLL24]. supervisory [ZZZL25]. Supplying [ZWLP24]. Support [ASK^{+}23, Ano23a, aBWH^{+}22, HZW21, KS21, LHH23, MS23, MAEK23, RJ23, SA24, SDW24, SS20, WSA22, ZZF25].

Supporting [LZG^{+}21b, HMG25].

Suppression [OFMH22]. surface [ZSH25]. Surpassing [KTK21]. Surrounding [SIT20]. Surveillance [MSH22, SK24a, VKM21, WSA22]. Survey [ATZ^{+}21, AAWX24, AYW^{+}22, CK21, GLLZ21]. Survivability [SRL20, SRLM22]. survivable [PL25]. Survival [MRVC23]. Susceptibility [MG22]. Suspicious [Ala23]. SVD [TYY^{+}21]. SVM [MK20, NBTB20, SK24b]. SVMs [SHJA24]. Swarm [Ano23d, BÖI22, CKA21, GGD22, Ken20, KB22, LdXZ^{+}21, LdXZ^{+}22, LQCM23, MP22, NCRS22, PK21, SMK23, WZ24]. Swin [LYL24]. Switching [LYK^{+}20]. Symmetric [HWW^{+}24, L^{+}23e, WZQ^{+}23, YVCC23]. Symptoms [MKB23]. Syndrome [aBWH^{+}22]. Synthesis [B^{+}23, XLL24]. Synthesizing [MJQ23]. Synthetic [BT23b, ZHL^{+}24a]. Synthetic-Aperture [BT23b]. System [AEZ20, ASK^{+}23, AP22, AKA^{+}21, BS24a, BS23, CR22, DLM20, DFS^{+}21a, FSN21, FNLW23, GKK22, Gök21, GÖ23, HSC^{+}24, HZKF23, HSL^{+}22, LZL20, LHL21, LKA24, MS20, MJQ23, MUA23, Oks24, PAO20, RD20, RM22, RKA^{+}22, SS22b, TYTZSE24, WSA22, YX24, ZOLX23, ZWG^{+}20, HBI^{+}25, PFGW25]. System-Based [BS23, RD20]. Systems [ATZ^{+}21, ACK^{+}24, AK20, AKD^{+}21, AYW^{+}22, AKI20, AS24b, ALS23b, BYYS21, BHJ20, CvDRV24, CDCN21, DFS^{+}21a, ECE20, HFT22, Liu24b, MNN20, RD20, RNJRLL^{+}22, RW23a, RJ23, SA24, VG23, WWSW24, WDC^{+}23, XZYZ24, YLX^{+}24, GSL25, QWHC25, XZXY25, XLL24]. Systems-of-Systems [MNN20].

T [AAL24]. Tab [XWCZ24]. Table [HLJW22, KSS^{+}24, YLZ20]. Tables [ACLA23]. Tabu [MSZ^{+}20]. tactical [PL25]. Tactics [YWTL23]. Tactile [SZL22, WSCL22]. Tag [GSFS21]. Tagging [VSAR24]. tailored [ICAAFSS25]. TaintAttack [PZL^{+}25]. Take [AV22]. Tallying [ZHYH22, ZHYH23]. Tamil [VSAR24]. Tamper [SJHL21]. Tamper-Proof [SJHL21]. Target [HSL^{+}22, LDFD23, XCC20, ZWM22, Z^{+}23b, YLW^{+}25b]. Target-specific [LDFD23]. Task [ASSA21, AS24a, BFM22, DY21, GS22, LLGC22, LZL^{+}24, PFT24, PAG^{+}22, SWJkL20, SK24b, LZ25, LXW^{+}25]. Task-Graph [SWJkL20]. Task-Scheduling [GS22]. Taxi [LLDX23]. Taylor [RG23]. TBFL [ZWL^{+}25a]. Technique [ANA^{+}22, AM20, BG24, CP22, GR21, KK23a, PG22, RJ23, RAD20, SS24a, SSB23, ZHL^{+}24a]. Techniques [AYW^{+}22, MPP23, Meg20, RMSMAH^{+}20, SDK^{+}22, VS21, YWTL23, VTZX25]. technologies [AAL25]. Technology [KAMA22, AAL24, WLZ^{+}25]. Template [LMMR22, ZY23]. Templates [ANG20]. Temporal [BG24, L^{+}23a, LLDX23, MPL21, ZWY^{+}24, BATN24, MLX^{+}25, YZS25]. Temporary [LSY^{+}20]. Tensor [Ano24f, LGX^{+}24, YTZ^{+}24]. Tenuous [LSH^{+}22]. Terahertz [WWY20a]. Term [BT24, Çal22, HHB24, LH22, MGB23, Tan24]. term-weighting [Tan24]. Terminating [DLP^{+}21]. Tertiary [MUA23]. Test [LSX^{+}21, LZG^{+}21b, LMH^{+}21, YMD^{+}24a]. Testbed [ARDP21]. Testing [CDCN21, LZC^{+}21, SB20, ZMWL20]. Tetrahedrons [LMMR22]. Text [BFM22, DK22, DFS^{+}21a, HAWA^{+}22, KF24, OS23, VB21, CGQ^{+}25, CLH^{+}25, ÇA25, JWJ24, Tan24]. text-based [ÇA25]. Text-to-Speech [DFS^{+}21a]. Texts [ZZC^{+}22]. Textual [Ano24a, SYC^{+}23, ZLL^{+}22]. Texture [DV22, XXW^{+}24]. TF [OEAA23]. TF-IDF [OEAA23]. th [CLX^{+}24, WX25]. Their [AEZ21, BÖI22, HCZ^{+}21, KCLH21, Sha22]. Thematic [Ang24, Ant24, Kam24, Man24a, Man24b]. Theme [HLML21]. Theoretic [HZS^{+}23]. Theoretical [ZY21]. Theory [CKR22, K.24, RAD20, WCD21, Eld25]. Theory-Based [RAD20]. Thermal [LHL21, NB21]. Things [TTCCMR^{+}23, Ant24, AYW^{+}22, Far20, GCR^{+}22, Mou22, M^{+}23, MBA23, SD23, WCC^{+}22, YX23, YNZ^{+}22, ZZL^{+}22]. Third [MHAR20]. Thoughts [CK22]. Threat [GW24, YWTL23, YJY^{+}24]. Three [KCLH21, WZ20]. Three-Ary [WZ20]. Threshold [CLX^{+}24, CCY^{+}24, KKM21, XTW^{+}23, QXX^{+}25]. Thresholding [XZLZ22]. Throughput [RG22]. thwart [WLZ^{+}25]. Tightly [WZG^{+}20]. Tile [RLIK21]. TIMA [YJY^{+}24]. Time [AS24a, Ano24g, CLH22, EC22, ECE20, GYY^{+}20, HWS21, HH24, KK23c, LZQL22, LTH21, LDW^{+}24, LZL^{+}24, MYB20, MKI20, RAD20, SCH^{+}20, SDM23, SJL20, TAA23, VG23, WSA22, WKW^{+}24, WHL22, ZWM22, ZYG23, DCV25, MKB23, MFSH^{+}25, ZYLW25]. Time-Aware [LZL^{+}24]. Time-Based [LZQL22]. time-controlled [DCV25]. Time-Critical [VG23]. time-dependent [ZYLW25]. Time-Domain [ZYG23]. Time-Series [SCH^{+}20]. Time-Weighted [KK23c]. Timed [CFRS24]. timeouts [TY23]. Timer [GHZ24]. TimeRider [VKM21]. TimeRider-Based [VKM21]. Times [Ano24d, QDZZ20, ZZ21, ZXL^{+}24]. Timeslot [NJ21]. Timing [JWR24]. TinyThunder [CL25]. Titanium [CRA23]. TLBO [KYAN21]. tokens [LWL^{+}25]. Tolerance [HLCX23, LC24a, SM21, SWFW24, CL25, DWLF25b, LFDW24]. Tolerant [MWQ^{+}24, XDL23, ZL24a]. Tomo [HHL^{+}24]. Tomography [JCG23, HHL^{+}24]. Tongue [NGS22, SSS^{+}22]. Too [BRL24]. Toolbox [ZSM24]. Tools [AASG23, MFHhK21]. Top [P^{+}23, RS22, TBH21, YÖ23]. Top- [YÖ23].

Top-2 [P^{+}23]. Top-Level [RS22]. Topic [Ano24a, SC22, SJL20, ZZC^{+}22, ZLL^{+}22, uHLH22]. Topological [MPL21]. Topology [YXLZ23]. Tor [XWCZ24]. tori [SKE^{+}25]. Total [ÇA23]. Touchscreens [WSCL22]. Tourism [SHL22]. Tourist [BA23a, SHL22]. Touristic [PCMPCA^{+}20]. Toxicity [GR21]. TPM [CEN24]. TPRT [TZwSX24]. Trace [XWCZ24]. Traceable [LHW21, ZC23b]. tracer [LUW^{+}25]. Tracing [Nar21]. Tracking [BGHG22, GYY^{+}20, SDK^{+}22, WWW^{+}22, Z^{+}23b, PZL^{+}25]. Tracy [SRC20]. Trade [RG22]. Trade-Off [RG22]. Trading [CZLY21, LLW22, YXZ23, ZSQS24, LSF25]. Traditional [LLY^{+}21]. Traffic [AKA^{+}21, ESC24, GLW^{+}24, HDK24, HJH^{+}24, HH24, JWR24, L^{+}23a, LJLQ24, LJ23, YFW^{+}23, YZX^{+}24, Zar20, ZWY^{+}24, ZCW^{+}24, MLX^{+}25]. Traffic-Aware [HH24]. Traffic-Centric [Zar20]. Trails [JZD21]. Train [CBEBPPZM20]. trained [OB21, WPF^{+}25]. Training [GK23, JZ23, Sae24, YLY24a, HTJ25, YTXW24, Zha25]. Traitors [LHW21]. Trajectory [LOGC22, WL20, TZwSX24]. Transaction [YZW23]. Transactional [AGZ24]. Transactions [AGZ24]. Transfer [BS24d, Far24, HJ20, LFDF24, TT24]. Transferred [DFS^{+}21a]. Transferring [HAWA^{+}22]. Transform [Gök21, LRP22, WWY^{+}20b]. Transformation [WG21, ZYS^{+}23, WZG25]. Transformations [SLY^{+}24]. Transformed [GCD22]. Transformer [AMA^{+}24, KC23, LUW^{+}25, LYL24]. transformers [ZLC^{+}25]. Transient [VD20, VC23]. transition [AAL24, AAL25]. Transitions [MMK22]. Translation [C^{+}23a, GCG^{+}22]. Translation-Invariant [C^{+}23a]. Transmission [LGFD24, LC24b, M^{+}23, TKV23, VSS^{+}22, WLTY25, Wan25]. Transparency [LZM^{+}20]. Transportation [HLML21]. Transposition [XDL23].

Trapdoor [HYZ^{+}20]. Trapezoidal [GZL^{+}21, L^{+}23b]. Travel [HZY21]. Traveling [DLW24]. Tree [BPNdQ22, FGC22, HM23, dAJS22, KA20, SWJkL20, SB20, WZQ^{+}23, TZwSX24]. Tree-Based [HM23]. Trees [BGGBN24, CC24, ESC24, SDW24, WCF^{+}22, XDL23, ZYWH23, HNR24, ICAAFSS25]. Trending [KTK23]. Trends [uHLH22]. Tri [LRP22]. Tri-directional [LRP22]. Triangulation [YLG^{+}24]. Tricyclic [YCZ^{+}22]. Triple [ZC23a, CLH^{+}25]. Triple-Point [ZC23a]. Trojan [WLZ^{+}25]. Truncated [GZY^{+}23, HWW^{+}24, Lem24, MA20]. Truncated-Differential [MA20]. Trust [ADF22, BS23, DY21, PP22, TEHG22]. Trust-Aware [BS23, DY21]. Trust-Based [PP22]. trusted [ZWL^{+}25a]. Tsunami [JMR23]. TTP [MWQ^{+}24, PCMPCNHR22]. Tumor [Ano23d, CP23, CKA21, KB22]. Tumour [RPP24]. tuned [WPF^{+}25]. Tunicate [GGD22]. Turkish [ÇÖİ21]. Tutoring [RNJRLL^{+}22]. Tweakable [LC22]. Tweet [GKK22, BA23a]. twin [ZZF25]. Twisted [KCLH21, qLlHG20]. Twitter [DH21, GKK22, WLY24, YJY^{+}24]. Two [BGGBN24, CPN^{+}21, JCG23, KCLH21, LOGC22, OMG22, PPR23, PJ20, RLIK21, Sar20, SLY^{+}24, TZ23, TEGB22, WZ20, LYZ^{+}25]. Two-Dimensional [BGGBN24]. Two-Disjoint-Cycle-Cover [KCLH21]. Two-Factor [CPN^{+}21]. Two-Forest [SLY^{+}24]. Two-Good-Neighbor [WZ20]. Two-Hop [LOGC22]. Two-Level [PJ20, RLIK21]. Two-Party [TZ23, LYZ^{+}25]. Two-Stage [JCG23]. Two-Step [Sar20]. Type [AA23, JQ24, LLCL24, MS20, CHQM25]. Type-2 [MS20]. Types [KCLH21].

U [BTT^{+}22, SP23]. U-Net [BTT^{+}22]. UAV [ASSA21, Hsu25]. UAV-Aided [ASSA21]. UAVs [CT23]. Ubiquitous [Man24b]. uBlock [ZZW^{+}24a]. UE [GAK20]. UeLP [LLS^{+}25]. ULDC [YZX^{+}24]. Ultrasound [BSM21]. unbalanced [LDCX25]. Unbiased [Zha25]. Uncertain [CCC^{+}23]. Uncertainty [PJ24, PL25]. Unclaimable [H^{+}23c]. Unconstrained [AM20]. Underground [SSB23]. Understanding [Rat20, Y^{+}23a]. Underwater [Gan20, PK21, RS21, SLC^{+}20, YLG^{+}24, YLW^{+}25b]. Unexpected [LCX^{+}22]. UnfairDuelMerge [Mer20]. Unified [GC23]. Uniform [HLX22, RLIK21]. unifying [LXWD24]. unit [WJHL25]. Universe [LSX^{+}21]. University [AHI22]. Unknown [GLW^{+}24]. unlearning [GSL25]. Unlinkable [BS24d, TLMY21]. unmanned [ZSH25]. Unobtrusive [RW23a]. Unravelling [ZLC^{+}25]. Unreachable [ZLWZ24]. Unsegmented [SDM23]. unsorted [BATN24]. Unsupervised [FS24, PS20, WCC^{+}22, YZX^{+}24]. Untrusted [SZX20]. Updatable [HYZ^{+}20]. Update [WZMZ23, XCC20, ZC23b, BS24b, SLL25]. Update-Free [ZC23b]. Updated [Ano23d, KB22, Nar22]. Updates [LC24b]. Uplift [SD20]. Upper [ZZW^{+}24a]. Urban [DLM20, HDK24, LHH23]. Urdu [HZKF23, KNWT24, RKA^{+}22]. Urdu-Like [RKA^{+}22]. URL [PCS25]. Usage [BCS22]. Use [ASYLAM25, GGŞ23, HYTG24]. User [ADF22, Ano24a, GKK22, HXLX18, HZY21, K.24, LOGC22, MYB20, TLD^{+}24, WtZLS20, XZYZ24, ZLL^{+}22, ANCLHD25, LLS^{+}25, HXLX22]. User-Experience-Oriented [HXLX18, HXLX22]. User-Generated [Ano24a, HZY21, ZLL^{+}22]. Users [ÇÖİ21, LFDF24, ÇA25]. Uses [Thi24]. Using [ASPB22, AKA^{+}22, AASG23, AT24, AP22, AV20, AR21, AAJ^{+}20, AKA^{+}21, BTT^{+}22, BS24c, BT23b, BC22, B^{+}23, BIZA21, Çal22, CK21, CR22, CLX^{+}22, CLH22, CCZ^{+}22, CFR23, DV22, DAR22, DFS^{+}21b, EI21, FSN21, Fat22, GAK20, Gök21, GG22, GK23, GCD22, HZKF23, HFT22, HH22, JCG23, JMATQ23, JMR23, K.24, KSD22, KSG20b, Ken20, KTK23, KM23, KASV23, KK23c, LMMR22, LQCM23, Liu24b, LJ23, LKA24, LHH23, MRVC23, MAB^{+}24, MP22, NJ21, Oks24, OMG22, OEAA23, PCMPCNHR22, PVFP22a, PVFP22b, PP22, PS22a, PGV^{+}22, PS22b, PCLZ23, PK21, QAA^{+}22, RA21, RA22a, RFAY22, RJ23, RA22b, RMSMAH^{+}20, SHJA24, Sae24, SRL20, SCH^{+}20, SJ22, SSS^{+}22, SWW^{+}22, SJ23, SJHL21, SS22d, SE23, SST23, SDM23, SA22, SR23, SZL22, SMK23, SSB23, TR24, TADD23, TK22, TT24, VD20, VKM21, VSS^{+}22]. Using [WSA22, WCD21, Wan21, WSCL22, WPY^{+}23, WZ24, XLLG22, XQYA23, YHX^{+}24, Y^{+}23c, ZWM22, ZYC^{+}24, ZKQH24, uHLH22, ATZ^{+}21, AK24, Alm25, ASCJ25, BSM21, BA23a, BGHG22, CKA21, Eld25, HZS^{+}25, KSG20a, KK23b, MJQ23, MPL21, SKB^{+}22, VC23, WLZ22, WZ25, ZYL^{+}25, ZWL^{+}25b, ZYWH23, ZZZL25]. UTC [GLW^{+}24]. Utilization [TEHG22]. utilizing [SS25]. UWANs [ZYG23]. UWSN [YLG^{+}24]. UWSN-DT [YLG^{+}24].

V [JLH20, SJ23]. v2 [WLLM23].

Validation [DM23, WLZ20]. Validity [LSX^{+}21]. Value [ErEE20, L^{+}23d, WZPZ24]. Values [TYY^{+}21]. VANETs [Zar20]. VAPR [Gan20]. Variational [CLYZ25, QWC24]. Variations [JYW^{+}25, LGC^{+}24]. various [WX25]. Vector [Ano23a, aBWH^{+}22, CC22, GZL^{+}21, HZW21, KS21, LZ23, LHH23, MS23, Sha24, SS20, WSA22, WZXX20, ZLZ22, ZZF25]. Vectorization [QGL^{+}22]. Vegetation [KK23c]. Vehicle [LL24a, RAD20, SDK^{+}22, LZ25]. Vehicles [HDK24, ZSH25, L^{+}23f, TZwSX24, YZS25]. Vehicular [TBH21]. Vein [WZQ^{+}22].

Verifiable [CX22, CZPL25, L^{+}23e, WZQ^{+}23, ZKQH24, ZCZ^{+}24]. Verification [BDFP23, DK22, FYH20, ZLZS22]. Verifier [MH21, SJHL21]. Verifier-Based [SJHL21]. Verify [MHG22]. veritable [QXX^{+}25]. Verse [Ano23d, KB22, RCK22]. Version [MG21]. Versus [TEGB22]. Vertex [HLX22, LX20, YXLZ23]. Vessel [CLY^{+}25, XDZ22]. Via [GW24, GGŞ23, HAWA^{+}22, LFDF24, WZH^{+}22, XZZ^{+}24, Als25, CGY22, LW23a, LX23, SZY^{+}25, SYC^{+}23, ZLZ^{+}24, ZL24b, zLB^{+}25]. Vibration [SZL22]. Video [C^{+}23a, DAR22, KKBK24, MK20, MMK22, MSH22, PK20, PVFP22b, RCK22, TCY^{+}20, VKM21, YTZ^{+}24, GPR20, ZLM^{+}24, LYL24, PVFP22a]. Videos [KTK23, QAA^{+}22, SK24a]. View [BG21, L^{+}23a, PDRC22, PKLK21, ZHT22]. Violence [LYL24, MSH22]. Virtual [AT24, CHT20, LL21, LLY^{+}21, LL24b, QMR^{+}20, YLZ20]. Virtual-Point [CHT20]. virtualization [FMX^{+}25]. virus [BY24]. Visibility [ZC23a]. Visible [QGL^{+}22, SLWL25]. Vision [AMA^{+}24]. Visit [CT23]. Visual [Ano24e, BGHG22, FLZ^{+}22, Gok22, NCL22, WWW^{+}22, YHW^{+}24, ZWM22, ZOLX23]. Visualization [PDRC22, TADD23]. Visually [NB21]. VoD [AK20]. Void [Gan20]. VoIP [GP21]. Voting [AA20, H^{+}23a, RFAY22, ZHYH22, ZHYH23]. VQCNN [QWC24]. VR [CDCN21]. VR-Based [CDCN21]. VSA [XXM^{+}22]. Vulnerability [GHZ24, WCD^{+}24, ZWC^{+}24, SK25, XJWX24]. Vulnerable [Y^{+}23a].

Walks [Ano24d, LZ20a, LDW^{+}24, QDZZ20, ZXL^{+}24]. Wallet [MQL23]. warehousing [zLB^{+}25]. Warning [Ano24g, L^{+}23f, TAA23]. WARP [HWW^{+}24]. Warping [KK23c, ZWM22]. Watching [MK20]. Water [BÖI22, SSB23]. Watermark [Nar22]. Watermarking [ANG20, HAWA^{+}22, QGL^{+}22, XWY^{+}24, QLH24]. Watermarks [XZZ^{+}24]. Watson [MMR22]. Wave [RM22, KA20]. Wavelet [Gök21, KSG22, MPL21, KSG20a]. Way [CPN^{+}21]. Weak [CRRY23]. Weakly [XXM^{+}22]. Wearable [HYTG24, TYTZSE24]. weather [SS24b]. Web [BS23, HDK24, JDH23, OCPM20, PLY^{+}23, PK22, PJ20, WtZLS20, WZPZ24]. Weblogs [PJ20]. Webpages [SP20b]. website [PCS25, SWZW25, LS23, XWCZ24, ZWZ^{+}24]. Weighed [GPR20, VB21]. Weight [LWZR24, RA22b, SB22, SK24a, S^{+}23, YXLZ23]. Weighted [GCD22, HB23, KK23c, LC20, NJ21, ZHL^{+}24a, ZZF25, HNR24, WZ25]. weighting [Tan24]. Weights [YCZ^{+}22]. Welded [CRA23]. Whale [JDH23, Rat20, SRC20, LXW^{+}25]. WHDA [KRR20]. WHDA-FCM [KRR20]. Wheel [LX20]. Whirlpool [WWZZ24]. White [BTT^{+}22, CLL^{+}24, LLCL24]. White-Box [LLCL24, CLL^{+}24]. Who [ZZJZ20]. WHOIS [CCZ^{+}22]. Wi [PFGW25]. Wi-Crack [PFGW25]. Width [Sha22]. Widths [Sha22]. Wildcard [LZQ^{+}23]. Wildcards [Wil22]. Wildfire [KS20b]. Window [RAD20, XZLZ22]. Windows [SJL20]. Winkler [KASV23]. WiPoTS [Far24]. Wireless [ASSA21, BRA21, CHT20, CLH22, EC22, Far24, Gan20, LOGC22, Q^{+}23, RS21, SB22, YQL24, Y^{+}23b, YLG^{+}24, Wan25]. Withdrawn [Ano22]. Within [KS24]. Without [CNL^{+}23, Kam24, PCMPCNHR22, XWQ24, ZYW^{+}20, MWQ^{+}24, YMD^{+}24a]. WLAN [WCLY24]. Wolf [KRR20, PP22, VB21]. Word [GZY^{+}23, HFT22, LDFD23]. word-masking [LDFD23]. Word-Oriented [GZY^{+}23]. Words [LH21]. Work [OFMH22]. Work-Efficient [OFMH22]. workload [LDCX25]. works [LWL^{+}25]. World [ASM^{+}21, Kam24, QXZZ21, ANCLHD25, SZZ^{+}25]. Worrying [Shu23]. Worst [NJ21, Suc24]. Worst-Case [Suc24]. Wrapper [GK21]. Wrapper-Enabled [GK21]. Wrinkles [AEZ20]. Writing [LLY^{+}21]. WS [PJ20]. WS-BD-Based [PJ20]. WSN [KYAN21, SMK23].

X [MST23]. X-Ray [MST23]. XGBoost [LQJ^{+}24, OEAA23]. Xoodoo [ZLD^{+}20]. Xoodoo-AE [ZLD^{+}20]. Xoodyak [ZLD^{+}20]. Xoofff [ZZD^{+}21]. XOR [HXW22].

YOLO [SS24b, ZBQ^{+}25, ZSH25]. YOLO-based [ZBQ^{+}25]. YOLO-fall [ZBQ^{+}25]. Yolo-like [ZSH25]. YOLOv9 [YLW^{+}25b]. YOLOX [RMW24]. Young [GCD21, GCD22]. Youtube [KTK23].

Zero [CCY^{+}24, SLW24b, SZH^{+}24, WZXX20, YD21, ZLZZ25]. Zero-Knowledge [CCY^{+}24, SLW24b, SZH^{+}24, WZXX20, YD21]. Zhaoqing [HJ20]. zk [ZLZZ25]. zk-DCIAExchange [ZLZZ25]. Zone [LYW^{+}22]. Zoom [WSCL22].

References

Aydin:2020:CMV

[AA20] Fatih Aydın and Zafer Aslan. The construction of a majority-voting ensemble based on the interrelation and amount of information of features. The Computer Journal, 63(11):1756–1774, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1756/5625929.

Albahli:2023:CTP

[AA23] Saleh Albahli and Waleed Albattah. Crime type prediction in Saudi Arabia based on intelligence gathering. The Computer Journal, 66(8):1936–1948, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1936/6572652.

Asad:2020:DDD

[AAJ^{+}20] Muhammad Asad, Muhammad Asim, Talha Javed, Mirza O. Beg, Hasan Mujtaba, and Sohail Abbas. DeepDetect: Detection of distributed denial of service attacks using deep learning. The Computer Journal, 63(7):983– 994, July 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/983/5525444.

Al-Azzawi:2024:SAM

[AAL24] Ameen Al-Azzawi and Gábor Lencse. Security analysis of the MAP-T IPv6 transition technology. The Computer Journal, 67(10):2945–2958, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2945/7708712.

Al-Azzawi:2025:MSA

[AAL25] Ameen Al-Azzawi and Gábor Lencse. Methodology for the security analysis of IPv4-as-a-service IPv6 transition technologies. The Computer Journal, 68(10):1450–1462, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1450/8124674.

Asif:2022:LMD

[AAM^{+}22] Muhammad Asif, Maaz Bin Ahmad, Shiza Mushtaq, Khalid Masood, Toqeer Mahmood, and Arfan Ali Nagra. Long multi-digit number recognition from images empowered by deep convolutional neural networks. The Computer Journal, 65(10):2815–2827, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2815/6369317.

Atwan:2024:EDF

[AAR^{+}24] Jaffar Atwan, Nedaa Almansour, Mohammad Hashem Ryalat, Shahnorbanun Sahran, Hamza Aldabbas, and Dheeb Albashish. Ensemble of deep features for breast cancer histopathological image classification. The Computer Journal, 67(6): 2126–2136, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2126/7536050.

Acharya:2023:ICH

[AASG23] Amitabha Acharya, Aman Aryan, Sujay Saha, and Anupam Ghosh. Impact of COVID-19 on the human personality: an analysis based on document modeling using machine learning tools. The Computer Journal, 66(4):963– 969, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/963/6509041.

Alhosaini:2024:ARM

[AAWX24] Hadeel Alhosaini, Sultan Alharbi, Xianzhi Wang, and Guandong Xu. API recommendation for mashup creation: a comprehensive survey. The Computer Journal, 67(5):1920– 1940, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1920/7456151.

Bi:2022:CER

[aBWH^{+}22] Xia an Bi, Hao Wu, Xi Hu, Yu Fu, and Shaoliang Peng. Clusteringevolutionary random support vector machine ensemble for fMRI-based Asperger syndrome diagnosis. The Computer Journal, 65(2):251–260, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/251/5821299.

Abid:2024:SCB

[ACK^{+}24] Amal Abid, Saoussen Cheikhrouhou, Slim Kallel, Zahir Tari, and Mohamed Jmaiel. A smart contractbased access control framework for smart healthcare systems. The Computer Journal, 67(2):407– 422, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/407/6965939.

Avoine:2023:RTH

[ACLA23] Gildas Avoine, Xavier Carpent, and Diane Leblanc-Albarel. Rainbow tables: How far can CPU go? The Computer Journal, 66(12):3029–3037, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3029/6777260.

Albert:2022:FHG

[ADF22] I. Edwin Albert, A. J. Deepa, and A. Lenin Fred. Fidelity homogeneous genesis recommendation model for user trust with item ratings. The Computer Journal, 65(6):1639–1652, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1639/6572852.

Angelini:2024:LPG

[ADFS24] Patrizio Angelini, Giordano Da Lozzo, Henry Förster, and Thomas Schneck. 2-layer k-planar graphs density, crossing lemma, relationships and pathwidth. The Computer Journal, 67(3):1005–1016, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1005/7135863.

Altisen:2023:SSS

[ADJ23] Karine Altisen, Stéphane Devismes, and Erwan Jahier. sasa: a SimulAtor of Self-stabilizing Algorithms. The Computer Journal, 66(4):796– 814, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/796/6499477.

Ataie:2022:HML

[AEGA22] Ehsan Ataie, Athanasia Evangelinou, Eugenio Gianniti, and Danilo Ardagna. A hybrid machine learning approach for performance modeling of cloud-based big data applications. The Computer Journal, 65(12):3123–3140, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3123/6372951.

See [AEGA23].

Ataie:2023:EHM

[AEGA23] Ehsan Ataie, Athanasia Evangelinou, Eugenio Gianniti, and Danilo Ardagna. Erratum to: A hybrid machine learning approach for performance modeling of cloud-based big data applications. The Computer Journal, 66(2): 524, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/524/6412923.

See [AEGA22].

Afdhal:2020:ERH

[AEZ20] Rim Afdhal, Ridha Ejbali, and Mourad Zaied. Emotion recognition by a hybrid system based on the features of distances and the shapes of the wrinkles. The Computer Journal, 63(3): 351–363, March 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/351/5466467.

Afdhal:2021:PER

[AEZ21] Rim Afdhal, Ridha Ejbali, and Mourad Zaied. Primary emotions and recognition of their intensities. The Computer Journal, 64(12):1848–1860, December 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1848/5717857.

Amin-Ghafari:2024:APC

[AGA24] Vahid Amin-Ghafari and Mahmoud Ahmadian Attari. An attack on a proposed construction of smallstate stream ciphers and proposals for new constructions. The Computer Journal, 67(1):169–178, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/169/6847608.

See correction [Ano24b].

Arboleda:2024:QOT

[AGZ24] Francisco Javier Moreno Arboleda, Georgia Garani, and Carlos Daniel Bolivar Zapata. Query operators for transactional data: Detecting similar and periodic transactions. The Computer Journal, 67(2):437– 446, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/437/6955255.

Alajlan:2022:EAC

[AHI22] Norah Alajlan, Mohammed Hadwan, and Dina M. Ibrahim. Effectiveness of adopting cloud-based elearning at Qassim University. The Computer Journal, 65(5):1098–1106, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1098/6012868.

Alaya:2020:QEV

[AK20] Bechir Alaya and Rehanullah Khan. QoS enhancement in VoD systems: Load management and replication policy optimization perspectives. The Computer Journal, 63(10):1547– 1563, October 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1547/5865453.

Aboluhom:2024:FRU

[AK24] Abdulatif Ahmed Ali Aboluhom and Ismet Kandilli. Face recognition using deep learning on Raspberry Pi. The Computer Journal, 67(10):3020–3030, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/10/3020/7754455.

Ata:2021:AIE

[AKA^{+}21] Ayesha Ata, Muhammad Adnan Khan, Sagheer Abbas, Muhammad Saleem Khan, and Gulzar Ahmad. Adaptive IoT empowered smart road traffic congestion control system using supervised machine learning algorithm. The Computer Journal, 64(11):1672– 1679, November 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1672/5838271.

Abbas:2022:ESC

[AKA^{+}22] Sagheer Abbas, Muhammad Adnan Khan, Atifa Athar, Syed Ali Shan, Anwar Saeed, and Tahir Alyas. Enabling smart city with intelligent congestion control using hops with a hybrid computational approach. The Computer Journal, 65(3):484– 494, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/484/5866138.

Ali:2021:FLA

[AKD^{+}21] Waqar Ali, Rajesh Kumar, Zhiyi Deng, Yansong Wang, and Jie Shao. A federated learning approach for privacy protection in context-aware recommender systems. The Computer Journal, 64(7): 1016–1027, July 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/1016/6259634.

Alyari:2020:SNR

[AKI20] Robab Alyari, Jaber Karimpour, and Habib Izadkhah. Specifying a new requirement model for secure adaptive systems. The Computer Journal, 63(8):1148–1167, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1148/5645559.

Alawatugoda:2020:PKE

[Ala20] Janaka Alawatugoda. Publickey encryption in the standard model against strong leakage adversary. The Computer Journal, 63(12):1904–1914, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/12/1904/5850745.

Alam:2023:LSL

[Ala23] Shahid Alam. LAM: Scrutinizing leading APIs for detecting suspicious call sequences. The Computer Journal, 66(11):2638–2655, November 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2638/6649277.

Almalki:2025:DIF

[Alm25] Faris A. Almalki. Developing an intelligent framework with Blockchain capabilities for environmental monitoring using a Cube-Sat. The Computer Journal, 68(8):968–984, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/8/968/8051282.

AlShaikh:2023:NRR

[AlS23a] Muath AlShaikh. A novel reduced reference image quality assessment based on formal concept analysis. The Computer Journal, 66(7):1749–1760, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1749/6566844.

Arellanes:2023:DDF

[ALS23b] Damian Arellanes, Kung-Kiu Lau, and Rizos Sakellariou. Decentralized data flows for the functional scalability of service-oriented IoT systems. The Computer Journal, 66(6):1477– 1506, June 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1477/6552276.

Alserhani:2025:ASM

[Als25] Faeiz Alserhani. Advanced social media crime prevention via deep learning and cryptographic data encryption. The Computer Journal, 68(9):1150–1162, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1150/8152776.

Al-Muhammed:2020:LGN

[AM20] Muhammed Jassem Al-Muhammed. Lookbackguess-next optimizer: Feedbackguided random search technique with biased mapping for solving unconstrained optimization problems. The Computer Journal, 63(5): 791–816, May 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/5/791/5487028.

Ashur:2023:SIF

[AM23] Tomer Ashur and Chris J. Mitchell. Special issue on failed approaches and insightful losses in cryptology — foreword. The Computer Journal, 66(6):1311, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1311/7152931.

Ahmed:2024:EAB

[AMA^{+}24] Hunar Abubakir Ahmed, Jafar Majidpour, Mohammed Hussein Ahmed, Samer Kais Jameel, and Amir Majidpour. Enhancing auditory brainstem response classification based on vision transformer. The Computer Journal, 67(5): 1872–1878, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1872/7335816.

Ashur:2023:HDE

[AMT23] Tomer Ashur, Mohammad Mahzoun, and Dilara Toprakhisar. How not to design an efficient FHE-friendly block cipher: Seljuk. The Computer Journal, 66(6):1312–1319, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1312/6986901.

Abidi:2022:HMC

[ANA^{+}22] Afef Abidi, Ibtihel Nouira, Ines Assali, Mohamed Ali Saafi, and Mohamed Hedi Bedoui. Hybrid multichannel EEG filtering method for ocular and muscular artifact removal based on the 3D spline interpolation technique. The Computer Journal, 65(5):1257– 1271, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1257/6066623.

Al-Naffakh:2025:RWC

[ANCLHD25] N. Al-Naffakh, N. Clarke, F. Li, and P. Haskell-Dowland. Real-world continuous smartwatch-based user authentication. The Computer Journal, 68(7): 717–733, July 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/717/7953397.

Abdul:2020:CWH

[ANG20] Wadood Abdul, Ohoud Nafea, and Sanaa Ghouzali. Combining watermarking and hyper-chaotic map to enhance the security of stored biometric templates. The Computer Journal, 63(3):479–493, March 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/479/5510728.

Angelides:2024:TEM

[Ang24] Marios C. Angelides. Thematic editorial: Mostly artificial intelligence (AI) or machine learning (ML) now in the engine room, in pursuit of a green agenda. The Computer Journal, 67(1):1– 2, January 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/1/7462043.

Ahmed-Nacer:2022:MDS

[ANKZ^{+}22] Mehdi Ahmed-Nacer, Slim Kallel, Faiez Zalila, Philippe Merle, and Walid Gaaloul. Model-driven simulation of elastic OCCI cloud resources. The Computer Journal, 65(5):1144–1166, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1144/6043303.

Anonymous:2022:WE

[Ano22] Anonymous. Withdrawn:

Editorial. The Computer Journal, 65(5):??, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/bxr040/2609372.

A blank Editorial was assigned to this DOI in error, which has been withdrawn.

Anonymous:2023:CAL

[Ano23a] Anonymous. Correction to:

Adaptive Laplacian support vector machine for semi-supervised learning. The Computer Journal, 66 (8):2075, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/2075/6563405.

See [HZW21].

Anonymous:2023:CIF

[Ano23b] Anonymous. Correction to:

Intelligent forecast of stock markets to handle COVID-19 economic crisis by modified generative adversarial networks. The Computer Journal, 66(10):2593, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2593/6691910.

See [SAK22].

Anonymous:2023:CLD

[Ano23c] Anonymous. Correction to: Learning disjunctive multiplicity expressions and disjunctive generalize multiplicity expressions from both positive and negative examples. The Computer Journal, 66(9):2329, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/9/2329/6665424.

See [LCC23].

Anonymous:2023:CMO

[Ano23d]  Anonymous.  Corrigen-

dum to: A multi-objective randomly updated beetle swarm and multi-verse optimization for brain tumor segmentation and classification. The Computer Journal, 66(6):1564, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1564/6527554.

See [KB22].

Anonymous:2023:CID

[Ano23e] Anonymous. Corrigendum to: IoT data quality issues and potential solutions: a literature review. The Computer Journal, 66(6):1563, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1563/6529197.

See [MMMZ23].

Anonymous:2024:CSE

[Ano24a] Anonymous. Correction to: A semantic embedding enhanced topic model for user-generated textual content modeling in social ecosystems. The Computer Journal, 67(4):1604, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1604/7238254.

See [ZLL^{+}22].

Anonymous:2024:CAP

[Ano24b] Anonymous. Correction to: An attack on a proposed construction of smallstate stream ciphers and proposals for new constructions. The Computer Journal, 67(2):806, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/806/7051110.

See [AGA24].

Anonymous:2024:CAD

[Ano24c] Anonymous. Correction to: Automatic diagnosis of diabetic retinopathy from retinal abnormalities: Improved jaya-based feature selection and recurrent neural network. The Computer Journal, 67(5):2007, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/2007/7342308.

See [RK22].

Anonymous:2024:CHT

[Ano24d] Anonymous. Correction to: Hitting times of random walks on edge corona product graphs. The Computer Journal, 67(3):1210, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1210/7158425.

See [ZXL^{+}24].

Anonymous:2024:CME

[Ano24e] Anonymous. Correction to: Metaheuristic-enabled artificial neural network framework for multimodal biometric recognition with local fusion visual features. The Computer Journal, 67 (4):1603, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1603/7188730.

See [Gok22].

Anonymous:2024:CPI

[Ano24f] Anonymous. Correction to: Perceptual image hashing based on canny operator and tensor for copy– move forgery detection. The Computer Journal, 67(3): 1209, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1209/7075504.

See [LGX^{+}24].

Anonymous:2024:CPE

[Ano24g] Anonymous. Correction to:

Performance evaluation of FPGA-based LSTM neural networks for pulse signal detection on real-time radar warning receivers. The Computer Journal, 67(2): 807, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/807/7055812.

See [TAA23].

Anta:2024:TEE

[Ant24] Antonio Fernández Anta.

Thematic editorial: edge computing, fog computing, and internet of things. The Computer Journal, 67(9):2721–2724, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2721/7811253.

Abdalla:2021:SAF

[AÖ21] Ghazi Abdalla and Fatih Özyurt. Sentiment analysis of fast food companies with deep learning models. The Computer Journal, 64(3): 383–390, March 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/383/5926765.

Alrahhal:2022:CDS

[AP22] Maher Alrahhal and Supreethi K P. COVID-19 diagnostic system using medical image classification and retrieval: a novel method for image analysis. The Computer Journal, 65(8):2146–2163, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/2146/6278158.

Aprem:2021:BOA

[AR21] Anup Aprem and Stephen Roberts. A Bayesian optimization approach to compute Nash equilibrium of potential games using bandit feedback. The Computer Journal, 64(12):1801–1813, December 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1801/5650974.

AlShboul:2022:SNA

[ARAA^{+}22] Bashar Al Shboul, Abdullateef Rabab’ah, Mahmoud Al-Ayyoub, Yaser Jararweh, and Thar Baker. Social network analysis of the Panama Papers concentrating on the MENA region. The Computer Journal, 65(9):2493–2505, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2493/6308677.

Adiwal:2021:RPD

[ARDP21] Sanjay Adiwal, Balaji Rajendran, Pushparaj Shetty D, and Gopinath Palaniappan. Revisiting the performance of DNS queries on a DNS hierarchy testbed over dual-stack. The Computer Journal, 64(6):843– 859, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/843/6169062.

Ali:2024:MSA

[AS24a] Hala S. Ali and R. Sridevi.

Mobility and security aware real-time task scheduling in fog–cloud computing for IoT devices: a fuzzylogic approach. The Computer Journal, 67(2):782– 805, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/782/7107926.

Antony:2024:BBP

[AS24b] Amalan Joseph Antony and Kunwar Singh. A blockchain-based public key infrastructure for IoTbased healthcare systems. The Computer Journal, 67 (4):1531–1537, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/67/4/1531/
                                        7235532.

Aruchamy:2025:ESM

[ASCJ25] Prasanth Aruchamy, M. Sundarrajan, Mani Deepak Choudhry, and Akshya Jothi. Enhancing security of medical images using code-based intermittent encryption and convolutional neural network. The Computer Journal, 68(5):539– 551, May 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/5/539/7921447.

Alam:2023:FIB

[ASK^{+}23] Talha Mahboob Alam, Kamran Shaukat, Adel Khelifi, Hanan Aljuaid, Malaika Shafqat, Usama Ahmed, Sadeem Ahmad Nafees, and Suhuai Luo. A fuzzy inference-based decision support system for disease diagnosis. The Computer Journal, 66(9):2169– 2180, September 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2169/6603446.

Alam:2021:CBP

[ASM^{+}21] Talha Mahboob Alam, Kamran Shaukat, Mubbashar Mushtaq, Yasir Ali, Matloob Khushi, Suhuai Luo, and Abdul Wahab. Corporate bankruptcy prediction: an approach towards better corporate world. The Computer Journal, 64(11):1731–1746, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/11/1731/5856206.

Alam:2022:MLA

[ASM^{+}22] Talha Mahboob Alam, Kamran Shaukat, Haris Mahboob, Muhammad Umer Sarwar, Farhat Iqbal, Adeel Nasir, Ibrahim A. Hameed, and Suhuai Luo. A machine learning approach for identification of malignant mesothelioma etiological factors in an imbalanced dataset. The Computer Journal, 65(7):1740– 1751, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/7/1740/6259657.

A:2022:LCD

[ASPB22] Selvapandian A, Nagendra Prabhu S, Sivakumar P, and Jagannadha Rao D B. Lung cancer detection and severity level classification using sine cosine sail fish optimization based generative adversarial network with CT images. The Computer Journal, 65(6):1611– 1630, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1611/6399871.

Al-Share:2021:CLB

[ASSA21] Rama Al-Share, Mohammad Shurman, and Abdallah Alma’aitah. A collaborative learning-based algorithm for task offloading in UAV-aided wireless sensor networks. The Computer Journal, 64(10):1575–1583, October 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/10/1575/6326765.

Al-Shamiri:2025:UDM

[ASYLAM25] Abdulkawi Yahya Radman Al-Shamiri, Dong-Jun Yu, Peipei Li, and Balqis Yahya Ali Abdullah Al-Mahweeti. Use of data mining in identifying the risk factors of optic neuropathy. The Computer Journal, 68(9):1181–1196, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1181/8242813.

Ajmera:2024:DVM

[AT24] Kashav Ajmera and Tribhuwan Kumar Tewari. Dynamic virtual machine scheduling using residual optimum power-efficiency in the cloud data center. The Computer Journal, 67(3):1099–1110, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1099/7161852.

Abduljabbar:2021:SPS

[ATZ^{+}21] Tamara Abdulmunim Abduljabbar, Xiaohui Tao, Ji Zhang, Xujuan Zhou, Lin Li, and Yi Cai. A survey of privacy solutions using blockchain for recommender systems: Current status, classification and open issues. The Computer Journal, 64(7):1104– 1129, July 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/7/1104/6289879.

Ameen:2020:PNN

[AV20] Salem Ameen and Sunil Vadera. Pruning neural networks using multi-armed bandits. The Computer Journal, 63(7):1099–1108, July 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/7/1099/5574718.

Albert:2022:HMP

[AV22] Michael Albert and Vincent Vatter. How many popstacks does it take to sort a permutation? The Computer Journal, 65(10):2610– 2614, October 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/10/2610/6316088.

Ali:2021:BBF

[AVA21] Muhammad Salek Ali, Massimo Vecchio, and Fabio Antonelli. A blockchainbased framework for IoT data monetization services. The Computer Journal, 64(2):195–210, February 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/2/195/5911075.

Aouar:2024:DPS

[AYS^{+}24] Aissam Aouar, Saïd Yahiaoui, Lamia Sadeg, Nadia Nouali-Taboudjemat, and Kadda Beghdad Bey. Distributed partial simulation for graph pattern matching. The Computer Journal, 67(1): 110–126, January 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/67/1/110/
6832432.

Altulyan:2022:SRS

[AYW^{+}22] May Altulyan, Lina Yao, Xianzhi Wang, Chaoran Huang, Salil S. Kanhere, and Quan Z. Sheng. A survey on recommender systems for Internet of Things: Techniques, applications and future directions. The Computer Journal, 65(8):2098–2132, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2098/6278157.

Bu:2023:RCS

[B^{+}23] Dengli Bu et al. Reversible circuit synthesis method using sub-graphs of shared functional decision diagrams. The Computer Journal, 66(10):2574–2592, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/10/2574/6652787.

Bharathi:2023:SCT

[BA23a] G. Bharathi and G. Anandharaj. Sentiment classification of tourist’s opinion on tourist places of interest in South India using Tweet reviews. The Computer Journal, 66(4):815– 825, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/4/815/6463427.

Bhutta:2023:EBC

[BA23b] Areeb Ahmed Bhutta and Hasnain Ahmed. Edgebased congestion-aware datacenter load balancing with smart probing. The Computer Journal, 66(12):2908–2920, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2908/6768050.

Badawi:2024:DLB

[Bad24]  Soran Badawi.  Deep

learning-based cyberbullying detection in Kurdish language. The Computer Journal, 67(7):2548–2558, July 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2548/7618088.

Brito:2024:DCD

[BATN24] Luiz Fernando Afra Brito, Marcelo Keese Albertini, Bruno Augusto Nassif Travençolo, and Gonzalo Navarro. Dynamic compact data structure for temporal reachability with unsorted contact insertions. The Computer Journal, 67(10):2984–2994, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2984/7717688.

Booher:2024:FHS

[BBD^{+}24] Jeremy Booher, Ross Bowden, Javad Doliskani, Tako Boris Fouotsa, Steven D. Galbraith, Sabrina Kunzweiler, Simon-Philipp Merz, Christophe Petit, Benjamin Smith, Katherine E. Stange, Yan Bo Ti, Christelle Vincent, José Felipe Voloch, Charlotte Weitkämper, and Lukas Zobernig. Failing to hash into supersingular isogeny graphs. The Computer Journal, 67(8):2702– 2719, August 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2702/7681095.

Boinski:2022:ODA

[BC22] Tomasz Boiński and Paweł Czarnul. Optimization of data assignment for parallel processing in a hybrid heterogeneous environment using integer linear programming. The Computer Journal, 65(6):1412– 1433, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1412/6127452.

Blundo:2022:RMH

[BCS22] Carlo Blundo, Stelvio Cimato, and Luisa Siniscalchi. Role mining heuristics for permissionrole-usage cardinality constraints. The Computer Journal, 65(6):1386–1411, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1386/6134261.

Bernardeschi:2023:CSF

[BDFP23] Cinzia Bernardeschi, Andrea Domenici, Adriano Fagiolini, and Maurizio Palmieri. Co-simulation and formal verification of co-operative drone control with logic-based specifications. The Computer Journal, 66(2):295–317, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/295/6408792.

Bekos:2021:HAT

[BFG^{+}21] Michael A. Bekos, Henry Förster, Christian Geckeler, Lukas Holländer, Michael Kaufmann, Amadäus M. Spallek, and Jan Splett. A heuristic approach towards drawings of graphs with high crossing resolution. The Computer Journal, 64(1):7–26, January 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/7/5645631.

Bidoki:2022:TSM

[BFM22] M. Bidoki, M. Fakhrahmad, and M. R. Moosavi. Text summarization as a multi-objective optimization task: Applying harmony search to extractive multi-document summarization. The Computer Journal, 65(5):1053–1072, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1053/6018134.

Boyd:2021:MVF

[BG21] Colin Boyd and Kai Gellert.

A modern view on forward security. The Computer Journal, 64(4):639– 652, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/4/639/5896207.

Brahmam:2024:NNO

[BG24] M. Veera Brahmam and S. Gopikrishnan. NOD-STAC: Novel outlier detection technique based on spatial, temporal and attribute correlations on IoT bigdata. The Computer Journal, 67(3):947– 960, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/947/7110816.

Behera:2025:SFD

[BG25] Rakesh Ranjan Behera and Debasis Gountia. A secure fault detection for digital microfluidic biochips. The Computer Journal, 68(3): 217–227, March 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/3/217/7874331.

Brisaboa:2024:TDB

[BGGBN24] Nieves R. Brisaboa, Travis Gagie, Adrián Gómez-Brandón, and Gonzalo Navarro. Two-dimensional block trees. The Computer Journal, 67(1):391– 406, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/391/6955257.

Biswas:2022:SBV

[BGHG22] Biswajit Biswas, Swarup Kr Ghosh, Moumita Hore, and Anupam Ghosh. SIFTbased visual tracking using optical flow and belief propagation algorithm. The Computer Journal, 65(1):1– 17, January 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/1/5686229.

Burns:2020:DSC

[BHJ20] Alan Burns, Ian J. Hayes, and Cliff B. Jones. Deriving specifications of control programs for cyber physical systems. The Computer Journal, 63(5):774– 790, May 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/774/5481807.

Bulbul:2021:IHA

[BIZA21] Mohammad Farhad Bulbul, Saiful Islam, Yatong Zhou, and Hazrat Ali. Improving human action recognition using hierarchical features and multiple classifier ensembles. The Computer Journal, 64(11):1633–1655, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1633/5625673.

Bradley:2022:BAS

[BJ22] Patrick Erik Bradley and Markus Wilhelm Jahn. On the behaviour of p-adic scaled space filling curve indices for high-dimensional data. The Computer Journal, 65(2):310–330, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/310/5869134.

Bonab:2023:ERA

[BK23] Mohammad Jalilvand Aghdam Bonab and Ramin Shaghaghi Kandovan. Effective resource allocation and load balancing in hierarchical HetNets: Toward QoSaware multi-access edge computing. The Computer Journal, 66(1):229– 244, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/229/6411851.

Benza:2021:SCS

[BKS21] Ekaterina Benza, Shmuel T. Klein, and Dana Shapira.

Smaller compressed suffix arrays. The Computer Journal, 64(5):721– 730, May 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/721/5819212.

Baruch:2022:ECS

[BKS22] Gilad Baruch, Shmuel T.

Klein, and Dana Shapira. Enhanced context sensitive flash codes. The Computer Journal, 65(5):1200– 1210, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1200/6071858.

Bindel:2023:NBE

[BM23] Nina Bindel and Sarah Mc-Carthy. The need for being explicit: Failed attempts to construct implicit certificates from lattices. The Computer Journal, 66(6): 1320–1334, June 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1320/6779419.

Bilal:2025:CBG

[BMA^{+}25] Ahmad Bilal, Hamid Turab Mirza, Adnan Ahmad, Ibrar Hussain, Ali Raza, Helena Garay, Josep Alemany Iturriaga, and Imran Ashraf. On the correlation between Google Play Store application icons and downloads. The Computer Journal, 68(10):1579–1593, October 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1579/8163570.

Bouillaguet:2022:CME

[BMV22] Charles Bouillaguet, Florette Martinez, and Damien Vergnaud. Cryptanalysis of modular exponentiation outsourcing protocols. The Computer Journal, 65(9):2299–2314, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2299/6289878.

Bulbul:2022:OCC

[BÖI22] Mehmet Akif Bülbül, Celal Öztürk, and Mehmet Fatih Işık. Optimization of climatic conditions affecting determination of the amount of water needed by plants in relation to their life cycle with particle swarm optimization, and determining the optimum irrigation schedule. The Computer Journal, 65(10):2654–2663, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2654/6324867.

BorgesSantos:2022:DTE

[BPNdQ22] Veruska Borges Santos, Carlos Eduardo S. Pires, Dimas Cassimiro Nascimento, and Andreza Raquel M. de Queiroz. A decision tree ensemble model for predicting bus bunching. The Computer Journal, 65(8):2044–2062, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2044/6269132.

B:2021:PAP

[BRA21] Nithya B., Naveen Ranjan, and Justin Gopinath A. Performance analysis of prioritization and contention control algorithm in wireless body area networks. The Computer Journal, 64(2):211–223, February 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/211/6032260.

Boudgoust:2024:OTL

[BRL24] Katharina Boudgoust and Adeline Roux-Langlois. Overfull: Too large aggregate signatures based on lattices. The Computer Journal, 67(2):719– 727, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/719/7079712.

Baliyan:2021:RBS

[BS21] Niyati Baliyan and Aarti Sharma. Related blogs’ summarization with natural language processing. The Computer Journal, 64(3):347–357, March 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/347/5901650.

Brahmi:2023:CSB

[BS23] Zaki Brahmi and Afef Selmi. Coordinate systembased trust-aware Web services composition in edge and cloud environment. The Computer Journal, 66(9):2102–2117, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2102/6589511.

Behal:2024:IAM

[BS24a] Veerawali Behal and Ramandeep Singh. An intelligent air monitoring system for pollution prediction: a predictive healthcare perspective. The Computer Journal, 67(5):1763– 1782, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1763/7292002.

Bel:2024:IRB

[BS24b] K. Nalini Sujantha Bel and I. Shatheesh Sam. Image retrieval based on autoencoder and clustering with centroid update. The Computer Journal, 67(11):3031– 3041, November 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/11/3031/7717690.

Beniwal:2024:HBC

[BS24c] Rohit Beniwal and Pavi Saraswat. A hybrid BERT-CNN approach for depression detection on social media using multimodal data. The Computer Journal, 67(7):2453– 2472, July 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2453/7614266.

Borges:2024:CBC

[BS24d] Ricard Borges and Francesc Sebé. An e-coin based construction for unlinkable priced oblivious transfer. The Computer Journal, 67(3):933–940, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/933/7098670.

Bedi:2021:UID

[BSM21] Anterpreet Kaur Bedi, Ramesh Kumar Sunkaria, and Deepti Mittal. Ultrasound image despeckling and enhancement using modified multiscale anisotropic diffusion model in non-subsampled shearlet domain. The Computer Journal, 64(12):1785–1800, December 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/12/1785/5670504.

Bergstra:2023:ACM

[BT23a] Jan A. Bergstra and John V. Tucker. On the axioms of common meadows: Fracterm calculus, flattening and incompleteness. The Computer Journal, 66(7):1565–1572, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1565/6562943.

Bhatt:2023:ODB

[BT23b] Abhishek Bhatt and Vandana Thakur. An optimized deep belief network for land cover classification using syntheticaperture radar images and Landsat images. The Computer Journal, 66(8):2043– 2058, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/2043/6649281.

Bergstra:2024:ETR

[BT24] Jan A. Bergstra and John V. Tucker. Eager term rewriting for the fracterm calculus of common meadows. The Computer Journal, 67(5):1866–1871, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1866/7335813.

Bao:2022:WMG

[BTT^{+}22] Pham The Bao, Tran Anh Tuan, Le Nhi Lam Thuy, Jin Young Kim, and João Manuel R. S. Tavares. White matter, gray matter and cerebrospinal fluid segmentation from brain magnetic resonance imaging using adaptive U-net and local convolutional neural network. The Computer Journal, 65(12):3081–3090, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3081/6370299.

Bao:2022:BDP

[BXZ22] Qi Bao, Wanyue Xu, and Zhongzhi Zhang. Benchmark for discriminating power of edge centrality metrics. The Computer Journal, 65(12):3141–3155, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3141/6370298.

Burukanli:2024:CVM

[BY24] Mehmet Burukanli and Nejat Yumuşak. COVID-19 virus mutation prediction with LSTM and attention mechanisms. The Computer Journal, 67(10):2934– 2944, October 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/10/2934/7708788.

Botangen:2021:ICP

[BYYS21] Khavee Agustus Botangen, Jian Yu, Wai Kiang Yeap, and Quan Z. Sheng. Integrating context to preferences and goals for goaloriented adaptability of software systems. The Computer Journal, 64(5): 675–706, May 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/675/5709019.

Ba:2022:SSC

[BZ22] Lina Ba and Heping Zhang.

The star-structure connectivity and star-substructure connectivity of hypercubes and folded hypercubes. The Computer Journal, 65(12):3156–3166, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3156/6370844.

Ba:2025:SCF

[BZ25] Lina Ba and Heping Zhang.

Structure connectivity of folded cross cubes. The Computer Journal, 68(5): 502–509, May 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/502/7930079.

Ba:2023:PSC

[BZZ23] Lina Ba, Yaxian Zhang, and Heping Zhang. The path-structure connectivity of augmented k-ary ncubes. The Computer Journal, 66(12):3119–3128, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3119/7075448.

Chen:2023:RHR

[C^{+}23a] Lv Chen et al. RTIM hashing: Robust and compact video hashing with a rotation- and translationinvariant model. The Computer Journal, 66(11):2741– 2757, November 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2741/6691913.

Chen:2023:NAS

[C^{+}23b] Yi Chen et al. Neural-aided statistical attack for cryptanalysis. The Computer Journal, 66(10):2480–2498, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/10/2480/6645489.

Ciftci:2023:DTD

[ÇA23] Canan Çiftçi and Vecd˙

Aytaç. Disjunctive total domination in Harary graphs. The Computer Journal, 66(12):2990–2999, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2990/6768048.

Coban:2025:ATB

[ÇA25] Önder Çoban and Şeyma Yücel Altay. Arming text-based gender inference with partition membership filtering and feature selection for online social network users. The Computer Journal, 68(9):1208–1224, September 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1208/8116799.

Caliskan:2022:NEA

[Çal22] Abidin Çalışkan. A new ensemble approach for congestive heart failure and arrhythmia classification using shifted one-dimensional local binary patterns with long short-term memory. The Computer Journal, 65(9):2535–2546, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2535/6643451.

Castro-Bleda:2020:NDC

[CBEBPPZM20] M. J. Castro-Bleda, S. España-Boquera, J. Pastor-Pellicer, and F. Zamora-Martínez. The NoisyOffice database: a corpus to train supervised machine learning filters for image processing. The Computer Journal, 63(11):1658–1667, November 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1658/5648014.

Chakraborty:2024:GHM

[CBM24] Abhinav Chakraborty, Subhash Bhagat, and Krishnendu Mukhopadhyaya. Gathering over heterogeneous meeting nodes. The Computer Journal, 67(5): 1794–1813, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1794/7292000.

Chen:2022:ACV

[CC22] Wenbin Chen and Jianer Chen. Approximating closest vector problem in ℓ_{∞}- norm revisited. The Computer Journal, 65(12):3100– 3105, December 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3100/6369598.

Chang:2024:PSP

[CC24] Yeim-Kuan Chang and Hsin-Mao Chen. Partitioned 2D set-pruning segment trees with compressed buckets for multidimensional packet classification. The Computer Journal, 67(6):2189–2207, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2189/7596736.

Cai:2023:MRP

[CCC^{+}23] Saihua Cai, Jinfu Chen, Haibo Chen, Chi Zhang, Qian Li, Dengzhou Shi, and Wei Lin. Minimal rare pattern-based outlier detection approach for uncertain data streams under monotonic constraints. The Computer Journal, 66(1): 16–34, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/16/6381823.

Cui:2024:BEZ

[CCY^{+}24] Handong Cui, Kwan Yin Chan, Tsz Hon Yuen, Xin Kang, and Cheng-Kang Chu. Bandwidth-efficient zero-knowledge proofs for threshold ECDSA. The Computer Journal, 67(4): 1265–1278, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1265/7188842.

Cheng:2022:DMD

[CCZ^{+}22] Yanan Cheng, Tingting Chai, Zhaoxin Zhang, Keyu Lu, and Yuejin Du. Detecting malicious domain names with abnormal WHOIS records using feature-based rules.

The Computer Journal, 65(9):2262–2275, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2262/6281305.

Correa:2021:STA

[CDCN21] Cléber G. Corrêa, Márcio E. Delamaro, Marcos L. Chaim, and Fátima L. S. Nunes. Software testing automation of VR-based systems with haptic interfaces. The Computer Journal, 64(5): 826–841, May 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/826/5856158.

Chen:2024:CBE

[CDR^{+}24] Zhiwei Chen, Lunzhi Deng, Yu Ruan, Shuai Feng, Tao Wang, and Bo Wang. Certificateless broadcast encryption with authorization suitable for storing personal health records. The Computer Journal, 67(2):617– 631, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/617/7024504.

Chen:2024:HBT

[CEN24] Liqun Chen, Nada El Kassem, and Christopher J. P. Newton. How to bind a TPM’s attestation keys with its endorsement key. The Computer Journal, 67(3):988–1004, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/988/7130967.

Cristia:2023:DPI

[CFR23] Maximiliano Cristiá, Andrea Fois, and Gianfranco Rossi. Declarative programming with intensional sets in Java using JSetL. The Computer Journal, 66(3):763– 784, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/763/6521749.

Cavalcanti:2024:LTS

[CFRS24] Ana Cavalcanti, Madiel Conserva Filho, Pedro Ribeiro, and Augusto Sampaio. Laws of timed state machines. The Computer Journal, 67(6):2066–2107, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2066/7492011.

Chen:2025:CBS

[CGQ^{+}25] Ruidong Chen, Shuiping Guo, Baohua Qiang, Xianyi Yang, Pingping Sun, Shihao Zhang, and Yuan Xie. CLIP-based semantic refinement method for imagetext retrieval. The Computer Journal, 68(10):1377– 1385, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1377/8120571.

Cai:2022:LAM

[CGY22] Binlei Cai, Qin Guo, and Junfeng Yu. LraSched: Admitting more longrunning applications via auto-estimating container size and affinity. The Computer Journal, 65(9):2377– 2391, September 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2377/6289887.

Chen:2025:CES

[CHQM25] Ligeng Chen, Zhongling He, Yi Qian, and Bing Mao. CATI++: empirical study and evaluation for adjacent instruction enhanced type inference. The Computer Journal, 68(7):788– 803, July 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/788/8019595.

Chang:2020:MMV

[CHT20] Hong-Yi Chang, Zih-Huan Hang, and Yih-Jou Tzang. MNVPCS: Multinode virtualpoint–based charging scheme to prolong the lifetime of wireless rechargeable sensor networks. The Computer Journal, 63(2):283– 294, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/283/5618853.

Chen:2021:PAI

[CHWM21] Haixia Chen, Xinyi Huang, Wei Wu, and Yi Mu. Privacy-aware image authentication from cryptographic primitives. The Computer Journal, 64(8): 1178–1192, August 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/64/8/1178/
5943716.

Chadha:2021:SPS

[CK21] Akshma Chadha and Baijnath Kaushik. A survey on prediction of suicidal ideation using machine and ensemble learning. The Computer Journal, 64(11):1617–1632, November 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1617/5612727.

Chadha:2022:PEL

[CK22] Akshma Chadha and Baijnath Kaushik. Performance evaluation of learning models for identification of suicidal thoughts. The Computer Journal, 65(1):139– 154, January 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/1/139/6281304.

Cristin:2021:SLC

[CKA21] R. Cristin, K. Suresh Kumar, and P. Anbhazhagan. Severity level classification of brain tumor based on MRI images using fractional-chicken swarm optimization algorithm. The Computer Journal, 64(10):1514–1530, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/10/1514/6287331.

Cristia:2022:PAT

[CKR22] Maximiliano Cristiá, Ricardo D. Katz, and Gianfranco Rossi. Proof automation in the theory of finite sets and finite set relation algebra. The Computer Journal, 65(7):1891– 1903, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1891/6262250.

Chen:2025:TEA

[CL25] Xiaorui Chen and Xiangxue Li. TinyThunder: enabling asynchronous Byzantine fault tolerance with optimal communication efficiency. The Computer Journal, 68(4):407– 418, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/407/7945155.

Cheng:2019:MBO

[CLC^{+}19] Binlin Cheng, Jinjun Liu, Jiejie Chen, Shudong Shi, Xufu Peng, Xingwen Zhang, and Haiqing Hai. MoG: Behavior-obfuscation resistance malware detection. The Computer Journal, 62(12):1734–1747, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1734/5510727.

See erratum [CLJ^{+}22].

Chen:2020:PAB

[CLG^{+}20] Yang Chen, Wenmin Li, Fei Gao, Kaitai Liang, Hua Zhang, and Qiaoyan Wen. Practical attribute-based conjunctive keyword search scheme. The Computer Journal, 63(8):1203–1215, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1203/5667449.

Cheng:2022:CAS

[CLH22] Rung-Shiang Cheng, Yin-Ming Li, and Chung-Ming Huang. The collision avoidance and situation-aware media access scheme using the registered-backoff-time method for the IEEE 802.11ah-based IoT wireless networks. The Computer Journal, 65(8):1977– 1997, August 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/1977/6262252.

Chen:2025:RTS

[CLH^{+}25] Xinyi Chen, Bo Liu, Xueshu Hong, Wangqun Chen, Houjie Qiu, and Xingkong Ma. Relation as text: a semantic-preserving method for relational triple extraction. The Computer Journal, 68(4):346– 354, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/346/7994664.

Cheng:2022:EMB

[CLJ^{+}22] Binlin Cheng, Jinjun Liu, Chen Jiejie, Shi Shudong, Peng Xufu, Zhang Xingwen, and Haiqing Hai. Erratum to: MoG: Behaviorobfuscation resistance malware detection. The Computer Journal, 65(10):2846, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2846/6154265.

See [CLC^{+}19].

Chen:2024:WBI

[CLL^{+}24] Jie Chen, Yinuo Luo, Jun Liu, Chao Wang, Yueyu Zhang, and Xiaoli Dong. A white-box implementation of SM4 with self-equivalence encoding. The Computer Journal, 67(3):1087–1098, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1087/7190927.

Chen:2020:BPP

[CLM^{+}20] Lixia Chen, Jian Li, Ruhui Ma, Haibing Guan, and Hans-Arno Jacobsen. Balancing power and performance in HPC clouds. The Computer Journal, 63(6): 880–899, June 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/6/880/5681778.

Cao:2022:EDL

[CLM^{+}22] Jin Cao, Liwei Lin, Ruhui Ma, Haibing Guan, Mengke Tian, and Yong Wang. An efficient deep learning approach to IoT intrusion detection. The Computer Journal, 65(11):2870–2879, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/11/2870/6732677.

Chen:2024:GBB

[CLWH24] Xiuyuan Chen, Chao Lin, Wei Wu, and Debiao He. A general blockchain-based automatic audit scheme for proofs of retrievability. The Computer Journal, 67(6): 2219–2229, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2219/7590677.

Chen:2022:EDO

[CLX^{+}22] Peng Chen, Hongyun Liu, Ruyue Xin, Thierry Carval, Jiale Zhao, Yunni Xia, and Zhiming Zhao. Effectively detecting operational anomalies in large-scale IoT data infrastructures by using a GAN-based predictive model. The Computer Journal, 65(11):2909–2925, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2909/6741149.

Chen:2024:TMT

[CLX^{+}24] Zhenhua Chen, Ting Li, Junrui Xie, Ni Li, and Jingjing Nie. K th min threshold encryption for privacy-preserving data evaluation. The Computer Journal, 67(5):1941–1950, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1941/7477782.

Cao:2025:VRI

[CLY^{+}25] Yuetian Cao, Jin Liu, Zijun Yu, Xingye Li, Lai Wei, and Zhongdai Wu. Vessel re-identification by a hierarchical perceptual aggregation network with inclination-aware attention. The Computer Journal, 68(6):617–634, June 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/617/7933710.

Chen:2025:FEC

[CLYZ25] Zhijin Chen, Nankai Lin, Aimin Yang, and Dong Zhou. Filter-enhanced Contrast Variational AutoEncoders for sequential recommendation. The Computer Journal, 68(7):775– 787, July 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/775/8019596.

Chen:2024:PPB

[CLZ^{+}24] Shuangquan Chen, Jinguo Li, Kai Zhang, Aoran Di, and Mengli Lu. Privacypreserving breast cancer prediction based on logistic regression. The Computer Journal, 67(8):2667–2676, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2667/7673534.

Chen:2023:GSE

[CNL^{+}23] Zhenhua Chen, Jingjing Nie, Zhanli Li, Chunpeng Ge, and Willy Susilo. Geometric searchable encryption without false positive and its applications. The Computer Journal, 66(9):2155–2168, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2155/6608476.

Coban:2021:DLB

[ÇÖİ21] Önder Çoban, Selma Ayşe Özel, and Ali İnan. Deep learning-based sentiment analysis of Facebook data: The case of Turkish users. The Computer Journal, 64 (3):473–499, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/64/3/473/
6095851.

Chaurasia:2022:ETP

[CP22] Vikas Chaurasia and Saurabh Pal. Ensemble technique to predict breast cancer on multiple datasets. The Computer Journal, 65(10):2730–2740, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2730/6334045.

Chakre:2023:PRM

[CP23] Rahul Ramesh Chakre and Dipak V. Patil. Particle rider mutual information and dendritic-squirrel search algorithm with artificial immune classifier for brain tumor classification. The Computer Journal, 66(3):743–762, March 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/3/743/6474501.

Chen:2025:SMP

[CPLW25] Kai Chen, Bo Pang, Yongqiang Li, and Mingsheng Wang. Secure multiparty shuffling with optimal communication. The Computer Journal, 68(10):1364– 1376, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1364/8133220.

Cui:2021:TFD

[CPN^{+}21] Hui Cui, Russell Paulet, Surya Nepal, Xun Yi, and Butrus Mbimbi. Two-factor decryption: a better way to protect data security and privacy. The Computer Journal, 64(4):550– 563, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/4/550/5868155.

Chandol:2022:BCC

[CR22] Mohan Kumar Chandol and M. Kameswara Rao. Border collie cat optimization for intrusion detection system in healthcare IoT network using deep recurrent neural network. The Computer Journal, 65(12):3181–3198, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3181/6421947.

Chandran:2023:LBW

[CRA23] Sudhin Chandran, R. Rajesh, and M. Dev Anand. Laser beam welded aluminumtitanium dissimilar sheet metals: Neural network based strength and hardness prediction model. The Computer Journal, 66(5): 1053–1068, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1053/6513724.

Chuengsatiansup:2023:RRR

[CRRY23] Chitchanok Chuengsatiansup, Eyal Ronen, Gregory G. Rose, and Yuval Yarom. Row, row, row your boat: How to not find weak keys in Pilsung. The Computer Journal, 66(6):1335– 1341, June 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1335/6645991.

Chen:2023:NND

[CSYY23] Yi Chen, Yantian Shen, Hongbo Yu, and Sitong Yuan. A new neural distinguisher considering features derived from multiple ciphertext pairs. The Computer Journal, 66(6):1419– 1433, June 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1419/6545266.

Calamoneri:2023:MAV

[CT23] Tiziana Calamoneri and Daniele Tavernelli. Modeling and approximating the visit of a set of sites with a fleet of UAVs. The Computer Journal, 66(7):1586– 1594, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1586/6555407.

Cheng:2025:FSS

[CTL^{+}25] Wen Cheng, Shengxia Tu, Yi Liu, Lingfang Zeng, Yang Wang, and André Brinkmann. FlashFox: a secret-sharing approach to securing data deletion for flash-based SSD. The Computer Journal, 68(7):734– 748, July 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/7/734/8034846.

Castaneda:2024:PMD

[CvDRV24] Armando Castañeda, Hans van Ditmarsch, David A. Rosenblueth, and Diego A. Velázquez. Pattern models: a dynamic epistemic logic for distributed systems. The Computer Journal, 67(7):2421–2440, July 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2421/7609837.

Cai:2023:QAA

[CWD^{+}23] Bin-Bin Cai, Yusen Wu, Jing Dong, Su-Juan Qin, Fei Gao, and Qiao-Yan Wen. Quantum attacks on 1K-AES and PRINCE. The Computer Journal, 66(5): 1102–1110, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1102/6518264.

Cao:2022:IPC

[CX22] Shujiao Cao and Rui Xue.

The (im)Possibility on constructing verifiable random functions. The Computer Journal, 65(7):1826–1845, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1826/6259651.

Chen:2020:PIR

[CYH^{+}20] Haishan Chen, Junying Yuan, Wien Hong, Jiangqun Ni, and Tung-Shou Chen. On performance improvement of reversible data hiding with contrast enhancement. The Computer Journal, 63(10):1584–1596, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1584/5875042.

Chen:2022:NCB

[CZC22] Siyuan Chen, Peng Zeng, and Kim-Kwang Raymond Choo. A new code-based blind signature scheme. The Computer Journal, 65(7):1776–1786, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1776/6236090.

Chen:2021:CEA

[CZH^{+}21] Shuhui Chen, Jincheng Zhong, Teng Huang, Ziling Wei, and Shuang Zhao. CMT: an efficient algorithm for scalable packet classification. The Computer Journal, 64(6):941– 959, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/941/6185121.

Cai:2021:PMM

[CZLY21] Hui Cai, Yanmin Zhu, Jie Li, and Jiadi Yu. A profitmaximizing mechanism for query-based data trading with personalized differential privacy. The Computer Journal, 64(2):264– 280, February 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/264/6043302.

Chen:2025:VAB

[CZPL25] Yang Chen, Chunlu Zhao, Jin Pan, and Yang Liu. Verifiable attribute-based multi-keyword ranked search scheme in blockchain. The Computer Journal, 68(5): 510–519, May 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/510/7933712.

Cao:2023:KDC

[CZZ23] Wenqin Cao, Wentao Zhang, and Xuefeng Zhao. The key-dependent capacity in multidimensional linear cryptanalysis. The Computer Journal, 66(2):269– 279, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/269/6397156.

Cheng:2021:HIC

[CZZW21] Wen Cheng, Yuqi Zou, Lingfang Zeng, and Yang Wang. Hercules: Intelligent coupling of dual-mode flash memory and hard disk drive. The Computer Journal, 64(2):224–235, February 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/2/224/6032259.

Junior:2022:FRM

[dAJS22] José Wagner de Andrade Júnior and Rodrigo Duarte Seabra. Fully retroactive minimum spanning tree problem. The Computer Journal, 65(4): 973–982, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/4/973/6012628.

Dangalchev:2024:CCL

[Dan24] Chavdar Dangalchev. Closeness centralities of lollipop graphs. The Computer Journal, 67(6):2020–2029, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2020/7477783.

Davoodi:2022:AEC

[DAR22] Omid Davoodi, Mehrdad Ashtiani, and Morteza Rajabi. An approach for the evaluation and correction of manually designed video game levels using deep neural networks. The Computer Journal, 65(3):495– 515, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/495/5868153.

Diaz-Caro:2025:CTC

[DCV25] Alejandro Díaz-Caro and Marcos Villagra. Classically time-controlled quantum automata: definition and properties. The Computer Journal, 68(1):23–31, January 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/23/7762872.

Dong:2023:HPD

[DFC^{+}23] Hui Dong, Jianxi Fan, Baolei Cheng, Yan Wang, and Li Xu. Hamiltonian properties of the data center network HSDC with faulty elements. The Computer Journal, 66(8):1965– 1981, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1965/6585620.

Dong:2021:CCS

[DFS^{+}21a] Yuhong Dong, Zetian Fu, Stevan Stankovski, Yaoqi Peng, and Xinxing Li. A call center system based on expert systems for the acquisition of agricultural knowledge transferred from text-to-speech in China. The Computer Journal, 64(6):895– 908, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/6/895/6139149.

Dong:2021:CDD

[DFS^{+}21b] Yuhong Dong, Zetian Fu, Stevan Stankovski, Yaoqi Peng, and Xinxing Li. A cotton disease diagnosis method using a combined algorithm of casebased reasoning and fuzzy logic. The Computer Journal, 64(2):155–168, February 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/2/155/5880022.

Drif:2021:OSP

[DH21] Ahlem Drif and Khalil Hadjoudj. An opinion spread prediction model with Twitter emotion analysis during Algeria’s Hirak. The Computer Journal, 64(3):358–368, March 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/358/5903958.

Dong:2024:AAK

[DJD24] Dong Dong, Hongxu Jiang, and Boyu Diao. AKGF:

Automatic kernel generation for DNN on CPU-FPGA. The Computer Journal, 67(5):1619–1627, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1619/7241314.

Devi:2022:RIV

[DK22] P. Suthanthira Devi and S. Karthika. Rumor identification and verification for text in social media content. The Computer Journal, 65(2):436–455, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/436/6363561.

DAndrea:2020:SMP

[DLM20] Eleonora D’Andrea, Beatrice Lazzerini, and Francesco Marcelloni. A system for multi-passenger urban ridesharing recommendations with ordered multiple stops. The Computer Journal, 63(5):657– 687, May 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/657/5363891.

Devismes:2021:TEG

[DLP^{+}21] Stéphane Devismes, Anissa Lamani, Franck Petit, Pascal Raymond, and Sébastien Tixeuil. Terminating exploration of a grid by an optimal number of asynchronous oblivious robots. The Computer Journal, 64(1):132– 154, January 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/132/5808802.

Dong:2024:HIA

[DLW24] Xueshi Dong, Qing Lin, and Wei Wang. Hybrid ITÖ algorithm for maximum scatter colored traveling salesman problem. The Computer Journal, 67(6):2172– 2188, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2172/7536051.

Daud:2023:CVE

[DM23] Marriam Daud and Ali Afzal Malik. Construction and validation of early software size estimation models based on ADAF-adjusted ACD metrics. The Computer Journal, 66(9):2123– 2137, September 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2123/6606050.

Datta:2023:RCC

[DMG^{+}23] Subrata Datta, Kalyani Mali, Udit Ghosh, Subrata Bose, Sourav Das, and Sourav Ghosh. Rare correlated coherent association rule mining with CLS-MMS. The Computer Journal, 66(2):342–359, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/342/6397157.

Devi:2023:PMM

[DMP^{+}23] K. Durga Devi, P. Uma Maheswari, Phani Kumar Polasi, R. Preetha, and M. Vidhyalakshmi. Pattern matching model for recognition of stone inscription characters. The Computer Journal, 66(3):554– 564, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/554/6424463.

Doan:2022:SMC

[DO22] Ha Thi Thu Doan and Kazuhiro Ogata. Specifying and model checking distributed control algorithms at meta-level. The Computer Journal, 65(12):2998– 3019, December 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/2998/6374066.

Duong:2020:MBR

[DST20] Dung Hoang Duong, Willy Susilo, and Ha Thanh Nguyen Tran. A multivariate blind ring signature scheme. The Computer Journal, 63(8):1194–1202, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/8/1194/5643521.

Dung:2024:CCA

[DTP24] Luong The Dung, Nguyen Ngoc Toan, and Tran Nghi Phu. CAIMP: Cross-architecture IoT malware detection and prediction based on static feature. The Computer Journal, 67(9):2763–2776, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2763/7680263.

Daniya:2022:DNN

[DV22] T. Daniya and S. Vigneshwari. Deep neural network for disease detection in rice plant using the texture and deep features. The Computer Journal, 65(7):1812– 1825, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1812/6236074.

Dong:2023:NMM

[DWGC23] Xiaoli Dong, Yongzhuang Wei, Wen Gao, and Jie Chen. New meet-in-the-middle attacks on FOX block cipher. The Computer Journal, 66(5):1195– 1212, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1195/6535018.

Dong:2025:DPC

[DWLF25a] Hui Dong, Huaqun Wang, Mengjie Lv, and Weibei Fan. Disjoint paths construction algorithm in the data center network DP-Cell. The Computer Journal, 68(1):47–57, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/47/7762873.

Dong:2025:FTA

[DWLF25b] Hui Dong, Huaqun Wang, Mengjie Lv, and Weibei Fan. Fault tolerance assessment of the data center network DPCell based on g-good-neighbor conditions. The Computer Journal, 68(8):985–997, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/8/985/8046863.

Ding:2024:QGD

[DWZS24] Lin Ding, Zheng Wu, Guixian Zhang, and Tairong Shi. Quantum guess and determine attack on stream ciphers. The Computer Journal, 67(1):292–303, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/292/6880884.

Donglai:2021:TAT

[DY21] Fu Donglai and Liu Yanhua. Trust-aware task allocation in collaborative crowdsourcing model. The Computer Journal, 64(6): 929–940, June 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/6/929/6146549.

Dong:2024:CLH

[DYZY24] Songshou Dong, Yanqing Yao, Yihua Zhou, and Yuguang Yang. A certificateless linearly homomorphic signature scheme based on lattice for network coding. The Computer Journal, 67(9):2739–2748, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2739/7668037.

ElGhor:2022:ORT

[EC22] Hussein El Ghor and Maryline Chetto. Optimal real-time scheduling algorithm for wireless sensors with regenerative energy. The Computer Journal, 65(8):2087–2097, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2087/6272096.

ElOsta:2020:OSS

[ECE20] Rola El Osta, Maryline Chetto, and Hussein El Ghor. Optimal slack stealing servicing for real-time energy harvesting systems. The Computer Journal, 63(10):1537–1546, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1537/5867669.

Elmezain:2021:RSI

[EI21] Mahmoud Elmezain and Hani M. Ibrahem. Retrieving semantic image using shape descriptors and latent-dynamic conditional random fields. The Computer Journal, 64(12):1876– 1885, December 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/12/1876/5904330.

Eldakhly:2025:OAI

[Eld25] Nabil M. Eldakhly. Optimizing arrhythmia identification using chaos theory and deep learning analysis. The Computer Journal, 68(8):859–871, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/8/859/8010274.

Estrada-Moreno:2021:KMD

[EMYRV21] A. Estrada-Moreno, I. G. Yero, and J. A. Rodríguez-Velázquez. On the (k, t)- metric dimension of graphs. The Computer Journal, 64 (5):707–720, May 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/64/5/707/
                                        5808798.

Essiz:2021:ABC

[EO21] Esra Sarac Essiz and Murat Oturakci. Artificial bee colony-based feature selection algorithm for cyberbullying. The Computer Journal, 64(3):305– 313, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/305/5860626.

Er-rajy:2020:NSR

[ErEE20] Latifa Er-rajy, My Ahmed El Kiram, and Mohamed El Ghazouani. New security risk value estimate method for Android applications. The Computer Journal, 63(4):593– 603, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/593/5618854.

Erdem:2024:PDT

[ESC24] Oğuzhan Erdem, Tuncay Soylu, and Aydın Carus. Pipelined decision trees for online traffic classification on FPGAs. The Computer Journal, 67(3):825– 839, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/825/7076641.

ElKhediri:2022:NDM

[ETA22] Salim El Khediri, Adel Thaljaoui, and Fayez Alfayez. A novel decisionmaking process for COVID-19 fighting based on association rules and Bayesian methods. The Computer Journal, 65(9):2360–2376, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2360/6291513.

Farooq:2020:RRP

[Far20] Muhammad Omer Farooq. RIoT: a routing protocol for the Internet of Things. The Computer Journal, 63(6): 958–973, June 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/6/958/5808800.

Farooq:2024:WAL

[Far24] Muhammad Omer Farooq. WiPoTS: an application layer protocol with network protocol stack enhancements for wireless power transfer networks. The Computer Journal, 67(6): 2108–2117, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2108/7491970.

Fati:2022:DCA

[Fat22] Suliman Mohamed Fati. Detecting cyberbullying across social media platforms in Saudi Arabia using sentiment analysis: a case study. The Computer Journal, 65(7):1787–1794, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1787/6232811.

Faghihi:2022:HLA

[FFH22] Hossein Rajaby Faghihi, MohammadAmin Fazli, and Jafar Habibi. Hybrid learning approach toward situation recognition and handling. The Computer Journal, 65(5):1293– 1305, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1293/6103953.

Feng:2022:BTE

[FGC22] Shengyuan Feng, Junqing Gong, and Jie Chen. Binary tree encryption with constant-size public key in the standard model. The Computer Journal, 65(6): 1489–1511, June 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1489/6154510.

Fulber-Garcia:2021:LDF

[FGM21] Vinicius Fulber-Garcia and Sérgio Luis Sardi Mergen. LUISA: Decoupling the frequency model from the context model in prediction-based compression. The Computer Journal, 64(9):1437–1450, September 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/9/1437/5868170.

Finocchi:2024:SDC

[FGS24] Irene Finocchi, Renan Leon Garcia, and Blerina Sinaimeri. From stars to diamonds: Counting and listing almost complete subgraphs in large networks. The Computer Journal, 67(6):2151– 2161, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2151/7492702.

Fang:2024:DPS

[FHWH24] Dong Fang, Guifang Huang, Mengfan Wang, and Lei Hu. Decreasing proof size of BLS scheme. The Computer Journal, 67(3):1030– 1040, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1030/7135776.

Faro:2023:LCC

[FLPS23] Simone Faro, Thierry Lecroq, Kunsoo Park, and Stefano Scafiti. On the longest common Cartesian substring problem. The Computer Journal, 66(4): 907–923, April 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/907/6500728.

Fan:2022:MEV

[FLZ^{+}22] Zhen-Bao Fan, Yi-Na Li, Kang Zhang, Jinhui Yu, and Mao Lin Huang. Measuring and evaluating the visual complexity of Chinese ink paintings. The Computer Journal, 65(8):1964–1976, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/1964/6278154.

Feng:2025:BEP

[FMX^{+}25] Yanchang Feng, Junchao Ma, Haojun Xia, Da Zhang, and Bibo Tu. BLFair: enabling proportional I/O sharing for NVMe SSD in SPDK para-virtualization architecture. The Computer Journal, 68(8):953– 967, August 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/953/8069015.

Feng:2023:SAC

[FNLW23] Li Feng, Xiaoling Ni, Zhen Ling, and Liangmin Wang. Strong anonymous communication system based on segment routing over SDN. The Computer Journal, 66(12):3092–3106, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/3092/6814393.

Fiori:2022:ASM

[FPT22] Fernando J. Fiori, Waltteri Pakalén, and Jorma Tarhio. Approximate string matching with SIMD. The Computer Journal, 65(6):1472– 1488, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1472/6134013.

Feng:2024:SBA

[FQZL24] Le Feng, Zhenxing Qian, Xinpeng Zhang, and Sheng Li. Stealthy backdoor attacks on deep point cloud recognization networks. The Computer Journal, 67(5):1879–1891, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1879/7462048.

Farhadi:2024:LML

[FS24] Amirfarhad Farhadi and Arash Sharifi. Leveraging meta-learning to improve unsupervised domain adaptation. The Computer Journal, 67(5):1838–1850, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1838/7335818.

Farzadnia:2021:NID

[FSN21] Ehsan Farzadnia, Hossein Shirazi, and Alireza Nowroozi. A new intrusion detection system using the improved dendritic cell algorithm. The Computer Journal, 64(8):1193–1214, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1193/6015901.

Feng:2024:IFD

[FWC24] Hao Feng, Jiong Wu, and Lin Chen. Intermittent fault diagnosis of product network based on PMC model. The Computer Journal, 67(9):2777–2786, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2777/7676522.

Fang:2020:EEA

[FYH20] Yong Fang, Yue Yang, and Cheng Huang. EmailDetective: an email authorship identification and verification model. The Computer Journal, 63(11):1775–1787, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1775/5870579.

Fang:2021:CCE

[FZH21] Yong Fang, Yuchi Zhang, and Cheng Huang. CyberEyes: Cybersecurity entity recognition model based on graph convolutional network. The Computer Journal, 64(8):1215– 1225, August 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/8/1215/6012869.

G:2020:NHU

[GAK20] Balaji C. G., Anu Monisha A., and Murugan K. A novel hybrid UE selection scheme for efficient data offloading using D2D communication. The Computer Journal, 63(10):1513–1523, October 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1513/5855740.

Ganesh:2020:PED

[Gan20] N. Ganesh. Performance evaluation of depth adjustment and void aware pressure routing (DA-VAPR) protocol for underwater wireless sensor networks. The Computer Journal, 63(2):193–202, February 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/193/5613613.

Ganapathi:2022:PDC

[GC22] Pramod Ganapathi and Rezaul Chowdhury. Parallel divide-and-conquer algorithms for bubble sort, selection sort and insertion sort. The Computer Journal, 65(10):2709–2719, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2709/6334046.

Ganapathi:2023:UFD

[GC23] Pramod Ganapathi and Rezaul Chowdhury. A unified framework to discover permutation generation algorithms. The Computer Journal, 66(3):603– 614, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/603/6429743.

Guruvammal:2021:OFS

[GCD21] S. Guruvammal, T. Chellatamilan, and L. Jegatha Deborah. Optimal feature selection and hybrid classification for autism detection in young children.

The Computer Journal, 64(11):1760–1774, November 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1760/6041173.

Guruvammal:2022:ADA

[GCD22] S. Guruvammal, T. Chellatamilan, and L. Jegatha Deborah. Automatic detection of autism in young children using weighted logarithmic transformed data with optimized deep learning. The Computer Journal, 65(10):2678–2692, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/10/2678/6323598.

Georgiou:2022:LTE

[GCG^{+}22] Kyriakos Georgiou, Zbigniew Chamski, Andres Amaya Garcia, David May, and Kerstin Eder. Lost in translation: Exposing hidden compiler optimization opportunities. The Computer Journal, 65(3):718– 735, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/718/5880035.

Gu:2020:CCE

[GCH20] Mei-Mei Gu, Jou-Ming Chang, and Rong-Xia Hao. On computing component (edge) connectivities of balanced hypercubes. The Computer Journal, 63(9):1311–1320, September 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1311/5554332.

Gu:2021:SMC

[GCH21] Mei-Mei Gu, Jou-Ming Chang, and Rong-Xia Hao. Strong Menger connectedness of augmented kary n-cubes. The Computer Journal, 64(5):812– 825, May 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/812/5830802.

Gomez:2024:CFM

[GCMRGM24] Juan Manuel Castelo Gómez, Javier Carrillo-Mondéjar, José Roldán-Gómez, and José Luis Martínez Martínez. A concept forensic methodology for the investigation of IoT cyberincidents. The Computer Journal, 67(4): 1324–1345, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1324/7208679.

Gao:2022:SAI

[GCR^{+}22] Wei Gao, Liqun Chen, Chunming Rong, Kaitai Liang, Xianghan Zheng, and Jiangshan Yu. Security analysis and improvement of a redactable consortium blockchain for industrial Internet-of-Things. The Computer Journal, 65(9):2430–2438, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2430/6299203.

Gnouma:2024:DHS

[GEZ24] Mariem Gnouma, Ridha Ejbali, and Mourad Zaied. Deep hashing and sparse representation of abnormal events detection. The Computer Journal, 67(1):3–17, January 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/3/6827723.

Gupta:2022:CPC

[GG22] Surbhi Gupta and Manoj K. Gupta. Computational prediction of cervical cancer diagnosis using ensemblebased classification algorithm. The Computer Journal, 65(6):1527–1539, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1527/6153484.

Gupta:2023:CMP

[GG23] Surbhi Gupta and Manoj Kumar Gupta. Computational model for prediction of malignant mesothelioma diagnosis. The Computer Journal, 66(1):86–100, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/86/6385801.

Gupta:2022:TSM

[GGD22] Shefali Gupta, Ankit Goel, and Meenu Dave. Tunicate swarm magnetic optimization with deep convolution neural network for collaborative filter recommendation. The Computer Journal, 65(10):2664–2677, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2664/6323602.

Gausseran:2021:SRM

[GGJM21] Adrien Gausseran, Frederic Giroire, Brigitte Jaumard, and Joanna Moulierac. Be scalable and rescue my slices during reconfiguration. The Computer Journal, 64(10):1584–1599, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/10/1584/6337892.

Gokcen:2023:PLI

[GGŞ23] Ahmet Gökçen, Alkım Gökçen, and Savaş Şahin. Prediction of Li-ion battery discharge patterns in IoT devices under random use via machine learning algorithms. The Computer Journal, 66(6):1541–1548, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1541/6627277.

Gu:2020:NAL

[GHC20] Mei-Mei Gu, Rong-Xia Hao, and Eddie Cheng. Note on applications of linearly many faults. The Computer Journal, 63 (9):1406–1416, September 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/9/1406/5614858.

Gu:2021:RAA

[GHC21] Mei-Mei Gu, Rong-Xia Hao, and Jou-Ming Chang. Reliability analysis of alternating group graphs and split-stars. The Computer Journal, 64(9):1425–1436, September 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1425/5872128.

Guo:2024:LAH

[GHZ24] Kaiwen Guo, Kexin Hu, and Zhenfeng Zhang. Liveness attacks on Hot-Stuff: The vulnerability of timer doubling mechanism. The Computer Journal, 67(8):2586–2600, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2586/7634133.

Gandhmal:2021:WEF

[GK21] Dattatray P. Gandhmal and K. Kumar. Wrapperenabled feature selection and CPLM-based NARX model for stock market prediction. The Computer Journal, 64(2):169– 184, February 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/169/5894133.

Gupta:2023:PPB

[GK23] Niharika Gupta and Baij Nath Kaushik. Prognosis and prediction of breast cancer using machine learning and ensemble-based training model. The Computer Journal, 66(1):70–85, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/70/6384777.

Geetha:2022:WRT

[GKK22] R. Geetha, S. Karthika, and Ponnurangam Kumaraguru. I regret for this tweet?’ — Twitter user’s behavior analysis system for private data disclosure. The Computer Journal, 65(2):275–296, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/275/5824617.

Gross:2025:CC

[GKO^{+}25] Yoav Gross, Shmuel T. Klein, Elina Opalinsky, Rivka Revivo, and Dana Shapira. Compression cryptosystem. The Computer Journal, 68(8):1062–1073, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/8/1062/8090573.

Gan:2021:NNM

[GLLZ21] Jiangzhang Gan, Tong Liu, Li Li, and Jilian Zhang. Non-negative matrix factorization: a survey. The Computer Journal, 64(7):1080– 1092, July 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/1080/6323682.

Gao:2024:CUC

[GLW^{+}24] Zhiyuan Gao, Jinguo Li, Liangliang Wang, Yin He, and Peichun Yuan. CM-UTC: a cost-sensitive matrix based method for unknown encrypted traffic classification. The Computer Journal, 67(7):2441– 2452, July 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2441/7614263.

Guo:2025:NEH

[GLX25] Chen Guo, Yaoyao Luo, and Zhifang Xiao. Nodeedge hybrid diagnosability and conditional nodeedge hybrid diagnosability under the HPMC^{∗} model. The Computer Journal, 68(10):1535–1545, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1535/8178806.

Guo:2023:DIN

[GLXP23] Chen Guo, Qiuming Liu, Zhifang Xiao, and Shuo Peng. The diagnosability of interconnection networks with missing edges and broken-down nodes under the PMC and MM^{∗} models. The Computer Journal, 66(8):2000–2010, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/2000/6581999.

Gong:2024:GBM

[GLY^{+}24] Yunchao Gong, Xueqiang Lv, Zhu Yuan, Xindong You, Feng Hu, and Yuzhong Chen. GNN-based multimodal named entity recognition. The Computer Journal, 67(8):2622–2632, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/8/2622/7641748.

Gozukara:2023:IHC

[GÖ23] Furkan Gözükara and Selma Ayşe Özel. An incremental hierarchical clustering based system for record linkage in e-commerce domain. The Computer Journal, 66(3): 581–602, March 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/3/581/6425234.

Gokcen:2021:CAD

[Gök21] Ahmet Gökçen. Computeraided diagnosis system for chronic obstructive pulmonary disease using empirical wavelet transform on auscultation sounds. The Computer Journal, 64(11):1775–1783, November 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/11/1775/6132390.

Gokulkumari:2022:MEA

[Gok22] G. Gokulkumari. Metaheuristicenabled artificial neural network framework for multimodal biometric recognition with local fusion visual features. The Computer Journal, 65(6):1586– 1597, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1586/6167838.

See correction [Ano24e].

Gandotra:2021:SSD

[GP21] Rahil Gandotra and Levi Perigo. SDVoIP — a software-defined VoIP framework for SIP and dynamic QoS. The Computer Journal, 64(2):254– 263, February 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/254/6032271.

Ghuge:2020:WQS

[GPR20] C. A. Ghuge, V. Chandra Prakash, and Sachin D. Ruikar. Weighed queryspecific distance and hybrid NARX neural network for video object retrieval. The Computer Journal, 63(11):1738–1755, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1738/5618962.

Guo:2020:IPR

[GQL^{+}20] Wei Guo, Sujuan Qin, Jun Lu, Fei Gao, Zhengping Jin, and Qiaoyan Wen. Improved proofs of retrievability and replication for data availability in cloud storage. The Computer Journal, 63(8):1216–1230, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/8/1216/5686324.

Gupta:2021:ETT

[GR21] Vishan Kumar Gupta and Prashant Singh Rana. Ensemble technique for toxicity prediction of small drug molecules of the antioxidant response element signalling pathway. The Computer Journal, 64(12):1861–1875, December 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1861/5739941.

Gavade:2022:DMH

[GR22] Anil B. Gavade and Vijay S. Rajpurohit. S-DolLion-MSVNN: a hybrid model for developing the super-resolution image from the multispectral satellite image. The Computer Journal, 65(4):757– 772, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/757/5894718.

Gokuldhev:2022:LPB

[GS22] M. Gokuldhev and G. Singaravel. Local pollinationbased moth search algorithm for task-scheduling heterogeneous cloud environment. The Computer Journal, 65(2):382– 395, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/382/5860965.

Gharibi:2021:CBM

[GSFS21] Reza Gharibi, Atefeh Safdel, Seyed Mostafa Fakhrahmad, and Mohammad Hadi Sadreddini. A contentbased model for tag recommendation in software information sites. The Computer Journal, 64(11):1680– 1691, November 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1680/5682398.

Guo:2025:EUD

[GSL25] Enting Guo, Chunhua Su, and Peng Li. Efficient unlearning for data security in deep learning systems. The Computer Journal, 68(9):1197–1207, September 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/9/1197/8159132.

Gao:2022:SIA

[GT22] Honghao Gao and Zhiyuan Tan. Special issue on adversarial AI to IoT security and privacy protection: Attacks and defenses. The Computer Journal, 65(11):2847–2848, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/11/2847/6732676.

Ge:2024:SBE

[GW24] Wenhan Ge and Junfeng Wang. SeqMask: Behavior extraction over cyber threat intelligence via multi-instance learning. The Computer Journal, 67(1):253–273, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/253/6852690.

Gong:2020:DCF

[GYY^{+}20] Faming Gong, Hanbing Yue, Xiangbing Yuan, Wenjuan Gong, and Tao Song. Discriminative correlation filter for long-time tracking. The Computer Journal, 63(3):460–468, March 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/460/5498234.

Guo:2025:RSI

[GZ25] Yuanzhao Guo and Jun Zhang. Remote sensing image dehazing method in mountaineering equipment. The Computer Journal, 68(4):397–406, April 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/397/7903423.

Guo:2025:PIB

[GZF25] Maohua Guo, Yuefei Zhu, and Jinlong Fei. ProInfer: inference of binary protocol keywords based on probabilistic statistics. The Computer Journal, 68(2):109– 125, February 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/2/109/7789868.

Gao:2021:EDC

[GZL^{+}21] Rui Gao, Yuanyuan Zhu, Mingye Li, Shoufeng Li, and Xiaohu Shi. Encoder– decoder couplet generation model based on ‘trapezoidal context’ character vector. The Computer Journal, 64(3):286– 295, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/286/5860617.

Guo:2023:NMF

[GZY^{+}23] Hao Guo, Zhiyu Zhang, Qianqian Yang, Lei Hu, and Yiyuan Luo. A new method to find all the high-probability wordoriented truncated differentials: Application to Midori, SKINNY and CRAFT. The Computer Journal, 66(5):1069–1082, May 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1069/6522735.

Harn:2023:MBS

[H^{+}23a] Lein Harn et al. Multiple blind signature for e-voting and e-cash. The Computer Journal, 66(10):2331–2338, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2331/6618068.

Hou:2023:PARb

[H^{+}23b] Zezhou Hou et al. Practical attacks of round-reduced SIMON based on deep learning. The Computer Journal, 66(10):2517–2534, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2517/6654842.

Hu:2023:ILB

[H^{+}23c] Mingxing Hu et al. An improved lattice-based ring signature with unclaimable anonymity in the standard model. The Computer Journal, 66(10):2542–2553, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2542/6652013.

Hilal:2022:HIT

[HAWA^{+}22] Anwer Mustafa Hilal, Fahd N.

Al-Wesabi, Abdelzahir Abdelmaboud, Manar Ahmed Hamza, Mohammad Mahzari, and Abdulkhaleq Q. A. Hassan. A hybrid intelligent text watermarking and natural language processing approach for transferring and receiving an authentic English text via Internet. The Computer Journal, 65(2):423–435, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/423/6309918.

Honnikoll:2023:MER

[HB23] Nagaraj Honnikoll and Ishwar Baidari. Mean error rate weighted online boosting method. The Computer Journal, 66(1):1–15, January 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/1/6381247.

Hajlaoui:2025:HAS

[HBI^{+}25] Nasreddine Hajlaoui, Chaima Bejaoui, Tayssir Ismail, Houaida Ghanmi, and Haifa Touati. A hybrid architecture for secure data sharing in multiclouds system. The Computer Journal, 68(1):58–73, January 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/1/58/7789870.

Huang:2021:CAC

[HCZ^{+}21] Rubing Huang, Haibo Chen, Yunan Zhou, Tsong Yueh Chen, Dave Towey, Man Fai Lau, Sebastian Ng, Robert Merkel, and Jinfu Chen. Covering array constructors: an experimental analysis of their interaction coverage and fault detection. The Computer Journal, 64(5):762–788, May 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/762/5822659.

Hou:2023:PARa

[HCZ23] Tao Hou, Ting Cui, and Jiyan Zhang. Practical attacks on reduced-round 3D and Saturnin. The Computer Journal, 66(2):479– 495, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/479/6419808.

Hireche:2024:TNR

[HDK24] Samia Hireche, Abdeslem Dennai, and Boufeldja Kadri. Toward a novel RESTFUL big data-based urban traffic incident data Web service for connected vehicles. The Computer Journal, 67(2):557–580, February 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/557/7021934.

Hasanzadeh:2022:RBR

[HFT22] S. Hasanzadeh, S. M. Fakhrahmad, and M. Taheri. Review-based recommender systems: a proposed rating prediction scheme using word embedding representation of reviews. The Computer Journal, 65(2):345– 354, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/345/5837819.

Huynh:2022:NAD

[HH22] Hieu Trung Huynh and Tran Minh Hoang. A novel approach for determining meal plan for gestational diabetes mellitus using artificial intelligence. The Computer Journal, 65(5):1088– 1097, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/5/1088/6023092.

Huang:2024:TAR

[HH24] Chung-Ming Huang and Shu-Hang Huang. Trafficaware re-grouping for load balance in IEEE 802.11ah IoT network based on the registered backoff time mechanism. The Computer Journal, 67(3):884– 898, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/884/7079713.

Hongjin:2024:NFS

[HHB24] Zhang Hongjin, Wei Hui, and Wang Bo. A new fuzzy smoothing term model for stereo matching. The Computer Journal, 67(2):746– 761, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/746/7075447.

Hsieh:2022:IES

[HHC22] Ping-Chi Hsieh, Der-Juinn Horng, and Hong-Yi Chang. International expansion selection model by machine learning — a proprietary model. The Computer Journal, 65(2):217–236, February 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/217/5824920.

Hou:2024:DTD

[HHL^{+}24] Changsheng Hou, Bingnan Hou, Xionglve Li, Tongqing Zhou, Yingwen Chen, and Zhiping Cai. DRL-Tomo: a deep reinforcement learning-based approach to augmented data generation for network tomography. The Computer Journal, 67(10):2995–3008, October 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/10/2995/7717689.

Hao:2022:AEC

[HHLL22] Wangli Hao, Meng Han, Shancang Li, and Fuzhong Li. An attention enhanced cross-modal image– sound mutual generation model for birds. The Computer Journal, 65(2):410– 422, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/410/6124656.

Huang:2020:DRR

[HHX^{+}20] Longji Huang, Jianbin Huang, Yueshen Xu, Zhiqiang Zhao, and Zhenghao Zhang. Driving route recommendation with profit maximization in ride sharing. The Computer Journal, 63(11):1607–1623, November 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1607/5628026.

Hsu:2023:CLA

[HHX^{+}23] Chingfang Hsu, Lein Harn, Zhe Xia, Jianqun Cui, and Jingxue Chen. Construction of lightweight authenticated joint arithmetic computation for 5G IoT networks. The Computer Journal, 66(1):208–220, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/208/6398683.

Huang:2020:SDT

[HJ20] Tianwen Huang and Fei Jiao. Study on data transfer in meteorological forecast of small and medium-sized cities and its application in Zhaoqing City. The Computer Journal, 63(7):1076– 1083, July 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/1076/5560310.

Hou:2024:SFF

[HJH^{+}24] Changsheng Hou, Chunbo Jia, Bingnan Hou, Tongqing Zhou, Yingwen Chen, and Zhiping Cai. A sketch framework for fast, accurate and fine-grained analysis of application traffic. The Computer Journal, 67(6):2039–2053, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2039/7484594.

Haddad:2020:PSS

[HJK20] Mohammed Haddad, Colette Johnen, and Sven Köhler. Polynomial silent self-stabilizing p-star decomposition. The Computer Journal, 63(2):253– 266, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/253/5618961.

Hamid:2024:EEB

[HKA24] Naila Hamid, Nazar Khan, and Arbish Akram. ELSM: Evidence-based line segment merging. The Computer Journal, 67(7):2498– 2514, July 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/7/2498/7634131.

Hua:2023:LDC

[HLC^{+}23] Jialiang Hua, Tai Liu, Yulong Cui, Lingyue Qin, Xiaoyang Dong, and Huiyong Cui. Low-data cryptanalysis on SKINNY block cipher. The Computer Journal, 66(4):970–986, April 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/4/970/6519795.

Huang:2023:CFD

[HLCX23] Yanze Huang, Limei Lin, Eddie Cheng, and Li Xu. Component fault diagnosis and fault tolerance of alternating group graphs. The Computer Journal, 66(5): 1184–1194, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1184/6528296.

Hu:2022:NDP

[HLJW22] Xichao Hu, Yongqiang Li, Lin Jiao, and Mingsheng Wang. New division property propagation table: Applications to block ciphers with large S-boxes. The Computer Journal, 65(6): 1560–1573, June 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1560/6134263.

Han:2021:TIC

[HLML21] Gang Han, Menggang Li, Yiduo Mei, and Deming Li. Transportation index computation: a development theme miningbased approach. The Computer Journal, 64(3):337– 346, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/337/5905458.

Huang:2022:SRA

[HLX22] Yanze Huang, Limei Lin, and Li Xu. Subgraph reliability of alternating group graph with uniform and nonuniform vertex faultfree probabilities. The Computer Journal, 65(3): 589–605, March 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/589/5879957.

Hougaard:2023:ATB

[HM23] Hector Hougaard and Atsuko Miyaji. Authenticated tree-based R-LWE group key exchange. The Computer Journal, 66(2):360– 372, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/360/6409535.

Huang:2025:BBB

[HMG25] Longxia Huang, Xue Mao, and Chengzhi Ge. BDEA: blockchain-based secure deduplication for lowentropy data supporting public auditing. The Computer Journal, 68(10):1424– 1435, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1424/8120572.

Hakim:2024:AAM

[HNR24] Sheikh Azizul Hakim, Rahnuma Islam Nishat, and Md Saidur Rahman. Approximation algorithms for maximum weighted internal spanning trees in regular graphs and subdivisions of graphs. The Computer Journal, 67(10):2898–2905, October 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2898/7713287.

Hellings:2021:RAS

[HPV^{+}21] Jelle Hellings, Catherine L.

Pilachowski, Dirk Van Gucht, Marc Gyssens, and Yuqing Wu. From relation algebra to semi-join algebra: an approach to graph query optimization. The Computer Journal, 64(5): 789–811, May 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/789/5827080.

Han:2024:PPC

[HSC^{+}24] Jinguang Han, Willy Susilo, Liquan Chen, Jianchang Lai, and Ge Wu. Privacypreserving confidential reporting system with designated reporters. The Computer Journal, 67(5):1951– 1962, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1951/7457339.

Huang:2022:CSB

[HSL^{+}22] Jianbin Huang, Heli Sun, He Li, Longji Huang, Ao Li, and Xiangyu Wang. Central station-based demand prediction for determining target inventory in a bikesharing system. The Computer Journal, 65(3):573– 588, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/573/5879956.

Hsu:2023:DDB

[Hsu23] Ching-Kuo Hsu. A dueling DQN-based computational offloading method in MEC-enabled IIoT network. The Computer Journal, 66(12):2887–2896, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2887/6751160.

Hsu:2024:GAB

[Hsu24] Ching-Kuo Hsu. A genetic algorithm based scheduling method with carrier aggregation in 5G networks. The Computer Journal, 67(4): 1279–1285, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1279/7192130.

Hsu:2025:ESR

[Hsu25]  Ching-Kuo Hsu.  En-

hanced SAC-RNN-based collision-avoidance methods in UAV networks with leader–follower scenario. The Computer Journal, 68(10):1546–1554, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1546/8131476.

Harish:2025:GTE

[HTJ25] Avanthika Vineetha Harish, Kimberly Tam, and Kevin Jones. Generating training events for building cyberphysical security skills. The Computer Journal, 68(5): 445–459, May 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/445/7921452.

Huang:2024:CDA

[HWC^{+}24] Jiale Huang, Jigang Wu, Long Chen, Yalan Wu, and Yidong Li. Coalitional double auction for ridesharing with desired benefit and QoE constraints. The Computer Journal, 67(5):1674– 1686, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1674/7280724.

Hao:2021:LSS

[HWS21] Shijie Hao, Zhonghao Wang, and Fuming Sun. LEDet: a single-shot realtime object detector based on low-light image enhancement. The Computer Journal, 64(7):1028–1038, July 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/1028/6275475.

Hou:2024:TDA

[HWW^{+}24] Shiqi Hou, Baofeng Wu, Shichang Wang, Hao Guo, and Dongdai Lin. Truncated differential attacks on symmetric primitives with linear key schedule: WARP and Orthros. The Computer Journal, 67(4):1483– 1500, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1483/7238220.

Hsu:2024:ELC

[HXCH24] Chingfang Hsu, Zhe Xia, Tianshu Cheng, and Lein Harn. Extremely lightweight constant-round membershipauthenticated group key establishment for resourceconstrained smart environments toward 5G. The Computer Journal, 67(3): 840–850, March 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/840/7079711.

Hou:2018:UEO

[HXLX18] Yonghong Hou, Lin Xue, Shuo Li, and Jiaming Xing. User-experience-oriented fuzzy logic controller for adaptive streaming. The Computer Journal, 61(7):1064–1074, July 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/1064/4843991.

See corrigendum [HXLX22].

Hou:2022:CUE

[HXLX22] Yonghong Hou, Lin Xue, Shuo Li, and Jiaming Xing. Corrigendum: Userexperience-oriented fuzzy logic controller for adaptive streaming. The Computer Journal, 65(7):1937, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1937/5498465.

See [HXLX18].

Huang:2022:PAS

[HXW22] Mingjiang Huang, Zhen Xu, and Liming Wang. On the probability and automatic search of rotational-XOR cryptanalysis on ARX ciphers. The Computer Journal, 65(12):3062–3080, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3062/6373560.

He:2024:CRR

[HY24a] Le He and Hongbo Yu. Cryptanalysis of reducedround SipHash. The Computer Journal, 67(3):875– 883, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/875/7080914.

Huo:2024:PHP

[HY24b] Jin Huo and Weihua Yang. Pancyclic and Hamiltonian properties of dragonfly networks. The Computer Journal, 67(3):1201–1208, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1201/7161851.

He:2025:MKG

[HY25] Jinyan He and Haitong Yang. Multilingual knowledge graph completion based on structural features. The Computer Journal, 68(9):1252–1262, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1252/8122431.

Hatkeposhti:2024:PAE

[HYTG24] Rahman Keramati Hatkeposhti, Meisam Yadollahzadeh-Tabari, and Mehdi Golsorkhtabariamiri. Providing an approach for early prediction of fall in human activities based on wearable sensor data and the use of deep learning algorithms. The Computer Journal, 67(2):658– 673, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/658/7071572.

Huang:2020:ULT

[HYZ^{+}20] Meijuan Huang, Bo Yang, Mingwu Zhang, Lina Zhang, and Hongxia Hou. Updatable lossy trapdoor functions under consecutive leakage. The Computer Journal, 63(4):648– 656, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/648/5667451.

Huang:2022:CLR

[HYZH22] Meijuan Huang, Bo Yang, Yanwei Zhou, and Xuewei Hu. Continual leakageresilient hedged public-key encryption. The Computer Journal, 65(6):1574–1585, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1574/6134264.

Hua:2025:FBC

[HZ25] Xiaohui Hua and Qin Zhao. The h-faulty-block connectivity of k-ary ncubes. The Computer Journal, 68(2):126–134, February 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/2/126/7828053.

Haq:2023:UNE

[HZKF23] Rafiul Haq, Xiaowang Zhang, Wahab Khan, and Zhiyong Feng. Urdu named entity recognition system using deep learning approaches. The Computer Journal, 66(8):1856–1869, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/1856/6572657.

Huang:2024:DSL

[HZLW24] Zhuojie Huang, Shuping Zhao, Zien Liang, and Jigang Wu. Discriminative subspace learning with adaptive graph regularization. The Computer Journal, 67(9):2823–2837, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2823/7690287.

Hao:2023:ITC

[HZS^{+}23] Qiaohong Hao, Qi Zhao, Mateu Sbert, Qinghe Feng, Cosmin Ancuti, Miquel Feixas, Marius Vila, and Jiawan Zhang. Informationtheoretic channel for multiexposure image fusion. The Computer Journal, 66(1):114–127, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/114/6384779.

Han:2025:DPU

[HZS^{+}25] Meng Han, Yan Zeng, Hao Shu, Lupeng Yue, Jilin Zhang, Jian Wan, and Yongjian Ren. Device placement using Laplacian PCA and graph attention networks. The Computer Journal, 68(2):175– 186, February 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/175/7819324.

Hu:2021:ALS

[HZW21] Rongyao Hu, Leyuan Zhang, and Jian Wei. Adaptive Laplacian support vector machine for semisupervised learning. The Computer Journal, 64(7): 1005–1015, July 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/7/1005/6259636.

See correction [Ano23a].

Huang:2021:DIR

[HZY21] Weidong Huang, Shuting Zhu, and Xinkai Yao. Destination image recognition and emotion analysis: Evidence from user-generated content of online travel communities. The Computer Journal, 64(3):296– 304, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/296/5866111.

He:2020:HMM

[HZZ20] Binjie He, Dong Zhang, and Chang Zhao. Hidden Markov model-based load balancing in data center networks. The Computer Journal, 63(10):1449–1462, October 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1449/5666159.

Irribarra-Cortes:2025:HBC

[ICAAFSS25] Ayleen Irribarra-Cortés, Roberto Asín-Achá, José Fuentes-Sepúlveda, and Diego Seco. Heuristicbased computation of tailored spanning trees to speed up compact planar graphs. The Computer Journal, 68(10):1476–1484, October 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1476/8131475.

Ito:2022:ERC

[IOS^{+}22] Sohei Ito, Kenji Osari, Masaya Shimakawa, Shigeki Hagihara, and Naoki Yonezaki. Efficient realizability checking by modularization of LTL specifications. The Computer Journal, 65(10):2801– 2814, October 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2801/6363569.

J:2023:CSO

[JB23] Sheela J. and Janet B.

Caviar-sunflower optimization algorithm-based deep learning classifier for multidocument summarization. The Computer Journal, 66 (3):727–742, March 2023.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/66/3/727/
                                        6474500.

Jain:2023:PLN

[JCG23] Sweta Jain, Pruthviraj Choudhari, and Mahesh Gour. Pulmonary lung nodule detection from computed tomography images using two-stage convolutional neural network. The Computer Journal, 66(4): 785–795, April 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/4/785/6464409.

Jiang:2020:MBC

[JCY^{+}20] Lili Jiang, Xiaolin Chang, Runkai Yang, Jelena Mišić, and Vojislav B. Mišić. Model-based comparison of cloud-edge computing resource allocation policies. The Computer Journal, 63(10):1564–1583, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1564/5865371.

Ju:2023:HSI

[JDH23] Chuanxiang Ju, Hangqi Ding, and Benjia Hu. A hybrid strategy improved whale optimization algorithm for Web service composition. The Computer Journal, 66(3):662– 677, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/662/6455660.

Jafri:2024:HSC

[JGJ^{+}24] Syed Irteza Hussain Jafri, Rozaida Ghazali, Irfan Javid, Yana Mazwin Mohmad Hassim, and Mubashir Hayat Khan. A hybrid solution for the cold start problem in recommendation. The Computer Journal, 67(5): 1637–1644, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1637/7252293.

Jiao:2022:GDA

[JLD22] Lin Jiao, Yongqiang Li, and Shaoyu Du. Guessand-determine attacks on AEGIS. The Computer Journal, 65(8):2221–2230, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2221/6280579.

Jiao:2020:GDA

[JLH20] Lin Jiao, Yongqiang Li, and Yonglin Hao. A guessand-determine attack on SNOW-V stream cipher. The Computer Journal, 63(12):1789–1812, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/12/1789/5739948.

Jin:2024:SCE

[JM24] Yaoan Jin and Atsuko Miyaji. Secure and compact elliptic curve scalar multiplication with optimized inversion. The Computer Journal, 67(2):474– 484, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/474/6966047.

Jameel:2023:FIU

[JMATQ23] Samer Kais Jameel, Jafar Majidpour, Abdulbasit K. Al-Talabani, and Jihad Anwar Qadir. Face identification using conditional generative adversarial network. The Computer Journal, 66(7):1687– 1697, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1687/6565998.

Jiji:2023:AMS

[JMR23] W. Jiji, G. Merlin, and A. Rajesh. Assessment and monitoring of salinity of soils after tsunami in Nagapattinam area using fuzzy logic-based classification. The Computer Journal, 66(2):333–341, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/333/6423536.

Jain:2022:FFG

[JMV22] Amita Jain, Kanika Mittal, and Kunwar Singh Vaisla. FLAKE: Fuzzy graph centrality-based automatic keyword extraction. The Computer Journal, 65(4):926–939, April 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/926/6015899.

Jia:2024:MTC

[JQ24] Huanshen Jia and Jianguo Qian. Menger-type connectivity of line graphs of generalized hypercubes with faulty edges. The Computer Journal, 67(6):2118– 2125, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2118/7510968.

Joseph:2024:RSE

[JSB24] Mabin Joseph, Gautham Sekar, and R. Balasubramanian. Revisiting the software-efficient stream ciphers RCR-64 and RCR-32. The Computer Journal, 67(4):1590–1602, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1590/7242318.

Joseph:2022:SSC

[JSBV22] Mabin Joseph, Gautham Sekar, R. Balasubramanian, and G. Venkiteswaran. On the security of the stream ciphers RCR-64 and RCR-32. The Computer Journal, 65(12):3091–3099, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3091/6369316.

Jiang:2020:SCR

[JTGJ20] Jiafu Jiang, Linyu Tang, Ke Gu, and WeiJia Jia. Secure computing resource allocation framework for open fog computing. The Computer Journal, 63(4): 567–592, April 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/4/567/5717858.

Jia:2024:GGR

[JWJ24] Zeguang Jia, Jianming Wang, and Rize Jin. GR-Net: a graph reasoning network for enhanced multimodal learning in scene text recognition. The Computer Journal, 67(12):3239–3250, December 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/12/3239/7760133.

Jia:2024:CLB

[JWR24] Hairong Jia, Suying Wang, and Zelong Ren. CNN-LSTM base station traffic prediction based on dual attention mechanism and timing application. The Computer Journal, 67(6): 2246–2256, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2246/7595363.

Jiang:2022:SCP

[JXZ^{+}22] Che Jiang, Wanyue Xu, Xiaotian Zhou, Zhongzhi Zhang, and Haibin Kan. Some combinatorial problems in power-law graphs. The Computer Journal, 65 (7):1679–1691, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/65/7/1679/
6189767.

Jia:2025:VTE

[JYW^{+}25] Yaxun Jia, Zhu Yuan, Haoyang Wang, Yunchao Gong, Haixiang Yang, and Zuo lin Xiang. Variations towards an efficient drug– drug interaction. The Computer Journal, 68(5):552– 564, May 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/552/7921450.

Jin:2023:EDL

[JZ23] Chengbin Jin and Yongbin Zhou. Enhancing deeplearning based side-channel analysis through simultaneously multi-byte training. The Computer Journal, 66(11):2674–2704, November 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2674/6691911.

Ji:2021:IMS

[JZD21] Fulei Ji, Wentao Zhang, and Tianyou Ding. Improving Matsui’s search algorithm for the best differential/linear trails and its applications for DES, DESL and GIFT. The Computer Journal, 64(4):610– 627, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/610/5880463.

Jiang:2023:FSP

[JZWN23] Zhe Jiang, Kai Zhang, Liangliang Wang, and Jianting Ning. Forward secure public-key authenticated encryption with conjunctive keyword search. The Computer Journal, 66(9):2265–2278, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2265/6611470.

Kalkar:2023:BOP

[K^{+}23] Oznur Kalkar et al. On the batch outsourcing of pairing computations. The Computer Journal, 66 (10):2437–2446, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2437/6649280.

K:2024:UNB

[K.24] Ashwini K. User namebased compression and encryption of images using chaotic compressive sensing theory. The Computer Journal, 67(1):304– 322, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/304/6918745.

Karaata:2020:SSM

[KA20] Mehmet Hakan Karaata and Anwar Nais AlMutairi. A snap-stabilizing mwave algorithm for tree networks. The Computer Journal, 63(2):220–238, February 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/220/5721330.

Kalpana:2023:CSO

[Kal23] Parsi Kalpana. Chronological sailfish optimizer for preserving privacy in cloud based on Khatri–Rao product. The Computer Journal, 66(1):101–113, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/101/6394996.

Kamareddine:2024:TEI

[Kam24] Fairouz Kamareddine. Thematic editorial, it is hard to imagine a world without algorithms and data science. The Computer Journal, 67(5):1605–1606, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1605/7686451.

Kazmi:2022:SCA

[KAMA22] Syed Saqib Ali Kazmi, Mehreen Ahmed, Rafia Mumtaz, and Zahid Anwar. Spatiotemporal clustering and analysis of road accident hotspots by exploiting GIS technology and kernel density estimation. The Computer Journal, 65(2):155–176, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/155/5709697.

Kumar:2023:FEB

[KASV23] B. Mathan Kumar, Bharati S.

Ainapure, Suryabhan Pratap Singh, and Sumit Vyas. Feature extraction based deep indexing by deep fuzzy clustering for image retrieval using Jaro Winkler distance. The Computer Journal, 66(9):2191–2207, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2191/6633645.

Kathavate:2023:RML

[Kat23] Pravin Narayan Kathavate.

Role of machine learning on key extraction for data privacy preservation of health care sectors in IoT environment. The Computer Journal, 66(6):1549–1562, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1549/7077265.

Kumar:2022:MOR

[KB22] Katukuri Arun Kumar and Ravi Boda. A multiobjective randomly updated beetle swarm and multi-verse optimization for brain tumor segmentation and classification. The Computer Journal, 65(4): 1029–1052, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/4/1029/6414419.

See corrigendum [Ano23d].

Kumar:2023:TBE

[KC23] Shobhan Kumar and Arun Chauhan. A transformer based encodings for detection of semantically equivalent questions in cQA. The Computer Journal, 66(5): 1139–1155, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1139/6534390.

Kung:2021:TTT

[KCLH21] Tzu-Liang Kung, Hon-Chan Chen, Chia-Hui Lin, and Lih-Hsing Hsu. Three types of two-disjoint-cycle-cover pancyclicity and their applications to cycle embedding in locally twisted cubes. The Computer Journal, 64(1):27–37, January 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/27/5628025.

Khaire:2021:ERF

[KD21] Utkarsh Mahadeo Khaire and R. Dhanalakshmi. Effects of random forest parameters in the selection of biomarkers. The Computer Journal, 64(12):1840–1847, December 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1840/5808799.

Kenny:2020:HFM

[Ken20] Ian Kenny. Hydrographical flow modelling of the River Severn using particle swarm optimization. The Computer Journal, 63(11):1713– 1726, November 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1713/5618958.

Khalifa:2022:DMS

[KESZ22] Intissar Khalifa, Ridha Ejbali, Raimondo Schettini, and Mourad Zaied. Deep multi-stage approach for emotional body gesture recognition in job interview. The Computer Journal, 65(7):1702–1716, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1702/6236092.

Karaoglan:2024:EAC

[KF24] Kursat Mustafa Karaoglan and Oguz Findik. Enhancing aspect category detection through hybridised contextualised neural language models: a case study in multi-label text classification. The Computer Journal, 67(6):2257–2269, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2257/7590676.

Kammoun:2022:MBS

[KGA22] Hager Kammoun, Imen Gabsi, and Ikram Amous. MeSH-based semantic indexing approach to enhance biomedical information retrieval. The Computer Journal, 65(3):516– 536, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/516/5869138.

Kumar:2023:FSO

[KGA23] Penta Anil Kumar, R. Gunasundari, and R. Aarthi. Fractional sailfish optimizer with deep convolution neural network for compressive sensing based magnetic resonance image reconstruction. The Computer Journal, 66(2):280–294, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/280/6398686.

K:2023:MFS

[KK23a] Poongodi K. and Dhananjay Kumar. Mining frequent serial positioning episode rules with forward and backward search technique from event sequences. The Computer Journal, 66(7):1622–1643, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1622/6561439.

Kaur:2023:ACA

[KK23b] Jaspreet Kaur and Prabhpreet Kaur. Automated computer-aided diagnosis of diabetic retinopathy based on segmentation and classification using Knearest neighbor algorithm in retinal images. The Computer Journal, 66(8):2011– 2032, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/2011/6585081.

Kumawat:2023:DAT

[KK23c] Manisha Kumawat and Arti Khaparde. Development of adaptive timeweighted dynamic time warping for time series vegetation classification using satellite images in Solapur District. The Computer Journal, 66(8):1982–1999, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/1982/6585938.

Kushwaha:2024:HAR

[KKBK24] Arati Kushwaha, Manish Khare, Reddy Mounika Bommisetty, and Ashish Khare. Human activity recognition based on video summarization and deep convolutional neural network. The Computer Journal, 67(8):2601–2609, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/8/2601/7634135.

Keles:2021:PAB

[KKK21] Mumine Kaya Keles, Umit Kilic, and Abdullah Emre Keles. Proposed artificial bee colony algorithm as feature selector to predict the leadership perception of site managers. The Computer Journal, 64(3):408– 417, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/408/6046268.

Kubilay:2021:KEP

[KKM21] Murat Yasin Kubilay, Mehmet Sabir Kiraz, and Haci Ali Mantar. KORGAN: an efficient PKI architecture based on PBFT through dynamic threshold signatures. The Computer Journal, 64(4): 564–574, April 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/564/5890396.

Krishan:2023:MCL

[KM23] A. Krishan and D. Mittal. Multi-class liver cancer diseases classification using CT images. The Computer Journal, 66(3):525– 539, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/525/6398685.

Khan:2024:RUC

[KNWT24] Marwa Khan, Asma Naseer, Aamir Wali, and Maria Tamoor. A Roman Urdu corpus for sentiment analysis. The Computer Journal, 67(9):2864–2876, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2864/7695664.

Kumar:2020:AIS

[KP20] B. Mathan Kumar and R. PushpaLakshmi. An approach for image search and retrieval by clusterbased indexing of binary MKSIFT codes. The Computer Journal, 63(6):857– 879, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/6/857/5728652.

Kumar:2020:WFW

[KRR20] J. Thrisul Kumar, Y. Mallikarjuna Reddy, and B. Prabhakara Rao. WHDA-FCM: Wolf hunting-based dragonfly with fuzzy C-mean clustering for change detection in SAR images. The Computer Journal, 63(2):308– 321, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/308/5667455.

Kaur:2020:AIB

[KS20a] Amandeep Kaur and Sandeep K.

Sood. Artificial intelligencebased model for drought prediction and forecasting. The Computer Journal, 63(11):1704–1712, November 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1704/5618957.

Kaur:2020:SDM

[KS20b] Harkiran Kaur and Sandeep K.

Sood. A smart disaster management framework for wildfire detection and prediction. The Computer Journal, 63(11):1644–1657, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1644/5612729.

K:2021:PPB

[KS21] Reshma V K and Vinod Kumar R S. Pixel prediction-based image steganography by support vector neural network. The Computer Journal, 64(5):731– 748, May 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/731/5819403.

Kumar:2024:SNA

[KS24] Ashutosh Kumar and Aakanksha Sharaff. SnorkelPlus: a novel approach for identifying relationships among biomedical entities within abstracts. The Computer Journal, 67(3):1187–1200, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1187/7152254.

Karacay:2020:IDE

[KSA20] Leyli Karaçay, Erkay Savaş, and Halit Alptekin. Intrusion detection over encrypted network data. The Computer Journal, 63(4): 604–619, April 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/4/604/5618960.

Kansal:2022:EMS

[KSD22] Meenakshi Kansal, Amit Kumar Singh, and Ratna Dutta. Efficient multisignature scheme using lattice. The Computer Journal, 65(9):2421–2429, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2421/6289877.

Kashyap:2020:DSF

[KSG20a] Abhishek Kashyap, B. Suresh, and Hariom Gupta. Detection of splicing forgery using differential evolution and wavelet decomposition. The Computer Journal, 63(11):1727–1737, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1727/5618852.

Kaur:2020:RIA

[KSG20b] Parampreet Kaur, Ashima Singh, and Sukhpal Singh Gill. RGIM: an integrated approach to improve QoS in AODV, DSR and DSDV routing protocols for FANETS using the chain mobility model. The Computer Journal, 63(10):1500– 1512, October 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1500/5827627.

Kashyap:2022:RDC

[KSG22] Abhishek Kashyap, B. Suresh, and Hariom Gupta. Robust detection of copy-move forgery based on wavelet decomposition and firefly algorithm. The Computer Journal, 65(4):983– 996, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/983/6012624.

Kobayashi:2023:SSD

[KSKM23] Hisaki Kobayashi, Yuichi Sudo, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. A self-stabilizing distributed algorithm for the generalized dominating set problem with safe convergence. The Computer Journal, 66(6):1452– 1476, June 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1452/6554131.

Khate:2024:FPE

[KSN24] Kevisino Khate, Bam Bahadur Sinha, and Arambam Neelima. Flower pollination-enhanced CNN for lung disease diagnosis. The Computer Journal, 67(11):3080–3093, November 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/11/3080/7723329.

Kaur:2024:EPI

[KSS^{+}24] Ravneet Kaur, Amrinderpreet Singh, Aekamjot Singh, Amit Goyal, Amritpal Singh, and Shalini Batra. An efficient pending interest table content search in NDN through stable Bloom filter. The Computer Journal, 67(3):941– 946, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/941/7100602.

Kamei:2024:AMI

[KT24] Sayaka Kamei and Sébastien Tixeuil. An asynchronous maximum independent set algorithm by myopic luminous robots on grids. The Computer Journal, 67(1): 57–77, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/57/6834177.

Khanam:2021:AIS

[KTK21] Sana Khanam, Safdar Tanweer, and Syed Khalid. Artificial intelligence surpassing human intelligence: Factual or hoax. The Computer Journal, 64(12):1832– 1839, December 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1832/5688168.

Khanam:2023:YTV

[KTK23] Sana Khanam, Safdar Tanweer, and Syed Sibtain Khalid. Youtube trending videos: Boosting machine learning results using exploratory data analysis. The Computer Journal, 66(1):35–46, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/35/6404045.

Kuang:2025:RBB

[KXLL25] Jian Kuang, Bailin Xie, Jianzhen Luo, and Jianhong Li. Reliable broadcastingbased content acquisition for named data MANETs. The Computer Journal, 68 (1):84–96, January 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/68/1/84/
7762870.

Krishnan:2021:EEC

[KYAN21] Kannan Krishnan, B. Yamini, Wael Mohammad Alenazy, and M. Nalini. Energyefficient cluster-based routing protocol for WSN based on hybrid BSO– TLBO optimization model. The Computer Journal, 64(10):1477–1493, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1477/6287332.

Khurshid:2022:BRS

[KZH^{+}22] Faisal Khurshid, Yan Zhu, Jie Hu, Muqeet Ahmad, and Mushtaq Ahmad. Battering review spam through ensemble learning in imbalanced datasets. The Computer Journal, 65(7):1666– 1678, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1666/6189766.

Li:2023:MVS

[L^{+}23a] He Li et al. Multi-view spatial–temporal graph neural network for traffic prediction. The Computer Journal, 66(10):2393–2408, October 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2393/6631438.

Li:2023:TSS

[L^{+}23b] Ning Li et al. Trapezoidal sketch: a sketch structure for frequency estimation of data streams. The Computer Journal, 66(11):2656– 2673, November 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2656/6658859.

Li:2023:EFC

[L^{+}23c] Yao Li et al. Ensemble framework combining family information for Android malware detection. The Computer Journal, 66(11):2721–2740, November 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2721/6672798.

Liqin:2023:CBV

[L^{+}23d] Xiong Liqin et al. Characterbased value factorization for MADRL. The Computer Journal, 66(11):2782– 2793, November 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2782/6702392.

Lu:2023:VCD

[L^{+}23e] Haitang Lu et al. Verifiable conjunctive dynamic searchable symmetric encryption with forward and backward privacy. The Computer Journal, 66(10):2379–2392, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/10/2379/6628786.

Luo:2023:CCB

[L^{+}23f] Xuanhao Luo et al. A clothoid curve-based intersection collision warning scheme in Internet of Vehicles. The Computer Journal, 66(10):2447–2461, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2447/6651323.

Luo:2022:PDP

[LAKL^{+}22] Fucai Luo, Saif Al-Kuwari, Changlu Lin, Fuqun Wang, and Kefei Chen. Provable data possession schemes from standard lattices for cloud computing. The Computer Journal, 65 (12):3223–3239, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3223/6457160.

Lei:2020:NEL

[LC20] Zhendong Lei and Shaowei Cai. NuDist: an efficient local search algorithm for (weighted) partial MaxSAT. The Computer Journal, 63(9):1321–1337, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1321/5544159.

Li:2022:IMM

[LC22] Manman Li and Shaozhen Chen. Improved meetin-the-middle attacks on reduced-round tweakable block cipher Deoxys-BC. The Computer Journal, 65(9):2411–2420, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2411/6291059.

Liu:2024:SFT

[LC24a] Heqin Liu and Dongqin Cheng. Structure fault tolerance of exchanged hypercube. The Computer Journal, 67(2):527–536, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/527/6995426.

Luo:2024:PPN

[LC24b] Qin Luo and Xin Cao. PNCTS: a prediction and network coding-based transmission scheme for efficient screen updates delivery in DaaS. The Computer Journal, 67(1):153–168, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/153/6847609.

Li:2023:LDM

[LCC23] Yeting Li, Haiming Chen, and Zixuan Chen. Learning disjunctive multiplicity expressions and disjunctive generalize multiplicity expressions from both positive and negative examples. The Computer Journal, 66(7):1733–1748, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1733/6569539.

See correction [Ano23c].

Lin:2024:RDD

[LCC^{+}24] Tu-Liang Lin, Hong-Yi Chang, Yuan-Yao Chiang, Shu-Cheng Lin, Tsung-Yen Yang, Chun-Jun Zhuang, Wha-Lee Tseng, and Bo-Hao Zhang. Ransomware detection by distinguishing API call sequences through LSTM and BERT models. The Computer Journal, 67(2):632–641, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/632/7067524.

Lv:2021:CRE

[LCF^{+}21] Mengjie Lv, Baolei Cheng, Jianxi Fan, Xi Wang, Jingya Zhou, and Jia Yu. The conditional reliability evaluation of data center network BCDC. The Computer Journal, 64(9):1451– 1464, September 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1451/5901609.

Lu:2022:RUD

[LCX^{+}22] Keyu Lu, Tingting Chai, Haiyan Xu, Shitala Prasad, Jianen Yan, and Zhaoxin Zhang. Research on unexpected DNS response from open DNS resolvers. The Computer Journal, 65(9):2276–2298, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2276/6281324.

Lv:2022:APL

[LCZ^{+}22] Zhuo Lv, Hongbo Cao, Feng Zhang, Yuange Ren, Bin Wang, Cen Chen, Nuannuan Li, Hao Chang, and Wei Wang. AWFC: Preventing label flipping attacks towards federated learning for intelligent IoT. The Computer Journal, 65(11):2849–2859, November 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/11/2849/6741165.

Liu:2021:PPS

[LDC^{+}21] Yuying Liu, Shaoyi Du, Wenting Cui, Xijing Wang, Qingnan Mou, Jiamin Zhao, Yucheng Guo, and Yong Zhang. Precise point set registration based on feature fusion. The Computer Journal, 64(7):1039– 1055, July 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/1039/6348008.

Liu:2025:RAW

[LDCX25] Hangda Liu, Boyu Diao, Wenxin Chen, and Yongjun Xu. A resource-aware workload scheduling method for unbalanced GEMMs on GPUs. The Computer Journal, 68(3):273– 282, March 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/273/7841854.

Liu:2023:TSS

[LDFD23] Xiaoyang Liu, Shanghong Dai, Giacomo Fiumara, and Pasquale De Meo. Targetspecific sentiment analysis method combining wordmasking data enhancement and adversarial learning. The Computer Journal, 66(9):2138–2154, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2138/6603445.

Liu:2024:DTQ

[LDW^{+}24] Xiaoyang Liu, Nan Ding, Yudie Wu, Giacomo Fiumara, and Pasquale De Meo. Discrete-time quantum walks community detection in multi-domain networks. The Computer Journal, 67(6):2379–2389, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2379/7603800.

Li:2021:MSC

[LdXZ^{+}21] Juan Li, Dan dan Xiao, Ting Zhang, Chun Liu, Yuan xiang Li, and Gai ge Wang. Multi-swarm cuckoo search algorithm with Q-learning model. The Computer Journal, 64(1):108–131, January 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/108/5802861.

See corrigendum [LdXZ^{+}22].

Li:2022:CMS

[LdXZ^{+}22] Juan Li, Dan dan Xiao, Ting Zhang, Chun Liu, Yuan xiang Li, and Gai ge Wang. Corrigendum to: Multi-swarm cuckoo search algorithm with Qlearning model. The Computer Journal, 65(7):1938, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/

article/65/7/1938/5910021.

See [LdXZ^{+}21].

Leeke:2020:SFM

[Lee20] Matthew Leeke. Simultaneous fault models for the generation and location of efficient error detection mechanisms. The Computer Journal, 63(5):758– 773, May 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/758/5481796.

Lemire:2024:ESP

[Lem24] Daniel Lemire. Exact short products from truncated multipliers. The Computer Journal, 67(4):1514–1520, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1514/7306807;
https://arxiv.org/abs/
2303.14321v1.

Liu:2024:CDR

[LFDF24] Xiaoyang Liu, Xiaoyang Fu, Pasquale De Meo, and Giacomo Fiumara. Crossdomain recommendation to cold-start users via categorized preference transfer. The Computer Journal, 67(8):2610–2621, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2610/7641747.

Lv:2024:FTH

[LFDW24] Mengjie Lv, Weibei Fan, Hui Dong, and Guijuan Wang. Fault tolerance of hierarchical cubic networks based on cluster fault pattern. The Computer Journal, 67(10):2890–2897, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2890/7709855.

Lv:2022:REK

[LFZ^{+}22] Mengjie Lv, Jianxi Fan, Jingya Zhou, Jia Yu, and Xiaohua Jia. The reliability of k-ary n-cube based on component connectivity. The Computer Journal, 65(8):2197–2208, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2197/6279277.

Liu:2024:FIC

[LFZ24] Changan Liu, Changjun Fan, and Zhongzhi Zhang. Finding influencers in complex networks: an effective deep reinforcement learning approach. The Computer Journal, 67(2):463– 473, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/463/6965541.

Liu:2024:AMC

[LGC^{+}24] Weihua Liu, Bing Gong, Kai Che, Jieming Ma, and Yushan Pan. Anomaly metrics on class variations for face anti-spoofing. The Computer Journal, 67(9):2725–2738, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2725/7663420.

Liu:2024:KNI

[LGFD24] Xiaoyang Liu, Luyuan Gao, Giacomo Fiumara, and Pasquale De Meo. Key node identification method integrating information transmission probability and path diversity in complex network. The Computer Journal, 67(1):127– 141, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/127/6832434.

Liu:2024:PIH

[LGX^{+}24] Mengqi Liu, Hang Gao, Xiaofan Xia, Suying Gui, and Tiegang Gao. Perceptual image hashing based on canny operator and tensor for copy–move forgery detection. The Computer Journal, 67(2):447– 462, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/447/6955258.

See correction [Ano24f].

Liu:2020:IPP

[LH20] Xiaoyang Liu and Daobing He. Information propagation and public opinion evolution model based on artificial neural network in online social network. The Computer Journal, 63(11):1689–1703, November 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1689/5612249.

Lu:2021:LRC

[LH21] Guangquan Lu and Jihong Huang. Learning representation from concurrencewords graph for aspect sentiment classification. The Computer Journal, 64(7): 1069–1079, July 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/7/1069/6326035.

Larkin:2022:IGD

[LH22] Andrew Larkin and Perry Hystad. Integrating geospatial data and social media in bidirectional longshort term memory models to capture human nature interactions. The Computer Journal, 65(3):667– 678, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/667/5893915.

Lu:2023:DUA

[LHH23] Cong Lu, Jianbin Huang, and Longji Huang. Detecting urban anomalies using factor analysis and one class support vector machine. The Computer Journal, 66(2):373–383, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/373/6404044.

Lai:2022:PSO

[LHHW22] Jianchang Lai, Xinyi Huang, Debiao He, and Wei Wu. Provably secure online/ offline identity-based signature scheme based on SM9. The Computer Journal, 65(7):1692–1701, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1692/6189769.

Li:2021:RFC

[LHL21] Qing Yun Li, Jie Han, and Lin Lu. A random forest classification algorithm based personal thermal sensation model for personalized conditioning system in office buildings. The Computer Journal, 64(3):500– 508, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/500/6120296.

Liu:2020:NAB

[LHLY20] Guangfeng Liu, Xianying Huang, Xiaoyang Liu, and Anzhi Yang. A novel aspect-based sentiment analysis network model based on multilingual hierarchy in online social network. The Computer Journal, 63(3):410– 424, March 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/410/5485130.

Liu:2021:EAB

[LHW21] Zhen Liu, Qiong Huang, and Duncan S. Wong. On enabling attribute-based encryption to be traceable against traitors. The Computer Journal, 64(4):575– 598, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/575/5874143.

Liu:2024:OFC

[Liu24a] Jia-Jie Liu. The orbits of folded crossed cubes. The Computer Journal, 67(5): 1719–1726, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1719/7306810.

Liu:2024:AAN

[Liu24b] Tzong-Jye Liu. Adversarial attacks on network intrusion detection systems using flow containers. The Computer Journal, 67(2):728–745, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/728/7075446.

Long:2023:NTC

[LJ23] Zhang Long and Wang Jinsong. Network traffic classification based on a deep learning approach using NetFlow data. The Computer Journal, 66(8):1882– 1892, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1882/6582000.

Li:2024:DHG

[LJLQ24] He Li, Duo Jin, XueJiao Li, and Shaojie Qiao. A dynamic heterogeneous graph convolution network for traffic flow prediction. The Computer Journal, 67(1): 31–44, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/31/6832431.

Louati:2024:ISS

[LKA24] Faten Louati, Farah Barika Ktata, and Ikram Amous. An intelligent security system using enhanced anomaly-based detection scheme. The Computer Journal, 67(6):2317–2330, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2317/7606361.

Li:2021:VNE

[LL21] Meng Li and MeiLian Lu. A virtual network embedding algorithm based on double-layer reinforcement learning. The Computer Journal, 64(6):973– 989, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/973/6280580.

Liu:2022:SNL

[LL22] Xiaoyang Liu and Xiang Li. A social network link prediction method based on stacked generalization. The Computer Journal, 65(10):2693–2708, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2693/6330818.

Liu:2023:CRD

[LL23] Pingshan Liu and Shuyue Lv. Chinese RoBERTa distillation for emotion classification. The Computer Journal, 66(12):3107–3118, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3107/6821147.

Lin:2024:FFL

[LL24a] Zhaoliang Lin and Jinguo Li. FedEVCP: Federated learning-based anomalies detection for electric vehicle charging pile. The Computer Journal, 67(4):1521– 1530, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1521/7237947.

Lu:2024:MQM

[LL24b] Meilian Lu and Meng Li. A multiple QoS metrics-aware virtual network embedding algorithm. The Computer Journal, 67(3):1171–1186, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1171/7147952.

Lu:2022:EKE

[LLAL22] Jinyu Lu, Yunwen Liu, Tomer Ashur, and Chao Li. On the effect of the key-expansion algorithm in Simon-like ciphers. The Computer Journal, 65(9):2454–2469, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2454/6314722.

Lu:2024:CTW

[LLCL24] Jiqiang Lu, Jingyu Li, Zexuan Chen, and Yanan Li. Cryptanalysis of a type of white-box implementations of the SM4 block cipher. The Computer Journal, 67(5):1663–1673, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1663/7277137.

Lv:2023:DDS

[LLDX23] Zhiqiang Lv, Jianbo Li, Chuanhao Dong, and Zhihao Xu. DeepSTF: a deep spatial–temporal forecast model of taxi flow. The Computer Journal, 66(3): 565–580, March 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/565/6428744.

Li:2021:RBE

[LLF^{+}21] Xiaoyan Li, Cheng-Kuan Lin, Jianxi Fan, Xiaohua Jia, Baolei Cheng, and Jingya Zhou. Relationship between extra connectivity and component connectivity in networks. The Computer Journal, 64(1):38–53, January 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/38/5644395.

Lin:2024:CGE

[LLF^{+}24] Nankai Lin, Xiaotian Lin, Yingwen Fu, Shengyi Jiang, and Lianxi Wang. A Chinese grammatical error correction model based on grammatical generalization and parameter sharing. The Computer Journal, 67(5): 1628–1636, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1628/7258762.

Li:2020:NBF

[LLG^{+}20] Zhidan Li, Wenmin Li, Fei Gao, Ping Yu, Hua Zhang Zhengping Jin, and Qiaoyan Wen. New blind filter protocol: an improved privacy-preserving scheme for location-based services. The Computer Journal, 63(12):1886–1903, December 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/12/1886/5856151.

Li:2020:RFE

[LLGC20] Yiming Li, Shengli Liu, Dawu Gu, and Kefei Chen. Reusable fuzzy extractor based on the LPN assumption. The Computer Journal, 63(12):1826–1834, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/12/1826/5854798.

Li:2022:FBC

[LLGC22] Zhiyao Li, Wei Liu, Xiaofeng Gao, and Guihai Chen. Friends-based crowdsourcing: Algorithms for task dissemination over social groups. The Computer Journal, 65(10):2615–2630, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2615/6324868.

Li:2025:SOD

[LLGP25] Sumin Li, Jinhua Lin, Yijin Gang, and Xiuqin Pan. Small object detection in remote sensing images through multi-scale feature fusion. The Computer Journal, 68(9):1329– 1340, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1329/8113253.

Li:2025:SGA

[LLL^{+}25] Qiuyu Li, Zuhe Li, Weihua Liu, Xiaojiang He, and Yushan Pan. Sarcasm-GPT: advancing sarcasm detection with large language models. The Computer Journal, 68(10):1523– 1534, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1523/8133221.

Li:2024:PSP

[LLM^{+}24] Nan Li, Yingjiu Li, Mark Manulis, Yangguang Tian, and Guomin Yang. Practical and secure policybased chameleon hash for redactable blockchains. The Computer Journal, 67 (11):3128–3139, November 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/11/3128/7723759.

Lu:2024:IRK

[LLS^{+}24] Jinyu Lu, Guoqiang Liu, Bing Sun, Chao Li, and Li Liu. Improved (relatedkey) differential-based neural distinguishers for SI-MON and SIMECK block ciphers. The Computer Journal, 67(2):537–547, February 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/537/6995424.

Liu:2025:UAU

[LLS^{+}25] Ruiting Liu, Ruixiang Li, Wenqi Shi, Shaoyong Du, Yimin Liu, and Xiangyang Luo. UeLP: accurate user linkage across social platforms against location errors. The Computer Journal, 68(3):228– 240, March 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/3/228/7833421.

Liu:2024:CAB

[LLT^{+}24] Kang Liu, Yang Lu, Shiyi Tan, Wei Liang, Huiping Sun, and Zhong Chen. CFAuditChain: Audit BlockChain based on cuckoo filter. The Computer Journal, 67(6):2208– 2218, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2208/7615717.

Li:2022:BBF

[LLW22] Taotao Li, Dequan Li, and Mingsheng Wang. Blockchain-based fair and decentralized data trading model. The Computer Journal, 65(8):2133–2145, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2133/6273372.

Li:2024:EPQ

[LLW24] Lingyun Li, Xianhui Lu, and Kunpeng Wang. eBiBa: a post-quantum hash-based signature with small signature size in the continuous communication of large-scale data. The Computer Journal, 67(4):1405– 1424, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1405/7229173.

Lu:2021:ETL

[LLY^{+}21] Li Lu, Jian Liu, Jiadi Yu, Yingying Chen, Yanmin Zhu, Linghe Kong, and Minglu Li. Enable traditional laptops with virtual writing capability leveraging acoustic signals. The Computer Journal, 64(12):1814–1831, December 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/12/1814/5699819.

Li:2023:IHC

[LLY^{+}23] He Li, Yanna Liu, Shuqi Yang, Yishuai Lin, Yi Yang, and Jaesoo Yoo. An improved hill climbing algorithm for graph partitioning. The Computer Journal, 66(7):1761–1776, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1761/6566842.

Ling:2021:EGI

[LMH^{+}21] Yunhao Ling, Sha Ma, Qiong Huang, Ximing Li, Yijian Zhong, and Yunzhi Ling. Efficient group ID-based encryption with equality test against insider attack. The Computer Journal, 64(4):661– 674, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/661/5910102.

Lahmidi:2022:FTP

[LMMR22] Ayoub Lahmidi, Khalid Minaoui, Chouaib Moujahdi, and Mohammed Rziza. Fingerprint template protection using irreversible minutiae tetrahedrons. The Computer Journal, 65(10):2741–2754, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2741/6341635.

Li:2022:THR

[LOGC22] Zhiyao Li, Siru Ouyang, Xiaofeng Gao, and Guihai Chen. Two-hop relay deployment based on user trajectory in wireless networks. The Computer Journal, 65(12):3106–3122, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3106/6372941.

Liu:2021:EDQ

[LPH21] Jun Liu, Yicheng Pan, and Qifu Hu. Exact distance query in large graphs through fast graph simplification. The Computer Journal, 64(1):93–107, January 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/1/93/5734741.

Liu:2021:CRA

[LPP21] Zhijun Liu, Xuefeng Pan, and Yuan Peng. Character recognition algorithm based on fusion probability model and deep learning. The Computer Journal, 64(11):1705–1714, November 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1705/5819224.

Li:2023:CSU

[LQCM23] Yue Li, Jianfang Qi, Xiaoquan Chu, and Weisong Mu. Customer segmentation using K-means clustering and the hybrid particle swarm optimization algorithm. The Computer Journal, 66(4):941– 962, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/4/941/6501352.

Li:2024:IGX

[LQJ^{+}24] Yue Li, Jianfang Qi, Haibin Jin, Dong Tian, Weisong Mu, and Jianying Feng. An improved genetic-XGBoost classifier for customer consumption behavior prediction. The Computer Journal, 67(3):1041–1059, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1041/7136737.

Lei:2024:MSA

[LQZ^{+}24] Yu Lei, Keshuai Qu, Yifan Zhao, Qing Han, and Xuguang Wang. Multimodal sentiment analysis based on composite hierarchical fusion. The Computer Journal, 67(6):2230– 2245, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2230/7595364.

Liu:2024:BBC

[LRG^{+}24] Shuai Liu, Jizhou Ren, Jie Guan, Bin Hu, Sudong Ma, and Hao Bai. A break of barrier to classical differential fault attack on the nonce-based authenticated encryption algorithm. The Computer Journal, 67(4): 1370–1380, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1370/7217095.

Lokku:2022:OSI

[LRP22] Gurukumar Lokku, G. Harinatha Reddy, and M. N. Giri Prasad. Optimized scaleinvariant feature transform with local tri-directional patterns for facial expression recognition with deep learning model. The Computer Journal, 65(9):2506– 2527, September 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2506/6311549.

Li:2023:RRB

[LS23] Jie Li and GuoYing Sun. RSCOEWR: Radical-based sentiment classification of online education Website reviews. The Computer Journal, 66(12):3000–3014, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3000/6793818.

Luo:2022:CFG

[LSC^{+}22] Yiqin Luo, Yanpeng Sun, Liang Chang, Tianlong Gu, Chenzhong Bin, and Long Li. Considering fine-grained and coarse-grained information for context-aware recommendations. The Computer Journal, 65(3): 679–688, March 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/679/5880461.

Liao:2025:BPM

[LSF25] Gengjian Liao, Shiyu Shao, and Zhenni Feng. Balancing privacy and model performance in federated learning through contract-based data trading. The Computer Journal, 68(9):1285– 1299, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1285/8110712.

Liu:2020:IMM

[LSG^{+}20] Ya Liu, Bing Shi, Dawu Gu, Fengyu Zhao, Wei Li, and Zhiqiang Liu. Improved meet-in-the-middle attacks on reduced-round Deoxys-BC-256. The Computer Journal, 63(12):1859–1870, December 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/12/1859/5860022.

Lv:2023:ADS

[LSG^{+}23] Yin Lv, Danping Shi, Yi Guo, Qiu Chen, Lei Hu, and Zihui Guo. Automatic Demirci–Selçuk meet-in-the-middle attack on SI-MON. The Computer Journal, 66(12):3052–3068, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3052/6772525.

Li:2022:QTG

[LSH^{+}22] Yang Li, Heli Sun, Liang He, Jianbin Huang, Jiyin Chen, Hui He, and Xiaolin Jia. Querying tenuous group in attributed networks. The Computer Journal, 65(4):858– 873, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/4/858/5898662.

Lv:2024:ILC

[LSH^{+}24] Yin Lv, Danping Shi, Lei Hu, Zihui Guo, Yi Guo, and Caibing Wang. Improved linear cryptanalysis of block cipher BORON. The Computer Journal, 67(1):210– 219, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/210/6955889.

Lin:2020:LFI

[LSQ20] Xi-Jun Lin, Lin Sun, and Haipeng Qu. Leakage-free ID-based signature, revisited. The Computer Journal, 63(8):1263–1270, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/8/1263/5716157.

Lou:2024:RAL

[LSS24] Paul Lou, Amit Sahai, and Varun Sivashankar. Relinearization attack on LPN over large fields. The Computer Journal, 67(4):1438– 1442, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1438/7224395.

Lin:2023:SLR

[LSW^{+}23] Hao Lin, Shi-Feng Sun, Mingqiang Wang, Joseph K. Liu, and Weijia Wang. Shorter linkable ring signature based on middleproduct learning with errors problem. The Computer Journal, 66(12):2974– 2989, December 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/2974/6768047.

Li:2021:LUC

[LSX^{+}21] Cong Li, Qingni Shen, Zhikang Xie, Xinyu Feng, Yuejian Fang, and Zhonghai Wu. Large universe CCA2 CP-ABE with equality and validity test in the standard model. The Computer Journal, 64(4):509– 533, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/4/509/5872129.

Lin:2020:SCS

[LSY^{+}20] Xi-Jun Lin, Lin Sun, Zhen Yan, Xiaoshuai Zhang, and Haipeng Qu. On the security of a certificateless signcryption with known session-specific temporary information security in the standard model. The Computer Journal, 63(8):1259– 1262, August 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/8/1259/5699818.

Liu:2021:DLN

[LTH21] Xiaoyang Liu, Ting Tang, and Daobing He. Doublelayer network negative public opinion information propagation modeling based on continuous-time Markov chain. The Computer Journal, 64(9):1315– 1325, September 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/9/1315/5828288.

Liu:2022:PKA

[LTT^{+}22] Zi-Yuan Liu, Yi-Fan Tseng, Raylin Tso, Masahiro Mambo, and Yu-Chi Chen. Public-key authenticated encryption with keyword search: a generic construction and its quantumresistant instantiation. The Computer Journal, 65 (10):2828–2844, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2828/6364551.

Liu:2025:MGM

[LUW^{+}25] Shu Liu, Zain Ul Abideen, Tongming Wan, Inzamam Shahzad, Abbas Waseem, and Yushan Pan. MuST-GAN MFAS: Multisemantic spoof tracer GAN with transformer layers for multi-modal face antispoofing. The Computer Journal, 68(8):891– 907, August 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/891/8010272.

Li:2023:FLC

[LW23a] Kaiju Li and Hao Wang.

Federated learning communicationefficiency framework via corset construction. The Computer Journal, 66 (9):2077–2101, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2077/6589512.

Li:2023:CBM

[LW23b] Xiaodan Li and Wenling Wu. Constructing binary matrices with good implementation properties for low-latency block ciphers based on Lai–Massey structure. The Computer Journal, 66(1):160–173, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/160/6381817.

Li:2025:OAS

[LWJ25] Yuke Li, Yong Wang, and Bin Jiang. Offset attention with seed generation for point cloud completion. The Computer Journal, 68(6):605– 616, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/6/605/7930080.

Liu:2022:EAS

[LWL^{+}22] Tonglai Liu, Jigang Wu, Jiaxing Li, Jingyi Li, and Zikai Zhang. Efficient algorithms for storage load balancing of outsourced data in blockchain network. The Computer Journal, 65(6): 1512–1526, June 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1512/6146586.

Liu:2025:SRS

[LWL^{+}25] Jianan Liu, Yongjuan Wang, Siqi Lu, Gang Yu, Xiangyu Wang, and Peixuan Feng. SAAChain: release and storage platform of digital works based on non-fungible tokens. The Computer Journal, 68(2):187–200, February 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/187/7960058.

Lin:2022:SOD

[LWM^{+}22] Xiao Lin, Zhi-Jie Wang, Lizhuang Ma, Renjie Li, and Mei-E Fang. Salient object detection based on multiscale segmentation and fuzzy broad learning. The Computer Journal, 65(4): 1006–1019, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/1006/6042246.

Lin:2021:SAF

[LWS^{+}21] Xi-Jun Lin, Qihui Wang, Lin Sun, Zhen Yan, and Peishun Liu. Security analysis of the first certificateless proxy signature scheme against malicious-but-passive KGC attacks. The Computer Journal, 64(4):653–660, April 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/653/5880730.

Li:2022:TBP

[LWX22] Shimin Li, Xin Wang, and Rui Xue. Toward both privacy and efficiency of homomorphic MACs for polynomial functions and its applications. The Computer Journal, 65(4):1020– 1028, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/1020/6281302.

Lin:2025:NCL

[LWZ^{+}25] Nankai Lin, Hongyan Wu, Peijian Zeng, Qifeng Bai, Dong Zhou, and Aimin Yang. A novel curriculum learning framework for multi-label emotion classification. The Computer Journal, 68(8):1050–1061, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/1050/8080214.

Liu:2024:BCP

[LWZR24] Ya Liu, Yuanhang Wu, Fengyv Zhao, and Yanli Ren. Balanced off-chain payment channel network routing strategy based on weight calculation. The Computer Journal, 67(3): 907–922, March 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/907/7085478.

Luo:2020:KCV

[LX20] Zuwen Luo and Liqiong Xu.

A kind of conditional vertex connectivity of Cayley graphs generated by wheel graphs. The Computer Journal, 63(9):1372–1384, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1372/5614857.

Li:2023:PCE

[LX23] Kaiju Li and Chunhua Xiao. PBFL: Communicationefficient federated learning via parameter predicting. The Computer Journal, 66(3):626–642, March 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/626/6438106.

Li:2025:FRA

[LXP25] Sumin Li, Zhen Xie, and Xiuqin Pan. Feature refinement and attention enhancement for click-through rate prediction. The Computer Journal, 68(6):697–705, June 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/6/697/7945687.

Luo:2025:HAT

[LXW^{+}25] Qianhua Luo, Bo Xie, Jiahuan Wang, Jiaqi Shuai, and Haixia Cui. A hybrid approach to task offloading optimization: integrating hybrid whale genetic algorithm and reinforcement learning. The Computer Journal, 68(9):1225–1236, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/9/1225/8110711.

Lin:2024:GLG

[LXWD24] Keao Lin, Xiaozhu Xie, Wei Weng, and Xiaofeng Du. Global-local graph attention: unifying global and local attention for node classification. The Computer Journal, 67(10):2959–2969, October 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/10/2959/7710998.

Lin:2020:LHS

[LXY^{+}20] Cheng-Jun Lin, Rui Xue, Shao-Jun Yang, Xinyi Huang, and Shimin Li. Linearly homomorphic signatures from lattices. The Computer Journal, 63 (12):1871–1885, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/12/1871/5824921.

Li:2020:ASB

[LYK^{+}20] Xian-Shu Li, Su-Kyung Yoon, Jeong-Geun Kim, Bernd Burgstaller, and Shin-Dug Kim. Algorithmswitching-based last-level cache structure with hybrid main memory architecture. The Computer Journal, 63(1):123–136, January 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/1/123/5310129.

Li:2024:KGV

[LYL24] Chenghao Li, Xinyan Yang, and Gang Liang. Keyframeguided Video Swin Transformer with multi-path excitation for violence detection. The Computer Journal, 67(5):1826–1837, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1826/7325339.

Liu:2022:RBB

[LYW^{+}22] Yan Liu, Haisheng Yu, Wenyong Wang, Sai Zou, Dong Liu, Daobiao Gong, and Zhen Li. A robust blockchain-based distribution master for distributing root zone data in DNS. The Computer Journal, 65(11):2880–2893, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/11/2880/6768049.

Li:2025:FFS

[LYZ^{+}25] Jinguo Li, Peichun Yuan, Jin Zhang, Sheng Shen, Yin He, and Ruyang Xiao. F2PQNN: a fast and secure two-party inference on quantized convolutional neural networks. The Computer Journal, 68(8):998– 1012, August 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/8/998/8102979.

Lin:2020:EEC

[LZ20a] Yuan Lin and Zhongzhi Zhang. Effects of edge centrality on random walks on graphs. The Computer Journal, 63(1):25–40, January 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/1/25/5259180.

Liu:2020:ASF

[LZ20b] Muhua Liu and Ping Zhang. An adaptively secure functional encryption for randomized functions. The Computer Journal, 63(8):1247–1258, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1247/5699816.

Li:2022:ESA

[LZ22] Sujuan Li and Futai Zhang.

eCK-secure authenticated key exchange against auxiliary input leakage. The Computer Journal, 65(8): 2063–2072, August 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/65/8/2063/
6269133.

Liu:2023:RIH

[LZ23] Shuai Liu and Yan Zhao.

Robust image hashing combining 3D space contour and vector angle features. The Computer Journal, 66(11):2844–2859, November 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2844/6694816.

Liu:2025:JTO

[LZ25] Jinlei Liu and Linbo Zhai. Joint task offloading and resource allocation for vehicle edge computing. The Computer Journal, 68(10):1510– 1522, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1510/8131473.

Lian:2021:PHD

[LZC^{+}21] Guanqin Lian, Shuming Zhou, Eddie Cheng, Jiafei Liu, and Gaolin Chen. Persistence of hybrid diagnosability of regular networks under testing diagnostic model. The Computer Journal, 64(9):1401– 1411, September 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/9/1401/5872132.

Liang:2021:MSM

[LZG21a] Caixia Liang, Bo Zhou, and Haiyan Guo. Minimum status, matching and domination of graphs. The Computer Journal, 64(9):1384– 1392, September 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/9/1384/5848717.

Lin:2021:LPK

[LZG^{+}21b] Hao Lin, Zhen Zhao, Fei Gao, Willy Susilo, Qiaoyan Wen, Fuchun Guo, and Yijie Shi. Lightweight public key encryption with equality test supporting partial authorization in cloud storage. The Computer Journal, 64(8):1226–1238, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1226/6025507.

Lian:2020:CDB

[LZH^{+}20] Guanqin Lian, Shuming Zhou, Sun-Yuan Hsieh, Gaolin Chen, Jiafei Liu, and Zhendong Gu. Characterization of diagnosabilities on the bounded PMC model. The Computer Journal, 63(9):1397–1405, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1397/5680451.

Li:2020:ARC

[LZL20] Chunlin Li, YiHan Zhang, and Youlong Luo. Adaptive replica creation and selection strategies for latencyaware application in collaborative edge–cloud system. The Computer Journal, 63(9):1338–1354, September 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/9/1338/5618850.

Liu:2021:OMA

[LZL^{+}21] Meiling Liu, Beixian Zhang, Xi Li, Weidong Tang, and GangQiang Zhang. An optimized k-means algorithm based on information entropy. The Computer Journal, 64(7):1130– 1143, July 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/7/1130/6291942.

Li:2023:REC

[LZL^{+}23] Xiao-Yan Li, Yufang Zhang, Ximeng Liu, Xiangke Wang, and Hongju Cheng. Reliability evaluation of clustered faults for regular networks under the probabilistic diagnosis model. The Computer Journal, 66(2):441–462, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/441/6408790.

Lu:2024:TAD

[LZL^{+}24] SenXing Lu, Mingming Zhao, Chunlin Li, Quanbing Du, and Youlong Luo. Time-aware data partition optimization and heterogeneous task scheduling strategies in spark clusters. The Computer Journal, 67(2):762–776, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/762/7077270.

Liu:2024:RAC

[LZLZ24] Xiaoqing Liu, Shuming Zhou, Jiafei Liu, and Hong Zhang. Reliability analysis of the cactus-based networks based on subsystem. The Computer Journal, 67(1):142–152, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/142/6847199.

Li:2020:NTO

[LZM^{+}20] Huizhong Li, Yongbin Zhou, Jingdian Ming, Guang Yang, and Chengbin Jin. The notion of transparency order, revisited. The Computer Journal, 63(12):1915–1938, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/12/1915/5866139.

Li:2022:RCC

[LZM^{+}22] Xiaowang Li, Shuming Zhou, Tianlong Ma, Xia Guo, and Xiangyu Ren. The h-restricted connectivity of a class of hypercubebased compound networks. The Computer Journal, 65(9):2528–2534, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2528/6330819.

Li:2024:LPB

[LZN^{+}24] Tongfeng Li, Ruisheng Zhang, Bojuan Niu, Yabing Yao, Jun Ma, Jing Jiang, and Zhili Zhao. Link prediction based on local structure and node information along local paths. The Computer Journal, 67(1): 45–56, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/45/6843556.

Liu:2023:MKR

[LZQ^{+}23] Jinlu Liu, Bo Zhao, Jing Qin, Xi Zhang, and Jixin Ma. Multi-keyword ranked searchable encryption with the wildcard keyword for data sharing in cloud computing. The Computer Journal, 66(1):184–196, January 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/184/6384775.

Li:2022:CTB

[LZQL22] Chunlin Li, Yihan Zhang, Xiaomei Qu, and Youlong Luo. Cost- and time-based data deployment for improving scheduling efficiency in distributed clouds. The Computer Journal, 65(4):874–889, April 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/874/5911726.

Liu:2022:NLN

[LZX^{+}22] Zeyi Liu, Weijuan Zhang, Ji Xiang, Daren Zha, and Lei Wang. NP-LFA: Nonprofiled leakage fingerprint attacks against improved rotating S-box masking scheme. The Computer Journal, 65(6):1598–1610, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1598/6178973.

Li:2020:RAB

[LZY20] Juan Li, Yanmin Zhu, and Jiadi Yu. Redundancyaware and budget-feasible incentive mechanism in crowd sensing. The Computer Journal, 63(1):66–79, January 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/66/5288327.

Lin:2025:SAS

[LZY^{+}25] Yuming Lin, Yinghao Zhang, Yan Yang, You Li, and Jingwei Zhang. Self-adaptive smoothing model for cardinality estimation. The Computer Journal, 68(4):360– 371, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/4/360/7889121.

Li:2025:RRD

[LZYT25] Deshun Li, Xinlong Zhao, Qiuling Yang, and Yuyin Tan. Research on reliability of DPCell structure. The Computer Journal, 68(10):1555–1564, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1555/8202950.

Liu:2021:CGG

[LZZZ21] Yufeng Liu, Xiaoqin Zeng, Kang Zhang, and Yang Zou. Coordinate graph grammar for the specification of spatial graphs. The Computer Journal, 64(5): 749–761, May 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/5/749/5819257.

Munagala:2023:BBI

[M^{+}23] N. V. L. M. Krishna Munagala et al. Blockchainbased Internet-of-Things for secure transmission of medical data in rural areas. The Computer Journal, 66(11):2705–2720, November 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2705/6671204.

Moghaddam:2020:NAS

[MA20] AmirHossein Ebrahimi Moghaddam and Zahra Ahmadian. New automatic search method for truncateddifferential characteristics application to Midori, SKINNY and CRAFT. The Computer Journal, 63(12):1813–1825, December 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/12/1813/5855734.

Ma:2025:LOB

[Ma25] Wenmin Ma. Learning to optimize based on rate decay. The Computer Journal, 68(8):908–925, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/908/8019597.

Marandi:2024:LBH

[MAAJ24] Ali Marandi, Pedro Geraldo M. R. Alves, Diego F. Aranha, and Rune Hylsberg Jacobsen. Latticebased homomorphic encryption for privacy-preserving smart meter data analytics. The Computer Journal, 67(5):1687–1698, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1687/7280723.

Manzali:2024:PSP

[MAB^{+}24] Youness Manzali, Yassine Akhiat, Khalidou Abdoulaye Barry, Elyazid Akachar, and Mohamed El Far. Prediction of student performance using random forest combined with Naïve Bayes. The Computer Journal, 67(8):2677–2689, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2677/7661614.

Mosa:2023:NHS

[MAEK23] Mona Mosa, Nedaa Agami, Ghada Elkhayat, and Mohamed Kholief. A novel hybrid segmentation approach for decision support: a case study in banking. The Computer Journal, 66(5): 1228–1240, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1228/6531899.

Manolopoulos:2024:TES

[Man24a] Yannis Manolopoulos. Thematic editorial: sentiment analysis. The Computer Journal, 67(7):2403–2407, July 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2403/7709771.

Manolopoulos:2024:TEU

[Man24b] Yannis Manolopoulos. Thematic editorial: The ubiquitous network. The Computer Journal, 67(3):809– 811, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/809/7641752.

Muppidi:2023:SSP

[MBA23] Satish Muppidi, Kishore Bhamidipati, and Sajeev Ram Arumugam. SPDPOA: Student psychology dragonfly political optimizer algorithm-based soil moisture and heat-level prediction for plant health monitoring in Internet of Things. The Computer Journal, 66(8):2059–2074, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/2059/6652001.

Meghanathan:2020:RAI

[Meg20] Natarajan Meghanathan. Relative assortativity index: a quantitative metric to assess the impact of link prediction techniques on assortativity of complex networks. The Computer Journal, 63(9):1417–1437, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/9/1417/5601486.

Mergen:2020:UME

[Mer20] Sergio L. S. Mergen. Unfair-DuelMerge: Merging with even fewer moves. The Computer Journal, 63(5): 701–708, May 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/5/701/5381951.

Mohammed:2021:ADM

[MFHhK21] Sabah Mohammed, Wai Chi Fang, Aboul Ella Hassanien, and Tai hoon Kim. Advanced data mining tools and methods for social computing. The Computer Journal, 64(3):281– 285, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/281/6212710.

Munoz:2025:CBC

[MFSH^{+}25] Martita Muñoz, José Fuentes-Sepúlveda, Cecilia Hernández, Gonzalo Navarro, Diego Seco, and Fernando Silva-Coira. Clustering-based compression for raster time series. The Computer Journal, 68(1):32–46, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/32/7789869.

Ma:2021:IKR

[MG21] Sudong Ma and Jie Guan.

Improved key recovery attacks on simplified version of K2 stream cipher. The Computer Journal, 64(8):1253–1263, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1253/6042244.

Mutlu:2022:ESH

[MG22] Alev Mutlu and Furkan Goz. SkySlide: a hybrid method for landslide susceptibility assessment based on landslideoccurring data only. The Computer Journal, 65(3): 473–483, March 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/473/5874134.

Muppidi:2023:DPO

[MGB23] Satish Muppidi, Om Prakash P G, and Kishore B. Dragonfly political optimizer algorithm-based rider deep long short-term memory for soil moisture and heat level prediction in IoT. The Computer Journal, 66(6): 1350–1365, June 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1350/6518263.

Ma:2020:NFI

[MH20] Sha Ma and Qiong Huang. A new framework of IND-CCA secure public key encryption with keyword search. The Computer Journal, 63(12):1849–1858, December 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/12/1849/5854458.

Ma:2021:CAF

[MH21] Sha Ma and Qiong Huang. CCA-almost-full anonymous group signature with verifier local revocation in the standard model. The Computer Journal, 64(8):1239–1252, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1239/6029314.

Mahdavi-Hezavehi:2020:EFT

[MHAR20] S. Mahdavi-Hezavehi, Y. Alimardani, and R. Rahmani. An efficient framework for a third party auditor in cloud computing environments. The Computer Journal, 63(9):1285–1297, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1285/5488731.

Mammar:2022:EBB

[MHG22] Amel Mammar, Lazhar Hamel, and Mohamed Graiet. An event-b-based approach to model and verify behaviors for component-based applications. The Computer Journal, 65(10):2780–2800, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2780/6367626.

Majidpour:2023:FIS

[MJQ23] Jafar Majidpour, Samer Kais Jameel, and Jihad Anwar Qadir. Face identification system based on synthesizing realistic image using edge-aided GANs. The Computer Journal, 66(1): 61–69, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/61/6381821.

Manshouri:2020:CAV

[MK20] Negin Manshouri and Temel Kayikcioglu. A comprehensive analysis of 2D&3D video watching of EEG signals by increasing PLSR and SVM classification results. The Computer Journal, 63(3):425–434, March 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/425/5485565.

Manocha:2023:IAI

[MKB23] Ankush Manocha, Gulshan Kumar, and Munish Bhatia. IoT analytics-inspired real-time monitoring for early prediction of COVID-19 symptoms. The Computer Journal, 66(1):144– 159, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/144/6384778.

Mohagheghi:2020:PMA

[MKI20]  Mohammadsadegh  Mo-

hagheghi, Jaber Karimpour, and Ayaz Isazadeh. Prioritizing methods to accelerate probabilistic model checking of discrete-time Markov models. The Computer Journal, 63(1):105– 122, January 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/1/105/5307039.

Masood:2022:SAI

[MKS^{+}22] Khalid Masood, Muhammad Adnan Khan, Usman Saeed, Mohammed A. Al Ghamdi, Muhammad Asif, and Muhammad Arfan. Semantic analysis to identify students’ feedback. The Computer Journal, 65(4): 918–925, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/918/5921713.

Ma:2020:SAR

[ML20] Xuecheng Ma and Dongdai Lin. Server-aided revocable IBE with identity reuse. The Computer Journal, 63(4):620– 632, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/620/5625927.

Ma:2025:STT

[MLX^{+}25] Zhaobin Ma, Zhiqiang Lv, Zhihao Xu, Rongkun Ye, and Jianbo Li. Spatiotemporal traffic flow forecasting based on secondorder continuous graph neural network. The Computer Journal, 68(9):1300– 1314, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1300/8117844.

Maisuria:2020:BBS

[MM20] Jemish V. Maisuria and Saurabh N. Mehta. Bayesianbased spectrum sensing and optimal channel estimation for MAC layer protocol in cognitive radio sensor networks. The Computer Journal, 63(6):942– 957, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/942/5739949.

Mondal:2021:PHD

[MM21] Safikureshi Mondal and Nandini Mukherjee. Pruning of health data in mobileassisted remote healthcare service delivery. The Computer Journal, 64(10):1549– 1564, October 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1549/6308678.

Manshouri:2022:DVT

[MMK22] Negin Manshouri, Mesut Melek, and Temel Kayıkcıoglu. Detection of 2D and 3D video transitions based on EEG power. The Computer Journal, 65(2):396– 409, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/396/5902223.

Mansouri:2023:IDQ

[MMMZ23] Taha Mansouri, Mohammad Reza Sadeghi Moghadam, Fatemeh Monshizadeh, and Ahad Zareravasan. IoT data quality issues and potential solutions: a literature review. The Computer Journal, 66(3):615– 625, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/615/6438757.

See corrigendum [Ano23e].

Mishra:2022:WCJ

[MMR22] U. K. Mishra, K. Mahalingam, and R. Rama. Watson–Crick jumping finite automata: Combination, comparison and closure. The Computer Journal, 65(5):1178–1188, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1178/6056295.

Manzano:2020:DSA

[MNN20] Wallace Manzano, Valdemar Vicente Graciano Neto, and Elisa Yumi Nakagawa. Dynamic-SoS: an approach for the simulation of systems-of-systems dynamic architectures. The Computer Journal, 63(5): 709–731, May 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/709/5448912.

Moulahi:2022:JFC

[Mou22]  Tarek Moulahi.  Join-

ing formal concept analysis to feature extraction for data pruning in Cloud of Things. The Computer Journal, 65(9):2484–2492, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2484/6301527.

Moorthy:2022:AAU

[MP22] Rajalakshmi Shenbaga Moorthy and P. Pabitha. Accelerating analytics using improved binary particle swarm optimization for discrete feature selection. The Computer Journal, 65(10):2547–2569, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2547/6316090.

Mo:2021:ELP

[MPL21] Xian Mo, Jun Pang, and Zhiming Liu. Effective link prediction with topological and temporal information using wavelet neural network embedding. The Computer Journal, 64(3):325– 336, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/325/5875044.

Meena:2023:CAO

[MPP23] Pawan Meena, Mahesh Pawar, and Anjana Pandey. Comparative analysis of overlap community detection techniques on social media platform. The Computer Journal, 66(8):1893– 1912, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1893/6575154.

Ma:2023:RGF

[MQL23] Jie Ma, Bin Qi, and Kewei Lv. Reusable group fuzzy extractor and group-shared bitcoin wallet. The Computer Journal, 66(3):643– 661, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/3/643/6484508.

Majji:2023:SSD

[MRVC23] Ramachandro Majji, R. Rajeswari, Ch Vidyadhari, and R. Cristin. Squirrel search deer hunting-based deep recurrent neural network for survival prediction using PAN-cancer gene expression data. The Computer Journal, 66(1):245– 266, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/1/245/6419809.

Mahalakshmi:2020:AFT

[MS20] T. Mahalakshmi and Alluri Sreenivas. Adaptive filter with type-2 fuzzy system and optimizationbased kernel interpolation for satellite image denoising. The Computer Journal, 63(6):913–926, June 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/6/913/5739947.

Mohamed:2023:PML

[MS23] Zahraa E. Mohamed and Hussein H. Saleh. Potential of machine learning based support vector regression for solar radiation prediction. The Computer Journal, 66(2):399–415, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/399/6406702.

Monti:2025:EGO

[MS25] Angelo Monti and Blerina Sinaimeri. Effects of graph operations on star pairwise compatibility graphs. The Computer Journal, 68(10):1355–1363, October 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1355/8116801.

Mumtaz:2022:FLT

[MSH22] Aqib Mumtaz, Allah Bux Sargano, and Zulfiqar Habib. Fast learning through deep multi-net CNN model for violence recognition in video surveillance. The Computer Journal, 65(3):457–472, March 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/457/5867750.

Magneta:2023:LLS

[MST23] S. Christina Magneta, C. Sundar, and M. S. Thanabal. Lung lobe segmentation and feature extraction-based hierarchical attention network for COVID-19 prediction from chest X-ray images. The Computer Journal, 66(2):508–522, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/508/6763329.

Ma:2020:DTS

[MSZ^{+}20] Zongjie Ma, Abdul Sattar, Jun Zhou, Qingliang Chen, and Kaile Su. Dropout with tabu strategy for regularizing deep neural networks. The Computer Journal, 63(7):1031–1038, July 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/7/1031/5541821.

Marjan:2023:EDM

[MUA23] Md Abu Marjan, Md Palash Uddin, and Masud Ibn Afjal. An educational data mining system for predicting and enhancing tertiary students’ programming skill. The Computer Journal, 66(5):1083–1101, May 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1083/6521709.

Ma:2021:DES

[MW21] Fei Ma and Ping Wang.

Determining exact solutions for structural parameters on hierarchical networks with density feature. The Computer Journal, 64(9):1412–1424, September 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1412/5878416.

Ma:2024:KSA

[MWH^{+}24] Dongchao Ma, Dongmei Wang, Xiaofu Huang, Yuekun Hu, and Li Ma. KEFSAR: a solar-aware routing strategy for rechargeable IoT based on highaccuracy prediction. The Computer Journal, 67(4): 1467–1482, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1467/7233669.

Mei:2024:EEF

[MWQ^{+}24] Xianyun Mei, Liangliang Wang, Baodong Qin, Kai Zhang, and Yu Long. EFTA: an efficient and fault-tolerant data aggregation scheme without TTP in smart Grid. The Computer Journal, 67(6):2368– 2378, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2368/7612814.

Ma:2025:RNF

[MXH^{+}25] Zengxie Ma, Yao Xin, Ning Hu, Tong Li, Zhaorui Zhang, and Feng Zhang. REFS: a novel framework for accelerated receive encrypted flow steering. The Computer Journal, 68(10):1565–1578, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1565/8145615.

Mahmoudi:2020:NRT

[MYB20] Amin Mahmoudi, Mohd Ridzwan Yaakub, and Azuraliza Abu Bakar. A new real-time link prediction method based on user community changes in online social networks. The Computer Journal, 63(3): 448–459, March 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/448/5498232.

Nofal:2021:CGE

[NAD21] Samer Nofal, Katie Atkinson, and Paul E. Dunne.

Computing grounded extensions of abstract argumentation frameworks. The Computer Journal, 64(1): 54–63, January 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/54/5644190.

Narayanan:2021:CTS

[Nar21] Hari T. S. Narayanan. Contact tracing solution for global community. The Computer Journal, 64 (10):1565–1574, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1565/6323600.

Narasimhulu:2022:NBW

[Nar22] C. Venkata Narasimhulu. A new blind watermark embedding model: Spiral updated rider optimization algorithm. The Computer Journal, 65(6):1365–1385, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1365/6124657.

Nancy:2021:TIB

[NB21] V. Nancy and G. Balakrishnan. Thermal imagebased object classification for guiding the visually impaired. The Computer Journal, 64(11):1747–1759, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1747/5879254.

Neffati:2020:ESK

[NBTB20] Syrine Neffati, Khaoula Ben Abdellafou, Okba Taouali, and Kais Bouzrara. Enhanced SVM–KPCA method for brain MR image classification. The Computer Journal, 63(3):383– 394, March 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/383/5480372.

Nouri:2022:GBC

[NCL22] Anass Nouri, Christophe Charrier, and Olivier Lézoray. A genetically based combination of visual saliency and roughness for FR 3D mesh quality assessment: a statistical study. The Computer Journal, 65(3):606– 620, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/606/5893202.

Narayana:2022:ACS

[NCRS22] Satyala Narayana, Suresh Babu Chandanapalli, Mekala Srinivasa Rao, and Kalyanapu Srinivas. Ant cat swarm optimization-enabled deep recurrent neural network for big data classification based on map reduce framework. The Computer Journal, 65(12):3167–3180, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3167/6409693.

Naveed:2022:EDD

[NGS22] Safia Naveed, Geetha G., and Leninisha S. Early diabetes discovery from tongue images. The Computer Journal, 65(2):237– 250, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/237/5834386.

Nawaz:2021:DQA

[NJ21] Falak Nawaz and Naeem Khalid Janjua. Dynamic QoSaware cloud service selection using best-worst method and timeslot weighted satisfaction scores. The Computer Journal, 64 (9):1326–1342, September 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1326/5850581.

Niu:2023:NGM

[NLD^{+}23] Weina Niu, Yuheng Luo, Kangyi Ding, Xiaosong Zhang, Yanping Wang, and Beibei Li. A novel generation method for diverse privacy image based on machine learning. The Computer Journal, 66(3):540– 553, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/540/6424462.

Nabipour:2023:NAA

[NM23] Mohammad Nabipour and Amir Reza Momen. A novel approach to adaptive resource allocation for energy saving in reconfigurable heterogeneous networks. The Computer Journal, 66(1):128–143, January 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/128/6410185.

Noor:2023:BAB

[Noo23] Talal H. Noor. Behavior analysis-based IoT services for crowd management. The Computer Journal, 66(9):2208–2219, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2208/6608474.

Nelson:2025:CDM

[NPF^{+}25] Jordan Nelson, Michalis Pavlidis, Andrew Fish, Stelios Kapetanakis, and Nikolaos Polatidis. ChatGPTdriven machine learning code generation for Android malware detection. The Computer Journal, 68(4): 331–345, April 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/331/7876388.

Nasir:2023:SMA

[NRA23] Muhammad Umar Nasir, Urva Rehmat, and Imran Ahmad. Social media analysis of customer emotions in pizza industry. The Computer Journal, 66(7):1777– 1783, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1777/6565162.

Ozcan:2021:PIP

[OB21] Tayyip Ozcan and Alper Basturk. Performance improvement of pre-trained convolutional neural networks for action recognition. The Computer Journal, 64(11):1715–1730, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1715/5856117.

ODwyer:2020:PSP

[OCPM20] Karl J. O’Dwyer, Eoin Creedon, Mark Purcell, and David Malone. Power saving proxies for Web servers. The Computer Journal, 63(2):179–192, February 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/179/5614859.

Ozogur:2023:AMD

[OEAA23] Gokhan Ozogur, Mehmet Ali Erturk, Zeynep Gurkas Aydin, and Muhammed Ali Aydin. Android malware detection in bytecode level using TF-IDF and XG-Boost. The Computer Journal, 66(9):2317–2328, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/9/2317/6995421.

Oro:2022:WEP

[OFMH22] David Oro, Carles Fernández, Xavier Martorell, and Javier Hernando. Workefficient parallel non-maximum suppression kernels. The Computer Journal, 65(4): 773–787, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/773/5894719.

Oksuz:2024:SSA

[Oks24] Ozgur Oksuz. A system for storing anonymous patient healthcare data using blockchain and its applications. The Computer Journal, 67(1):18–30, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/18/6832430.

Osmani:2022:SCU

[OMG22] Amjad Osmani, Jamshid Bagherzadeh Mohasefi, and Farhad Soleimanian Gharehchopogh. Sentiment classification using two effective optimization methods derived from the artificial bee colony optimization and imperialist competitive algorithm. The Computer Journal, 65(1): 18–66, January 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/18/5743591.

Oswald:2023:SMC

[OS23] C. Oswald and B. Sivaselvan. Smart multimedia compressor — intelligent algorithms for text and image compression. The Computer Journal, 66(2):463– 478, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/463/6414805.

Park:2023:PIE

[P^{+}23] Wooyoung Park et al. Practical implementation of encoding range top-2 queries. The Computer Journal, 66(11):2794–2809, November 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2794/6695274.

Pradeep:2022:CHA

[PAG^{+}22] K. Pradeep, L. Javid Ali, N. Gobalakrishnan, C. J. Raman, and N. Manikandan. CWOA: Hybrid approach for task scheduling in cloud environment. The Computer Journal, 65(7): 1860–1873, July 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/7/1860/6262249.

Periola:2020:ASD

[PAO20] A. A. Periola, A. A. Alonge, and K. A. Ogudo. Architecture and system design for marine cloud computing assets. The Computer Journal, 63(6):927– 941, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/927/5721769.

Prasad:2022:PIP

[PC22] Shitala Prasad and Tingting Chai. Palmprint for individual’s personality behavior analysis. The Computer Journal, 65(2):355– 370, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/355/5860494.

Prasad:2023:CLI

[PCLZ23] Shitala Prasad, Tingting Chai, Jiahui Li, and Zhaoxin Zhang. CR loss: Improving biometric using ClassRoom learning approach. The Computer Journal, 66(12):2897–2907, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/2897/6761065.

Payeras-Capella:2020:IEM

[PCMPCA^{+}20] M. Magdalena Payeras-Capella, Macia Mut-Puigserver, Pau Conejero-Alberola, Jordi Castella-Roca, and Llorenç Huguet-Rotger. Implementation and evaluation of the mCityPASS protocol for secure and private access to associated touristic services.

The Computer Journal, 63(8):1168–1193, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1168/5670506.

Payeras-Capella:2022:BBC

[PCMPCNHR22] M. Magdalena Payeras-Capellà, Macià Mut-Puigserver, Miquel À. Cabot-Nadal, and Llorenç Huguet-Rotger. Blockchain-based confidential multiparty contract signing protocol without TTP using elliptic curve cryptography. The Computer Journal, 65(10):2755– 2768, October 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2755/6334042.

Prabhakaran:2025:PAR

[PCS25] Manoj Kumar Prabhakaran, Abinaya Devi Chandrasekar, and Parvathy Meenakshi Sundaram. PHISH ATTENTION: achieving robust phishing website detection with balanced datasets and advanced URL features. The Computer Journal, 68(9):1263–1284, September 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1263/8113024.

Phyo:2022:DCA

[PDO22] Yati Phyo, Canh Minh Do, and Kazuhiro Ogata. A divide & conquer approach to leads-to model checking. The Computer Journal, 65(6):1353–1364, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1353/6125355.

Pradhan:2022:MLM

[PDRC22] Nitesh Pradhan, Vijaypal Singh Dhaka, Geeta Rani, and Himanshu Chaudhary. Machine learning model for multi-view visualization of medical images. The Computer Journal, 65(4):805–817, April 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/805/5896997.

Perez:2024:HAF

[Per24] J. Ayuso Perez. Hardware addition over finite fields based on Booth–Karatsuba algorithm. The Computer Journal, 67(8):2643–2666, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2643/7667243.

Peng:2025:WCS

[PFGW25] Min Peng, Xianxin Fu, Benling Ge, and Lusheng Wang. Wi-Crack: a smartphone keystroke recognition system based on multi-dimensional information. The Computer Journal, 68(1):1–12, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/1/7758128.

Pourian:2024:DLM

[PFT24] Reza Ebrahim Pourian, Mehdi Fartash, and Javad Akbari Torkestani. A deep learning model for energyaware task scheduling algorithm based on learning automata for fog computing. The Computer Journal, 67(2):508–518, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/508/6972877.

Puri:2022:IHB

[PG22] Arjun Puri and Manoj Kumar Gupta. Improved hybrid bag-boost ensemble with K-Means-SMOTE-ENN technique for handling noisy class imbalanced data. The Computer Journal, 65(1):124– 138, January 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/124/6262253.

Prasad:2022:PSP

[PGV^{+}22] Venkata Vara Prasad, Srinivas Gumparthi, Lokeswari Y. Venkataramana, S. Srinethe, R. M. Sruthi Sree, and K. Nishanthi. Prediction of stock prices using statistical and machine learning models: a comparative analysis. The Computer Journal, 65(5):1338– 1351, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1338/6317640.

Popova:2021:CAS

[PiKT21] Diana Popova, Ken ichi Kawarabayashi, and Alex Thomo. CutTheTail: an accurate and spaceefficient heuristic algorithm for influence maximization. The Computer Journal, 64(9):1343–1357, September 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1343/5860966.

Prakash:2020:WBB

[PJ20] Pg Om Prakash and A. Jaya. WS-BD-based two-level match: Interesting sequential patterns and Bayesian fuzzy clustering for predicting the Web pages from weblogs. The Computer Journal, 63(2):322–336, February 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/322/5644196.

Paik:2024:MFP

[PJ24] Joon-Young Paik and Rize Jin. Malware family prediction with an awareness of label uncertainty. The Computer Journal, 67(1):376–390, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/376/6927254.

Pantula:2020:MAR

[PK20] Muralidhar Pantula and K. S. Kuppusamy. A metric to assess the readability of video closed captions for the persons with low literacy skills. The Computer Journal, 63(7):1063– 1075, July 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/1063/5560308.

Prasath:2021:UIE

[PK21] R. Prasath and T. Kumanan. Underwater image enhancement with optimal histogram using hybridized particle swarm and dragonfly. The Computer Journal, 64(10):1494–1513, October 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1494/6287339.

Pantula:2022:MLB

[PK22] Muralidhar Pantula and K. S. Kuppusamy. A machine learning-based model to evaluate readability and assess grade level for the Web pages. The Computer Journal, 65(4):831– 842, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/831/5902220.

Peng:2021:RDG

[PKLK21] Liang Peng, Fei Kong, Chongzhi Liu, and Ping Kuang. Robust and dynamic graph convolutional network for multi-view data classification. The Computer Journal, 64(7):1093– 1103, July 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/1093/6299204.

Pan:2025:AON

[PL25] Ying Pan and Na Lyu. Attack-Oblivious Node Protection Strategy for survivable routing in airborne tactical networks under uncertainty. The Computer Journal, 68(6):675– 683, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/675/7942796.

Pang:2023:CCW

[PLY^{+}23] Bo Pang, Gang Liang, Jin Yang, Yijing Chen, Xinyi Wang, and Wenbo He. CWSOGG: Catching Web Shell obfuscation based on genetic algorithm and generative adversarial network. The Computer Journal, 66(5):1295–1309, May 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1295/6566841.

Pattanayak:2022:ACM

[PMMS22] Debasish Pattanayak, Kaushik Mondal, Partha Sarathi Mandal, and Stefan Schmid. Area convergence of monoculus robots with additional capabilities. The Computer Journal, 65(5):1306– 1319, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1306/6125293.

Pol:2022:QSE

[PP22] Pooja Shashank Pol and Vinod K. Pachghare. Quality of service estimation enabled with trust-based resource allocation in collaborative cloud using improved grey wolf optimization. The Computer Journal, 65(12):3209–3222, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3209/6396791.

Papan:2023:IPC

[PPR23] Bishal Basak Papan, Protik Bose Pranto, and Md Saidur Rahman. On 2-interval pairwise compatibility properties of two classes of grid graphs. The Computer Journal, 66(5): 1256–1267, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1256/6536120.

Pavithra:2020:ISP

[PS20] L. K. Pavithra and T. Sree Sharmila. An improved seed point selection-based unsupervised color clustering for content-based image retrieval application. The Computer Journal, 63(3): 337–350, March 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/337/5420616.

Prasad:2022:CSS

[PS22a] M. Krishna Siva Prasad and Poonam Sharma. Corrigendum to: Similarity of sentences with contradiction using semantic similarity measures. The Computer Journal, 65(10):2845, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/10/2845/5924456.

See [PS22b].

Prasad:2022:SSC

[PS22b] M Krishna Siva Prasad and Poonam Sharma. Similarity of sentences with contradiction using semantic similarity measures. The Computer Journal, 65(3):701– 717, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/701/5893603.

See corrigendum [PS22a].

Phu:2021:EAE

[PTH^{+}21] Tran Nghi Phu, Nguyen Dai Tho, Le Huy Hoang, Nguyen Ngoc Toan, and Nguyen Ngoc Binh. An efficient algorithm to extract control flow-based features for IoT malware detection. The Computer Journal, 64(4):599– 609, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/599/5940626.

Philip:2022:CVA

[PVFP22a] Felix M. Philip, Jayakrishnan V, Ajesh F, and Haseena P. Corrigendum to: Video anomaly detection using the optimizationenabled deep convolutional neural network. The Computer Journal, 65(5):1352, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1352/6224791.

See [PVFP22b].

Philip:2022:VAD

[PVFP22b] Felix M. Philip, Jayakrishnan V, Ajesh F, and Haseena P. Video anomaly detection using the optimizationenabled deep convolutional neural network. The Computer Journal, 65(5):1272– 1292, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/

                                        article/65/5/1272/6090337.

See corrigendum [PVFP22a].

Peng:2022:DBN

[PXWL22] Tao Peng, Thomas Canhao Xu, Yihuai Wang, and Fanzhang Li. Deep belief network and closed polygonal line for lung segmentation in chest radiographs. The Computer Journal, 65(5):1107–1128, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/5/1107/6029261.

Pan:2025:TRA

[PZL^{+}25] Yuedong Pan, Lixin Zhao, Tao Leng, Chaofei Li, Aimin Yu, and Lijun Cai. TaintAttack: rapid attack investigation based on information flow tracking. The Computer Journal, 68(7):813–829, July 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/7/813/7994526.

Qiao:2023:NCC

[Q^{+}23]  Zirui Qiao et al.  A

novel construction of certificateless aggregate signature scheme for healthcare wireless medical sensor networks. The Computer Journal, 66(11):2810–2824, November 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2810/6694815.

Qananwah:2022:DCF

[QAA^{+}22] Qasem Qananwah, Ali Mohammad Alqudah, Moh’d Alodat, Ahmad Dagamseh, and Oliver Hayden. Detecting cognitive features of videos using EEG signal. The Computer Journal, 65(1):105–123, January 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/105/6077136.

Qi:2020:HTR

[QDZZ20] Yi Qi, Yuze Dong, Zhongzhi Zhang, and Zhang Zhang. Hitting times for random walks on Sierpiński graphs and hierarchical graphs. The Computer Journal, 63(9):1385–1396, September 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1385/5612725.

Qi:2022:RRV

[QGL^{+}22] Wenfa Qi, Sirui Guo, Yuxin Liu, Xiang Wang, and Zongming Guo. Research on reversible visible watermarking algorithms based on vectorization compression method. The Computer Journal, 65(5):1320– 1337, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/5/1320/6120302.

Qi:2024:RBR

[QLH24] Wenfa Qi, Cheng Li, and Xinhui Han. Research on blind reversible database watermarking algorithm based on dual embedding strategy. The Computer Journal, 67(12):3169–3179, December 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/12/3169/7746723.

Liu:2020:GNC

[qLlHG20] Hui qing Liu, Xiao lan Hu, and Shan Gao. The ggood-neighbor conditional diagnosability of locally exchanged twisted cubes. The Computer Journal, 63(1): 80–90, January 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/80/5288326.

Qiu:2024:NMI

[QLZ24] Liqing Qiu, Yuying Liu, and Jianyi Zhang. A new method for identifying influential spreaders in complex networks. The Computer Journal, 67(1):362– 375, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/362/6927253.

Qadeer:2020:VIO

[QMR^{+}20] Arslan Qadeer, Asad Waqar Malik, Anis Ur Rahman, Hamayun Mian Muhammad, and Arsalan Ahmad. Virtual infrastructure orchestration for cloud service deployment. The Computer Journal, 63(2):295– 307, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/295/5631925.

Qian:2024:FFD

[QSC^{+}24] Wenjun Qian, Qingni Shen, Xiaoyi Chen, Cong Li, Yuejian Fang, and Zhonghai Wu. FDP-FL: differentially private federated learning with flexible privacy budget allocation. The Computer Journal, 67(12):3180–3195, December 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/12/3180/7762888.

Qi:2024:VVQ

[QWC24] Han Qi, Jingtong Wang, and Yufan Cui. VQCNN: variational quantum convolutional neural networks based on quantum filters and fully connected layers. The Computer Journal, 67(10):2970–2983, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2970/7721727.

Qi:2025:ILP

[QWHC25] Du Qi, Feng Wang, Hui Huang, and Jinlin Chen. Improving LAMMPS performance for molecular dynamic simulation on largescale HPC systems. The Computer Journal, 68(6): 706–716, June 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/706/7942797.

Qin:2025:PVT

[QXX^{+}25] Shihan Qin, Yuting Xiao, Yansen Xin, Birou Gao, and Rui Zhang. Practical and veritable threshold multi-factor authentication for mobile devices. The Computer Journal, 68(7): 749–762, July 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/749/8015436.

Qi:2021:RWN

[QXZZ21] Yi Qi, Wanyue Xu, Liwang Zhu, and Zhongzhi Zhang. Real-world networks are not always fast mixing. The Computer Journal, 64(2):236–244, February 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/236/6032261.

Qi:2025:FDP

[QYH^{+}25] Qingqiang Qi, Xingye Yang, Chengyu Hu, Peng Tang, Zhiyuan Su, and Shanqing Guo. FLDS: differentially private federated learning with double shufflers. The Computer Journal, 68(4):431– 443, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/431/7918703.

Qiao:2021:NPK

[QYZ^{+}21] Zirui Qiao, Qiliang Yang, Yanwei Zhou, Zhe Xia, and Mingwu Zhang. Novel public-key encryption with continuous leakage amplification. The Computer Journal, 64(8):1163–1177, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1163/5921729.

Qiu:2021:IIN

[QZTZ21] Liqing Qiu, Jianyi Zhang, Xiangbo Tian, and Shuang Zhang. Identifying influential nodes in complex networks based on neighborhood entropy centrality. The Computer Journal, 64(10):1465–1476, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1465/6284270.

Rui:2023:DLC

[R^{+}23] Lanlan Rui et al. Doublelead content search and producer location prediction scheme for producer mobility in named data networking. The Computer Journal, 66(11):2825–2843, November 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2825/6711415.

Raghu:2021:ROD

[RA21] A. Francis Alexander Raghu and J. P. Ananth. Robust object detection and localization using semantic segmentation network. The Computer Journal, 64(10):1531–1548, October 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1531/6295807.

Raghu:2022:ODL

[RA22a] A. Francis Alexander Raghu and J. P. Ananth. Object detection and localization using sparse-FCM and optimization-driven deep convolutional neural network. The Computer Journal, 65(5):1225–1241, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1225/6124655.

Rezaeipanah:2022:BCD

[RA22b] Amin Rezaeipanah and Gholamreza Ahmadi. Breast cancer diagnosis using multi-stage weight adjustment in the MLP neural network. The Computer Journal, 65(4):788– 804, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/788/5894134.

Rizkallah:2020:SLN

[RAD20] Lydia W. Rizkallah, Mona F.

Ahmed, and Nevin M. Darwish. SMT-LH: a new satisfiability modulo theorybased technique for solving vehicle routing problem with time window constraints. The Computer Journal, 63(1):91–104, January 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/91/5288325.

Ratre:2020:SGD

[Rat20] Avinash Ratre. Stochastic gradient descent–whale optimization algorithm-based deep convolutional neural network to crowd emotion understanding. The Computer Journal, 63(2):267– 282, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/267/5627747.

Rao:2022:HCS

[RCK22] B. Janardhana Rao, Y. Chakrapani, and S. Srinivas Kumar. Hybridized cuckoo search with multi-verse optimization-based patch matching and deep learning concept for enhancing video inpainting. The Computer Journal, 65(9):2315–2338, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2315/6291065.

Raj:2020:ANF

[RD20] Ajay Amrit Raj and Dejey. Adaptive neuro-fuzzy inference system-based nonlinear equalizer for CO-OFDM systems. The Computer Journal, 63(2):169– 178, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/169/5627775.

Rakhshaninejad:2022:EBC

[RFAY22] Morteza Rakhshaninejad, Mohammad Fathian, Babak Amiri, and Navid Yazdanjue. An ensemble-based credit card fraud detection algorithm using an efficient voting strategy. The Computer Journal, 65(8):1998– 2015, August 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/1998/6281303.

Ramchandran:2022:MAH

[RG22] M. Ramchandran and E. N.

Ganesh. MBSO algorithm for handling energythroughput trade-off in cognitive radio networks. The Computer Journal, 65(7): 1717–1725, July 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1717/6287333.

R:2023:TSF

[RG23] Sekar R. and Ravi G. Taylor sun flower optimizationbased compressive sensing for image compression and recovery. The Computer Journal, 66(4):873– 887, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/873/6503939.

Ratra:2024:GNN

[RGB^{+}24] Siftee Ratra, Mohona Ghosh, Niyati Baliyan, Jinka Rashmitha Mohan, and Sanjana Singh. Graph neural network based phishing account detection in Ethereum. The Computer Journal, 67(12):3160–3168, December 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/12/3160/7746724.

Ravishankar:2023:API

[RJ23] K. Ravishankar and C. Jothikumar. Analysis performance of image processing technique its application by decision support systems on Covid-19 disease prediction using convolution neural network. The Computer Journal, 66(4):1030– 1039, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/1030/6845388.

Rafiee:2021:PSO

[RK21] Mojtaba Rafiee and Shahram Khazaei. Private set operations over encrypted cloud dataset and applications. The Computer Journal, 64(8):1145–1162, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1145/5921321.

Ravala:2022:ADD

[RK22] Lavanya Ravala and Rajini G. K. Automatic diagnosis of diabetic retinopathy from retinal abnormalities: Improved jaya-based feature selection and recurrent neural network. The Computer Journal, 65(7): 1904–1922, July 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1904/6295808.

See correction [Ano24c].

Rizvi:2022:DEL

[RKA^{+}22] Syed Saqib Raza Rizvi, Muhammad Adnan Khan, Sagheer Abbas, Muhammad Asadullah, Nida Anwer, and Areej Fatima. Deep extreme learning machine-based optical character recognition system for Nastalique Urdu-like script languages. The Computer Journal, 65(2):331– 344, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/331/5860094.

Reyhan:2021:ITP

[RLIK21] Zahra Abdi Reyhan, Shahriar Lotfi, Ayaz Isazadeh, and Jaber Karimpour. Intratile parallelization for twolevel perfectly nested loops with non-uniform dependences. The Computer Journal, 64(9):1358–1383, September 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/9/1358/5846191.

Rezgui:2020:TSF

[RM20] Kalthoum Rezgui and Hédia Mhiri. Towards a semantic framework for lifelong integrated competency management and development. The Computer Journal, 63(7):1004–1016, July 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/7/1004/5543654.

Rao:2022:CEM

[RM22] Y. Srinivasa Rao and R. Madhu. Channel estimation for millimeter wave massive MIMO system: Proposed hybrid optimization with heuristicenabled precoding and combining. The Computer Journal, 65(5):1211–1224, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1211/6090335.

Rosales-Morales:2020:INA

[RMSMAH^{+}20] Viviana Yarel Rosales-Morales, Laura Nely Sánchez-Morales, Giner Alor-Hernández, Jorge Luis Garcia-Alcaraz, José Luis Sánchez-Cervantes, and Lisbeth Rodriguez-Mazahua. ImagIngDev: a new approach for developing automatic crossplatform mobile applications using image processing techniques. The Computer Journal, 63(5):732– 757, May 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/732/5476716.

Ren:2024:BDR

[RMW24] Dawei Ren, Lingwei Meng, and Rui Wang. Borehole depth recognition based on improved YOLOX detection. The Computer Journal, 67(7):2408–2420, July 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/7/2408/7606363.

Ramirez-Noriega:2022:MBQ

[RNJRLL^{+}22] Alan Ramírez-Noriega, Reyes Juárez-Ramírez, Juan Carlos Leyva-López, Samantha Jiménez, and J. Francisco Figueroa-Pérez. A method for building the quantitative and qualitative part of Bayesian networks for intelligent tutoring systems. The Computer Journal, 65(12):3035–3048, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3035/6369386.

R:2022:RCO

[RP22a] Vinoth R and Ananth J P. Rider chicken optimization algorithm-based recurrent neural network for big data classification in spark architecture. The Computer Journal, 65(8):2183–2196, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/2183/6278159.

Rani:2022:OEB

[RP22b] A. Jaya Mabel Rani and A. Pravin. Optimization enabled black hole entropic fuzzy clustering approach for medical data. The Computer Journal, 65(7):1795– 1811, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1795/6236075.

Ramtekkar:2024:CRB

[RPP24] Praveen Kumar Ramtekkar, Anjana Pandey, and Mahesh Kumar Pawar. A comprehensive review of brain tumour detection mechanisms. The Computer Journal, 67(3):1126–1152, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1126/7147953.

Roy:2021:MOC

[RS21] Alak Roy and Nityananda Sarma. Multichannel ordered contention MAC protocol for underwater wireless sensor networks. The Computer Journal, 64(2):185–194, February 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/185/5894897.

Rastegari:2022:ROF

[RS22] Parvin Rastegari and Willy Susilo. On random-oracle-free top-level secure certificateless signature schemes. The Computer Journal, 65(12):3049–3061, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3049/6369345.

Radha:2023:DBD

[RSSJ23] P. Radha, N. Selvakumar, J. Raja Sekar, and J. V. Johnsonselva. Densitybased dynamically selfparameterized clustering for material inspection. The Computer Journal, 66(2):416–428, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/416/6406701.

Ren:2020:DCG

[RW20] Yunxia Ren and Shiying Wang. Diagnosability of the Cayley graph generated by complete graph with missing edges under the MM—^{∗} — model. The Computer Journal, 63(9):1438–1447, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/9/1438/5612728.

Rasheed:2023:UAE

[RW23a] Fareeha Rasheed and Abdul Wahid. An unobtrusive approach to emotion detection in e-learning systems. The Computer Journal, 66(8):1840–1855, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1840/6569538.

Ren:2023:LDC

[RW23b] Yunxia Ren and Shiying Wang. The local diagnosability of a class of Cayley graphs with conditional faulty edges under the PMC model. The Computer Journal, 66(8):1913–1921, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1913/6574618.

Shi:2023:MCS

[S^{+}23] Juanjuan Shi et al. On the maximum cliques of the subgraphs induced by binary constant weight codes in powers of hypercubes. The Computer Journal, 66(10):2535–2541, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2535/6651999.

Sohane:2022:KMF

[SA22] Anurag Sohane and Ravinder Agarwal. Knee muscle force estimating model using machine learning approach. The Computer Journal, 65(5):1167–1177, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/5/1167/6032301.

Silva:2024:OSS

[SA24] Timóteo G. Silva and Fernanda Alencar. An ontology as support for specification of non-functional requirements of AAL systems considering compliance aspects. The Computer Journal, 67(4):1211– 1225, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1211/7169136.

Saeed:2024:DSO

[Sae24] Mozamel M. Saeed. Designing scenarios for inorganization training using the CyberCIEGE game. The Computer Journal, 67(1):338–346, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/338/6931872.

Sornavalli:2022:IFS

[SAK22] G. Sornavalli, Gladston Angelin, and Nehemiah H.

Khanna. Intelligent forecast of stock markets to handle COVID-19 economic crisis by modified generative adversarial networks. The Computer Journal, 65(12):3250–3264, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3250/6576181.

See correction [Ano23b].

Sarkar:2020:TSK

[Sar20] Bikash Kanti Sarkar. A two-step knowledge extraction framework for improving disease diagnosis. The Computer Journal, 63(3): 364–382, March 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/364/5476714.

Sajid:2022:CDL

[SAR^{+}22a] Muhammad Sajid, Nouman Ali, Naeem Iqbal Ratyal, Muhammad Usman, Faisal Mehmood Butt, Imran Riaz, Usman Musaddiq, Mirza Jabbar Aziz Baig, Shahbaz Baig, and Umair Ahmad Salaria. Corrigendum to: Deep Learning in Ageinvariant Face Recognition: a Comparative Study. The Computer Journal, 65(8): 2245, August 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/2245/6190175.

See [SAR^{+}22b].

Sajid:2022:DLA

[SAR^{+}22b] Muhammad Sajid, Nouman Ali, Naeem Iqbal Ratyal, Muhammad Usman, Faisal Mehmood Butt, Imran Riaz, Usman Musaddiq, Mirza Jabbar Aziz Baig, Shahbaz Baig, and Umair Ahmad Salaria. Deep learning in age-invariant face recognition: a comparative study. The Computer Journal, 65(4):940– 972, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/4/940/6023086.

See corrigendum [SAR^{+}22a].

Soucha:2020:OTA

[SB20] Michal Soucha and Kirill Bogdanov. Observation tree approach: Active learning relying on testing. The Computer Journal, 63(9):1298–1310, September 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/9/1298/5525443.

S:2022:SBL

[SB22] Smrithy G. S. and Ramadoss Balakrishnan. A statistical-based light-weight anomaly detection framework for wireless body area networks. The Computer Journal, 65(7):1752–1759, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1752/6236077.

Strappa:2024:ECI

[SB24] Jan Strappa and Facundo Bromberg. Efficient comparison of independence structures of log-linear models. The Computer Journal, 67(4):1226–1252, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1226/7192129.

Saraswat:2022:ETC

[SC22] Mala Saraswat and Shampa Chakraverty. Enriching topic coherence on reviews for cross-domain recommendation. The Computer Journal, 65(1):80–90, January 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/80/5827875.

Sarker:2020:CML

[SCH^{+}20] Iqbal H. Sarker, Alan Colman, Jun Han, A. S. M.

Kayes, and Paul Watters. CalBehav: a machine learning-based personalized calendar behavioral model using time-series smartphone data. The Computer Journal, 63(7):1109– 1123, July 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/1109/5610858.

Seth:2024:CDB

[SCS24] Sugandh Seth, Kuljit Kaur Chahal, and Gurvinder Singh. Concept driftbased intrusion detection for evolving data stream classification in IDS: Approaches and comparative study. The Computer Journal, 67(7):2529–2547, July 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/7/2529/7618465.

Sinha:2020:BFL

[SD20] Bam Bahadur Sinha and R. Dhanalakshmi. Building a fuzzy logic-based artificial neural network to uplift recommendation accuracy. The Computer Journal, 63(11):1624–1632, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/11/1624/5612726.

Srivastava:2023:MBP

[SD23] Vikas Srivastava and Sumit Kumar Debnath. A multivariatebased provably secure certificateless signature scheme with applications to the Internet Of Medical Things. The Computer Journal, 66(10):2499–2516, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2499/6649938.

Sepulvene:2022:PEM

[SDK^{+}22] Luis Sepulvene, Isabela Drummond, Bruno Kuehne, Rafael Frinhani, Dionisio Leite Filho, Maycon Peixoto, Stephan Reiff-Marganiec, and Bruno Batista. Performance evaluation of machine learning techniques for fault diagnosis in vehicle fleet tracking modules. The Computer Journal, 65(8):2073–2086, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2073/6275474.

Singh:2023:IUP

[SDM23] Sinam Ajitkumar Singh, Ningthoujam Dinita Devi, and Swanirbhar Majumder. An improved unsegmented phonocardiogram classification using nonlinear time scattering features. The Computer Journal, 66(6): 1525–1540, June 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1525/6561444.

Simpson:2024:FAT

[SDW24] Andrew Simpson, Matthias Dellago, and Daniel Woods. Formalizing attack trees to support economic analysis. The Computer Journal, 67(1):220–235, January 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/220/6855081.

Shyla:2023:GDU

[SE23] N. S. Jeya Shyla and W. R. Sam Emmanuel. Glaucoma detection using multiple feature set with recurrent neural network. The Computer Journal, 66(10):2426–2436, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/10/2426/6648787.

Sun:2023:RBC

[SFC^{+}23] Xueli Sun, Jianxi Fan, Baolei Cheng, Jingya Zhou, and Yan Wang. Relationship between component connectivity and component diagnosability of some regular networks. The Computer Journal, 66(8):2033–2042, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/2033/6585062.

Sun:2023:ASK

[SG23] Keshuo Sun and Haiying Gao. Adaptively secure KP-ABE for circuits with fan-in n and fan-out 1. The Computer Journal, 66(10):2554–2573, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2554/6649279.

Sasikala:2021:PFC

[SGGM21] S. Sasikala, S. Gomathi, V. Geetha, and L. Murali. A proposed framework for cloud-aware multimodal multimedia big data analysis toward optimal resource allocation. The Computer Journal, 64(6): 880–894, June 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/880/6132478.

Shaswat:2021:HBD

[Sha21] Kumar Shaswat. Hybridbased deep belief network model for cement compressive strength prediction. The Computer Journal, 64(6):909–920, June 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/6/909/6134262.

Sharmin:2022:EAE

[Sha22] Sadia Sharmin. An experimental approach to exact and random Booleanwidths and their comparison with other width parameters. The Computer Journal, 65(9):2392–2399, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2392/6300205.

Shang:2024:RVC

[Sha24] Yilun Shang. Resilient vector consensus over random dynamic networks under mobile malicious attacks. The Computer Journal, 67(3):1076–1086, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1076/7135860.

S:2024:CDC

[SHJA24] Navin K S, Khanna Nehemiah H, Nancy Y. Jane, and Kannan Arputharaj. Clinical dataset classification using feature ranking and satin bower bird optimized SVMs. The Computer Journal, 67(5):1993– 2006, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1993/7477781.

Sun:2022:TDF

[SHL22] Qiong Sun, Xiankai Huang, and Zheng Liu. Tourists’ digital footprint: Prediction method of tourism consumption decision preference. The Computer Journal, 65(6):1631–1638, June 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1631/6513723.

Shumow:2023:IGR

[Shu23] Daniel Shumow. Incorrectly generated RSA keys: How I learned to stop worrying and recover lost plaintexts. The Computer Journal, 66(6):1342–1349, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1342/6995423.

Salmanian:2022:ASR

[SII22] Zolfaghar Salmanian, Habib Izadkhah, and Ayaz Isazadeh. Auto-scale resource provisioning in IaaS clouds. The Computer Journal, 65(2):297–309, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/297/5836052.

Shen:2020:DBS

[SIT20] Bojie Shen, Saiful Islam, and David Taniar. Direction-based spatial skyline for retrieving arbitraryshaped surrounding objects. The Computer Journal, 63(11):1668–1688, November 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1668/5625928.

Sugave:2020:MEM

[SJ20] Shounak Sugave and Balaso Jagdale. Monarch-EWA: Monarch-earthworm-based secure routing protocol in IoT. The Computer Journal, 63(6):817–831, June 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/817/5645634.

Selvi:2022:LNF

[SJ22] M. Tamil Selvi and B. Jaison. Lemuria: a novel future crop prediction algorithm using data mining. The Computer Journal, 65(3):655–666, March 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/655/5877531.

Shi:2023:LAS

[SJ23] Zhen Shi and Chenhui Jin.

Linear attacks on SNOW 3G and SNOW-V using automatic search. The Computer Journal, 66(5):1268– 1278, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1268/6540627.

Smitha:2023:FAN

[SJC23] J. C. Smitha, Ambily Jane, and Lekshmi Chandran. 3D flattering amplified neural network-based segmentation of amygdala and hippocampus. The Computer Journal, 66(8):1949–1964, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1949/6585622.

Shin:2021:VBP

[SJHL21] Ji Sun Shin, Minjae Jo, Jung Yeon Hwang, and Jaehwan Lee. A verifier-based password-authenticated key exchange using tamperproof hardware. The Computer Journal, 64(8):1293– 1302, August 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/8/1293/6064819.

Su:2020:RTT

[SJL20] Na Su, Shujuan Ji, and Jimin Liu. Real-time topic detection with dynamic windows. The Computer Journal, 63(3):469– 478, March 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/469/5510726.

Sumesh:2022:SAC

[SK22] Sreenithya Sumesh and Aneesh Krishna. Sensitivity analysis of conflicting goals in the i∗ goal model. The Computer Journal, 65(6): 1434–1460, June 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/6/1434/6132358.

Sandhya:2024:LWD

[SK24a] Sandhya and Abhishek Kashyap. A light weight depthwise separable layer optimized CNN architecture for object-based forgery detection in surveillance videos. The Computer Journal, 67(6):2270– 2285, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2270/7606362.

Subramaniam:2024:HOE

[SK24b] Erana Veerappa Dinesh Subramaniam and Valarmathi Krishnasamy. Hybrid optimal ensemble SVM forest classifier for task offloading in mobile cloud computing. The Computer Journal, 67(4):1286–1297, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1286/7192131.

Safaei:2025:EAV

[SK25] Farshad Safaei and Mohammad Mahdi Emadi Kouchak. Error and attack vulnerability of Apollonian networks. The Computer Journal, 68(3):283– 305, March 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/283/7839395.

Selmi:2023:SBH

[SKA23] Wided Selmi, Hager Kammoun, and Ikram Amous. Semantic-based hybrid query reformulation for biomedical information retrieval. The Computer Journal, 66(9):2296–2316, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/9/2296/6627272.

Soykan:2022:CAA

[SKB^{+}22] Elif Ustundag Soykan, Leyli KaraÇay, Zeki Bilgin, Emrah Tomur, Mehmet Akif Ersoy, Ferhat Karakoç, and Pinar Çomak. Contextaware authentication with dynamic credentials using electricity consumption data. The Computer Journal, 65(10):2631–2640, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/10/2631/6316087.

Shibata:2025:PGM

[SKE^{+}25] Masahiro Shibata, Naoki Kitamura, Ryota Eguchi, Yuichi Sudo, Junya Nakamura, and Yonghwan Kim. Partial gathering of mobile agents in dynamic tori. The Computer Journal, 68(10):1485–1501, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1485/8127192.

Sun:2020:NCB

[SLC^{+}20] Jianqiu Sun, Xingguang Li, Kang Chen, Wei Cui, and Ming Chu. A novel CMA + DD LMS blind equalization algorithm for underwater acoustic communication. The Computer Journal, 63(6):974– 981, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/974/5821573.

Shi:2024:DAC

[SLL24] Jiali Shi, Chao Li, and Guoqiang Liu. Differential attack with constants on μ^{2} block cipher. The Computer Journal, 67(1):195– 209, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/195/6855082.

Shan:2025:AUS

[SLL25] Lingxiao Shan, Jian Li, and Guanjun Liu. Alternatingupdate-strategy based Graph Autoencoder for graph neural network. The Computer Journal, 68(7):830– 838, July 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/7/830/8001803.

Sengupta:2022:PBE

[SLT^{+}22] Binanda Sengupta, Yingjiu Li, Yangguang Tian, Robert H. Deng, and Zheng Yang. Policy-based editing-enabled signatures: Authenticating fine-grained and restricted data modification. The Computer Journal, 65(10):2570–2588, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/10/2570/6318170.

Shou:2024:IDM

[SLW^{+}24a] Dingyu Shou, Chao Li, Zhen Wang, Song Cheng, Xiaobo Hu, Kai Zhang, Mi Wen, and Yong Wang. An intrusion detection method based on attention mechanism to improve CNN-BiLSTM model. The Computer Journal, 67(5): 1851–1865, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1851/7335815.

Si:2024:CZK

[SLW24b] Shumin Si, Xiuhan Lin, and Puwen Wei. Compressed zero-knowledge proofs for lattice-based accumulator. The Computer Journal, 67(2):694–708, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/694/7075452.

Sun:2025:ACA

[SLWL25] Hongliang Sun, Zhihui Liu, Chao Wang, and Dan Li. Admission control algorithm for visible light communication random access network under delay and jitter constraints. The Computer Journal, 68(5): 479–486, May 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/479/7918701.

Sun:2024:EST

[SLY^{+}24] Daoqiang Sun, Hongbo Liu, Yu Yang, Long Li, Heng Zhang, and Asfand Fahad. Enumeration of subtrees of two families of self-similar networks based on novel two-forest dual transformations. The Computer Journal, 67(5):1652–1662, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1652/7280726.

Sabir:2021:SFT

[SM21] Eminjan Sabir and Jixiang Meng. Structure fault tolerance of recursive interconnection networks. The Computer Journal, 64(1): 64–75, January 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/1/64/5716153.

Shobana:2023:ISA

[SM23] J. Shobana and M. Murali. An improved self attention mechanism based on optimized BERT-BiLSTM model for accurate polarity prediction. The Computer Journal, 66(5):1279– 1294, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1279/6555411.

Suresh:2023:FEE

[SMK23] K. Suresh, S. S. Sreeja Mole, and A. Joseph Selva Kumar. F^{2}SO: an energy efficient cluster based routing protocol using fuzzy firebug swarm optimization algorithm in WSN. The Computer Journal, 66(5):1126– 1138, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1126/6527755.

S:2020:MOD

[SP20a] Oswalt Manoj S. and Ananth J. P. MapReduce and optimized deep network for rainfall prediction in agriculture. The Computer Journal, 63(6):900– 912, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/900/5721350.

Sankpal:2020:RRA

[SP20b] Lata Jaywant Sankpal and Suhas H. Patil. Rider-rank algorithm-based feature extraction for re-ranking the Webpages in the search engine. The Computer Journal, 63(10):1479–1489, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/10/1479/5855737.

Sekar:2023:NSU

[SP23] Aravindkumar Sekar and Varalakshmi Perumal. A novel SGD-U-network-based pixel-level road crack segmentation and classification. The Computer Journal, 66(7):1595–1608, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1595/6561441.

Shyamasundar:2020:MLM

[SR20] L. B. Shyamasundar and P. Jhansi Rani. A multiplelayer machine learning architecture for improved accuracy in sentiment analysis. The Computer Journal, 63(3):395–409, March 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/3/395/5480914.

Sudhakar:2023:DIA

[SR23] R. Sudhakar and P. V. Venkateswara Rao. Digital image anti-forensic model using exponential chaotic biogeography-based optimization algorithm. The Computer Journal, 66 (12):3038–3051, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/3038/6777261.

S:2020:TSP

[SRC20] Thanga Revathi S., N. Ramaraj, and S. Chithra. Tracy–Singh product and genetic whale optimization algorithm for retrievable data perturbation for privacy preserved data publishing in cloud computing. The Computer Journal, 63(2):239–253, February 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/2/239/5618959.

Salehi:2020:NDM

[SRL20] Mohsen Salehi, Jafar Razmara, and Shahriar Lotfi. A novel data mining on breast cancer survivability using MLP ensemble learners. The Computer Journal, 63(3):435– 447, March 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/3/435/5498230.

Salehi:2022:ODP

[SRLM22] Mohsen Salehi, Jafar Razmara, Shahriar Lotfi, and Farnaz Mahan. A onedimensional probabilistic convolutional neural network for prediction of breast cancer survivability. The Computer Journal, 65(10):2641–2653, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2641/6326762.

Shukla:2023:AEM

[SRS^{+}23] Piyush Kumar Shukla, Vandana Roy, Prashant Kumar Shukla, Anoop Kumar Chaturvedi, Aumreesh Kumar Saxena, Manish Maheshwari, and Parashu Ram Pal. An advanced EEG motion artifacts eradication algorithm. The Computer Journal, 66(2):429– 440, February 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/2/429/6469155.

Srinivas:2020:OBS

[SS20] Vasamsetti Srinivas and Ch Santhirani. Optimizationbased support vector neural network for speaker recognition. The Computer Journal, 63(1):151–167, January 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/151/5380618.

S:2022:BAL

[SS22a] Yoganand S. and Chithra S. Bat ant lion optimizationbased generative adversarial network for structural health monitoring in IoT. The Computer Journal, 65(9):2439–2453, September 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2439/6294164.

Sahil:2022:FAE

[SS22b] Sahil and Sandeep Kumar Sood. Fog-assisted energy efficient cyber physical system for panic-based evacuation during disasters. The Computer Journal, 65(6): 1540–1559, June 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1540/6249696.

Seddigh:2022:FAS

[SS22c] Milad Seddigh and Hadi Soleimany. Flush+Reload attacks on SEED. The Computer Journal, 65 (10):2769–2779, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2769/6341631.

Shoba:2022:AFR

[SS22d] Betcy Thanga Shoba and I. Shatheesh Sam. Aging facial recognition for feature extraction using adaptive fully recurrent deep neural learning. The Computer Journal, 65(7):1923– 1936, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1923/6518262.

Shaji:2024:ECO

[SS24a] C. Shaji and I. Shatheesh Sam. Encoding with combination orientation technique for RDH in dual stego images. The Computer Journal, 67(1):347– 361, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/347/6965539.

Shi:2024:DYR

[SS24b] Xiaolong Shi and Anjun Song. Defog YOLO for road object detection in foggy weather. The Computer Journal, 67(11):3115–3127, November 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/11/3115/7730391.

Sharma:2025:TSH

[SS25] Nikhil Sharma and Prashant Giridhar Shambharkar. Towards secure healthcare: SA-GBO-ODBN model utilizing blockchain and deep learning for data handling and diagnosis. The Computer Journal, 68(10):1386– 1423, October 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1386/8123022.

Sureshkumar:2023:NRP

[SSB23] Veluguri Sureshkumar, Rajasomashekar Somarajadikshitar, and B. Sarala Beeram. A novel representation and prediction initiative for underground water by using deep learning technique of remote sensing images. The Computer Journal, 66(7):1784– 1801, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1784/6649482.

Sood:2022:FCA

[SSMS22] Sandeep Kumar Sood, Vaishali Sood, Isha Mahajan, and Sahil. Fog–cloud assisted IoT-based hierarchical approach for controlling dengue infection. The Computer Journal, 65(1): 67–79, January 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/67/5817593.

Sarkar:2025:ICD

[SSN25] Nayan Kumar Sarkar, Moirangthem Marjit Singh, and Utpal Nandi. Image classification with deconvolution operation and augmentation. The Computer Journal, 68(2):135– 144, February 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/135/7841855.

Shamim:2022:ADO

[SSS^{+}22] Mohammed Zubair M.

Shamim, Sadatullah Syed, Mohammad Shiblee, Mohammed Usman, Syed Jaffar Ali qnd Hany S. Hussein, and Mohammed Farrag. Automated detection of oral pre-cancerous tongue lesions using deep learning for early diagnosis of oral cavity cancer. The Computer Journal, 65(1): 91–104, January 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/1/91/5985297.

Simpson:2023:SMM

[SST23] Michael Simpson, Venkatesh Srinivasan, and Alex Thomo. Scalable misinformation mitigation in social networks using reverse sampling. The Computer Journal, 66(9):2230–2253, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2230/6631439.

Song:2025:SGM

[SSY^{+}25] Qing Song, Wenfeng Sun, Donghan Yang, Mengjie Hu, and Chun Liu. Semantic Guided Matting Net. The Computer Journal, 68(3):322–330, March 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/322/7874330.

Suchenek:2024:WCA

[Suc24] Marek A. Suchenek. Worstcase analysis of Heapsort, exactly. The Computer Journal, 67(3):812– 824, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/812/7606364.

S:2022:MFD

[SV22a] Arul Antran Vijay S. and Jothi Prakash V. A modified firefly deep ensemble for microarray data classification. The Computer Journal, 65(12):3265–3274, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3265/6763328.

Singh:2022:RBM

[SV22b] Ravindra Kumar Singh and Harsh Kumar Verma. Redis-based messaging queue and cache-enabled parallel processing social media analytics framework. The Computer Journal, 65(4): 843–857, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/843/5956494.

Sultan:2024:NRA

[SVD^{+}24] Nazatul Haque Sultan, Vijay Varadharajan, Saurab Dulal, Seyit Camtepe, and Surya Nepal. NDN-RBE: an accountable privacy aware access control framework for NDN. The Computer Journal, 67(4):1572– 1589, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1572/7241751.

Shu:2024:SSF

[SWFW24] Chang Shu, Yan Wang, Jianxi Fan, and Guijuan Wang. Super structure fault-tolerance assessment of the generalized hypercube. The Computer Journal, 67(4):1457–1466, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1457/7233670.

Shi:2020:MCH

[SWJkL20] Wenjun Shi, Jigang Wu, Guiyuan Jiang, and Siew kei Lam. Multiple-choice hardware/software partitioning for tree task-graph on MPSoC. The Computer Journal, 63(5):688– 700, May 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/5/688/5363890.

Shen:2022:ICD

[SWW^{+}22] Yanzhao Shen, Ting Wu, Gaoli Wang, Xinfeng Dong, and Haifeng Qian. Improved collision detection of MD5 using sufficient condition combination. The Computer Journal, 65(10):2720–2729, October 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2720/6334047.

Shi:2025:MFR

[SWZW25] Jiadong Shi, Wenjin Wang, Qiang Zhou, and Liangmin Wang. Multiscale fingerprinting for robust website fingerprinting attack. The Computer Journal, 68(10):1436–1449, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1436/8123021.

Sun:2023:GDN

[SYC^{+}23] GuoYing Sun, Feng Ye, Tingting Chai, Zhaoxin Zhang, Xiaojun Tong, and Shitala Prasad. Gambling domain name recognition via certificate and textual analysis. The Computer Journal, 66(8):1829–1839, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/8/1829/6570684.

Shamsizadeh:2024:IFM

[SZA24] Marzieh Shamsizadeh, Mohammad Mehdi Zahedi, and Khadijeh Abolpour. Irreducible fuzzy multiset finite automaton. The Computer Journal, 67(2):519– 526, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/519/6995425.

Shi:2025:PCC

[SZCY25] Wenlong Shi, Jing Zhang, Xiaoping Chen, and Xiucai Ye. PCDP-CRLPPM: a classified regional location privacy-protection model based on personalized clustering with differential privacy in data management. The Computer Journal, 68(4):372– 396, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/372/7896302.

Song:2024:ACZ

[SZH^{+}24] Yongcheng Song, Jiang Zhang, Xinyi Huang, Wei Wu, and Haixia Chen. Analysis and construction of zero-knowledge proofs for the MinRank problem. The Computer Journal, 67(3):1060–1075, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1060/7136738.

Sun:2020:IFD

[SZL^{+}20] Xueli Sun, Shuming Zhou, Mengjie Lv, Jiafei Liu, and Guanqin Lian. Intermittent fault diagnosability of some general regular networks. The Computer Journal, 63(1):16–24, January 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/1/16/5255731.

Sun:2022:ITP

[SZL22] Xiaoying Sun, Chen Zhang, and Guohong Liu. Improved tactile perception of 3D geometric bumps using coupled electrovibration and mechanical vibration stimuli. The Computer Journal, 65(3):621– 630, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/621/5880725.

Sarbizhan:2023:GAS

[SZS23] Elham Raisi Sarbizhan, Mohammad Mehdi Zahedi, and Marzieh Shamsizadeh. L-graph automata and some applications. The Computer Journal, 66(7): 1698–1716, July 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1698/6572655.

Su:2020:SOA

[SZX20] Qianqian Su, Rui Zhang, and Rui Xue. Secure outsourcing algorithms for composite modular exponentiation based on single untrusted cloud. The Computer Journal, 63(8):1271, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1271/5823571.

Shang:2025:CLD

[SZY^{+}25] Shang Shang, Xinqiang Zhao, Zhongjiang Yao, Yepeng Yao, Liya Su, Zijing Fan, Xiaodan Zhang, and Zhengwei Jiang. Can LLMs deeply detect complex malicious queries? A framework for jailbreaking via obfuscating intent. The Computer Journal, 68(5):460– 478, May 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/460/7919595.

Sun:2025:PPP

[SZZ^{+}25] Han Sun, Yan Zhang, Huiping Zhuang, Jiatong Li, Zhen Xu, and Liji Wu. PEAR: privacy-preserving and effective aggregation for byzantine-robust federated learning in real-world scenarios. The Computer Journal, 68(9):1087–1104, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1087/8006389.

Tian:2023:RPB

[T^{+}23] Yangguang Tian et al. Revocable policy-based chameleon hash for blockchain rewriting. The Computer Journal, 66(10):2365–2378, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2365/6627275.

Toksari:2023:SSP

[TA23] M. Duran Toksari and Berrin Atalay. Some scheduling problems with job rejection and a learning effect. The Computer Journal, 66(4):866– 872, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/866/6494710.

Tekincan:2023:PEF

[TAA23] Erdoğan Berkay Tekincan, Tülin Erçelebi Ayyıldız, and Nizam Ayyıldız. Performance evaluation of FPGA-based LSTM neural networks for pulse signal detection on real-time radar warning receivers. The Computer Journal, 66(4): 1040–1052, April 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/4/1040/6918760.

See correction [Ano24g].

Teijeiro:2023:IVL

[TADD23] Diego Teijeiro, Margarita Amor, Ramón Doallo, and David Deibe. Interactive visualization of large point clouds using an autotuning multiresolution out-of-core strategy. The Computer Journal, 66(7):1802– 1816, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1802/6927252.

Tang:2024:GML

[Tan24]  Zhong Tang.  A gen-

eric multi-level framework for building term-weighting schemes in text classification. The Computer Journal, 67(11):3042–3055, November 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/11/3042/7720756.

Takrouni:2021:SID

[TBH21] Manel Takrouni, Rim Bouhouch, and Salem Hasnaoui. Simulink implementation of the data distribution service for vehicular controllers on top of GBE and AFDX. The Computer Journal, 64(6):860– 879, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/860/6124919.

Tang:2020:VHD

[TCY^{+}20] Zhenjun Tang, Lv Chen, Heng Yao, Xianquan Zhang, and Chunqiang Yu. Video hashing with DCT and NMF. The Computer Journal, 63(7):1017–1030, July 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/1017/5540176.

Tang:2024:ISD

[TDZ^{+}24] Fang Tang, Nengsheng Du, Zhong Zhengwei, Chunlin Li, and Youlong Luo. An improved sensing data cleaning scheme for object localization in edge computing environment.

The Computer Journal, 67(9):2838–2853, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2838/7713739.

Touati:2022:ATA

[TEGB22] Imen Touati, Mariem Ellouze, Marwa Graja, and Lamia Hadrich Belguith. Appraisal of two Arabic opinion summarization methods: Statistical versus machine learning. The Computer Journal, 65(2):192–202, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/192/5819213.

Trabay:2022:TFU

[TEHG22] Doaa Wagdy Trabay, Ibrahim El-Henawy, and Wajeb Gharibi. A trust framework utilization in cloud computing environment based on multi-criteria decisionmaking methods. The Computer Journal, 65(4): 997–1005, April 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/997/6090334.

Troudi:2021:MMD

[TGZ^{+}21] Abir Troudi, Leila Ghorbel, Corinne Amel Zayani, Salma Jamoussi, and Ikram Amous. MDER: Multidimensional event recommendation in social media context. The Computer Journal, 64(3):369– 382, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/369/5934818.

Truong:2023:NAC

[TH23] Hieu Minh Truong and Hieu Trung Huynh. A novel approach for CTbased COVID-19 classification and lesion segmentation based on deep learning. The Computer Journal, 66(6):1366–1375, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1366/6542227.

Thimbleby:2024:ISU

[Thi24] Harold Thimbleby. Improving science that uses code. The Computer Journal, 67(4):1381–1404, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1381/7235536.

Tian:2023:CCC

[TJB23] Junfeng Tian, Haoyi Jia, and Wenqing Bai. CCESHP: Causal consistency model of edge storage based on hash ring and partial georeplication. The Computer Journal, 66(12):2874–2886, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/2874/6702393.

Thethi:2022:DFS

[TK22] Sukhmani K. Thethi and Ravi Kumar. Dynamic frequency scaling of a singlecore processor using machine learning paradigms. The Computer Journal, 65 (3):631–654, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/65/3/631/
5880100.

Tomar:2023:PRB

[TKV23] Parul Tomar, Gyanendra Kumar, and Lal Pratap Verma. Path-rank-based data chunk scheduling for concurrent multipath transmission. The Computer Journal, 66(9):2254–2264, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2254/6611471.

Tian:2020:NCL

[TLD^{+}20] Yangguang Tian, Yingjiu Li, Robert H. Deng, Nan Li, Guomin Yang, and Zheng Yang. A new construction for linkable secret handshake. The Computer Journal, 63(4):536– 548, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/536/5612724.

Tian:2024:PBR

[TLD^{+}24] Yangguang Tian, Yingjiu Li, Robert H. Deng, Guomin Yang, and Nan Li. Policy-based remote user authentication from multibiometrics. The Computer Journal, 67(5):1814–1825, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1814/7323308.

Tian:2021:URS

[TLMY21] Yangguang Tian, Yingjiu Li, Yi Mu, and Guomin Yang. Unlinkable and revocable secret handshake. The Computer Journal, 64(8):1303–1314, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1303/6095852.

Taheri:2024:CIS

[TR24] Fatemeh Taheri and Kambiz Rahbar. Coverless image steganography using content-based image patch retrieval. The Computer Journal, 67(9):2799–2808, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2799/7681802.

Tiryaki:2024:BCM

[TT24] Volkan Müjdat Tiryaki and Nedim Tutkun. Breast cancer mass classification using machine learning, binarycoded genetic algorithms and an ensemble of deep transfer learning. The Computer Journal, 67(3):1111– 1125, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1111/7140287.

Tenorio-Trigoso:2023:ACR

[TTCCMR^{+}23] Alonso Tenorio-Trigoso, Manuel Castillo-Cara, Giovanny Mondragón-Ruiz, Carmen Carrión, and Blanca Caminero. An analysis of computational resources of event-driven streaming data flow for Internet of things: a case study. The Computer Journal, 66(1): 47–60, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/47/6381858.

Tripathy:2021:NAR

[TTPD21] A. K. Tripathy, S. K. Tripathy, S. R. Pattanaik, and S. K. Das. A new algorithm for reconstruction of a computer-generated hologram (CGH). The Computer Journal, 64(2):245– 253, February 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/2/245/6029221.

Tvardovskii:2023:DHS

[TY23] Aleksandr Tvardovskii and Nina Yevtushenko. Deriving homing sequences for finite state machines with timeouts. The Computer Journal, 66(9):2181–2190, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2181/6608475.

Tehrani:2024:WSB

[TYTZSE24] Amir Tehrani, Meisam Yadollahzadeh-Tabari, Aidin Zehtab-Salmasi, and Rasul Enayatifar. Wearable sensor-based human activity recognition system employing Bi-LSTM algorithm. The Computer Journal, 67(3):961– 975, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/961/7113990.

Tang:2021:RIH

[TYY^{+}21] Zhenjun Tang, Mengzhu Yu, Heng Yao, Hanyun Zhang, Chunqiang Yu, and Xianquan Zhang. Robust image hashing with singular values of quaternion SVD. The Computer Journal, 64(11):1656–1671, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1656/5670505.

Tang:2023:TPS

[TZ23] Guofeng Tang and Zhenfeng Zhang. Two-party signing for ISO/IEC Digital Signature Standards. The Computer Journal, 66(5): 1111–1125, May 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1111/6531895.

Tian:2023:DAD

[TZS^{+}23] Yuchen Tian, Weizhe Zhang, Andrew Simpson, Yang Liu, and Zoe Lin Jiang. Defending against data poisoning attacks: From distributed learning to federated learning. The Computer Journal, 66(3):711– 726, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/711/6469154.

Tian:2024:KNN

[TZW^{+}24] Gang Tian, Guangxin Zhao, Rui Wang, Jiachang Wang, and Cheng He. Knearest neighbor smart contract classification with semantic feature enhancement. The Computer Journal, 67(11):3067–3079, November 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/11/3067/7721728.

Tian:2024:TTP

[TZwSX24] Junfeng Tian, Qi Zhu, Jia wei Shen, and Tengfei Xu. TPRT: a trajectory publishing scheme for the Internet of Vehicles based on radix tree. The Computer Journal, 67(10):2920–2933, October 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2920/7708711.

ulHaq:2022:ART

[uHLH22] Muhammad Inaam ul Haq, Qianmu Li, and Jun Hou. Analyzing the research trends of IoT using topic modeling. The Computer Journal, 65(10):2589–2609, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/10/2589/6318691.

ulhaq:2025:BBD

[uhLM^{+}25] Muhammad Inaam ul haq, Qianmu Li, Khalid Mahmood, Ayesha Shafique, and Rizwan Ullah. BioElectra-BiLSTM-Dual Attention classifier for optimizing multilabel scientific literature classification. The Computer Journal, 68(5): 565–576, May 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/565/7924286.

Venkanna:2021:OTD

[VB21] Gugulothu Venkanna and K. F. Bharati. Optimal text document clustering enabled by weighed similarity oriented Jaya with grey wolf optimization algorithm. The Computer Journal, 64(6):960– 972, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/960/6259660.

Vanamoorthy:2023:CCF

[VC23] Muthumanikandan Vanamoorthy and Valliyammai Chinnaiah. Corrigendum to: Congestion Free Transient Plane (CFTP) using bandwidth sharing during link failures in SDN. The Computer Journal, 66(1):267, January 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/267/5866758.

See [VD20].

Vanamoorthy:2020:CFT

[VD20] Muthumanikandan Vanamoorthy and Valliyammai Devendran. Congestion-Free Transient Plane (CFTP) using bandwidth sharing during link failures in SDN. The Computer Journal, 63(6):832–843, June 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/6/832/5627746.

See corrigendum [VC23].

Verma:2023:RAC

[VG23] Anjushi Verma and Tirthankar Gayen. Reliability assessment of combined hardware– software non-repairable time-critical systems. The Computer Journal, 66(7): 1644–1663, July 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1644/6565999.

Veluchamy:2021:DLA

[VKM21] S. Veluchamy, L. R. Karlmarx, and K. Michael Mahesh. Detection and localization of abnormalities in surveillance video using TimeRider-based neural network. The Computer Journal, 64(12):1886–1906, December 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/12/1886/6178419.

Vinolin:2021:HCR

[VS21] V. Vinolin and M. Sucharitha. Hierarchical categorization and review of recent techniques on image forgery detection. The Computer Journal, 64(11):1692–1704, November 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/11/1692/5681102.

Visuwalingam:2024:DLM

[VSAR24] Hemakasiny Visuwalingam, Ratnasingam Sakuntharaj, Janaka Alawatugoda, and Roshan Ragel. Deep learning model for Tamil part-of-speech tagging. The Computer Journal, 67(8):2633– 2642, August 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2633/7641754.

Vijayakarthick:2022:MPA

[VSS^{+}22] M. Vijayakarthick, E. Sivaraman, S. Sathishbabu, N. Vinoth, and S. N. Sivaraj. Modelling and predictive analytics of COVID-19 transmission using Gustafson– Kessel fuzzy clustering approach. The Computer Journal, 65(12):3240–3249, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3240/6554132.

Voudouris:2025:GPC

[VTZX25] Anastassis Voudouris, Aristomenis Tressos, Apostolis Zarras, and Christos Xenakis. Game on: a performance comparison of interpolation techniques applied to Shamir’s secret sharing. The Computer Journal, 68(3):261– 272, March 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/261/7900561.

Wang:2021:PCM

[Wan21] Ming-Hung Wang. Positioning and categorizing mass media using reaction emojis on Facebook. The Computer Journal, 64(3): 451–461, March 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/451/6050691.

Wang:2025:EED

[Wan25] Haitao Wang. Energy efficient data transmission scheme integrating clustering and compressive sensing in wireless sensor networks. The Computer Journal, 68(10):1345–1354, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1345/8121018.

Wani:2021:GAC

[WBG21] Zahid Hussain Wani, Javaid Iqbal Bhat, and Kaisar Javeed Giri. A generic analogycentered software cost estimation based on differential evolution exploration process. The Computer Journal, 64(3):462– 472, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/3/462/6124918.

Wu:2022:AAN

[WCC^{+}22] Jiawei Wu, Huangfei Cheng, Bin Cao, Jiaxing Wang, and Jing Fan. ASCUE: an adversarial networkbased semantical conformance checking method for unsupervised event extraction in social Internet of Things. The Computer Journal, 65(11):2939–2952, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2939/6652556.

Wang:2021:IFI

[WCD21] Gaoli Wang, Zhenfu Cao, and Xiaolei Dong. Improved file-injection attacks on searchable encryption using finite set theory. The Computer Journal, 64(8):1264–1276, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1264/6048928.

Wu:2024:CCB

[WCD^{+}24] Tongshuai Wu, Liwei Chen, Gewangzi Du, Chenguang Zhu, Ningning Cui, and Gang Shi. CDNM: Clustering-based data normalization method for automated vulnerability detection. The Computer Journal, 67(4):1538–1549, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1538/7235534.

Wang:2022:ACC

[WCF^{+}22] Yifeng Wang, Baolei Cheng, Jianxi Fan, Yu Qian, and Ruofan Jiang. An algorithm to construct completely independent spanning trees in line graphs. The Computer Journal, 65(12):2979–2990, December 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/2979/6369599.

Wang:2024:IDM

[WCLY24] Haizhen Wang, Zhiqing Cui, Zuozheng Lian, and Jinying Yan. An improved DNN model for WLAN intrusion detection. The Computer Journal, 67(9):2854–2863, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2854/7695665.

Wang:2024:SMC

[WCW^{+}24] Yihong Wang, Baolei Cheng, Yan Wang, Jia Yu, and Jianxi Fan. Strongly Menger connectedness of a class of recursive networks. The Computer Journal, 67(6):2030–2038, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2030/7477779.

Wu:2023:SHL

[WDC^{+}23] Shusen Wu, Xiaoshe Dong, Heng Chen, Longxiang Wang, Qiang Wang, and Zhengdong Zhu. Simplified high level parallelism expression on heterogeneous systems through data partition pattern description. The Computer Journal, 66(6):1400–1418, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1400/6548383.

Wei:2023:DAP

[WDM^{+}23] Congming Wei, Xiaoyang Dong, Willi Meier, Lingyue Qin, and Ximing Fu. Differential-aided preimage attacks on round-reduced Keccak. The Computer Journal, 66(12):3069–3091, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3069/6814392.

Wang:2021:ICK

[WG21] Ai Wang and Xuedong Gao. Intelligent computing: Knowledge acquisition method based on the management scale transformation. The Computer Journal, 64(3):314– 324, March 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/3/314/5879955.

Wen:2025:IRF

[WG25] Xiaoyang Wen and Shaopeng Guan. Improving recommendation fairness with dependency-based graph collaborative filtering. The Computer Journal, 68(7): 839–858, July 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/839/8003476.

Wang:2022:SEB

[WGL^{+}22] Huan Wang, Qing Gao, Hao Li, Hao Wang, Liping Yan, and Guanghua Liu. A structural evolution-based anomaly detection method for generalized evolving social networks. The Computer Journal, 65(5):1189– 1199, May 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1189/6055137.

Wambura:2022:RAD

[WHL22] Stephen Wambura, Jianbin Huang, and He Li. Robust anomaly detection in feature-evolving time series. The Computer Journal, 65(5):1242–1256, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1242/6064823.

Wei:2023:HFD

[WhLY23] Yulong Wei, Rong hua Li, and Weihua Yang. Hybrid fault diagnosis capability analysis of highly connected graphs. The Computer Journal, 66(1):221– 228, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/221/6384774.

Wild:2022:CWC

[Wil22] Marcel Wild. Compression with wildcards: From CNFs to orthogonal DNFs by imposing the clauses one-by-one. The Computer Journal, 65(5):1073–1087, May 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1073/6012400.

Wang:2025:NDA

[WJHL25] Haizhen Wang, Na Jia, Yang He, and Zuozheng Lian. A new DDoS attack detection model based on improved stacked autoencoder and gated recurrent unit for software defined network. The Computer Journal, 68(8):1028–1049, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/1028/8069014.

Walsh:2024:NRT

[WKW^{+}24] James Walsh, Oluwafunmilola Kesa, Andrew Wang, Mihai Ilas, Patrick O’Hara, Oscar Giles, Neil Dhir, Mark Girolami, and Theodoros Damoulas. Near real-time social distance estimation in London. The Computer Journal, 67(1): 95–109, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/95/7071574.

Wang:2020:RIN

[WL20] Hao Wang and Kaiju Li. Resistance of IID noise in differentially private schemes for trajectory publishing. The Computer Journal, 63(4):549– 566, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/549/5625061.

Wang:2023:GCF

[WL23] Jing Wang and Fangmin Li. The generalized 3connectivity of the folded hypercube FQ_{n}. The Computer Journal, 66(12):2921– 2927, December 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/2921/6751165.

Wang:2023:GLS

[WLLM23] Shichang Wang, Meicheng Liu, Dongdai Lin, and Li Ma. On grain-like small state stream ciphers against fast correlation attacks: Cryptanalysis of Plantlet, Fruit-v2 and Fruit-80. The Computer Journal, 66(6): 1376–1399, June 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1376/6549846.

Wang:2025:DTO

[WLTY25] Cuiping Wang, Xiaole Li, Jinwei Tian, and Yilong Yin. Data transmission optimization based on multiobjective deep reinforcement learning. The Computer Journal, 68(2):201– 215, February 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/201/7828054.

Wang:2024:SAS

[WLW24] Xin Wang, Li Lin, and Yao Wang. Stealth address schemes with fast retrievability based on subgroup membership assumptions related to factoring. The Computer Journal, 67(4): 1253–1264, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1253/7188840.

Wei:2024:BTB

[WLY24] Chuancheng Wei, Gang Liang, and Kexiang Yan. BotGSL: Twitter bot detection with graph structure learning. The Computer Journal, 67(7):2486– 2497, July 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/7/2486/7618086.

Wang:2020:GGA

[WLZ20] Xiao-Yu Wang, Yu-Feng Liu, and Kang Zhang. A graph grammar approach to the design and validation of floor plans. The Computer Journal, 63(1):137– 150, January 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/137/5316176.

Wang:2022:DDC

[WLZ22] Weidong Wang, Xiaofeng Li, and He Zhao. DCAF: Dynamic cross-chain anchoring framework using smart contracts. The Computer Journal, 65(8):2164– 2182, August 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2164/6272097.

Wang:2025:DPT

[WLZ^{+}25] Zhen Wang, Jinfeng Lv, Yuhao Zhou, Jianhui Jiang, and Yong Wang. A dual protection technology to thwart hardware Trojan insertion based on observability. The Computer Journal, 68(8):1074–1085, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/1074/8154527.

Wang:2022:MPE

[WM22] Shiying Wang and Xiaolei Ma. The matching preclusion of enhanced hypercubes. The Computer Journal, 65(7):1874–1890, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1874/6259656.

Wang:2025:ACS

[WPF^{+}25] Yongshan Wang, Siyuan Pang, Zijing Fan, Shang Shang, Yepeng Yao, Zhengwei Jiang, and Baoxu Liu. Advanced code slicing with pre-trained model fine-tuned for open-source component malware detection. The Computer Journal, 68(9):1163–1180, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1163/8209415.

Wang:2023:ABS

[WPY^{+}23] Xiaodi Wang, Xiaoge Pan, Tian Yang, Jianhua Xie, and Mingwei Tang. Aspectbased sentiment analysis using interaction matrix and global attention neural network. The Computer Journal, 66(5):1167– 1183, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1167/6543490.

Wang:2022:CMB

[WQZ^{+}22] Lei Wang, Qing Qian, Qiang Zhang, Jishuai Wang, Wenbo Cheng, and Wei Yan. Classification model on big data in medical diagnosis based on semi-supervised learning. The Computer Journal, 65(2):177–191, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/177/5808795.

W:2022:SAR

[WSA22] Thamba Meshach W, Hemajothi S, and Mary Anita E A. Smart affect recognition system for real-time biometric surveillance using hybrid features and multilayered binary structured support vector machine. The Computer Journal, 65(4):897–917, April 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/4/897/5930867.

Wang:2022:EIP

[WSCL22] Qinglong Wang, Xiaoying Sun, Dekun Cao, and Guohong Liu. Enhanced interactive performance of zoom–in/out gestures using electrostatic tactile feedback on touchscreens. The Computer Journal, 65(2):261–274, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/261/5854799.

Wang:2025:CIN

[WTWZ25] Cenman Wang, Ye Tian, Yiwen Wu, and Xinming Zhang. Confluence: improving network monitoring accuracy on multi-pipeline data plane. The Computer Journal, 68(9):1315–1328, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/9/1315/8113252.

Wang:2020:EUP

[WtZLS20] Yan Wang, Jian tao Zhou, Xinyuan Li, and Xiaoyu Song. Effective user preference clustering in Web service applications. The Computer Journal, 63(11):1633–1643, November 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/11/1633/5603743.

Wang:2022:DPP

[WW22] Hao Wang and Huan Wang. Differentially private publication for correlated nonnumerical data. The Computer Journal, 65(7):1726– 1739, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/7/1726/6236085.

Wang:2024:CAE

[WWSW24] Zhuowei Wang, Hao Wang, Xiaoyu Song, and JiaHui Wu. Communication-aware energy consumption model in heterogeneous computing systems. The Computer Journal, 67(1):78–94, January 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/78/6835386.

Wang:2022:LSA

[WWW^{+}22] Yanbo Wang, Fasheng Wang, Chang Wang, Fuming Sun, and Jianjun He. Learning saliency-aware correlation filters for visual tracking. The Computer Journal, 65(7):1846– 1859, July 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1846/6272652.

Wang:2020:ICM

[WWY20a] Chao-Chao Wang, Wan-Liang Wang, and Xin-Wei Yao. Interference and coverage modeling for indoor terahertz communications with beamforming antennas. The Computer Journal, 63(10):1597–1606, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1597/5872130.

Wang:2020:MBC

[WWY^{+}20b] Haijiang Wang, Jingpu Wang, Fuqi Yao, Yongqiang Ma, Lihong Li, and Qinke Yang. Multi-band contourlet transform for adaptive remote sensing image denoising. The Computer Journal, 63(7):1084–1098, July 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/7/1084/5573574.

Wang:2024:NID

[WWZZ24] Bolin Wang, Wenling Wu, Yuhan Zhang, and Li Zhang. New integral distinguishers on permutation of Whirlpool. The Computer Journal, 67(3):899– 906, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/899/7086811.

Wu:2025:RAT

[WX25] Xuemin Wu and Liqiong Xu. Reliability analysis for the n-th Cartesian product of the balanced complete multipartite graph under various hypotheses. The Computer Journal, 68(10):1502–1509, October 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/10/1502/8134341.

Wen:2025:MMA

[WY25] Xin Wen and Xiaohui Yang.

MAPBFT: multilevel adaptive PBFT algorithm based on discourse and reputation models. The Computer Journal, 68(6):635– 648, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/635/7933711.

Wu:2023:EPS

[WYC^{+}23] Yalan Wu, Jiaquan Yan, Long Chen, Jigang Wu, and Yidong Li. Efficient parameter server placement for distributed deep learning in edge computing. The Computer Journal, 66(3):678– 691, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/3/678/6446972.

Wang:2022:MHO

[WYXZ22] Yucheng Wang, Yuhao Yi, Wanyue Xu, and Zhongzhi Zhang. Modeling higherorder interactions in complex networks by edge product of graphs. The Computer Journal, 65(9):2347– 2359, September 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/9/2347/6307490.

Wang:2020:TGN

[WZ20] Shiying Wang and Nan Zhao. The two-good-neighbor connectivity and diagnosability of the augmented three-ary n-cubes. The Computer Journal, 63 (1):1–15, January 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/63/1/1/
                                        5239646.

Wang:2023:RCM

[WZ23] Yanna Wang and Bo Zhou. Residual closeness, matching number and chromatic number. The Computer Journal, 66(5):1156–1166, May 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1156/6527283.

Wang:2024:IMS

[WZ24] Ping Wang and Ruisheng Zhang. Influence maximization in social networks using an improved multiobjective particle swarm optimization. The Computer Journal, 67(6):2137– 2150, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2137/7491975.

Wang:2025:OSS

[WZ25] Jinyong Wang and Ce Zhang.

Optimal selection of software reliability growth model for open-source software using weighted grey relational analysis method. The Computer Journal, 68(6):663–674, June 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/663/7934012.

Wang:2025:FJN

[WZB^{+}25] Yingjie Wang, Chengye Zhang, Fengbo Bai, Zumin Wang, and Jing Qin. FINB: a Japanese named entity recognition model based on multi-feature integration method. The Computer Journal, 68(4):419– 430, April 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/4/419/8002236.

Wen:2021:GLS

[WZC^{+}21] Guoqiu Wen, Yonghua Zhu, Linjun Chen, Mengmeng Zhan, and Yangcai Xie. Global and local structure preservation for nonlinear high-dimensional spectral clustering. The Computer Journal, 64(7):993–1004, July 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/993/6275471.

Wu:2020:GCT

[WZG^{+}20] Ge Wu, Zhen Zhao, Fuchun Guo, Willy Susilo, and Futai Zhang. On the general construction of tightly secure identity-based signature schemes. The Computer Journal, 63(12):1835– 1848, December 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/12/1835/5808792.

Wang:2024:LLS

[WZG^{+}24] Shurui Wang, Yuhe Zhang, Ruizhe Guo, Yaning Zhang, Yifei Xie, and Xinyu Zhou. 2L-LSH: a localitysensitive hash functionbased method for rapid point cloud indexing. The Computer Journal, 67 (9):2809–2818, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2809/7774503.

Wang:2025:TEP

[WZG25] Chuanyin Wang, Yifei Zhang, and Neng Gao. Towards effective privacy preservation in federated learning with automatic gradient clipping and gradient transformation perturbation. The Computer Journal, 68(8):939– 952, August 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/939/8051621.

Wang:2022:PBB

[WZH^{+}22] Minxian Wang, Zijian Zhang, Jialing He, Feng Gao, Meng Li, Shubin Xu, and Liehuang Zhu. Practical blockchain-based steganographic communication via adversarial AI: a case study in bitcoin. The Computer Journal, 65(11):2926–2938, November 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2926/6741202.

Wang:2024:KRA

[WZJ^{+}24] Ke Wang, Zhenfeng Zhang, Haodong Jiang, Huiqin Xie, Yanjun Li, Ying Sun, and Lidong Han. Key reuse attacks on postquantum cryptosystems, revisited. The Computer Journal, 67(1):323– 337, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/323/6918741.

Wang:2024:NIB

[WZL^{+}24] Xiwen Wang, Kai Zhang, Jinguo Li, Mi Wen, Shengmin Xu, and Jianting Ning. Non-interactive Boolean searchable asymmetric encryption with bilateral access control. The Computer Journal, 67(1):179– 194, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/179/6845389.

Wang:2023:FCA

[WZMZ23] Ti Wang, Yongbin Zhou, Hui Ma, and Rui Zhang. Flexible and controllable access policy update for encrypted data sharing in the cloud. The Computer Journal, 66(6):1507–1524, June 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/6/1507/6555409.

Wang:2024:KKV

[WZPZ24] Chiheng Wang, Shibin Zhao, Jianshan Peng, and Junhu Zhu. KVFL: Key-value-based persistent fuzzing for IoT Web servers. The Computer Journal, 67(5):1892–1909, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1892/7456153.

Wang:2022:MPP

[WZQ^{+}22] Lei Wang, Qiang Zhang, Qing Qian, Jishuai Wang, Yujun Pan, Renbing Yang, and Wenbo Cheng. Multispectral palm print and palm vein acquisition platform and recognition method based on convolutional neural network. The Computer Journal, 65(6):1461– 1471, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/6/1461/6178420.

Wang:2023:VSS

[WZQ^{+}23] Qing Wang, Xi Zhang, Jing Qin, Jixin Ma, and Xinyi Huang. A verifiable symmetric searchable encryption scheme based on the AVL tree. The Computer Journal, 66(1):174– 183, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/174/6384776.

Wang:2020:ZKV

[WZXX20] Qiang Wang, Fucai Zhou, Jian Xu, and Zifeng Xu.

A (zero-knowledge) vector commitment with sum binding and its applications. The Computer Journal, 63(4):633–647, April 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/633/5627774.

Xu:2023:MMK

[X^{+}23] Wenju Xu et al. Modified multi-key fully homomorphic encryption scheme in the plain model. The Computer Journal, 66(10):2355– 2364, October 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2355/6627274.

Xu:2020:DRL

[XCC20] Z. Xu, L. Cao, and X. Chen. Deep reinforcement learning with adaptive update target combination. The Computer Journal, 63(7): 995–1003, July 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/7/995/5543068.

Xu:2024:AFI

[XCHY24] Xintong Xu, Run Cui, Chanying Huang, and Kedong Yan. Anonymization of face images with contrastive learning. The Computer Journal, 67(5):1910– 1919, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1910/7456152.

Xue:2023:FTS

[XDL23] Shudan Xue, Qingying Deng, and Pingshan Li. Fault-tolerant strongly Hamiltonian laceability and hyper-Hamiltonian laceability of Cayley graphs generated by transposition trees. The Computer Journal, 66(2):384–398, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/384/6398684.

Xu:2022:SRB

[XDZ22] Dong Xu, Wei Dong, and Han Zhou. Sclera recognition based on efficient sclera segmentation and significant vessel matching. The Computer Journal, 65(2):371–381, February 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/2/371/5848604.

Xu:2022:ACC

[XG22] Liqiong Xu and Litao Guo.

Analysis on the component connectivity of enhanced hypercubes. The Computer Journal, 65(4):890– 896, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/4/890/5921320.

Xu:2024:FFB

[XJWX24] Shenglin Xu, Zhiyuan Jiang, Yongjun Wang, and Peidai Xie. FormatAEG: a framework for bypassing ASLR defense and automated exploitation of format string vulnerability. The Computer Journal, 67(11):3056–3066, November 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/11/3056/7819323.

Xie:2024:SSO

[XLL24] Huawei Xie, Jing Liu, and Na Li. Supervisor synthesis for opacity enforcement in partially observed discrete event systems. The Computer Journal, 67(12):3151– 3159, December 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/12/3151/7739096.

Xia:2022:GEU

[XLLG22] Yifan Xia, Baosheng Liang, Zhaotong Li, and Song Gao. Gaze estimation using neural network and logistic regression. The Computer Journal, 65(8):2034–2043, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2034/6269131.

Xu:2022:OSC

[XLW22] Zheng Xu, Yongqiang Li, and Mingsheng Wang. Observations on the security of COMET. The Computer Journal, 65(9):2247–2261, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2247/6280578.

Xu:2024:LRC

[XLZ24] Leyou Xu, Chengli Li, and Bo Zhou. Link residual closeness of graphs with fixed parameters. The Computer Journal, 67(6): 2286–2302, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2286/7600383.

Xu:2023:NAR

[XQYA23] Zhiwang Xu, Huibin Qin, Shengying Yang, and Seyedeh Maryam Arefzadeh. A new approach for resource recommendation in the fogbased IoT using a hybrid algorithm. The Computer Journal, 66(3):692– 710, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/3/692/6455666.

Xiang:2024:IRI

[XTLM24] Nan Xiang, Xiao Tang, Huiling Liu, and Xiaoxia Ma. Identifying and ranking influential spreaders in complex networks by localized decreasing gravity model. The Computer Journal, 67(5):1727–1746, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1727/7292003.

Xu:2023:THE

[XTW^{+}23] Kexin Xu, Benjamin Hong Meng Tan, Li-Ping Wang, Khin Mi Mi Aung, and Huaxiong Wang. Threshold homomorphic encryption from provably secure NTRU. The Computer Journal, 66(12):2861–2873, December 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2861/6695522.

Xu:2024:TTI

[XWCZ24] Yifan Xu, Liangmin Wang, Jie Chen, and Qiang Zhou. Tor trace in images: a novel multi-tab Website fingerprinting attack with object detection. The Computer Journal, 67(8):2690–2701, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2690/7664604.

Xu:2023:MKF

[XWH^{+}23] Wenju Xu, Baocang Wang, Yupu Hu, Pu Duan, Benyu Zhang, and Momeng Liu. Multi-key fully homomorphic encryption from additive homomorphism. The Computer Journal, 66(1):197– 207, January 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/1/197/6408803.

Xu:2023:NSI

[XWL23] Yaqi Xu, Baofeng Wu, and Dongdai Lin. New strategies to improve differentiallinear attacks with applications to Chaskey. The Computer Journal, 66(9):2279– 2295, September 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2279/6611469.

Xu:2024:QBC

[XWQ24] Lidong Xu, Mingqiang Wang, and Jing Qin.

Quantum bit commitment without quantum memory. The Computer Journal, 67(3):1163–1170, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/1163/7156518.

Xu:2023:ILR

[XWW23] Shiyou Xu, Jian Wang, and Liangliang Wang. Indistinguishable leakage-resilient circuit compiler. The Computer Journal, 66(7):1717– 1732, July 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1717/6599079.

Xiang:2024:SRB

[XWY^{+}24] Yuyuan Xiang, Hongxia Wang, Ling Yang, Mingze He, and Fei Zhang. SEDD: Robust blind image watermarking with single encoder and dual decoders. The Computer Journal, 67(6):2390–2402, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2390/7603799.

Xue:2022:WSS

[XXM^{+}22] Luoyang Xue, Ang Xu, Qirong Mao, Lijian Gao, and Jie Chen. Weakly supervised sentiment-specific region discovery for VSA. The Computer Journal, 65 (4):818–830, April 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/65/4/818/
5900572.

Xu:2024:CPE

[XXW^{+}24] Qiang Xu, Dongmei Xu, Hao Wang, Jianye Yuan, and Zhe Wang. Color patterns and enhanced texture learning for detecting computer-generated images. The Computer Journal, 67(6):2303–2316, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2303/7603802.

Xu:2024:OOM

[XZL^{+}24] Zhixiong Xu, Wei Zhang, Ailin Li, Feifei Zhao, Yuanyuan Jing, Zheng Wan, Lei Cao, and Xiliang Chen. Online optimization method of learning process for meta-learning. The Computer Journal, 67(5): 1645–1651, May 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1645/7457340.

Xiong:2022:NIT

[XZLZ22] Fusong Xiong, Jian Zhang, Yun Ling, and Zhiqiang Zhang. A novel image thresholding method combining entropy with Parzen window estimation. The Computer Journal, 65(8):2231–2244, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/2231/6438112.

Xia:2025:ASQ

[XZXY25] Qiqing Xia, Qianru Zhu, Huiqin Xie, and Li Yang. Algorithm for solving quantum linear systems of equations with coherent superposition and extended applications. The Computer Journal, 68(5):520– 538, May 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/5/520/7921455.

Xu:2024:DGS

[XZYZ24] Yishu Xu, Peng Zhang, Hongtao Yu, and Fuzhi Zhang. Detecting group shilling attacks in recommender systems based on user multi-dimensional features and collusive behaviour analysis. The Computer Journal, 67(2):604– 616, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/604/7024503.

Xue:2024:RWI

[XZZ^{+}24] Yuliang Xue, Yuhao Zhu, Zhiying Zhu, Sheng Li, Zhenxing Qian, and Xinpeng Zhang. Removing watermarks for image processing networks via referenced subspace attention. The Computer Journal, 67(2):498–507, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/498/6965543.

Yang:2023:UPV

[Y^{+}23a] Runkai Yang et al. Understanding performance of a vulnerable heterogeneous edge data center: a modeling approach. The Computer Journal, 66(10):2339– 2354, October 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2339/6627273.

Yao:2023:LCA

[Y^{+}23b] Mianyang Yao et al. Loading cost-aware model caching and request routing in edgeenabled wireless sensor networks. The Computer Journal, 66(10):2409–2425, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/10/2409/6627276.

Yuan:2023:MHB

[Y^{+}23c] Junying Yuan et al. Multiple histograms-based reversible data hiding using fast performance optimization and adaptive pixel distribution. The Computer Journal, 66(11):2623–2637, November 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2623/6654515.

Yousefi-Azar:2020:BMR

[YAHVC20] Mahmood Yousefi-Azar, Len Hamey, Vijay Varadharajan, and Shiping Chen. Byte2vec: Malware representation and feature selection for Android. The Computer Journal, 63(8):1125– 1138, August 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1125/5618685.

Yang:2024:HHL

[Yan24]  Yuxing Yang.  Hyper-

Hamiltonian laceability of Cartesian products of cycles and paths. The Computer Journal, 67(2):548– 556, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/548/6995422.

Yang:2020:SLC

[YCL^{+}20] Zhichao Yang, Rongmao Chen, Chao Li, Longjiang Qu, and Guomin Yang. On the security of LWE cryptosystem against subversion attacks. The Computer Journal, 63(4):495– 507, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/4/495/5560305.

Yang:2022:ABP

[YCZ^{+}22] Yu Yang, Beifang Chen, Guoping Zhang, Yongming Li, Daoqiang Sun, and Hongbo Liu. Algorithms based on path contraction carrying weights for enumerating subtrees of tricyclic graphs. The Computer Journal, 65(3):554– 572, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/554/5872830.

Yan:2021:NMZ

[YD21] Zhenbin Yan and Yi Deng.

Non-malleable zero-knowledge arguments with lower round complexity. The Computer Journal, 64(4):534– 549, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/4/534/5869147.

Yang:2020:HIB

[YDS^{+}20] Zhichao Yang, Dung H. Duong, Willy Susilo, Guomin Yang, Chao Li, and Rongmao Chen. Hierarchical identity-based signature in polynomial rings. The Computer Journal, 63(10):1490–1499, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1490/5826091.

Yazdanjue:2020:EAK

[YFA20] Navid Yazdanjue, Mohammad Fathian, and Babak Amiri. Evolutionary algorithms for k-anonymity in social networks based on clustering approach. The Computer Journal, 63(7): 1039–1062, July 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/7/1039/5550898.

Yi:2022:ECD

[YFC^{+}22] Yi Yi, Jianxi Fan, Baolei Cheng, Yan Wang, and Jia Yu. The 3-extra connectivity of the data center network BCube. The Computer Journal, 65(12):3199– 3208, December 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/12/3199/6381256.

Yang:2023:DEI

[YFW^{+}23] Luming Yang, Shaojing Fu, Yongjun Wang, Kaitai Liang, Fan Mo, and Bo Liu. DEV-ETA: an interpretable detection framework for encrypted malicious traffic. The Computer Journal, 66(5):1213–1227, May 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/5/1213/6543491.

Yang:2024:VID

[YHW^{+}24] Zhongjun Yang, Qing Huang, Qi Wang, Xuejun Zong, and Ran Ao. Visual intrusion detection based on CBAM-capsule networks. The Computer Journal, 67(6):2357–2367, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2357/7603801.

Yang:2024:TAS

[YHX^{+}24] Runze Yang, Baoqi Huang, Zhendong Xu, Bing Jia, and Gang Xu. Towards accurate smartphone localization using CSI measurements. The Computer Journal, 67(4):1361–1369, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1361/7209528.

You:2024:PTF

[YJY^{+}24] Yizhe You, Zhengwei Jiang, Peian Yang, Jun Jiang, Kai Zhang, Xuren Wang, Chenpeng Tu, and Huamin Feng. P-TIMA: a framework of Twitter threat intelligence mining and analysis based on a prompt-learning NER model. The Computer Journal, 67(12):3221–3238, December 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/12/3221/7791002.

Yao:2024:CBC

[YLC^{+}24] Shixiong Yao, Pei Li, Jing Chen, Yuexing Zeng, Jiageng Chen, and Donghui Wang. CD-BCM: Crossdomain batch certificates management based on blockchain. The Computer Journal, 67(3):864– 874, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/864/7080911.

Yuan:2024:UWS

[YLG^{+}24] Xin Yuan, Ning Li, Xiaobo Gong, Changli Yu, Xiaoteng Zhou, and José-Fernán Martínez Ortega. Underwater wireless sensor network-based Delaunay triangulation (UWSN-DT) algorithm for sonar map fusion. The Computer Journal, 67(5):1699– 1709, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1699/7306808.

Yuan:2021:RBG

[YLW21] Jun Yuan, Aixia Liu, and Xi Wang. The relationship between the g-extra connectivity and the g-extra diagnosability of networks under the MM* model. The Computer Journal, 64(6): 921–928, June 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/921/6153485.

Yan:2025:DFA

[YLW^{+}25a] Kexiang Yan, Gang Liang, Lei Wang, Mingxu Sun, and Kui Zhao. Dualfeature adaptive framework for multimodal disinformation detection. The Computer Journal, 68(9):1105– 1117, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/9/1105/8100540.

Yuan:2025:IYU

[YLW^{+}25b] Xin Yuan, Jiapeng Li, Weiwei Wang, Xiaoteng Zhou, Ning Li, and Changli Yu. Improved YOLOv9 for underwater side scan sonar target detection. The Computer Journal, 68(6):591– 604, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/6/591/7921449.

Yu:2025:GPR

[YLWL25] Song Yu, Wenlong Liu, Hannan Wu, and Zhifang Liao. GitHub project recommendation based on knowledge graph and developer similarity. The Computer Journal, 68(9):1128– 1136, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/9/1128/8106766.

Yuan:2024:CFG

[YLX^{+}24] Zhu Yuan, Xueqiang Lv, Ping Xie, Haojie Ge, and Xindong You. CSEA: a fine-grained framework of climate-season-based energy-aware in cloud storage systems. The Computer Journal, 67(2):423– 436, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/423/6965540.

Yang:2024:SAM

[YLY24a] Weiwei Yang, Shangsong Liang, and Jian Yin. Spatial-aware multi-directional autoencoder for pre-training. The Computer Journal, 67 (4):1346–1360, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/67/4/1346/
7208680.

Yang:2024:DDE

[YLY^{+}24b] Yifan Yang, Xionglve Li, Tao Yang, Bingnan Hou, Lingbin Zeng, Zhiping Cai, and Wenyuan Kuang. DGA domain embedding with deep metric learning. The Computer Journal, 67(11):3094–3104, November 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/11/3094/7723330.

Yu:2020:VBG

[YLZ20] Bin Yu, Xiaofeng Li, and He Zhao. Virtual block group: a scalable blockchain model with partial node storage and distributed hash table. The Computer Journal, 63(10):1524–1536, October 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1524/5828295.

Yang:2024:GCS

[YLZ24] Yang Yang, Wenhao Liu, and Guang Zeng. A general condition of structural distinguisher. The Computer Journal, 67(3):923– 932, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/3/923/7086814.

Yan:2022:HSL

[YLZY22] Di Yan, Hanlin Liu, Shuoyao Zhao, and Yu Yu. On the hardness of sparsely learning parity with noise. The Computer Journal, 65(8):1939–1947, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/1939/6275472.

Yang:2024:RPK

[YMD^{+}24a] Tian Yang, Sha Ma, Jiaojiao Du, Chengyu Jiang, and Qiong Huang. Revocable public key encryption with equality test without pairing in cloud storage. The Computer Journal, 67(2):642–657, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/642/7135773.

Yousaf:2024:CEE

[YMD24b] Tameem Yousaf, Ali Afzal Malik, and Marriam Daud. Customization effort estimation of Microsoft ERP projects. The Computer Journal, 67(11):3105–3114, November 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/11/3105/7723331.

Yu:2022:SDF

[YNZ^{+}22] Ping Yu, Wei Ni, Hua Zhang, Ren Ping Liu, Qiaoyan Wen, Wenmin Li, and Fei Gao. Secure and differentiated fog-assisted data access for Internet of Things. The Computer Journal, 65(8):1948–1963, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/1948/6275473.

Yilmazer:2023:IIP

[YÖ23] Hakan Yilmazer and Selma Ayşe Özel. ImposeSVD: Incrementing PureSVD for top-N recommendations for cold-start problems and sparse datasets. The Computer Journal, 66(11):2595– 2622, November 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2595/6653218.

Yu:2024:IRK

[YQDJ24] Qingyuan Yu, Lingyue Qin, Xiaoyang Dong, and Keting Jia. Improved relatedkey rectangle attacks on GIFT. The Computer Journal, 67(4):1443–1456, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1443/7224396.

Yan:2024:SNC

[YQL24] Zhen Yan, Haipeng Qu, and Xi-Jun Lin. On the security of a novel construction of certificateless aggregate signature scheme for healthcare wireless medical sensor networks. The Computer Journal, 67(9):2819–2822, September 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/9/2819/7690089.

Youbi:2022:AML

[YS22] Fatiha Youbi and Nesma Settouti. Analysis of machine learning and deep learning frameworks for opinion mining on drug reviews. The Computer Journal, 65(9):2470–2483, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2470/6311550.

Yang:2022:EES

[YSH^{+}22] Runze Yang, Jian Song, Baoqi Huang, Wuyungerile Li, and Guodong Qi. An energy-efficient step-counting algorithm for smartphones. The Computer Journal, 65(3):689– 700, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/3/689/5880469.

Yan:2024:SFP

[YSTD24] Guofeng Yan, Zhengwen Su, Hengliang Tan, and Jiao Du. Service function placement optimization for cloud service with end-to-end delay constraints. The Computer Journal, 67(7): 2473–2485, July 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/7/2473/7618464.

Yu:2023:RIH

[YTL^{+}23] Mengzhu Yu, Zhenjun Tang, Zhixin Li, Xiaoping Liang, and Xianquan Zhang. Robust image hashing with saliency map and sparse model. The Computer Journal, 66(5):1241– 1255, May 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/5/1241/6548378.

Yang:2024:ACS

[YTXW24] Hongjuan Yang, Gang Tian, Chengrui Xu, and Rui Wang. Algorithms for cold-start game recommendation based on GNN pretraining model. The Computer Journal, 67(9):2787– 2798, September 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2787/7681804.

Yu:2024:VHT

[YTZ^{+}24] Mengzhu Yu, Zhenjun Tang, Hanyun Zhang, Xiaoping Liang, and Xianquan Zhang. Video hashing with tensor robust PCA and histogram of optical flow for copy detection. The Computer Journal, 67(6):

2162–2171, June 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2162/7503684.

Yan:2023:OSR

[YVCC23] Hailun Yan, Serge Vaudenay, Daniel Collins, and Andrea Caforio. Optimal symmetric ratcheting for secure communication. The Computer Journal, 66(4): 987–1016, April 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/987/6524371.

Yang:2020:PSC

[YWHY20] Wenjie Yang, Jian Weng, Xinyi Huang, and Anjia Yang. A provably secure certificateless proxy signature scheme against malicious-but-passive KGC attacks. The Computer Journal, 63(8):1139–1147, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1139/5667454.

Yu:2023:TTC

[YWTL23] Zhongkun Yu, JunFeng Wang, BinHui Tang, and Li Lu. Tactics and techniques classification in cyber threat intelligence. The Computer Journal, 66(8):1870–1881, August 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/1870/6571525.

Yu:2021:ACL

[YWY21] Jianxing Yu, Shiqi Wang, and Jian Yin. Adaptive cross-lingual question generation with minimal resources. The Computer Journal, 64(7):1056–1068, July 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/64/7/1056/6323599.

Yao:2024:PSC

[YWZB24] Zhicheng Yao, Sa Wang, Jinbin Zhu, and Yungang Bao. Panoptic segmentation with convex object representation. The Computer Journal, 67(6):2009– 2019, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/6/2009/7484595.

Yin:2022:ECE

[YX22] Shanshan Yin and Liqiong Xu. On the g-extra connectivity of the enhanced hypercubes. The Computer Journal, 65(9):2339–2346, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2339/6311551.

Yao:2023:OOA

[YX23] Jiahui Yao and Lingling Xu.

Online/offline attributebased Boolean keyword search for Internet of Things. The Computer Journal, 66(12):2948–2960, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/2948/6779420.

Yu:2024:REM

[YX24] Zhecheng Yu and Liqiong Xu. Reliability evaluation of multiprocessor system based on the balanced complete multipartite graphs. The Computer Journal, 67(2):688– 693, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/688/7074334.

Yao:2025:OHI

[YXLS25] Maoqun Yao, Ziwei Xue, Haiwei Li, and Shanhu Shen. An optimized hardware implementation of SHA-256 round computation. The Computer Journal, 68(4):355–359, April 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/4/355/7899894.

Yan:2023:LSE

[YXLZ23] Jianen Yan, Haiyan Xu, Ning Li, and Zhaoxin Zhang. Large-scale emulation network topology partition based on community detection with the weight of vertex similarity. The Computer Journal, 66(8):1817– 1828, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/1817/6568471.

Yin:2023:IEC

[YXZ23] Lingyuan Yin, Jing Xu, and Zhenfeng Zhang. Interopera: an efficient crosschain trading protocol. The Computer Journal, 66(7): 1609–1621, July 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1609/6565625.

Yu:2020:RCE

[YYL^{+}20] Hui Yu, Jiejie Yang, Limei Lin, Yanze Huang, Jine Li, and Riqing Chen. {1, 2, 3}-restricted connectivity of |(n, k)|-enhanced hypercubes. The Computer Journal, 63(9):1355–1371, September 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/9/1355/5614860.

Yuan:2021:SOS

[YYLX21] Zhu Yuan, Xindong You, Xueqiang Lv, and Ping Xie. SS6: Online short-code RAID-6 scaling by optimizing new disk location and data migration. The Computer Journal, 64(10):1600– 1616, October 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/10/1600/6372943.

Yang:2024:CEC

[YZC^{+}24] Xin Yang, Peng Zeng, Kim-Kwang Raymond Choo, Chengju Li, and Yanzhao Yang. CoinFA: an efficient coin mixing scheme with flexible amounts. The Computer Journal, 67(12):3141– 3150, December 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/12/3141/7740901.

Yang:2025:IFL

[YZLL25] Xinyan Yang, Rongchuan Zhang, Shao Li, and Gang Liang. Image forgery localization integrating multiscale and boundary features. The Computer Journal, 68(3):306–321, March 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/306/7901221.

Yang:2023:ADE

[YZMC23] Yayu Yang, Mingzu Zhang, Jixiang Meng, and Rongda Chen. The a-average degree edge-connectivity of bijective connection networks. The Computer Journal, 66(9):2118–2122, September 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2118/6606041.

Yuan:2024:RDH

[YZN24] Junying Yuan, Huicheng Zheng, and Jiangqun Ni. Reversible data hiding with pattern adaptive prediction. The Computer Journal, 67(4):1564–1571, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1564/7235533.

Yang:2025:IVI

[YZS25] Zhongjun Yang, Jixue Zhang, and Beimin Su. Internet of Vehicles intrusion detection method based on CFS-COA feature selection and spatio-temporal feature extraction. The Computer Journal, 68(5):487– 501, May 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/5/487/7918702.

Yan:2023:IBT

[YZW23] Kailun Yan, Jilian Zhang, and Yongdong Wu. Improving bitcoin transaction propagation efficiency through local clique network. The Computer Journal, 66(2):318–332, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/318/6455661.

Yuan:2024:UUL

[YZX^{+}24] Qingjun Yuan, Yuefei Zhu, Gang Xiong, Yongjuan Wang, Wentao Yu, Bin Lu, and Gaopeng Gou. ULDC: Unsupervised learning-based data cleaning for malicious traffic with high noise. The Computer Journal, 67(3): 976–987, March 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/976/7113989.

Zeng:2023:AFL

[Z^{+}23a] Yan Zeng et al. Adaptive federated learning with non-IID data. The Computer Journal, 66(11):2758– 2772, November 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/11/2758/6732678.

Zhou:2023:AOT

[Z^{+}23b] Wei Zhou et al. Antiocclusion target tracking based on joint confidence. The Computer Journal, 66(10):2462–2479, October 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/10/2462/6649278.

Zarei:2020:TCM

[Zar20]  Mani Zarei.  Traffic-

centric mesoscopic analysis of connectivity in VANETs. The Computer Journal, 63(2):203– 219, February 2020. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/2/203/5618855.

Zhang:2025:YFY

[ZBQ^{+}25] Xiaoyang Zhang, Jian Bai, Guanchao Qiao, Xiao Xiao, Liwei Meng, and Shaogang Hu. YOLO-fall: a YOLObased fall detection model with high precision, shrunk size, and low latency. The Computer Journal, 68(7):

804–812, July 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/7/804/7994663.

Zarrabi:2023:SPT

[ZC23a] Mohammad Reza Zarrabi and Nasrollah Moghaddam Charkari. Single-point and triple-point queries visibility constrained minimum link paths in simple polygons. The Computer Journal, 66(2):496–507, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/2/496/6518261.

Zhang:2023:ETA

[ZC23b] Jixin Zhang and Jiageng Chen. Efficient traceable attribute-based signature with update-free revocation for blockchain. The Computer Journal, 66(4):842– 865, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/842/6489103.

Zhao:2023:IME

[ZCC^{+}23] Zhiruo Zhao, Lei Cao, Xiliang Chen, Jun Lai, and Legui Zhang. Improvement of MADRL equilibrium based on Pareto optimization. The Computer Journal, 66(7):1573–1585, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/7/1573/6567701.

Zhong:2023:FCA

[ZCH23] Jincheng Zhong, Shuhui Chen, and Biao Han. FPGA-CPU architecture accelerated regular expression matching with fast preprocessing. The Computer Journal, 66(12):2928–2947, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2928/6770084.

Zhao:2024:LSM

[ZCW^{+}24] Shuang Zhao, Shuhui Chen, Fei Wang, Ziling Wei, Jincheng Zhong, and Jianbing Liang. A large-scale mobile traffic dataset for mobile application identification. The Computer Journal, 67(4):1501–1513, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1501/7235535.

Zhang:2024:IIV

[ZCZ^{+}24] Xiaolei Zhang, Zhaoyu Chen, Xin Zhang, Qingni Shen, and Zhonghai Wu.

IPOD2: an irrecoverable and verifiable deletion scheme for outsourced data. The Computer Journal, 67(10):2877–2889, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2877/7701810.

Zhang:2024:KAD

[ZeWSL24] Hanwen Zhang, Li e Wang, Zhigang Sun, and Xianxian Li. Knowledge-aware dual-channel graph neural networks for denoising recommendation. The Computer Journal, 67(5):1607– 1618, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1607/7241315.

Zhang:2025:ISA

[ZFY^{+}25] Hongbin Zhang, Ya Feng, Meng Yuan, Jingyi Hou, Jin Zhang, and Guangli Li. Image sentiment analysis based on distillation and sentiment region localization network. The Computer Journal, 68(6):577– 590, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/577/7925233.

Zhao:2024:SEM

[Zha24]  Xuefeng Zhao.  Secure

and efficient masking of lightweight ciphers in software and hardware. The Computer Journal, 67(2): 581–603, February 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/67/2/581/
                                        7066354.

Zhang:2025:UTF

[Zha25] Huihui Zhang. Unbiased training framework on deep reinforcement learning. The Computer Journal, 68(6): 649–662, June 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/6/649/7942795.

Zhang:2024:SSS

[ZHL^{+}24a] Qiang Zhang, Junjiang He, Tao Li, Xiaolong Lan, Wenbo Fang, and Yihong Li. SPAW-SMOTE: Space partitioning adaptive weighted synthetic minority oversampling technique for imbalanced data set learning. The Computer Journal, 67(5):1747– 1762, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1747/7292004.

Zhou:2024:ADC

[ZHL^{+}24b] Shaohong Zhou, Junjiang He, Tao Li, Xiaolong Lan, Yunpeng Wang, Hui Zhao, and Yihong Li. Automating the deployment of cyber range with OpenStack. The Computer Journal, 67(3): 851–863, March 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/851/7126356.

Zhao:2021:RAS

[ZHP21] Shu-Li Zhao, Rong-Xia Hao, and Sheng-Lung Peng. Reliability assessment of some regular networks. The Computer Journal, 64(1):1– 6, January 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/1/5611267.

Zhang:2025:IDB

[ZHS^{+}25] Fengjun Zhang, Lisheng Huang, Kai Shi, Shengjie Zhai, Yunhai Lan, and Qinghua Li. Intrusion detection based on hybrid metaheuristic feature selection. The Computer Journal, 68(1):13–22, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/13/7758618.

Zhang:2022:MVD

[ZHT22] Zhenghao Zhang, Jianbin Huang, and Qinglin Tan. Multi-view dynamic heterogeneous information network embedding. The Computer Journal, 65(8):2016– 2033, August 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/8/2016/6262255.

Zeng:2022:STE

[ZHYH22] Gongxian Zeng, Meiqi He, Siu Ming Yiu, and Zhengan Huang. A self-tallying electronic voting based on blockchain. The Computer Journal, 65(12):3020–3034, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/3020/6375072.

See corrigendum [ZHYH23].

Zeng:2023:CST

[ZHYH23] Gongxian Zeng, Meiqi He, Siu Ming Yiu, and Zhengan Huang. Corrigendum to: A self-tallying electronic voting based on blockchain. The Computer Journal, 66(2):523, February 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/

article/66/2/523/6398682.

See [ZHYH22].

Zheng:2023:IPF

[ZJ23] Wanwan Zheng and Mingzhe Jin. Improving the performance of feature selection methods with low-sample-size data. The Computer Journal, 66(7):1664–1686, July 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/7/1664/6565626.

Zhang:2025:GSI

[ZJLM25] Yue Zhang, Chen Jue, Wanxiao Liu, and Yurui Ma. GRAN: a SDN intrusion detection model based on graph attention network and residual learning. The Computer Journal, 68(3): 241–260, March 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/3/241/7839364.

Zhang:2022:NEL

[ZJZ^{+}22] Pingyuan Zhang, Han Jiang, Zhihua Zheng, Hao Wang, and Qiuliang Xu. A new and efficient latticebased online/offline signature from perspective of abort. The Computer Journal, 65(9):2400–2410, September 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/9/2400/6289888.

Zhang:2024:PVO

[ZKQH24] Liang Zhang, Haibin Kan, Feiyang Qiu, and Feng Hao. A publicly verifiable optimistic fair exchange protocol using decentralized CP-ABE. The Computer Journal, 67(3):1017–1029, March 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1017/7135811.

Zhao:2024:CEF

[ZL24a] Shijie Zhao and Pingshan Li. On conditional edge-fault-tolerant strong Menger edge connectivity of folded hypercubes. The Computer Journal, 67(2):777–781, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/777/7075451.

Zhao:2024:DCF

[ZL24b] WeiDong Zhao and Xiao-Tong Liu. Detection of e-commerce fraud review via self-paced graph contrast learning. The Computer Journal, 67(6):2054– 2065, June 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2054/7477780.

zhang:2025:DTE

[zLB^{+}25] Weikang zhang, Zhi Liu, Tongxin Bai, Furong Zheng, Wenming Jin, and Yang Wang. DuoSQL: towards elastic data warehousing via separated data management and processing. The Computer Journal, 68(8): 926–938, August 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/926/8045384.

Zhang:2022:CDH

[ZLC^{+}22] Shurong Zhang, Dongyue Liang, Lin Chen, Ronghua Li, and Weihua Yang. The component diagnosability of hypercubes with largescale faulty nodes. The Computer Journal, 65(5): 1129–1143, May 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/5/1129/6032262.

Zhang:2025:USM

[ZLC^{+}25] Cheng Zhang, Jinxin Lv, Jingxu Cao, Jiachuan Sheng, Dawei Song, and Tiancheng Zhang. Unravelling the semantic mysteries of transformers layer by layer. The Computer Journal, 68(9):1237–1251, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1237/8119372.

Zhang:2024:SRF

[ZLCW24] Yun Zhang, Yuling Liu, Ge Cheng, and Jie Wang. Similarity regression of functions in different compiled forms with neural attentions on dual controlflow graphs. The Computer Journal, 67(5):1710– 1718, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1710/7306809.

Zhou:2020:PKR

[ZLD^{+}20] Haibo Zhou, Zheng Li, Xiaoyang Dong, Keting Jia, and Willi Meier. Practical key-recovery attacks on round-reduced Ketje Jr, Xoodoo-AE and Xoodyak. The Computer Journal, 63(8):1231–1246, August 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/63/8/1231/5709729.

Zhang:2022:SEE

[ZLL^{+}22] Peng Zhang, Baoxi Liu, Tun Lu, Hansu Gu, Xianghua Ding, and Ning Gu. A semantic embedding enhanced topic model for user-generated textual content modeling in social ecosystems. The Computer Journal, 65(11):2953–2968, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2953/6741145.

See correction [Ano24a].

Zeng:2024:PVD

[ZLM^{+}24] Junhao Zeng, Gang Liang, Yixin Ma, Xinyan Yang, and Cheng Chen. Pornographic video detection based on semantic and image enhancement. The Computer Journal, 67 (10):3009–3019, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/3009/7721726.

Zu:2024:GGM

[ZLWZ24] Shuodi Zu, Xiangyang Luo, Liang Wang, and Fan Zhang. GUI: a geolocation method for unreachable IP. The Computer Journal, 67(5):1963–1978, May 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/5/1963/7462049.

Zhang:2024:HSC

[ZLX24] Yaofang Zhang, Peixuan Li, and Ping Xie. A hybrid scheme combining duplications and LDPC decoding to reduce NAND flash. The Computer Journal, 67(4): 1425–1437, April 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/4/1425/7224394.

Zhu:2022:MMV

[ZLZ22] Lin Zhu, Wen Ming Li, and Liang Feng Zhang. On the modulus in matching vector codes. The Computer Journal, 65(12):2991–2997, December 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/12/2991/6372942.

Zhang:2024:EOD

[ZLZ^{+}24] Zihang Zhang, Yuling Liu, Zhili Zhou, Gaobo Yang, and Q. M. Jonathan Wu. Efficient object detector via dynamic prior and dynamic feature fusion. The Computer Journal, 67(12):3196– 3206, December 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/12/3196/7758619.

Zhang:2022:ERV

[ZLZS22] Zhaodi Zhang, Jing Liu, Min Zhang, and Haiying Sun. Efficient robustness verification of the deep neural networks for smart IoT devices. The Computer Journal, 65(11):2894–2908, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2894/6741092.

Zhan:2025:ZDS

[ZLZZ25] Jing Zhan, Bo Li, Jiang Zhao, and Yong Zhao. zk-DCIAExchange: SGX protected fair exchange with distributed zero knowledge proof for data confidentiality and authentication. The Computer Journal, 68(8): 872–890, August 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/872/8037982.

Zhang:2021:RDB

[ZM21] Hong Zhang and Jixiang Meng. Reliability of DQcube based on g-extra conditional fault. The Computer Journal, 64(9):1393– 1400, September 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/9/1393/5856207.

Zhou:2025:DAD

[ZMC25] Qi Zhou, Xuechun Mao, and Ying Chen. A DDoS attack detection method combining federated learning and hybrid deep learning in software-defined networking. The Computer Journal, 68(10):1463–1475, October 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/10/1463/8122430.

Zheng:2020:SGT

[ZMWL20] Lixiao Zheng, Shuai Ma, Yuanyang Wang, and Gang Lin. String generation for testing regular expressions. The Computer Journal, 63(1):41–65, January 2020. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/1/41/5288328.

Zhang:2024:DBD

[ZNMZ24] Peng Zhang, Zhendong Niu, Ru Ma, and Fuzhi Zhang. DHCL-BR: Dual hypergraph contrastive learning for bundle recommendation. The Computer Journal, 67(10):2906–2919, October 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/10/2906/7731071.

Zhang:2023:HVS

[ZOLX23] Cheng Zhang, Bo Ou, Xiaolong Li, and Jianqin Xiong. Human visual system guided reversible data hiding based on multiple histograms modification. The Computer Journal, 66(4):888– 906, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/888/6503940.

Zou:2025:CRG

[Zou25] Yang Zou. A comparison of reserved graph grammar and edge-based graph grammar. The Computer Journal, 68(2):145– 162, February 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/145/7820625.

Zhang:2022:CLA

[ZQY^{+}22] Wenzheng Zhang, Zirui Qiao, Bo Yang, Yanwei Zhou, and Mingwu Zhang. Continuous leakageamplified public-key encryption with CCA security. The Computer Journal, 65(7):1760–1775, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1760/6236091.

Zhong:2024:SBA

[ZQZ24] Nan Zhong, Zhenxing Qian, and Xinpeng Zhang. Sparse backdoor attack against neural networks. The Computer Journal, 67(5):1783– 1793, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/5/1783/7292001.

Zhou:2025:YLL

[ZSH25] Weina Zhou, Wei Shao, and Wenhua Hu. A Yolo-like lightweight ship detection network for unmanned surface vehicles. The Computer Journal, 68(9):1118– 1127, September 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1118/8102313.

Zhai:2024:TMB

[ZSM24] Zhonghao Zhai, Subin Shen, and Yanqin Mao. A toolbox for migrating the blockchain-based application from Ethereum to hyperledger fabric. The Computer Journal, 67(4):1309– 1323, April 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1309/7199521.

Zhang:2024:SSD

[ZSQS24] Xi Zhang, Ye Su, Jing Qin, and Jiameng Sun. SDTA: Secure decentralized trading alliance for electronic medical data. The Computer Journal, 67(8):2573– 2585, August 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2573/7634134.

Zhu:2021:PPM

[ZTK21] Xiaofeng Zhu, Kim Han Thung, and Minjeong Kim. Privacy-preserving multimedia data analysis. The Computer Journal, 64(7): 991–992, July 2021. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/7/991/6356799.

Zhong:2024:LCD

[ZTT24] Yi Zhong, Zihan Teng, and Mengjun Tong. LiteMixer: Cauliflower disease diagnosis based on a novel lightweight neural network. The Computer Journal, 67(6):2346–2356, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2346/7596734.

Zhu:2024:VLB

[ZWC^{+}24] Chenguang Zhu, Renzheng Wei, Liwei Chen, Tongshuai Wu, Gewangzi Du, and Gang Shi. Vulnerability localization based on intermediate code representation and feature fusion. The Computer Journal, 67(9):2749–2762, September 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/9/2749/7670955.

Zhao:2020:BBA

[ZWG^{+}20] Zhen Zhao, Ge Wu, Fuchun Guo, Willy Susilo, Yi Mu, Baocang Wang, and Yupu Hu. Black-box accountable authority identity-based revocation system. The Computer Journal, 63(4):525– 535, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/4/525/5618849.

Zhang:2025:TBE

[ZWL^{+}25a] Yan Zhang, Liangliang Wang, Yiyuan Luo, Kai Zhang, Yu Long, and Kefei Chen. TBFL: blockchainenabled trusted byzantinerobust federated learning framework for photovoltaic power generation forecasting. The Computer Journal, 68(6):684– 696, June 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/6/684/7934588.

Zhang:2025:FFR

[ZWL^{+}25b] Yuheng Zhang, Guojun Wang, Peiqiang Li, Wanyi Gu, and Houji Chen. FRACE: Front-Running Attack Classification on Ethereum using ensemble learning. The Computer Journal, 68(9):1137–1149, September 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/9/1137/8108227.

Zhang:2024:EES

[ZWLP24] Dinghua Zhang, Shihao Wu, Yang Li, and Quan Pan. An evaluation on the entropy supplying capability of smartphone sensors. The Computer Journal, 67(4):1550–1563, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1550/7241316.

Zhang:2022:STR

[ZWM22] Xinpeng Zhang, Jigang Wu, and Min Meng. Small target recognition using dynamic time warping and visual attention. The Computer Journal, 65(2):203– 216, February 2022. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/2/203/5860020.

Zhou:2023:IBE

[ZWQ^{+}23] Yanwei Zhou, Zhaolong Wang, Zirui Qiao, Ying Wang, Bo Yang, Yi Mu, and Mingwu Zhang. Identitybased encryption with continuous leakage-resilient CCA security from static complexity assumption. The Computer Journal, 66 (4):924–940, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://academic.oup.com/
                                        comjnl/article/66/4/924/
                                        6573319.

Zou:2024:IDP

[ZWXL24] Yujuan Zou, Zhijian Wang, Pengfei Xu, and Taizhi Lv. An improved density peaks clustering algorithm based on density ratio. The Computer Journal, 67(7):2515– 2528, July 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/7/2515/7618087.

Zhang:2024:HCS

[ZWY^{+}24] Xu Zhang, Shunjie Wen, Liang Yan, Jiangfan Feng, and Ying Xia. A hybridconvolution spatial–temporal recurrent network for traffic flow prediction. The Computer Journal, 67(1):236– 252, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/1/236/6855103.

Zhou:2024:JAN

[ZWZ^{+}24] Qiang Zhou, Liangmin Wang, Huijuan Zhu, Tong Lu, and Heping Song. Joint alignment networks for fewshot Website fingerprinting attack. The Computer Journal, 67(6):2331–2345, June 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/6/2331/7596735.

Zhang:2023:CRR

[ZWZW23] Li Zhang, Wenling Wu, YaFei Zheng, and Bolin Wang. Cryptanalysis on reduced-round 3D and Saturnin. The Computer Journal, 66(4):1017–1029, April 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/1017/6702391.

Zhao:2024:SCS

[ZXDL24] Xiaojun Zhao, Shudan Xue, Qingying Deng, and Pingshan Li. Structure connectivity and substructure connectivity of Möbius cubes. The Computer Journal, 67(12):3207–3220, December 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/12/3207/7758129.

Zhu:2024:HTR

[ZXL^{+}24] Mingzhe Zhu, Wanyue Xu, Wei Li, Zhongzhi Zhang, and Haibin Kan. Hitting times of random walks on edge corona product graphs. The Computer Journal, 67(2):485–497, February 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/485/6975872.

See correction [Ano24d].

Zhu:2023:RDS

[ZXZ^{+}23] Mingzhe Zhu, Wanyue Xu, Zhongzhi Zhang, Haibin Kan, and Guanrong Chen. Resistance distances in simplicial networks. The Computer Journal, 66(8):1922– 1935, August 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/8/1922/6572654.

Zhang:2021:TES

[ZY21] Hailong Zhang and Wei Yang. Theoretical estimation on the success rate of the asymptotic higher order optimal distinguisher. The Computer Journal, 64(8):1277–1292, August 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/8/1277/6062487.

Zhang:2023:TAA

[ZY23] Hailong Zhang and Wei Yang. Template attack assisted linear cryptanalysis on outer rounds protected DES implementations. The Computer Journal, 66(6): 1434–1451, June 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/6/1434/6548382.

Zhu:2022:FSE

[ZYA^{+}22] Fei Zhu, Xun Yi, Alsharif Abuadbba, Ibrahim Khalil, Surya Nepal, and Xinyi Huang. Forwardsecure edge authentication for graphs. The Computer Journal, 65(7):1653–1665, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1653/6178962.

Zhang:2024:CCS

[ZYC^{+}24] Jiaomei Zhang, Ayong Ye, Jianwei Chen, Yuexin Zhang, and Wenjie Yang. CSFL: Cooperative security aware federated learning model using the blockchain. The Computer Journal, 67(4):1298–1308, April 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/4/1298/7203816.

Zhao:2023:AMP

[ZYG23] Nan Zhao, Nianmin Yao, and Zhenguo Gao. An adaptive MAC protocol based on time-domain interference alignment for UWANs. The Computer Journal, 66(12):3015–3028, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/12/3015/6770085.

Zeng:2025:CEA

[ZYL^{+}25] Lingbin Zeng, Shixin Yao, Xinwang Liu, Liquan Xiao, and Yue Qian. A clustering ensemble algorithm for handling deep embeddings using cluster confidence. The Computer Journal, 68(2):163–174, February 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/2/163/7820669.

Zhang:2025:RCR

[ZYLW25] Li-Han Zhang, Shu-Han Yang, Dan-Yang Lv, and Ji-Bo Wang. Research on convex resource allocation scheduling with exponential time-dependent learning effects. The Computer Journal, 68(1):97–108, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/97/7764031.

Zeng:2023:DAE

[ZYS^{+}23] Xiao Zeng, Guowu Yang, Xiaoyu Song, Marek A. Perkowski, and Gang Chen. Detecting affine equivalence of Boolean functions and circuit transformation. The Computer Journal, 66(9):2220–2229, September 2023. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/9/2220/6618067.

Zhou:2020:CLR

[ZYW^{+}20] Yanwei Zhou, Bo Yang, Tao Wang, Zhe Xia, and Hongxia Hou. Continuous leakage-resilient certificate-based encryption scheme without bilinear pairings. The Computer Journal, 63(4):508– 524, April 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/4/508/5614856.

Zhao:2025:RHR

[ZYW25] Wenyang Zhao, Xiaoyao Yang, and Yong Wang. RADNet: a highly and robust dynamic network for object detection in complex road scenes. The Computer Journal, 68(7):763– 774, July 2025. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/68/7/763/8025940.

Zhao:2023:LBR

[ZYWH23] Yong Zhao, Shaojun Yang, Wei Wu, and Xinyi Huang. A lattice-based redactable signature scheme using cryptographic accumulators for trees. The Computer Journal, 66(12):2961– 2973, December 2023. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/66/12/2961/6770083.

Zeng:2021:SHT

[ZZ21] Yibo Zeng and Zhongzhi Zhang. Spectra, hitting times and resistance distances of q-subdivision graphs. The Computer Journal, 64(1):76–92, January 2021. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/1/76/5670503.

Zhang:2023:NOS

[ZZ23] Zheng Zhang and Fangguo Zhang. New obfuscation scheme for conjunctions. The Computer Journal, 66(11):2773–2781, November 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/11/2773/6695273.

Zhu:2024:SFP

[ZZ24] Lina Zhu and Zuochang Zhang. Software failure prediction based on program state and first-error characteristics. The Computer Journal, 67(8):2559–2572, August 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/8/2559/7634132.

Zhang:2022:IBC

[ZZC^{+}22] Kai Zhang, Yuan Zhou, Zheng Chen, Yufei Liu, Zhuo Tang, Li Yin, and Jihong Chen. Incorporating biterm correlation knowledge into topic modeling for short texts. The Computer Journal, 65(3):537– 553, March 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/3/537/5868156.

Zhang:2024:CDH

[ZZC24a] Hong Zhang, Shuming Zhou, and Eddie Cheng. The cyclic diagnosability of hypercubes under the PMC model and the MM^{∗} model. The Computer Journal, 67(2):709– 718, February 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/2/709/7086913.

Zhou:2024:NAE

[ZZC24b] Chunning Zhou, Wentao Zhang, and Weiwei Cao. A new approach of evaluating the security against differential and linear cryptanalysis and its applications to Serpent, NOEKEON and ASCON. The Computer Journal, 67(1):274– 291, January 2024. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/67/1/274/6862009.

Zhou:2021:IAR

[ZZD^{+}21] Haibo Zhou, Rui Zong, Xiaoyang Dong, Keting Jia, and Willi Meier. Interpolation attacks on roundreduced Elephant, Kravatte and Xoofff. The Computer Journal, 64(4):628– 638, April 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/4/628/5880072.

Zhang:2025:WTS

[ZZF25] Siyuan Zhang, Yixuan Zhang, and Jianying Feng. Weighted twin support vector machine with rescaled hinge loss. The Computer Journal, 68(8):1013–1027, August 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/8/1013/8069422.

Zhou:2024:PPI

[ZZH24] Fucai Zhou, Zongye Zhang, and Ruiwei Hou. Privacypreserving image retrieval with multi-modal query. The Computer Journal, 67 (5):1979–1992, May 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/67/5/1979/
7462028.

Zhang:2020:WFM

[ZZJZ20] Jianhui Zhang, Tianhao Zhang, Feilong Jiang, and Bei Zhao. Who face me? Neighbor identification with LED camera communication. The Computer Journal, 63(10):1463–1478, October 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/10/1463/5821147.

Zhao:2020:MRS

[ZZL20] Jumin Zhao, Liang Zhang, and Deng’ao Li. Multiantenna receiver signal detection in AmBC based on cluster analysis. The Computer Journal, 63(6):844– 856, June 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/63/6/844/5691261.

Zeng:2022:PEF

[ZZL^{+}22] Qingkui Zeng, Liwen Zhou, Zhuotao Lian, Huakun Huang, and Jung Yoon Kim. Privacy-enhanced federated generative adversarial networks for Internet of Things. The Computer Journal, 65(11):2860–2869, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/65/11/2860/6580518.

Zhou:2023:CCS

[ZZLY23] Qianru Zhou, Shuming Zhou, Xiaoqing Liu, and Zhengqin Yu. Cluster connectivity and super cluster connectivity of DQcube. The Computer Journal, 66(4):826– 841, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/66/4/826/6479784.

Zhang:2022:RKM

[ZZW22] Xue Zhang, Zhongxiang Zheng, and Anyu Wang. A refinement of key mismatch attack on NewHope. The Computer Journal, 65(8):2209–2220, August 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/8/2209/6291064.

Zhang:2024:EUB

[ZZW^{+}24a] Li Zhang, Yu Zhang, Wenling Wu, Yongxia Mao, and Yafei Zheng. Explicit upper bound of impossible differentials for AES-like ciphers: Application to uBlock and Midori. The Computer Journal, 67(2):674–687, February 2024. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/2/674/7074317.

Zhang:2024:CEA

[ZZW^{+}24b] Zongyang Zhang, Lingyue Zhang, Zhuo Wang, Yichen Li, Rongxing Lu, and Yong Yu. Chronos: an efficient asynchronous Byzantine ordered consensus. The Computer Journal, 67(3):1153– 1162, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/67/3/1153/7140286.

Zhang:2022:EMI

[ZZZ^{+}22] Zhaoxi Zhang, Leo Yu Zhang, Xufei Zheng, Bilal Hussain Abbasi, and Shengshan Hu. Evaluating membership inference through adversarial robustness. The Computer Journal, 65(11):2969–2978, November 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/11/2969/6746755.

Zheng:2025:MFF

[ZZZL25] Juhong Zheng, Lei Zhang, Jingbo Zhang, and Meibo Lv. Maximum faultfree enforcement in Petri nets using supervisory control. The Computer Journal, 68(1):74–83, January 2025. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/68/1/74/7758867.